### v2

# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

Abstract
We present masked Lagrange reconstruction, a technique that enables threshold ML-DSA (FIPS 204) with arbitrary thresholds $T$ while producing standard 3.3 KB signatures verifiable by unmodified FIPS 204 implementations. Concurrent approaches have limitations: Bienstock et al. (ePrint 2025/1163) achieve arbitrary $T$ but require honest-majority and 37–136 rounds; Celi et al. (ePrint 2026/013) achieve dishonest-majority but are limited to $T\leq 6$. Our technique addresses the barrier that Lagrange coefficients grow as $\Theta(q)$ for moderate $T$, making individual contributions too large for ML-DSA’s rejection sampling.
Unlike ECDSA threshold schemes where pairwise masks suffice for correctness, ML-DSA requires solving three additional challenges absent in prior work: (1) rejection sampling on $\|\mathbf{z}\|_{\infty}$ must still pass after masking, (2) the r0-check exposes $c\mathbf{s}_{2}$ enabling key recovery if unprotected, and (3) the resulting Irwin-Hall nonce distribution must preserve EUF-CMA security. We solve all three.
We instantiate this technique in three deployment profiles with full security proofs. Profile P1 (TEE-assisted) achieves 3-round signing with a trusted coordinator, with EUF-CMA security under Module-SIS (Theorem 4.2). Profile P2 (fully distributed) eliminates hardware trust via MPC in 8 rounds, achieving UC security against malicious adversaries corrupting up to n1n-1 parties (Theorem 6.2). Profile P3 (2PC-assisted) uses lightweight 2PC for the r0-check in 3–5 rounds, achieving UC security under a 1-of-2 CP honest assumption (Theorem 6.5) with the best empirical performance (249ms).
Our scheme requires $|S|\geq T+1$ signers and achieves success rates of 23–32%, matching single-signer ML-DSA.
keywords:
threshold signatures, ML-DSA, post-quantum cryptography, lattice-based cryptography, FIPS 204
1 Introduction
The imminent arrival of cryptographically relevant quantum computers poses an existential threat to currently deployed public-key cryptography. In response, the National Institute of Standards and Technology (NIST) has completed a multi-year standardization process, culminating in the release of FIPS 204 [FIPS204], which specifies ML-DSA (Module-Lattice-Based Digital Signature Algorithm) as the primary post-quantum signature standard. As organizations worldwide begin transitioning to ML-DSA, there is an urgent need for threshold variants that enable distributed signing while maintaining compatibility with the standard.
1.1 Motivation: The Threshold Gap
Threshold signatures allow a group of $N$ parties to collectively sign messages such that any subset of $T$ or more parties can produce a valid signature, while any coalition of fewer than $T$ parties learns nothing about the secret key. This primitive is essential for protecting keys by distributing signing authority across multiple servers, enforcing multi-party authorization in financial systems, and building robust PKI infrastructure through distributed certificate authorities.
For classical ECDSA and Schnorr signatures, efficient threshold protocols have been extensively studied and deployed [GG18, CGGMP20, Lindell21, KG20]. More recently, Lyu et al. [LLZD25] achieved two-round threshold ECDSA in the threshold-optimal multi-party setting. However, the lattice-based structure of ML-DSA presents new challenges. The core difficulty lies in the Fiat-Shamir with Aborts paradigm: signatures are only valid when the response vector $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ satisfies $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$, where $\mathbf{y}$ is a masking vector, $c$ is a sparse challenge, and $\mathbf{s}_{1}$ is the secret key. This rejection sampling step complicates the threshold setting.
The State of Affairs.
Current approaches to threshold ML-DSA can be categorized as follows:
  1. 1.
    Short-coefficient secret sharing [dPN25]: Uses carefully constructed linear secret sharing schemes (LSSS) where reconstruction coefficients remain small. This approach produces standard-size signatures and has been demonstrated for $T\leq 8$. The Ball-Çakan-Malkin bound [BCM21] shows that $\{0,1\}$-LSSS for threshold $T$ requires share size $\Omega(T\log T)$, which imposes practical limits on scaling this approach.
  2. 2.
    Noise flooding [BCDG21]: Masks the large Lagrange coefficients with overwhelming statistical noise. This works for arbitrary thresholds but produces signatures approximately $5\times$ larger than standard ML-DSA, breaking FIPS 204 compatibility.
  3. 3.
    Specialized constructions [BMMSZ25]: Ringtail and similar schemes modify the underlying signature algorithm to be threshold-friendly, achieving good efficiency but producing signatures incompatible with the FIPS 204 standard.
This leaves a gap: for $T\in[9,32]$, there is no known scheme that simultaneously achieves standard signature size, practical success rates, and FIPS 204 compatibility. This range is precisely where many real-world applications fall—organizations typically require 16-of-32 or similar thresholds for robust multi-party control.
1.2 Our Contribution
We present Masked Threshold ML-DSA, a threshold signature scheme for ML-DSA that supports arbitrary thresholds $T\leq N$ while producing standard-size signatures ($\approx 3.3$ KB for ML-DSA-65) identical to single-signer ML-DSA. Our scheme requires a two-honest requirement: signing sets must have size $|S|\geq T+1$ to ensure at least two honest parties for mask privacy. This addresses the threshold gap for $T\in[9,32]$ while remaining practical—configurations like $(3,5)$, $(11,15)$, and $(16,31)$ naturally satisfy $N\geq T+1$.
The scheme achieves practical success rates comparable to single-signer ML-DSA (theoretical baseline $\approx 20\%$, observed $3232\%$ due to the more concentrated Irwin-Hall nonce distribution) and is fully verifier-compatible with FIPS 204: produced signatures are accepted by unmodified FIPS 204 verifiers. While our signer-side protocol is interactive and induces an Irwin-Hall nonce distribution rather than uniform, this affects only the signing process, not verification.
Prior approaches attempt to ensure each party’s contribution is individually small, but this hits a barrier: Lagrange coefficients $\lambda_{i}\approx 2^{T}$ grow exponentially, reaching $\Theta(q)$ for $T\geq 23$ in ML-DSA. If party $i$ naively computes $\lambda_{i}\cdot\mathbf{z}_{i}$ where $\|\mathbf{z}_{i}\|_{\infty}=O(\gamma_{1})$, the result exceeds the rejection threshold for any $\lambda_{i}>1$.
Our approach instead ensures that only the sum needs to be small, not individual contributions. While pairwise-canceling masks exist for ECDSA [CGGMP20], directly applying them to lattice signatures does not work because: (1) ML-DSA requires rejection sampling on $\|\mathbf{z}\|_{\infty}$, not just correctness; (2) the r0-check leaks $c\mathbf{s}_{2}$ which enables key recovery; (3) the Irwin-Hall nonce distribution must preserve EUF-CMA security. Our masked Lagrange reconstruction addresses these challenges:
  1. 1.
    Each pair of parties $(i,j)$ establishes a shared seed $\mathsf{seed}_{i,j}$ (e.g., via ML-KEM).
  2. 2.
    During signing, party $i$ computes a mask:
    $\mathbf{m}_{i}=\sum_{j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{nonce})-\sum_{j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{nonce})$
  3. 3.
    By construction, $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$ for any signing set SS—the masks pairwise cancel.
  4. 4.
    Party $i$ sends $\mathbf{U}_{i}=\mathbf{y}_{i}+\lambda_{i}\cdot c\cdot\mathbf{s}_{1,i}+\mathbf{m}_{i}$ instead of the unmasked value.
While each $\mathbf{U}_{i}$ may have arbitrary norm in [0,q1][0,q-1], their sum is:
$\sum_{i\in S}\mathbf{U}_{i}=\sum_{i}\mathbf{y}_{i}+c\cdot\sum_{i}\lambda_{i}\mathbf{s}_{1,i}+\sum_{i}\mathbf{m}_{i}=\mathbf{y}+c\cdot\mathbf{s}_{1}+\mathbf{0}=\mathbf{z}$
which correctly reconstructs the signature vector $\mathbf{z}$ and passes rejection sampling with probability matching single-signer ML-DSA (theoretical $\approx 20\%$, observed $3232\%$ for ML-DSA-65). We note that while the aggregated nonce $\mathbf{y}=\sum_{i}\mathbf{y}_{i}$ follows an Irwin-Hall distribution rather than uniform (see Definition O.5), this does not affect EUF-CMA security and only introduces a quantifiable statistical difference that we analyze in Section 3.
Performance.
Our implementation achieves practical signing times: approximately 1 second average for $(3,5)$-threshold and 4 seconds for $(16,31)$-threshold on commodity hardware (including retries due to rejection sampling), using NTT-accelerated polynomial arithmetic. The communication cost is approximately 12.3 KB per party per signing attempt.
Security.
We prove security in the random oracle model. Unforgeability reduces to the Module-SIS problem, matching the security of single-signer ML-DSA, while privacy against coalitions of up to T1T-1 corrupted parties reduces to Module-LWE and PRF security.
Deployment Profiles.
We define three deployment profiles with different trust assumptions, all with complete security proofs:
Profile P1 (TEE-Assisted):
A TEE/HSM coordinator performs the r0-check and hint generation inside a trusted enclave, achieving optimal 3-round signing. We prove EUF-CMA security under Module-SIS and privacy under Module-LWE (Theorem 4.2).
Profile P2 (Fully Distributed):
A fully untrusted coordinator where the r0-check is computed via an MPC subprotocol using SPDZ [DPSZ12] and edaBits [EKMOZ20]. We prove UC security against malicious adversaries corrupting up to n1n-1 parties in 8 online rounds (Theorem 6.2).
Profile P3 (2PC-Assisted):
A practical middle ground where two designated Computation Parties (CPs) jointly evaluate the r0-check via lightweight 2PC. Motivated by recent round-reduction efforts in threshold ECDSA [LLZD25], we pursue a similar goal for the lattice setting; while their Non-Interactive Multiplication technique does not directly apply to ML-DSA’s r0-check structure, we achieve analogous round improvements via 2PC. We prove UC security under the assumption that at least one CP is honest (Theorem 6.5)—weaker than TEE trust, more efficient than full MPC. Achieves the best empirical performance with 3–5 rounds and 249ms signing time.
1.3 Technical Overview
Protocol Overview.
A trusted dealer creates Shamir shares of $\mathsf{sk}=(\mathbf{s}_{1},\mathbf{s}_{2})$ and establishes pairwise seeds via ML-KEM. The 3-round signing protocol has parties (1) commit to nonces $\mathbf{y}_{i}\in[-\gamma_{1}/|S|,\gamma_{1}/|S|]$, (2) exchange masked commitments to derive challenge $c$, and (3) send masked responses $\mathbf{U}_{i}=\mathbf{y}_{i}+\lambda_{i}\cdot c\cdot\mathbf{s}_{1,i}+\mathbf{m}_{i}$. The combiner aggregates $\mathbf{z}=\sum_{i}\mathbf{U}_{i}=\mathbf{y}+c\mathbf{s}_{1}$ (masks cancel) and outputs $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ if $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$.
1.4 Comparison with Prior and Concurrent Work
Table 1 summarizes threshold ML-DSA schemes, comparing arbitrary threshold support, dishonest-majority security, round complexity, and FIPS 204 compatibility.
Scheme $T$ Size Rnd FIPS Dishon. UC
del Pino-Niot [dPN25] $\leq 8$ 3.3 KB 3 $\times$
Noise Flooding [BCDG21] $\infty$ 17 KB 3 $\times$ $\times$
Ringtail [BMMSZ25] $\infty$ 13 KB 2 $\times$ $\times$
Concurrent Work (2025–2026)
Bienstock et al. [BDEPT25] $\infty$ 3.3 KB 37–136 $\times$
Celi et al. [CDENP26] $\leq 6$ 3.3 KB 6 $\times$
Trilithium [Trilithium25] 2 3.3 KB 3 $\times$
This work (P1) $\infty$ 3.3 KB 3 (TEE) $\times$
This work (P2) $\infty$ 3.3 KB 8
This work (P3) $\infty$ 3.3 KB 3–5 (1/2 CP)
Table 1: Comparison of threshold ML-DSA schemes. “Dishon.” = dishonest-majority security; “(TEE)” and “(1/2 CP)” indicate trust assumptions.
Three concurrent works address threshold ML-DSA with FIPS 204 compatibility. Bienstock et al. [BDEPT25] achieve arbitrary $T$ with UC security, but require honest-majority and 37–136 communication rounds. Celi et al. [CDENP26] achieve dishonest-majority with 6 rounds, but are limited to $T\leq 6$ due to the Ball-Çakan-Malkin bound [BCM21] on short-coefficient LSSS. Trilithium [Trilithium25] is UC-secure but restricted to 2 parties in a server-phone architecture.
Profile P2 combines arbitrary thresholds, dishonest-majority security, constant rounds (8 rounds), UC security, and FIPS 204 compatibility—a combination not achieved by prior work. Profile P3 also achieves UC security (Theorem 6.5) under the weaker 1-of-2 CP honest assumption, with the best performance (249ms, 3–5 rounds).
We contribute: (1) Masked Lagrange reconstruction, adapting pairwise-canceling masks [CGGMP20] to lattices while addressing rejection sampling, r0-check key leakage, and Irwin-Hall nonce security; (2) Three deployment profiles (P1-TEE, P2-MPC, P3-2PC) with different trust/performance trade-offs; (3) Security proofs reducing to Module-SIS/LWE, including UC security for P2 and P3; (4) A reference implementation with benchmarks from $(3,5)$ to $(16,31)$ threshold configurations.
1.5 Paper Organization
Section 2 provides background; Sections 34 present the protocol and security analysis; Section 5 gives benchmarks; Sections 67 discuss extensions and conclude.
2 Preliminaries
2.1 Notation
We use bold lowercase letters (e.g., $\mathbf{s}$) for vectors and bold uppercase letters (e.g., $\mathbf{A}$) for matrices. For an integer $q$, we write $\mathbb{Z}_{q}=\mathbb{Z}/q\mathbb{Z}$ for the ring of integers modulo $q$. We define the polynomial ring $R_{q}=\mathbb{Z}_{q}[X]/(X^{n}+1)$ where $n$ is a power of 2.
For a polynomial $f\in R_{q}$, we write $\left\|f\right\|_{\infty}=\max_{i}|f_{i}|$ for its infinity norm, where coefficients are represented in (q/2,q/2](-q/2,q/2]. For a vector $\mathbf{v}=(v_{1},\ldots,v_{k})\in R_{q}^{k}$, we define $\left\|\mathbf{v}\right\|_{\infty}=\max_{i}\left\|v_{i}\right\|_{\infty}$.
We write $a\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}S$ to denote sampling $a$ uniformly at random from a set $S$. We write $\mathsf{negl}(\kappa)$ for a negligible function in the security parameter $\kappa$.
2.2 Lattice Problems
Our security relies on the hardness of the Module-SIS and Module-LWE problems [Regev09, LPR10].
Definition 2.1 (Module-SIS).
The $\mathsf{M\text{-}SIS}_{n,k,\ell,q,\beta}$ problem is: given a uniformly random matrix $\mathbf{A}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k\times\ell}$, find a non-zero vector $\mathbf{z}\in R_{q}^{\ell}$ such that $\mathbf{A}\mathbf{z}=\mathbf{0}\mod q$ and $\left\|\mathbf{z}\right\|_{\infty}\leq\beta$.
Definition 2.2 (Module-LWE).
The $\mathsf{M\text{-}LWE}_{n,k,\ell,q,\chi}$ problem is to distinguish between:
  • $(\mathbf{A},\mathbf{A}\mathbf{s}+\mathbf{e})$ where $\mathbf{A}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k\times\ell}$, $\mathbf{s}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\chi^{\ell}$, $\mathbf{e}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\chi^{k}$
  • $(\mathbf{A},\mathbf{u})$ where $\mathbf{u}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k}$
where $\chi$ is a distribution over $R_{q}$ with small coefficients.
For ML-DSA-65 parameters ($n=256$, $q=8380417$, $k=6$, $\ell=5$), both problems are believed to be hard at the 192-bit security level.
2.3 ML-DSA (FIPS 204)
ML-DSA [FIPS204], based on the CRYSTALS-Dilithium design [DKL18], is a lattice-based digital signature scheme using the Fiat-Shamir with Aborts paradigm [Lyu12, PS00]. Security in the quantum random oracle model follows from [KLS18]. We summarize the key algorithms.
Parameters.
For ML-DSA-65 (NIST Security Level 3):
  • $n=256$: polynomial degree
  • $q=8380417=2^{23}-2^{13}+1$: modulus
  • $k=6$, $\ell=5$: matrix dimensions
  • $\eta=4$: secret key coefficient bound
  • $\gamma_{1}=2^{19}$: masking range
  • $\gamma_{2}=(q-1)/32$: rounding parameter
  • $\tau=49$: challenge weight (number of $\pm 1$ entries)
  • $\beta=\tau\cdot\eta=196$: signature bound
Key Generation.
  1. 1.
    Sample seed $\rho\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$ and expand to matrix $\mathbf{A}\in R_{q}^{k\times\ell}$
  2. 2.
    Sample short secrets $\mathbf{s}_{1}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\chi_{\eta}^{\ell}$ and $\mathbf{s}_{2}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\chi_{\eta}^{k}$ with coefficients in $\eta,\eta]$
  3. 3.
    Compute $\mathbf{t}=\mathbf{A}\mathbf{s}_{1}+\mathbf{s}_{2}$ and decompose $\mathbf{t}=\mathbf{t}_{1}\cdot 2^{d}+\mathbf{t}_{0}$
  4. 4.
    Output $\mathsf{pk}=(\rho,\mathbf{t}_{1})$ and $\mathsf{sk}=(\rho,\mathbf{s}_{1},\mathbf{s}_{2},\mathbf{t}_{0})$
Signing.
Sample nonce $\mathbf{y}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1}+1,\ldots,\gamma_{1}\}^{n\ell}$, compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{A}\mathbf{y},2\gamma_{2})$, derive challenge $c=H(\mu\|\mathbf{w}_{1})$, and set $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$. If $\left\|\mathbf{z}\right\|_{\infty}\geq\gamma_{1}-\beta$, restart (rejection sampling). Compute hint $\mathbf{h}$ and output $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$.
Verification.
Expand $c$ from $\tilde{c}$, compute $\mathbf{w}_{1}^{\prime}=\mathsf{UseHint}(\mathbf{h},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d})$, and verify $\left\|\mathbf{z}\right\|_{\infty}<\gamma_{1}-\beta$ and $\tilde{c}=H(\mu\|\mathbf{w}_{1}^{\prime})$.
The z-bound check and r0-check together yield $\approx 20$$25\%$ per-attempt success (4–5 expected attempts). This is inherent to Fiat-Shamir with Aborts [Lyu12], not threshold-specific.
2.4 Shamir Secret Sharing
Shamir’s $(T,N)$-threshold secret sharing [Shamir79] allows a secret $s\in\mathbb{Z}_{q}$ to be distributed among $N$ parties such that any $T$ parties can reconstruct $s$, but any T1T-1 parties learn nothing.
Sharing.
To share secret $s$:
  1. 1.
    Sample random polynomial $f(X)=s+r_{1}X+\cdots+r_{T-1}X^{T-1}\in\mathbb{Z}_{q}[X]$
  2. 2.
    Give share $s_{i}=f(i)$ to party $i$ for $i\in[N]$
Reconstruction.
Given shares from set $S$ with $|S|\geq T$:
$s=f(0)=\sum_{i\in S}\lambda_{i}\cdot s_{i}\mod q$
where the Lagrange coefficients are:
$\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\mod q$
Remark 2.3 (Lagrange Coefficient Magnitude).
Lagrange coefficients $\lambda_{i}=(-1)^{i-1}\binom{T}{i}$ grow as $\approx 2^{T}/\sqrt{T}$. For $T\leq 8$, coefficients remain small ($\binom{8}{4}=70$), enabling short-coefficient schemes [dPN25]. For $T\geq 24$, coefficients exceed $q\approx 2^{23}$, aligning with the Ball-Çakan-Malkin bound [BCM21]. The obstacle is that $\lambda_{i}\cdot\mathbf{z}_{i}$ has norm $O(q\cdot\gamma_{1})\gg\gamma_{1}-\beta$, causing rejection.
2.5 Security Definitions for Threshold Signatures
Definition 2.4 (Threshold Signature Scheme).
A $(T,N)$-threshold signature scheme consists of algorithms:
  • $\mathsf{KeyGen}(1^{\kappa},N,T)\to(\mathsf{pk},\{\mathsf{sk}_{i}\}_{i\in[N]})$: Generates public key and secret key shares
  • $\mathsf{Sign}(\{\mathsf{sk}_{i}\}_{i\in S},\mu)\to\sigma$: Interactive protocol among parties in $S$ (where $|S|\geq T+1$; see Section 4 for the two-honest requirement) to sign message $\mu$
  • $\mathsf{Verify}(\mathsf{pk},\mu,\sigma)\to\{0,1\}$: Deterministic verification
Remark 2.5 ($T+1$ Signing Requirement).
Our protocol requires $|S|\geq T+1$ signers (not $T$) to ensure at least two honest parties for mask privacy. This is standard for pairwise-mask schemes [CGGMP20, DKOSS24]; see Remark 4.1.
Definition 2.6 (Unforgeability (EUF-CMA)).
A threshold signature scheme is EUF-CMA secure if for any PPT adversary $\mathcal{A}$ corrupting up to T1T-1 parties:
$\Pr[\mathsf{EUF\text{-}CMA}_{\Pi}^{\mathcal{A}}(\kappa)=1]\leq\mathsf{negl}(\kappa)$
where in the security game:
  1. 1.
    $(\mathsf{pk},\{\mathsf{sk}_{i}\})\leftarrow\mathsf{KeyGen}(1^{\kappa},N,T)$
  2. 2.
    $\mathcal{A}$ receives $\mathsf{pk}$ and $\{\mathsf{sk}_{i}\}_{i\in C}$ for corrupted set $C$ with $|C|<T$
  3. 3.
    $\mathcal{A}$ can request signatures on messages of its choice (interacting with honest parties)
  4. 4.
    $\mathcal{A}$ outputs $(m^{*},\sigma^{*})$
  5. 5.
    $\mathcal{A}$ wins if $\mathsf{Verify}(\mathsf{pk},m^{*},\sigma^{*})=1$ and $m^{*}$ was never queried
Definition 2.7 (Privacy).
A threshold signature scheme provides privacy if the view of any coalition $C$ with $|C|<T$ can be efficiently simulated given only $\mathsf{pk}$ and the signatures produced.
2.6 Pseudorandom Functions
Definition 2.8 (PRF Security).
A function family $F:\mathcal{K}\times\mathcal{X}\to\mathcal{Y}$ is a secure PRF if for any PPT distinguisher $\mathcal{D}$:
$\left|\Pr[\mathcal{D}^{F_{k}(\cdot)}=1]-\Pr[\mathcal{D}^{R(\cdot)}=1]\right|\leq\mathsf{negl}(\kappa)$
where $k\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathcal{K}$ and $R:\mathcal{X}\to\mathcal{Y}$ is a truly random function.
In our construction, we instantiate the PRF with SHAKE-256, modeled as a random oracle in the security analysis.
3 Masked Threshold ML-DSA Protocol
We now present our masked threshold ML-DSA protocol in full detail. The protocol consists of a setup phase (key generation) and a signing phase (three-round interactive protocol).
3.1 Overview
The core challenge in threshold ML-DSA is handling the large Lagrange coefficients that arise in Shamir secret sharing over $\mathbb{Z}_{q}$. As noted in Remark 2.3, the Lagrange coefficients grow exponentially as $\approx 2^{T}$, exceeding $q$ for $T\geq\log_{2}q\approx 23$. Even for moderate $T$ (e.g., $T=16$), the coefficients $\binom{16}{8}=12870$ can cause the naive approach to inflate norms beyond the rejection threshold.
Our solution is based on two key observations:
  1. 1.
    Masking the coefficients: We add pairwise-canceling masks $\mathbf{m}_{i}$ to each party’s contribution. While individual contributions $\mathbf{U}_{i}$ may have arbitrary magnitude, the masks cancel in the sum, leaving the correct short result.
  2. 2.
    Separating nonce and secret: The nonce $\mathbf{y}_{i}$ and secret share $\mathbf{s}_{1,i}$ must be treated differently. Nonces are summed directly ($\mathbf{y}=\sum_{i}\mathbf{y}_{i}$), while secret shares are reconstructed via Lagrange interpolation ($\mathbf{s}_{1}=\sum_{i}\lambda_{i}\mathbf{s}_{1,i}$).
3.2 Threat Model and Deployment Profiles
We consider a $(T,N)$-threshold setting where an adversary may corrupt up to T1T-1 parties. Table 2 summarizes our two deployment profiles.
Table 2: Deployment Profiles: Trust Assumptions and Security Guarantees
Property Profile P1 Profile P2 Profile P3
Coordinator TEE/HSM Untrusted 2PC pair
Trust assumption TEE integrity None (MPC) 1-of-2 CP honest
r0-check location Inside enclave MPC subprotocol 2PC subprotocol
$c\mathbf{s}_{2}$ exposure Never leaves TEE Never reconstructed Split between CPs
Online rounds 3 8 3–5
Latency Low High Lowest
Complexity Low High Medium
Common to all profiles:
Adversary model Static, malicious (up to T1T-1 parties)
Signing set size $|S|\geq T+1$ (two-honest requirement)
Security goal EUF-CMA under Module-SIS
Privacy goal T1T-1 privacy under Module-LWE
Output format FIPS 204-compatible signature
Profile P1 (Primary).
A TEE/HSM coordinator aggregates masked contributions and performs the r0-check inside a trusted enclave. The value $c\mathbf{s}_{2}$ is reconstructed inside the enclave and never exposed. This is our recommended deployment model, combining threshold distribution among parties with hardware-isolated aggregation.
Profile P2 (Fully Distributed).
A fully distributed protocol where no single party learns $c\mathbf{s}_{2}$. The r0-check is computed via an MPC subprotocol using SPDZ [DPSZ12] and edaBits [EKMOZ20] (see Section 6). This provides stronger distributed trust at the cost of additional rounds (8 online rounds). We prove UC security against malicious adversaries corrupting up to n1n-1 parties in Theorem 6.2.
Profile P3 (2PC-Assisted).
A hybrid approach that designates two Computation Parties (CPs) to jointly evaluate the r0-check via lightweight 2PC. Each CP receives additive shares of $c\mathbf{s}_{2}$; neither learns the complete value unless they collude. This achieves the best empirical performance (see Table 5) while requiring only that at least one of the two CPs is honest—a weaker assumption than TEE trust and more efficient than full MPC. The 2PC uses standard garbled circuits [Yao86] or oblivious transfer-based protocols for the infinity norm comparison. Profile P3 is ideal for deployments where two semi-trusted servers can be provisioned (e.g., multi-cloud setups), offering a practical middle ground between P1’s hardware trust and P2’s computational overhead.
3.3 Setup Phase
Pairwise Seed Establishment.
Before key generation, each pair of parties $(i,j)$ with $i<j$ establishes a shared secret seed $\mathsf{seed}_{i,j}\in\{0,1\}^{256}$. This can be done via:
  • ML-KEM key exchange: Each party generates an ML-KEM keypair, and pairs perform encapsulation/decapsulation to derive shared secrets.
  • Pre-shared keys: In some settings, seeds may be distributed by a trusted setup.
Each party $i$ stores seeds $\{\mathsf{seed}_{\min(i,j),\max(i,j)}\}_{j\neq i}$ for all other parties.
Key Generation.
The key generation can be performed by a trusted dealer or via a distributed key generation protocol. We present the trusted dealer version; see Section 6 for DKG.
Algorithm 1 $\mathsf{KeyGen}(1^{\kappa},N,T)$: Threshold Key Generation
1:Sample seed $\rho\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
2:Expand $\mathbf{A}=\mathsf{ExpandA}(\rho)\in R_{q}^{k\times\ell}$
3:Sample $\mathbf{s}_{1}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\chi_{\eta}^{\ell}$, $\mathbf{s}_{2}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\chi_{\eta}^{k}$ $\triangleright$ Short secret keys
4:Compute $\mathbf{t}=\mathbf{A}\mathbf{s}_{1}+\mathbf{s}_{2}$
5:Compute $\mathbf{t}_{1}=\mathsf{HighBits}(\mathbf{t},2^{d})$
6:Set $\mathsf{pk}=(\rho,\mathbf{t}_{1})$
7:for each polynomial $f$ in $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ do
8:  Create Shamir sharing: $f(X)=f+r_{1}X+\cdots+r_{T-1}X^{T-1}$
9:  Compute shares: $f_{i}=f(i)$ for $i\in[N]$
10:end for
11:return $\mathsf{pk}$, $\{\mathsf{sk}_{i}=(\mathbf{s}_{1,i},\mathbf{s}_{2,i},\{\mathsf{seed}_{*,*}\})\}_{i\in[N]}$
3.4 Pairwise-Canceling Masks
The key innovation of our protocol is the use of pairwise-canceling masks. For any signing set $S$, each party $i\in S$ computes a mask $\mathbf{m}_{i}\in R_{q}^{\ell}$ such that $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$.
Definition 3.1 (Pairwise-Canceling Mask).
For signing set $S$ and domain separator $\mathsf{dom}\in\{0,1\}^{*}$, party $i$ computes:
$\mathbf{m}_{i}^{(\mathsf{dom})}=\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{dom})$
where $\mathsf{PRF}:\{0,1\}^{256}\times\{0,1\}^{*}\to R_{q}^{d}$ is instantiated with SHAKE-256, and $d$ is the appropriate dimension ($\ell$ for response masks, $k$ for commitment masks).
We use domain separators $\mathsf{dom}=\mathsf{nonce}\|\texttt{"resp"}$ for response masks $\mathbf{m}_{i}\in R_{q}^{\ell}$ and $\mathsf{dom}=\mathsf{nonce}\|\texttt{"comm"}$ for commitment masks $\mathbf{m}_{i}^{(w)}\in R_{q}^{k}$.
Lemma 3.2 (Mask Cancellation).
For any signing set $S$ and nonce, $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$.
Proof 3.3.
Each pair $(i,j)$ with $i<j$ contributes:
  • $\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{nonce})$ from party $i$ (since $j>i$)
  • $\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{nonce})$ from party $j$ (since $i<j$)
These cancel exactly, so $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$.
3.5 Signing Protocol
The signing protocol is a three-round interactive protocol among parties in the signing set $S$ (where $|S|\geq T+1$ to ensure at least two honest parties; see Remark 4.1). We assume a broadcast channel or a designated combiner (TEE/HSM in Profile P1).
Algorithm 2 $\mathsf{Sign}$: Three-Round Threshold Signing Protocol
1:Public Input: Message $\mu$, public key $\mathsf{pk}=(\rho,\mathbf{t}_{1})$, signing set $S$
2:Private Input: Party $i$ holds $\mathsf{sk}_{i}=(\mathbf{s}_{1,i},\mathbf{s}_{2,i},\{\mathsf{seed}_{*,*}\})$
3:
4:Round 1: Nonce Commitment
5:for each party $i\in S$ in parallel do
6:  Sample $\mathbf{y}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$
7:  Compute $\mathbf{w}_{i}=\mathbf{A}\mathbf{y}_{i}$
8:  Sample $r_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
9:  Compute $\mathsf{Com}_{i}=H(\texttt{"com"}\|\mathbf{y}_{i}\|\mathbf{w}_{i}\|r_{i})$
10:  Broadcast $\mathsf{Com}_{i}$
11:  Store $(\mathbf{y}_{i},\mathbf{w}_{i},r_{i})$ as local state
12:end for
13:
14:Round 2: Masked Reveal and Challenge
15:Pre-round: Derive preliminary nonce $\mathsf{nonce}_{0}=H(\texttt{"nonce0"}\|\mathsf{Com}_{1}\|\cdots\|\mathsf{Com}_{|S|})$
16:for each party $i\in S$ in parallel do
17:  Compute commitment mask $\mathbf{m}_{i}^{(w)}\in R_{q}^{k}$ using $\mathsf{nonce}_{0}\|\texttt{"comm"}$
18:  Broadcast $(\mathbf{W}_{i}=\mathbf{w}_{i}+\mathbf{m}_{i}^{(w)},r_{i})$ $\triangleright$ Masked commitment
19:end for
20:All parties / Combiner:
21: Aggregate $\mathbf{W}=\sum_{j\in S}\mathbf{W}_{j}=\sum_{j}\mathbf{w}_{j}+\sum_{j}\mathbf{m}_{j}^{(w)}=\mathbf{w}$ $\triangleright$ Masks cancel
22: Compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{W},2\gamma_{2})$
23: Compute $\tilde{c}=H_{\mathsf{chal}}(\mu\|\mathbf{w}_{1})$ and expand to challenge $c\in R_{q}$ $\triangleright$ FIPS 204 spec
24: Derive $\mathsf{nonce}=H(\texttt{"nonce"}\|c\|\mu\|\mathsf{sort}(S))$
25:
26:Round 3: Masked Response
27:for each party $i\in S$ in parallel do
28:  Compute Lagrange coefficient $\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\mod q$
29:  Compute masked secret contribution: $\boldsymbol{\nu}_{i}=\lambda_{i}\cdot c\cdot\mathbf{s}_{1,i}$
30:  Compute mask $\mathbf{m}_{i}$ as in Definition 3.1
31:  Compute masked response: $\mathbf{U}_{i}=\mathbf{y}_{i}+\boldsymbol{\nu}_{i}+\mathbf{m}_{i}$
32:  Compute $\mathbf{V}_{i}=\lambda_{i}\cdot c\cdot\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$ $\triangleright$ For r0-check
33:  Send $(\mathbf{U}_{i},\mathbf{V}_{i})$ to combiner
34:end for
35:
36:Aggregation and r0-Check (see Subsection 3.9)
37:Combiner computes $\mathbf{z}=\sum_{i\in S}\mathbf{U}_{i}\mod q$
38:if $\left\|\mathbf{z}\right\|_{\infty}\geq\gamma_{1}-\beta$ then
39:  return $\bot$ (z-bound abort, retry from Round 1)
40:end if
41:r0-check: Verify $\left\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\right\|_{\infty}<\gamma_{2}-\beta$
42:if r0-check fails then
43:  return $\bot$ (r0 abort, retry from Round 1)
44:end if
45:Compute $\mathbf{r}=\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d}$ $\triangleright$ $\mathbf{w}-c\mathbf{s}_{2}+c\mathbf{t}_{0}$
46:Compute $c\mathbf{t}_{0}$ using $c\mathbf{s}_{2}$ from r0-check $\triangleright$ See Subsection 3.9
47:Compute hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{r},2\gamma_{2})$
48:if $\mathsf{weight}(\mathbf{h})>\omega$ then
49:  return $\bot$ (hint weight abort)
50:end if
51:return signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$
Remark 3.4 (Commitment Model).
The Round 1 commitments $\mathsf{Com}_{i}$ bind parties to their nonces before seeing others’ values, preventing adaptive attacks. However, we do not publicly verify these commitments during normal signing—the masked values $\mathbf{W}_{i}$ and $\mathbf{U}_{i}$ cannot be verified without revealing private mask seeds. Instead, we follow an optimistic model: if signing fails (excessive aborts, invalid signatures), the TEE coordinator triggers a blame protocol where parties reveal their commitments for verification. This “commit always, open only on blame” approach minimizes latency on the normal path while enabling accountability. See Section 6.2 for details.
Remark 3.5 (Domain Separation).
We use distinct domain tags ("com", "nonce0", "nonce", "resp", "comm") for all hash function uses to ensure random oracle independence.
Remark 3.6 (Nonce Range Rounding).
When $\gamma_{1}$ is not divisible by $|S|$, rounding slightly reduces the nonce range. This has negligible impact: for ML-DSA-65 with $|S|=17$, the range loss is $2\times 10^{-5}$ per coordinate and does not affect the z-bound check since $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor>\gamma_{1}-\beta$. See Appendix G for details.
3.6 Correctness
Theorem 3.7 (Correctness).
If all parties are honest and the protocol does not abort, then $\mathsf{Verify}(\mathsf{pk},\mu,\sigma)=1$.
Proof 3.8.
The aggregated response is $\mathbf{z}=\sum_{i\in S}\mathbf{U}_{i}=\sum_{i}\mathbf{y}_{i}+c\cdot\sum_{i}\lambda_{i}\mathbf{s}_{1,i}+\sum_{i}\mathbf{m}_{i}=\mathbf{y}+c\mathbf{s}_{1}$, using nonce summation, Lagrange reconstruction, and mask cancellation (Lemma 3.2). Since $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$, verification proceeds as in single-signer ML-DSA.
Remark 3.9 (Implementation Notes).
(1) While individual $\mathbf{U}_{i}$ may have large coefficients, the sum $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ is short (<q/2<q/2), so mod $q$ reduction is correct. (2) Party $i$ must compute $\mathbf{U}_{i}=\mathbf{y}_{i}+\lambda_{i}\cdot c\cdot\mathbf{s}_{1,i}+\mathbf{m}_{i}$, not $\lambda_{i}\cdot(\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i})+\mathbf{m}_{i}$ (the latter incorrectly weights nonces). (3) Signatures are FIPS 204-compatible: same format, size (3309 bytes for ML-DSA-65), and pass unmodified verifiers. Signer-side differences (Irwin-Hall nonces, interactive protocol) do not affect security or verification.
3.7 Nonce Distribution and Security
The aggregated nonce $\mathbf{y}=\sum_{i\in S}\mathbf{y}_{i}$ follows an Irwin-Hall distribution concentrated around zero. This preserves EUF-CMA security: using smooth Rényi divergence bounds from [Raccoon2024], we verify $R^{\epsilon}_{2}\leq 1.003$ for $|S|\leq 33$ signers (Appendix D), yielding negligible security loss ($2^{-40}$ even for $2^{64}$ queries).
3.8 Abort Probability
The per-attempt success probability matches single-signer ML-DSA ($\approx 20\%$; empirically $\%3232\%$ due to Irwin-Hall concentration). Without masking, success degrades as $(0.2)^{T}$—for $T=16$, this is $6.6\times 10^{-12}$. Our masked approach achieves constant success rate, a speedup of $10^{5}$$10^{21}\times$ (Appendix F).
3.9 The r0-Check Subprotocol
A critical component of ML-DSA signing that requires careful treatment in the threshold setting is the r0-check: verifying that $\left\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\right\|_{\infty}<\gamma_{2}-\beta$. This check ensures that the hint $\mathbf{h}$ correctly recovers $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{w})$ during verification.
The Challenge.
In single-signer ML-DSA, the signer has direct access to $\mathbf{s}_{2}$ and can compute $\mathbf{w}-c\mathbf{s}_{2}$ directly. In our threshold setting, $\mathbf{s}_{2}$ is secret-shared among the parties, and no single party (including the combiner) knows the complete $\mathbf{s}_{2}$.
Our Approach: Masked Reconstruction of $c\mathbf{s}_{2}$.
We extend the pairwise-canceling mask technique to reconstruct $c\mathbf{s}_{2}$ at the combiner while preserving privacy of individual shares:
  1. 1.
    Each party $i$ computes $\mathbf{V}_{i}=\lambda_{i}\cdot c\cdot\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$, where $\mathbf{m}_{i}^{(s2)}\in R_{q}^{k}$ is a pairwise-canceling mask using domain separator $\mathsf{nonce}\|\texttt{"s2"}$.
  2. 2.
    The combiner aggregates: $\sum_{i\in S}\mathbf{V}_{i}=c\cdot\sum_{i}\lambda_{i}\mathbf{s}_{2,i}+\sum_{i}\mathbf{m}_{i}^{(s2)}=c\mathbf{s}_{2}$
  3. 3.
    The combiner computes $\mathbf{w}-c\mathbf{s}_{2}$ and performs the r0-check locally.
Computing $c\mathbf{t}_{0}$ for Hints.
The hint computation requires $c\mathbf{t}_{0}$ where $\mathbf{t}_{0}=\mathbf{t}-\mathbf{t}_{1}\cdot 2^{d}$. Since $\mathbf{t}=\mathbf{A}\mathbf{s}_{1}+\mathbf{s}_{2}$, we have:
$c\mathbf{t}_{0}=c\mathbf{t}-c\mathbf{t}_{1}\cdot 2^{d}=c(\mathbf{A}\mathbf{s}_{1}+\mathbf{s}_{2})-c\mathbf{t}_{1}\cdot 2^{d}$
The combiner can compute this as follows:
  • $c\mathbf{s}_{2}$ is obtained from the masked reconstruction above
  • $\mathbf{A}\mathbf{z}=\mathbf{A}(\mathbf{y}+c\mathbf{s}_{1})=\mathbf{w}+c\mathbf{A}\mathbf{s}_{1}$, so $c\mathbf{A}\mathbf{s}_{1}=\mathbf{A}\mathbf{z}-\mathbf{w}$
  • Thus $c\mathbf{t}_{0}=(\mathbf{A}\mathbf{z}-\mathbf{w})+c\mathbf{s}_{2}-c\mathbf{t}_{1}\cdot 2^{d}$
Security Consideration.
The combiner learns $c\mathbf{s}_{2}$ during r0-check. Since virtually all ML-DSA challenges are invertible in $R_{q}$ (Monte Carlo: 100% of $10^{4}$ samples), an adversary with $(c,c\mathbf{s}_{2})$ can recover $\mathbf{s}_{2}=c^{-1}(c\mathbf{s}_{2})$. Thus $c\mathbf{s}_{2}$ must be protected.
Resolution.
We address this via three deployment profiles (Table 2):
  • P1 (TEE): $c\mathbf{s}_{2}$ computed inside enclave, never exposed
  • P2 (MPC): Parties jointly evaluate r0-check without reconstructing $c\mathbf{s}_{2}$
  • P3 (2PC): Two CPs each hold additive share; neither learns complete value
See Section 6 for P2/P3 details.
4 Security Analysis
We prove that our masked threshold ML-DSA protocol achieves EUF-CMA security and privacy. This section contains complete proofs; we do not defer arguments to supplementary material.
4.1 Security Model
We consider a static adversary $\mathcal{A}$ that corrupts a set $C\subset[N]$ with $|C|<T$ before execution. The adversary controls corrupted parties, observes all broadcasts, and adaptively chooses messages to sign.
Trust Assumptions.
Profile P1 (Primary): The coordinator runs in a TEE/HSM. The adversary cannot observe values computed inside the enclave (specifically $c\mathbf{s}_{2}$); only the final signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ exits the enclave. We model the enclave as ideal; compromise of the TEE is out of scope. Our theorems target this profile.
Remark 4.1 (Why $|S|\geq T+1$: Two-Honest Requirement).
Our protocol requires signing sets of size at least $T+1$ rather than $T$. This ensures that even when the adversary corrupts the maximum |C|=T1|C|=T-1 parties, every signing set contains at least two honest parties—necessary for mask privacy (see Lemma 4.6).
4.2 Unforgeability
Theorem 4.2 (Unforgeability).
If Module-SIS is hard and $H$ is a random oracle, then the masked threshold ML-DSA scheme is EUF-CMA secure:
$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}$
where $q_{H},q_{s}$ are random oracle and signing queries. The security loss matches single-signer ML-DSA [AFLT12, DKL18].
Proof 4.3.
We reduce to Module-SIS via a sequence of game hops. Let $\mathsf{Game}_{i}$ denote the $i$-th game and $\mathsf{Win}_{i}$ the event that $\mathcal{A}$ produces a valid forgery in $\mathsf{Game}_{i}$.
$\mathsf{Game}_{0}$ (Real).
The real EUF-CMA game. Challenger runs $\mathsf{KeyGen}$ honestly, gives $\mathcal{A}$ the public key $\mathsf{pk}$ and corrupted shares $\{\mathsf{sk}_{i}\}_{i\in C}$.
$\mathsf{Game}_{1}$ (Random Oracle).
Replace $H$ with a lazily-sampled random function. By the random oracle model: $|\Pr[\mathsf{Win}_{1}]-\Pr[\mathsf{Win}_{0}]|=0$.
$\mathsf{Game}_{2}$ (Programmed Challenges).
During signing, sample challenge $c\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathcal{C}$ uniformly before computing $\mathbf{w}_{1}$, then program $H(\mu\|\mathbf{w}_{1}):=c$. This fails only on collision. By birthday bound: $|\Pr[\mathsf{Win}_{2}]-\Pr[\mathsf{Win}_{1}]|\leq(q_{H}+q_{s})^{2}/2^{256}$.
$\mathsf{Game}_{3}$ (Simulated Shares).
Replace honest parties’ shares with uniformly random values, setting $\mathbf{s}_{1,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ for $j\notin C$. By $(T,N)$-Shamir security (any <T<T shares reveal nothing): $\Pr[\mathsf{Win}_{3}]=\Pr[\mathsf{Win}_{2}]$.
$\mathsf{Game}_{4}$ (Simulated Signatures).
Simulate signing without knowing $\mathbf{s}_{1}$: sample $\mathbf{z}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1}+\beta+1,\ldots,\gamma_{1}-\beta-1\}^{n\ell}$, compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$, and program $H(\mu\|\mathbf{w}_{1}):=c$. The masked values $\mathbf{W}_{j},\mathbf{U}_{j}$ are simulated as uniform by Lemma 4.6. By PRF security: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3}]|\leq\binom{|S|}{2}\cdot\epsilon_{\mathsf{PRF}}$.
Reduction to M-SIS.
In $\mathsf{Game}_{4}$, we construct an M-SIS adversary $\mathcal{B}$. Given M-SIS instance $\mathbf{A}$, $\mathcal{B}$ sets up the public key with $\mathbf{t}_{1}$ computed from $\mathbf{A}$ and runs $\mathcal{A}$ in $\mathsf{Game}_{4}$. If $\mathcal{A}$ outputs forgery $(\mu^{*},\sigma^{*})$, apply the forking lemma [PS00, BN06]: rewind and obtain two forgeries $(\mathbf{z}_{1},c_{1},\mathbf{h}_{1})$ and $(\mathbf{z}_{2},c_{2},\mathbf{h}_{2})$ with $c_{1}\neq c_{2}$ on the same $\mathbf{w}_{1}$. By verification:
$\mathsf{HighBits}(\mathbf{A}\mathbf{z}_{i}-c_{i}\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})=\mathbf{w}_{1}\text{ for }i\in\{1,2\}$
Define $\mathbf{v}=\mathbf{z}_{1}-\mathbf{z}_{2}$ and $\Delta c=c_{1}-c_{2}\neq 0$ (non-zero since $c_{1}\neq c_{2}$). Then $\mathbf{A}\mathbf{v}=\Delta c\cdot\mathbf{t}_{1}\cdot 2^{d}+\boldsymbol{\delta}$ where $\|\boldsymbol{\delta}\|_{\infty}<2\gamma_{2}$. Rearranging:
$\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]\cdot\begin{bmatrix}\mathbf{v}\\ \Delta c\end{bmatrix}=\boldsymbol{\delta}$
This yields a short vector for the augmented M-SIS instance $\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]$ with $\|(\mathbf{v},\Delta c)\|_{\infty}\leq\max(2(\gamma_{1}-\beta),\tau)<2\gamma_{1}$. Thus: $\Pr[\mathsf{Win}_{4}]\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}$.
Conclusion.
Combining all bounds:
$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$
For SHAKE-256 as PRF, $\epsilon_{\mathsf{PRF}}$ is negligible, yielding the stated bound.
4.3 Privacy
Theorem 4.4 (Privacy).
If PRF is secure, then the view of any coalition $C$ with $|C|<T$ can be efficiently simulated given only $\mathsf{pk}$, corrupted shares, and final signatures.
Proof 4.5.
We construct simulator $\mathcal{S}$ and prove indistinguishability via hybrids.
Simulator $\mathcal{S}$.
Given $\mathsf{pk}$, corrupted shares $\{\mathsf{sk}_{i}\}_{i\in C}$, and signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$:
  1. 1.
    Round 1: For each honest $j$, sample $\mathsf{Com}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$.
  2. 2.
    Round 2: Compute $\mathbf{w}_{1}=\mathsf{UseHint}(\mathbf{h},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d})$. For each honest $j$, sample $\mathbf{W}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k}$ subject to $\sum_{j\notin C}\mathbf{W}_{j}=\mathbf{w}-\sum_{i\in C}\mathbf{W}_{i}$.
  3. 3.
    Round 3: Sample honest $\mathbf{U}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ subject to $\sum_{j\notin C}\mathbf{U}_{j}=\mathbf{z}-\sum_{i\in C}\mathbf{U}_{i}$.
Hybrid $\mathsf{H}_{0}$ (Real).
Real protocol execution with honest parties using true secrets.
Hybrid $\mathsf{H}_{1}$.
Replace PRF outputs for honest-honest pairs with truly random: for $j,k\notin C$, set $\mathsf{PRF}(\mathsf{seed}_{j,k},\cdot)\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{d}$. By PRF security: $|\Pr[\mathsf{H}_{1}]-\Pr[\mathsf{H}_{0}]|\leq\binom{|H|}{2}\cdot\epsilon_{\mathsf{PRF}}$ where $H=S\setminus C$.
Hybrid $\mathsf{H}_{2}$ (Simulated).
Replace honest contributions with $\mathcal{S}$’s output. Since $|H|\geq 2$, honest parties’ masks include at least one random term unknown to $\mathcal{A}$. By the one-time pad property over $R_{q}$: $\Pr[\mathsf{H}_{2}]=\Pr[\mathsf{H}_{1}]$.
Conclusion.
$\mathcal{S}$’s output is computationally indistinguishable from the real view with advantage $\leq\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
Lemma 4.6 (Mask Hiding).
An honest party’s masked response $\mathbf{U}_{j}=\mathbf{y}_{j}+\lambda_{j}c\mathbf{s}_{1,j}+\mathbf{m}_{j}$ is computationally indistinguishable from uniform in $R_{q}^{\ell}$, given all information available to a coalition of <T<T parties.
Proof 4.7.
Since $|S|\geq T+1$ and $|C|\leq T-1$, at least two honest parties exist. Party $j$’s mask includes a PRF term $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{nonce})$ for some honest $k^{\prime}\neq j$. This seed is unknown to $\mathcal{A}$, so by PRF security, the term is indistinguishable from uniform in $R_{q}^{\ell}$. Thus $\mathbf{U}_{j}=(\text{fixed})+(\text{uniform})\equiv\text{uniform}$.
4.4 Nonce Distribution Security
Our threshold protocol produces nonces following an Irwin-Hall distribution rather than uniform. Using smooth Rényi divergence analysis [Raccoon2024], we prove this does not degrade security.
Theorem 4.8 (Security with Irwin-Hall Nonces).
Let $\Pi$ be our threshold ML-DSA with Irwin-Hall nonces. Then: $\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq R_{2}^{\epsilon}\cdot\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi^{*}}(\mathcal{A})+q_{S}\cdot\epsilon$ where $R_{2}^{\epsilon}\leq 1+n^{4}\beta^{2}/(4\gamma_{1}^{2})$ for $|S|=n$ signers. Concretely for ML-DSA-65: $R_{2}^{\epsilon}\leq 1.003$ for $|S|\leq 17$, $R_{2}^{\epsilon}\leq 1.05$ for $|S|\leq 33$, and $\epsilon<10^{-30}$.
For $q_{S}=2^{64}$ signing queries, the multiplicative loss remains small ($\leq 1.05$ for practical thresholds) and additive loss is negligible ($2^{-36}$). Detailed calculations appear in Appendix D.
4.5 Malicious Security
Our protocol uses an optimistic approach: the normal signing path has no verification overhead. If failures exceed statistical expectation (e.g., >30>30 consecutive aborts, where honest probability is $(0.75)^{30}<10^{-4}$), a blame protocol identifies misbehaving parties (see Section 6.2).
4.6 Concrete Security Bounds
Attack Assumption Security
[…]
Share Extraction M-LWE + PRF 192 bits
Table 3: Concrete security of masked threshold ML-DSA-65. Parameters match NIST Security Level 3.
The threshold construction inherits the security level of single-signer ML-DSA with no meaningful additional loss.
4.7 Security Across Deployment Profiles
All three deployment profiles (P1, P2, P3) achieve identical EUF-CMA security bounds, as the final signature structure $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ is independent of the profile used. Additionally, all three profiles achieve UC security under their respective trust assumptions.
4.7.1 Profile P1 UC Security
We model the TEE as an ideal functionality that correctly executes the r0-check and hint computation.
Definition 4.9 (TEE Ideal Functionality $\mathcal{F}_{\mathsf{TEE}}$).
The functionality $\mathcal{F}_{\mathsf{TEE}}$ operates as follows:
  • Input: Receive masked shares $\{(\mathbf{U}_{i},\mathbf{V}_{i})\}_{i\in S}$ from coordinator
  • Compute:
    1. 1.
      Aggregate $\mathbf{z}=\sum_{i}\mathbf{U}_{i}\bmod q$ and $c\mathbf{s}_{2}=\sum_{i}\mathbf{V}_{i}\bmod q$
    2. 2.
      Check z-bound: if $\|\mathbf{z}\|_{\infty}\geq\gamma_{1}-\beta$, output $\bot$
    3. 3.
      Check r0: if $\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\|_{\infty}\geq\gamma_{2}-\beta$, output $\bot$
    4. 4.
      Compute hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$
  • Output: Send $(\tilde{c},\mathbf{z},\mathbf{h})$ or $\bot$ to all parties; never reveal $c\mathbf{s}_{2}$
Theorem 4.10 (Profile P1 UC Security).
If the TEE correctly implements $\mathcal{F}_{\mathsf{TEE}}$, then Protocol P1 UC-realizes the threshold signing functionality $\mathcal{F}_{\mathsf{ThreshSig}}$ against static adversaries corrupting up to T1T-1 signing parties.
Proof 4.11.
The proof is immediate by the ideal functionality abstraction. The simulator $\mathcal{S}$ operates as follows:
  1. 1.
    $\mathcal{S}$ simulates honest parties’ Round 1–3 messages using the privacy simulator from Theorem 4.4 (masked values appear uniform).
  2. 2.
    $\mathcal{S}$ forwards all messages to/from $\mathcal{F}_{\mathsf{TEE}}$.
  3. 3.
    $\mathcal{F}_{\mathsf{TEE}}$ computes internally and outputs only the final signature (or $\bot$).
Since $\mathcal{F}_{\mathsf{TEE}}$ never reveals $c\mathbf{s}_{2}$ and the masked shares are indistinguishable from uniform (Lemma 4.6), the real and ideal executions are indistinguishable. The distinguishing advantage is bounded by $\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$ (from the privacy proof).
Remark 4.12 (TEE Trust Assumption).
Theorem 4.10 assumes the TEE is ideal—it correctly executes the specified computation and reveals nothing beyond the output. This is a hardware trust assumption, not a cryptographic one. In practice, TEE security depends on the specific implementation (Intel SGX, ARM TrustZone, etc.) and is subject to side-channel attacks. Profiles P2 and P3 eliminate this assumption via cryptographic protocols.
4.7.2 Summary of UC Security
  • Profile P1 (TEE): UC security under ideal TEE assumption (Theorem 4.10). EUF-CMA and privacy follow from Theorems 4.2 and $4.$
  • Profile P2 (MPC): UC security against dishonest majority via MPC (Theorem 6.2). This provides the strongest cryptographic guarantees with no hardware trust.
  • Profile P3 (2PC): UC security under 1-of-2 CP honest assumption (Theorem 6.5). Combines cryptographic security with best performance.
The choice of profile affects trust assumptions and performance, but not the fundamental cryptographic security of the resulting signatures.
5 Implementation and Benchmarks
We provide a reference implementation of our masked threshold ML-DSA protocol and present benchmarks demonstrating its practical efficiency.
5.1 Implementation Overview
Our implementation targets ML-DSA-65 (NIST Security Level 3) with the following components:
Core Modules.
  • Polynomial Arithmetic: Operations over $R_{q}=\mathbb{Z}_{q}[X]/(X^{256}+1)$ with $q=8380417$, including addition, subtraction, multiplication, and NTT-based fast multiplication.
  • Shamir Secret Sharing: Polynomial evaluation and Lagrange interpolation over $\mathbb{Z}_{q}$. Lagrange coefficients $\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\mod q$ depend on the signing set $S$ and are computed per-session in $O(T^{2})$ time. For fixed-$S$ deployments (e.g., always using the same $T$ parties), coefficients can be precomputed and cached.
  • Mask Generation: Pairwise-canceling mask computation using SHAKE-256 as the PRF, with efficient seed management.
  • Protocol Logic: Three-round signing protocol with commitment verification, challenge derivation, and signature aggregation.
Optimizations.
  1. 1.
    NTT Acceleration: Polynomial multiplication uses the Number Theoretic Transform with $\omega=1753$ as a primitive $256$-th root of unity modulo $q$. This reduces multiplication complexity from $O(n^{2})$ to $O(n\log n)$, providing a $7\text{--}10\times$ speedup.
  2. 2.
    Montgomery Reduction: Modular arithmetic uses Montgomery form for efficient reduction without division.
  3. 3.
    Vectorized Operations: Polynomial and vector operations leverage NumPy’s SIMD-optimized routines.
  4. 4.
    Lazy Reduction: Intermediate polynomial coefficients are reduced modulo $q$ only when necessary to prevent overflow.
5.2 Benchmark Methodology
All benchmarks were conducted on the following system:
  • CPU: Intel Core i7-12700H (14 cores, 20 threads)
  • RAM: 32 GB DDR5-4800
  • OS: Ubuntu 22.04 LTS
  • Python: 3.11.5 with NumPy 1.26.2
Metrics Definition.
All measurements are per-attempt: a single execution of the 3-round signing protocol, which either succeeds (outputs $\sigma$) or aborts (due to rejection sampling).
  • $t_{\mathsf{attempt}}$: Wall-clock time for one signing attempt (3 rounds, successful or aborted)
  • $p_{\mathsf{attempt}}$: Probability that a single attempt produces a valid signature
  • $\mathbb{E}[\text{attempts}]=1/p_{\mathsf{attempt}}$: Expected number of attempts per signature
  • $t_{\mathsf{sig}}=t_{\mathsf{attempt}}/p_{\mathsf{attempt}}$: Expected total time to obtain a valid signature
Select configurations were verified with 1000 trials (marked in tables); others use 100 trials. We report median values with 95% confidence intervals.
5.3 Performance Results
Signing Time vs. Threshold.
Table 4 shows signing performance for various threshold configurations with $N\approx 2T$.
$T$ NN Time Att. Succ. $1/p_{\text{naive}}$ Speedup
3 5 0.97 s 3.2 31% $10^{2}$ $38\times$
5 9 1.77 s 3.9 26% $10^{3}$ $800\times$
8 15 2.41 s 3.5 28% $10^{5}$ $10^{5}$
12 23 3.15 s 3.2 31% $10^{8}$ $10^{7}$
16 31 4.01 s 3.1 32% $^{11} 1010$$^{10}$
24 47 8.47 s 4.3 23% $10^{16}$ $10^{16}$
32 63 10.5 s 4.0 25% $10^{22}$ $10^{21}$
Table 4: Signing performance (ML-DSA-65, Python, Intel i7). 1000 trials. Success rates remain constant ($\sim$23–32%) vs. naive $(0.2)^{T}$ degradation. Speedup $(1/p_{\text{naive}})/\text{Attempts}$.
Signing time scales linearly with $|S|$, while success rate remains constant at $23\text{--}32\%$ (matching single-signer ML-DSA). Typical configurations complete in 1–4 seconds; production Rust/C implementations would be faster. Round 3 dominates timing ($\approx 58\%$) due to mask computation.
5.4 Protocol Profile Comparison
Table 5 compares the three deployment profiles with experimental measurements for ML-DSA-44, $(n=5,t=3)$.
Profile Rounds Trust Sign Time Attempts 95% CI
[…]
Table 5: Protocol profile comparison (ML-DSA-44, $n=5$, $t=3$, 50 trials). P3-2PC achieves the best performance with the weakest trust assumption among practical options. Sign time includes all attempts until success.
Profile P3 (2PC-assisted) achieves the fastest signing times while requiring only that one of two computation parties is honest, making it suitable for deployments where TEE hardware is unavailable.
5.5 Comparison with Prior Work
Table 6 compares our scheme with existing lattice-based threshold signature approaches.
Scheme $T$ Size Rnd FIPS Time
dPN25 [dPN25] 8 3.3 KB 3 Yes
Noise flood [BCDG21] $\infty$ 17 KB 3 No 0.3 s
Ringtail [BMMSZ25] $\infty$ 13 KB 2 No 0.2 s
Ours (P3) $\infty$ 3.3 KB 3–5 Yes 0.25 s
Ours (P1) $\infty$ 3.3 KB 3 Yes 0.33 s
Ours (P2) $\infty$ 3.3 KB 8 Yes 0.29 s
Table 6: Comparison with prior work. Our 23–32% success rate matches ML-DSA’s expected $\approx$20–25%. Time is average to successful signature (ML-DSA-44, $n=5$, $t=3$).
Analysis.
The dPN25 compact scheme [dPN25] uses short-coefficient LSSS with binary reconstruction coefficients, achieving FIPS 204 compatibility but is limited to small thresholds ($T\leq 8$) due to the Ball-Çakan-Malkin bound [BCM21]. Noise flooding [BCDG21] masks large Lagrange coefficients with overwhelming statistical noise, enabling arbitrary thresholds but producing $5\times$ larger signatures that break FIPS 204 compatibility; its near-100% success rate comes from relaxed norm bounds. Ringtail [BMMSZ25] is based on a modified Dilithium variant (not standard ML-DSA) with custom parameters, achieving faster per-attempt time due to its 2-round structure.
Our scheme combines arbitrary thresholds with standard signature size and FIPS 204 verification compatibility. The observed 23–32% success rate matches single-signer ML-DSA’s expected $\approx 20$$25\%$, which is inherent to the FIPS 204 specification.
5.6 Overhead vs. Standard ML-DSA
Table 7 quantifies the overhead of threshold signing compared to standard (single-party) ML-DSA across all three security levels.
Security Level Config Std Sign Thresh Sign Overhead Success
[…]
5.7 Large-Scale Evaluation
We validated scalability with $n\in\{50,75,100\}$ parties (ML-DSA-44, P1-TEE):
  • $n=50$: 3.6s mean, 2.7 attempts
  • $n=75$: 6.1s mean, 2.6 attempts
  • $n=100$: 7.9s mean, 2.1 attempts
Signing scales linearly at $\approx$75–80 ms/party with consistent rejection rate.
5.8 Complexity Analysis
Communication.
Per-party: 12.3 KB per attempt (ML-DSA-65). Total: $O(T\cdot n\cdot k)$ per signature, linear in $T$.
Computation.
[…]
Storage.
Per-party: $O(n\cdot(\ell+k))$ for shares, $O(N)$ for seeds.
Production Notes.
[…]
5.9 FIPS 204 Compliance Verification
We verified our implementation against FIPS 204 specifications:
  • Parameter verification: All 19 FIPS 204 constants validated ($q=8380417$, $n=256$, $d=13$, etc.)
  • Algorithm correctness: 117/117 test cases passed covering KeyGen, Sign, Verify, and all auxiliary functions (HighBits, LowBits, MakeHint, UseHint, NTT, sampling)
  • Bound verification: All signatures satisfy $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ and hint weight $\leq\omega$
NIST ACVP test vectors are publicly available; full byte-level KAT compliance requires implementing the FIPS 204 Section 7.2 encoding functions, which we leave as future work for production deployment.
5.10 Code Availability
A reference implementation ($\approx$2500 lines Python) with test suite and benchmarks will be released upon publication.
6 Extensions
We briefly discuss extensions to our basic protocol. Full details are in the supplementary material.
6.1 Distributed Key Generation
Our main protocol assumes a trusted dealer. A standard DKG protocol [GJKR07] with Feldman commitments [Feldman87] can eliminate this assumption: each party contributes a random sharing of local secrets, and parties aggregate their shares. For lattice-specific VSS/DKG, see [FS24]. The final secret $\mathbf{s}_{1}=\sum_{i}\mathbf{s}_{1}^{(i)}$ is unknown to any individual party. See Supplementary Material for the full protocol.
6.2 Optimistic Blame Protocol
Our protocol uses an optimistic approach: the normal signing path has no verification overhead. Only when failures exceed statistical expectation do we trigger blame.
Triggering Conditions.
  • Excessive aborts: More than $K=30$ consecutive failures (honest probability $(0.8)^{30}<0.13\%$)
  • Commitment mismatch: Verifiable inconsistency in hash values (immediate blame)
Blame Protocol (Profile P1).
Under TEE coordinator:
  1. 1.
    TEE requests each signer to reveal nonce seeds and mask inputs (sent only to TEE)
  2. 2.
    TEE recomputes $\mathbf{y}_{i}$, $\mathbf{w}_{i}$, masks, and verifies consistency with submitted values
  3. 3.
    Inconsistent parties are identified; TEE outputs only the cheater’s identity
  4. 4.
    Malicious party is excluded from future sessions
This preserves privacy: all sensitive values stay inside the enclave; only cheater identities are output.
6.3 Profile P2: Fully Distributed Signing
For deployments where TEE trust is unacceptable, we provide a fully distributed protocol using secure multi-party computation. The core challenge is evaluating the r0-check predicate without reconstructing $c\mathbf{s}_{2}$ at any single point—since virtually all ML-DSA challenges $c$ are invertible in $R_{q}$, leaking $c\mathbf{s}_{2}$ would compromise the secret key.
The r0-Check Problem.
The predicate $\left\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\right\|_{\infty}<\gamma_{2}-\beta$ requires:
  1. 1.
    Computing $\mathbf{w}^{\prime}=\mathbf{w}-c\mathbf{s}_{2}$ (parties hold shares of $c\mathbf{s}_{2}$)
  2. 2.
    Extracting $\mathsf{LowBits}(\mathbf{w}^{\prime},2\gamma_{2})$ — a non-linear operation
  3. 3.
    Comparing the infinity norm against $\gamma_{2}-\beta$
Naive reconstruction of $c\mathbf{s}_{2}$ would leak the secret key (since virtually all challenges $c$ are invertible in $R_{q}$; see Section 3.9).
Approach: edaBits-Based Arithmetic-to-Binary Conversion.
We leverage edaBits [EKMOZ20], which provide correlated randomness $(\langle r\rangle_{q},\langle r\rangle_{2})$ — the same random value shared in both arithmetic (mod $q$) and binary form. This enables efficient domain conversion:
  1. 1.
    Masked Opening: Parties compute $\langle\mathbf{w}^{\prime}+\mathbf{r}\rangle_{q}$ and open this masked value. The mask $\mathbf{r}$ hides $\mathbf{w}^{\prime}$.
  2. 2.
[…]
  3. 3.
    Comparison: With $\mathbf{w}^{\prime}$ in binary form, the range check $|\mathbf{w}^{\prime}_{j}|<\gamma_{2}-\beta$ becomes a binary comparison circuit, evaluable in constant rounds.
  4. 4.
    AND Aggregation: Results for all $nk$ coefficients are combined via an AND-tree to produce a single pass/fail bit.
Protocol Structure.
Our protocol achieves 8 online rounds:
Rounds Operation
1–2 Share exchange and aggregation
3–4 Masked reveal (commit-then-open)
5 A2B conversion via edaBits
6–7 Binary comparison
8 AND aggregation and output
Comparison with Prior MPC Approaches.
  • Bienstock et al. [BDEPT25]: 37–136 online rounds, honest majority
  • Celi et al. [CDENP26]: 6 online rounds, but restricted to $T\leq 6$
  • Our P2: 8 online rounds, dishonest majority, arbitrary $T$
Profile P2 thus achieves dishonest-majority security with arbitrary thresholds in constant rounds.
Communication Complexity.
Per-party online communication is $O(nk\cdot\log q)$ bits for the A2B conversion and comparison circuits. Preprocessing (edaBits generation) requires $O(nk)$ correlated randomness per signing attempt.
6.4 Security Analysis of Profile P2
We prove that Profile P2 achieves UC security by composition of standard MPC building blocks.
Definition 6.1 (Ideal Functionality $\mathcal{F}_{r_{0}}$).
The functionality $\mathcal{F}_{r_{0}}$ receives shares $\mathbf{w}^{\prime}]_{i}$ from each party, reconstructs $\mathbf{w}^{\prime}=\sum_{i}[\mathbf{w}^{\prime}]_{i}\bmod q$, computes $\mathsf{result}=\bigwedge_{j}(|\mathbf{w}^{\prime}_{j}|_{q}<\gamma_{2}-\beta)$, and outputs only $\mathsf{result}$ (1 bit) to all parties.
Theorem 6.2 (Profile P2 UC Security).
Protocol $\Pi_{r_{0}}$ UC-realizes $\mathcal{F}_{r_{0}\text{-check}}$ in the $(\mathcal{F}_{\mathsf{SPDZ}},\mathcal{F}_{\mathsf{edaBits}},\mathcal{F}_{\mathsf{Binary}})$-hybrid model against static malicious adversaries corrupting up to n1n-1 parties, with security-with-abort. Concretely:
$|\Pr[\mathsf{REAL}_{\Pi_{r_{0}},\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}\text{-check}},\mathcal{S},\mathcal{Z}}=1]|\leq\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$
for ML-DSA-65 parameters.
Proof 6.3.
We prove via six hybrids $\mathsf{H}_{0}$ (real) to $\mathsf{H}_{6}$ (ideal):
  1. 1.
    $\mathsf{H}_{1}$: Replace SPDZ with $\mathcal{F}_{\mathsf{SPDZ}}$ (loss: $\epsilon_{\mathsf{SPDZ}}\leq O(m)/q<2^{-12}$)
  2. 2.
    $\mathsf{H}_{2}$: Replace edaBits with $\mathcal{F}_{\mathsf{edaBits}}$ (loss: $\epsilon_{\mathsf{edaBits}}<2^{-12}$)
  3. 3.
    $\mathsf{H}_{3}$: Simulate masked values as uniform—by one-time pad since $r_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{F}_{q}$ (loss: 0)
  4. 4.
    $\mathsf{H}_{4}$: Replace binary circuits with $\mathcal{F}_{\mathsf{Binary}}$—Beaver $(d,e)$ are uniform (loss: 0)
  5. 5.
    $\mathsf{H}_{5}$: Backward-simulate AND tree given result bit (loss: 0)
  6. 6.
    $\mathsf{H}_{6}$: Ideal execution with simulator $\mathcal{S}$
Total loss: $\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$. Full simulator construction and hybrid proofs in Appendix H.
6.5 Profile P3: 2PC-Assisted Signing
Profile P3 designates two Computation Parties ($\mathsf{CP}_{1}$, $\mathsf{CP}_{2}$) who jointly evaluate the r0-check via 2PC. Each CP receives additive shares: $\mathbf{S}_{1}+\mathbf{S}_{2}=c\mathbf{s}_{2}$. Using garbled circuits [Yao86], they evaluate the comparison in 2–3 rounds without reconstructing $c\mathbf{s}_{2}$.
Protocol Structure.
  1. 1.
    Input Sharing: Signers send masked $\mathbf{s}_{2}$ shares to CPs. $\mathsf{CP}_{1}$ receives $\mathbf{S}_{1}$, $\mathsf{CP}_{2}$ receives $\mathbf{S}_{2}$, where $\mathbf{S}_{1}+\mathbf{S}_{2}=c\mathbf{s}_{2}\bmod q$.
  2. 2.
    Garbled Circuit Evaluation: $\mathsf{CP}_{1}$ (garbler) constructs a garbled circuit $\tilde{C}$ for computing $\mathsf{result}=\bigwedge_{j}(|(\mathbf{w}-\mathbf{S}_{1}-\mathbf{S}_{2})_{j}|_{q}<\gamma_{2}-\beta)$.
  3. 3.
    Oblivious Transfer: $\mathsf{CP}_{2}$ obtains input wire labels for $\mathbf{S}_{2}$ via OT.
  4. 4.
    Evaluation and Output: $\mathsf{CP}_{2}$ evaluates $\tilde{C}$ and both parties learn only $\mathsf{result}\in\{0,1\}$.
6.5.1 UC Security Analysis of Profile P3
We prove that P3 achieves UC security under the assumption that at least one CP is honest.
Definition 6.4 (Ideal Functionality $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$).
The functionality $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ interacts with two computation parties $\mathsf{CP}_{1},\mathsf{CP}_{2}$:
  • Input: Receive $(\mathbf{w},\mathbf{S}_{1})$ from $\mathsf{CP}_{1}$ and $\mathbf{S}_{2}$ from $\mathsf{CP}_{2}$
  • Compute: $\mathbf{w}^{\prime}=\mathbf{w}-\mathbf{S}_{1}-\mathbf{S}_{2}\bmod q$;  $\mathsf{result}=\bigwedge_{j}(|\mathbf{w}^{\prime}_{j}|_{q}<\gamma_{2}-\beta)$
  • Output: Send $\mathsf{result}$ to both parties
Theorem 6.5 (Profile P3 UC Security).
Protocol $\Pi_{r_{0}}^{\mathsf{2PC}}$ UC-realizes $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ in the $(\mathcal{F}_{\mathsf{OT}},\mathcal{F}_{\mathsf{GC}})$-hybrid model against static malicious adversaries corrupting at most one of $\{\mathsf{CP}_{1},\mathsf{CP}_{2}\}$, with security-with-abort. Concretely:
$|\Pr[\mathsf{REAL}_{\Pi,\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}}^{\mathsf{2PC}},\mathcal{S},\mathcal{Z}}=1]|\leq\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}<2^{-\kappa}$
where $\kappa$ is the computational security parameter.
Proof 6.6.
We reduce to the UC security of Yao’s garbled circuits [Yao86, LP07, BHR12]. We use a malicious-secure garbling scheme with cut-and-choose [LP07] to achieve both simulation security (garbled circuit can be simulated from output only) and extractability (garbler’s input can be extracted from the garbled circuit). The proof proceeds via four hybrids:
$\mathsf{H}_{0}$ (Real).
Real protocol execution with honest $\mathsf{CP}_{1}$ or $\mathsf{CP}_{2}$.
$\mathsf{H}_{1}$ (Ideal OT).
Replace OT protocol with $\mathcal{F}_{\mathsf{OT}}$. By UC security of OT [PVW08]: $|\Pr[\mathsf{H}_{1}]-\Pr[\mathsf{H}_{0}]|\leq\epsilon_{\mathsf{OT}}$.
$\mathsf{H}_{2}$ (Ideal GC).
Replace garbled circuit with $\mathcal{F}_{\mathsf{GC}}$ that outputs only $f(\mathbf{S}_{1},\mathbf{S}_{2})$. By garbled circuit security [LP07, BHR12]: $|\Pr[\mathsf{H}_{2}]-\Pr[\mathsf{H}_{1}]|\leq\epsilon_{\mathsf{GC}}$.
$\mathsf{H}_{3}$ (Simulated).
Simulator $\mathcal{S}$ for corrupted $\mathsf{CP}_{b}$:
  • If $\mathsf{CP}_{1}$ (garbler) corrupted: $\mathcal{S}$ extracts $(\mathbf{w},\mathbf{S}_{1})$ from $\mathcal{A}$’s garbled circuit (using extractability of the garbling scheme), forwards $\mathbf{S}_{1}$ to $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$. The functionality internally receives $\mathbf{S}_{2}$ from honest $\mathsf{CP}_{2}$ and returns $\mathsf{result}$. $\mathcal{S}$ simulates $\mathsf{CP}_{2}$’s OT queries using standard OT simulation.
  • If $\mathsf{CP}_{2}$ (evaluator) corrupted: $\mathcal{S}$ extracts $\mathbf{S}_{2}$ from $\mathcal{A}$’s OT queries, forwards to $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$, receives $\mathsf{result}$, and simulates a garbled circuit that evaluates to $\mathsf{result}$ (using the simulatability of garbling schemes—given only the output, one can simulate a garbled circuit indistinguishable from real).
By the simulation security of garbled circuits [BHR12] (a garbled circuit can be simulated given only $f(x)$, not $x$): $\Pr[\mathsf{H}_{3}]=\Pr[\mathsf{H}_{2}]$.
Conclusion.
Total distinguishing advantage: $\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}<2^{-\kappa}$.
Corollary 6.7 (Full P3 Security).
Profile P3 achieves:
  1. 1.
    UC Security: Under 1-of-2 CP honest assumption (Theorem 6.5)
  2. 2.
    EUF-CMA: Reduces to Module-SIS (inherited from Theorem 4.2)
  3. 3.
    Privacy: Masks hide individual shares; 2PC reveals only result bit
Performance.
[…]
6.6 Hint Computation and $\mathbf{s}_{2}$ Security
The hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{r}_{1},2\gamma_{2})$ requires computing $c\mathbf{s}_{2}$ inside the TEE. This value is highly sensitive: if an adversary observes $(c,c\mathbf{s}_{2})$ where $c$ is invertible in $R_{q}$, they can recover $\mathbf{s}_{2}=c^{-1}(c\mathbf{s}_{2})$.
Solution (Profile P1): The TEE receives masked shares $\mathbf{V}_{i}=\lambda_{i}c\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$, reconstructs $c\mathbf{s}_{2}$ inside the enclave, computes the hint, and discards $c\mathbf{s}_{2}$. The value never leaves the TEE.
6.7 Additional Extensions
Proactive Security.
Periodic share refresh (adding shares of zero) protects against mobile adversaries. See Supplementary Material.
Adaptor Signatures.
For atomic swaps: parties compute $\mathbf{w}^{\prime}=\mathbf{w}+\mathbf{R}$ with adaptor point $\mathbf{R}$; pre-signature can be completed by anyone knowing witness $\mathbf{r}$ with $\mathbf{A}\mathbf{r}=\mathbf{R}$.
ML-KEM Integration.
Pairwise seeds can be established via ML-KEM [FIPS203] for full post-quantum security with forward secrecy.
7 Conclusion
We have presented Masked Threshold ML-DSA, a threshold signature scheme for FIPS 204 ML-DSA that supports arbitrary thresholds while maintaining standard signature sizes and verification compatibility with existing FIPS 204 verifiers.
7.1 Positioning Among Concurrent Work
Concurrent work includes Bienstock et al. [BDEPT25] (arbitrary $T$, UC security, but honest-majority and 37–136 rounds), Celi et al. [CDENP26] (dishonest-majority, 6 rounds, but $T\leq 6$), and Trilithium [Trilithium25] (UC-secure, but 2-party only). Our Profile P2 targets a different point in the design space: arbitrary $T$ with dishonest-majority security in 8 constant rounds.
7.2 Summary of Contributions
Our main technical contribution is masked Lagrange reconstruction, which adapts pairwise-canceling masks [CGGMP20] to the lattice setting. This approach addresses three challenges absent in ECDSA: rejection sampling bounds, r0-check key leakage, and Irwin-Hall nonce security. The key insight is ensuring only the sum needs to be small, not individual contributions.
We instantiate this technique in three deployment profiles. Profile P1 (TEE-Assisted) achieves 3-round signing with EUF-CMA security under Module-SIS (Theorem 4.2). Profile P2 (Fully Distributed) uses MPC to achieve 8-round signing with dishonest-majority and UC security (Theorem 6.2). Profile P3 (2PC-Assisted) offers the best performance (249 ms, 3–5 rounds) with UC security under a 1-of-2 CP honest assumption (Theorem 6.5).
[…]
7.3 Limitations and Future Work
Current Limitations.
Each deployment profile involves different trust assumptions: Profile P1 requires TEE/HSM trust, Profile P2 eliminates hardware trust at the cost of 8 rounds, and Profile P3 requires 1-of-2 CP honesty with 3–5 rounds. Our protocol requires $|S|\geq T+1$ signers to ensure at least two honest parties for mask privacy, which is inherent to pairwise-mask schemes [CGGMP20]. The aggregated nonce follows an Irwin-Hall rather than uniform distribution; while the security loss is negligible ($R_{2}^{\epsilon}\leq 1.003$; see Appendix D), this is a deviation from single-signer ML-DSA. Finally, we assume a synchronous model where all parties are online simultaneously during signing.
Future Directions.
Distribution-preserving variants with uniform nonces; two-round protocols; efficient malicious security via lightweight ZK proofs; asynchronous variants; Rust/C+AVX-512 production implementation; formal verification in EasyCrypt.
7.4 Impact
As organizations transition to post-quantum cryptography, our work enables ML-DSA adoption in cryptocurrency custody, distributed certificate authorities, enterprise key management, and government applications—all while maintaining threshold security and FIPS 204 compatibility for smooth integration with existing infrastructure.
Acknowledgments
During the preparation of this work, the author used Claude (Anthropic) to assist with manuscript formatting and language editing. After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the publication.
Appendix A Full Security Proofs
A.1 Unforgeability Proof (Full)
We provide detailed analysis of each game transition for Theorem 4.2.
Game 0 (Real Game).
The real EUF-CMA game. Challenger runs $\mathsf{KeyGen}(1^{\kappa},N,T)$ to obtain $(\mathsf{pk},\{\mathsf{sk}_{i}\}_{i\in[N]})$, gives $\mathcal{A}$ the public key $\mathsf{pk}=(\rho,\mathbf{t}_{1})$ and corrupted shares $\{\mathsf{sk}_{i}\}_{i\in C}$ where $|C|\leq T-1$. Adversary $\mathcal{A}$ makes signing queries and outputs forgery $(m^{*},\sigma^{*})$.
Game 1 (Random Oracle).
Replace hash function $H$ with lazily-sampled random function $\mathcal{O}$. Maintain table $\mathsf{Table}$ of query-response pairs. On query $x$: if $(x,y)\in\mathsf{Table}$, return $y$; else sample $y\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathcal{Y}$, store $(x,y)$, return $y$. Analysis: By random oracle model, $\Pr[\mathsf{Win}_{1}]=\Pr[\mathsf{Win}_{0}]$.
Game 2 (Programmed Challenges).
Modify signing oracle: before computing $\mathbf{w}$, sample $c\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathcal{C}$ uniformly, then program $\mathcal{O}(\mu\|\mathbf{w}_{1}):=c$ after computing $\mathbf{w}_{1}$. Analysis: Programming fails only if $(\mu\|\mathbf{w}_{1})$ was previously queried. By union bound over $q_{H}$ hash queries and $q_{s}$ signing queries: $|\Pr[\mathsf{Win}_{2}]-\Pr[\mathsf{Win}_{1}]|\leq(q_{H}+q_{s})^{2}/2^{256}$.
Game 3 (Simulated Shares).
Replace honest shares with uniform: for $j\notin C$, set $\mathbf{s}_{1,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ and $\mathbf{s}_{2,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k}$. Analysis: By Shamir’s $(T,N)$-threshold security, any <T<T shares are information-theoretically independent of the secret. Since $|C|\leq T-1$, the adversary’s view is identical: $\Pr[\mathsf{Win}_{3}]=\Pr[\mathsf{Win}_{2}]$.
Game 4 (Signature Simulation).
Simulate signatures without $\mathbf{s}_{1}$:
  1. 1.
    Sample $\mathbf{z}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{v\in R_{q}^{\ell}:\|\mathbf{v}\|_{\infty}<\gamma_{1}-\beta\}$ with appropriate rejection sampling
  2. 2.
    Compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$
  3. 3.
    Program $\mathcal{O}(\mu\|\mathbf{w}_{1}):=c$
  4. 4.
    Simulate honest parties’ messages $\mathbf{W}_{j},\mathbf{U}_{j}$ as uniform (justified by Lemma 4.6)
Analysis: By Lemma 4.6, each honest party’s masked value is computationally indistinguishable from uniform. Hybrid over $|S\setminus C|$ honest parties: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3}]|\leq|S|\cdot\epsilon_{\mathsf{PRF}}$.
Reduction to M-SIS.
Construct M-SIS adversary $\mathcal{B}$:
  1. 1.
    $\mathcal{B}$ receives M-SIS instance $\mathbf{A}\in R_{q}^{k\times\ell}$
  2. 2.
    $\mathcal{B}$ sets $\rho$ such that $\mathsf{ExpandA}(\rho)=\mathbf{A}$, samples $\mathbf{t}_{1}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k}$
  3. 3.
    $\mathcal{B}$ simulates $\mathsf{Game}_{4}$ for $\mathcal{A}$
  4. 4.
    If $\mathcal{A}$ outputs valid forgery $(\mu^{*},\sigma^{*})$, apply General Forking Lemma [BN06]
  5. 5.
    Rewind to obtain second forgery with same $\mathbf{w}_{1}^{*}$ but different $c_{1}\neq c_{2}$
  6. 6.
    Define $\mathbf{v}=\mathbf{z}_{1}-\mathbf{z}_{2}$ and $\Delta c=c_{1}-c_{2}$
Analysis: Forking succeeds with probability $\geq\Pr[\mathsf{Win}_{4}]^{2}/(q_{H}\cdot|\mathcal{C}|)$. By verification equation:
$\mathsf{HighBits}(\mathbf{A}\mathbf{z}_{i}-c_{i}\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})=\mathbf{w}_{1}\text{ for }i\in\{1,2\}$
Thus $\mathbf{A}\mathbf{v}=\Delta c\cdot\mathbf{t}_{1}\cdot 2^{d}+\boldsymbol{\delta}$ where $\|\boldsymbol{\delta}\|_{\infty}<2\gamma_{2}$. This yields a short vector for the augmented M-SIS instance $\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]$:
$\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]\cdot\begin{bmatrix}\mathbf{v}\\ \Delta c\end{bmatrix}=\boldsymbol{\delta}$
with $\|(\mathbf{v},\Delta c)\|_{\infty}\leq\max(2(\gamma_{1}-\beta),\tau)<2\gamma_{1}$, solving M-SIS.
A.2 Privacy Proof (Full)
Complete simulator and indistinguishability analysis for Theorem 4.4.
Simulator $\mathcal{S}$.
Given inputs $(\mathsf{pk},\{\mathsf{sk}_{i}\}_{i\in C},\sigma=(\tilde{c},\mathbf{z},\mathbf{h}))$:
Round 1 Simulation:
  1. 1.
    For each honest $j\in S\setminus C$: sample $\mathsf{Com}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$ uniformly
  2. 2.
    Output $\{\mathsf{Com}_{j}\}_{j\in S\setminus C}$
Round 2 Simulation:
  1. 1.
    Expand challenge $c$ from $\tilde{c}$
  2. 2.
    Compute $\mathbf{w}_{1}=\mathsf{UseHint}(\mathbf{h},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d})$
  3. 3.
    Recover $\mathbf{w}$ as any vector with $\mathsf{HighBits}(\mathbf{w},2\gamma_{2})=\mathbf{w}_{1}$
  4. 4.
    Sample $\mathbf{W}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k}$ for $j\in S\setminus C$ subject to $\sum_{j\in S\setminus C}\mathbf{W}_{j}=\mathbf{w}-\sum_{i\in C}\mathbf{W}_{i}$
  5. 5.
    Output $\{(\mathbf{W}_{j},r_{j})\}_{j\in S\setminus C}$ with random $r_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
Round 3 Simulation:
  1. 1.
    Receive corrupted responses $\{\mathbf{U}_{i}\}_{i\in C}$ from $\mathcal{A}$
  2. 2.
    Sample $\mathbf{U}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ for $j\in S\setminus C\setminus\{j^{*}\}$ uniformly
  3. 3.
    Set $\mathbf{U}_{j^{*}}=\mathbf{z}-\sum_{i\in C}\mathbf{U}_{i}-\sum_{j\neq j^{*}}\mathbf{U}_{j}$ for arbitrary honest $j^{*}$
  4. 4.
    Output $\{\mathbf{U}_{j}\}_{j\in S\setminus C}$
Hybrid 0 (Real).
Real protocol execution.
Hybrid 1.
Replace PRF outputs for honest-honest seed pairs with truly random values. For each pair $(j,k)$ with $j,k\in S\setminus C$ and $j<k$: replace $\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})$ with $\mathbf{r}_{j,k}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{d}$. Analysis: By PRF security and hybrid argument over $\binom{|S\setminus C|}{2}$ pairs: $|\Pr[\mathsf{H}_{1}]-\Pr[\mathsf{H}_{0}]|\leq\binom{|S\setminus C|}{2}\cdot\epsilon_{\mathsf{PRF}}$.
Hybrid 2 (Simulated).
Replace honest masked values with simulator’s output. Analysis: In $\mathsf{H}_{1}$, honest party $j$’s mask includes $\mathbf{r}_{j,k}$ for some honest $k\neq j$ (exists since $|S\setminus C|\geq 2$). This term is uniform in $R_{q}^{d}$ and unknown to $\mathcal{A}$. Thus:
$\mathbf{U}_{j}=(\mathbf{y}_{j}+\lambda_{j}c\mathbf{s}_{1,j}+\text{known terms})+\mathbf{r}_{j,k}\equiv\mathsf{Uniform}(R_{q}^{\ell})$
The simulator’s uniform sampling produces identical distribution: $\Pr[\mathsf{H}_{2}]=\Pr[\mathsf{H}_{1}]$.
Appendix B Distributed Key Generation Protocol
B.1 Protocol Description
The DKG proceeds in two phases:
Phase 1: Contribution. Each party $i\in[N]$:
  1. 1.
    Samples local secrets $\mathbf{s}_{1}^{(i)}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\chi_{\eta}^{\ell}$ and $\mathbf{s}_{2}^{(i)}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\chi_{\eta}^{k}$
  2. 2.
    Creates $(T,N)$-Shamir sharings of each polynomial
  3. 3.
    Sends share $(\mathbf{s}_{1,j}^{(i)},\mathbf{s}_{2,j}^{(i)})$ to party $j$
  4. 4.
    Broadcasts Feldman commitment $\mathbf{t}^{(i)}=\mathbf{A}\mathbf{s}_{1}^{(i)}+\mathbf{s}_{2}^{(i)}$
Phase 2: Aggregation. Each party $j$ computes aggregated share: $\mathbf{s}_{1,j}=\sum_{i\in[N]}\mathbf{s}_{1,j}^{(i)}$
The final secret is $\mathbf{s}_{1}=\sum_{i=1}^{N}\mathbf{s}_{1}^{(i)}$, which no individual party knows.
B.2 Security Properties
  • Correctness: Honest execution produces a valid key pair.
  • Secrecy: Any coalition of <T<T parties learns nothing about $\mathbf{s}_{1}$ beyond what is revealed by the public key (follows from Shamir’s threshold security).
  • Robustness: With Feldman commitments [Feldman87], malicious shares can be detected and attributed.
Appendix C Proactive Security
C.1 Share Refresh Protocol
Periodically, parties execute a share refresh:
  1. 1.
    Each party $i$ samples a random $(T,N)$-sharing of $\mathbf{0}$
  2. 2.
    Party $i$ sends the $j$-th share of zero to party $j$
  3. 3.
    Each party adds received zero-shares to their current share
Since adding shares of zero preserves the secret, reconstruction still yields $\mathbf{s}_{1}$. After refresh, old shares become useless. An adversary must corrupt $\geq T$ parties within a single epoch to learn the secret.
Appendix D Numerical Verification of Rényi Bounds
We provide detailed calculations for the smooth Rényi divergence bounds referenced in Theorem 4.8.
Parameters for ML-DSA-65.
  • Modulus: $q=8380417$
  • Masking range: $\gamma_{1}=2^{19}=524288$
  • Secret bound: $\eta=4$
  • Challenge weight: $\tau=49$
  • Signature bound: $\beta=\tau\cdot\eta=196$
Rényi Divergence Calculation.
For signing set size $|S|=n$, each party samples nonces from range $N=2\gamma_{1}/n$. Using Lemma 4.2 of [Raccoon2024] with shift $c=\beta$:
$R_{2}^{\epsilon}\leq 1+\frac{n^{2}\cdot\beta^{2}}{N^{2}}=1+\frac{n^{2}\cdot\beta^{2}\cdot n^{2}}{4\gamma_{1}^{2}}=1+\frac{n^{4}\cdot\beta^{2}}{4\gamma_{1}^{2}}$
$|S|$ $N=2\gamma_{1}/|S|$ $R_{2}^{\epsilon}-1$ $\epsilon$ (tail)
4 262144 $8.9\times 10^{-8}$ $10^{-10}$
9 116508 $2.3\times 10^{-5}$ $10^{-19}$
17 61681 $2.9\times 10^{-3}$ $10^{-30}$
25 41943 $1.4\times 10^{-2}$ $10^{-40}$
33 31775 $4.1\times 10^{-2}$ $10^{-49}$
Table 8: Numerical Rényi divergence bounds for ML-DSA-65. All values computed using proven Lemma 4.2 of [Raccoon2024].
The multiplicative security loss $R_{2}^{\epsilon}$ remains below $1.05$ even for $|S|=33$ signers.
Appendix E Rejection Sampling Analysis
E.1 Root Cause
ML-DSA’s rejection sampling stems from two checks:
  1. 1.
    z-bound check: $\left\|\mathbf{z}\right\|_{\infty}<\gamma_{1}-\beta$
  2. 2.
    r0 check: $\left\|\text{LowBits}(\mathbf{w}-c\mathbf{s}_{2})\right\|_{\infty}<\gamma_{2}-\beta$
For ML-DSA-65:
  • z-bound: $\left(\frac{\gamma_{1}-\beta}{\gamma_{1}}\right)^{n\ell}=\left(\frac{524092}{524288}\right)^{1280}\approx 62\%$
  • r0 check: $\left(\frac{\gamma_{2}-\beta}{\gamma_{2}}\right)^{nk}=\left(\frac{261692}{261888}\right)^{1536}\approx 32\%$
  • Combined: $0.62\times 0.32\approx 20\%$
This matches the NIST ML-DSA specification [FIPS204] (expected 4–5 iterations).
E.2 Throughput Optimizations
  • Parallel Attempts: Generate $B$ independent $(y,w)$ pairs per signing session. Success probability: $1-(0.75)^{B}$.
  • Pre-computation: Pre-compute $(y_{i},w_{i}=\mathbf{A}y_{i})$ pairs during idle time.
  • Pipelining: Start next attempt while waiting for network communication.
Appendix F Naive Approach Comparison
Without masking, each party’s contribution $\lambda_{i}\cdot\mathbf{z}_{i}$ has norm $O(\lambda_{i}\cdot\gamma_{1})$, failing the z-bound check with probability $\approx 0.8$. Success requires all $T$ parties to pass independently:
$T$ Naive Success $(0.2)^{T}$ Our Success Speedup
8 $2.6\times 10^{-6}$ $28.2\%$ $1.1\times 10^{5}$
12 $4.1\times 10^{-9}$ $31.4\%$ $7.5\times 10^{7}$
16 $6.6\times 10^{-12}$ $32.4\%$ $4.8\times 10^{10}$
24 $1.7\times 10^{-17}$ $23.1\%$ $1.4\times 10^{16}$
32 $4.3\times 10^{-23}$ $25.0\%$ $5.8\times 10^{21}$
Table 9: Improvement over naive threshold approach.
Appendix G Nonce Range Rounding Details
Each party samples $\mathbf{y}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$. When $\gamma_{1}$ is not divisible by $|S|$, the sum $\mathbf{y}=\sum_{i\in S}\mathbf{y}_{i}$ has range $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor,|S|\cdot\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$, slightly smaller than standard $\gamma_{1},\gamma_{1}]^{n\ell}$.
Concrete Impact.
For ML-DSA-65 with $\gamma_{1}=524288$ and $|S|=17$: $\lfloor 524288/17\rfloor=30840$, so the actual range is [524280,524280][-524280,524280]. The “missing” probability mass is $(524288-524280)/524288<2\times 10^{-5}$ per coordinate.
Effect on Security.
The z-bound check is $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ where $\gamma_{1}-\beta=524092$. Since $524280>524092$, the reduced nonce range does not affect the pass probability. The Irwin-Hall security analysis uses the actual sampled range $N=2\lfloor\gamma_{1}/|S|\rfloor$, so all bounds account for this rounding.
Appendix H UC Security Framework
We recall the Universal Composability (UC) framework of Canetti [Canetti01, Canetti20]. Our presentation follows the conventions of [CLOS02, DPSZ12].
H.1 Execution Model
The UC framework considers three types of entities:
  • Environment $\mathcal{Z}$: An interactive Turing machine that provides inputs to parties, receives their outputs, and attempts to distinguish real from ideal executions.
  • Adversary $\mathcal{A}$: Controls corrupted parties and the network (scheduling message delivery).
  • Parties $P_{1},\ldots,P_{n}$: Execute the protocol or interact with ideal functionalities.
Definition H.1 (Real Execution).
In the real execution $\mathsf{REAL}_{\Pi,\mathcal{A},\mathcal{Z}}$, parties execute protocol $\Pi$. The environment $\mathcal{Z}$ provides inputs and receives outputs. The adversary $\mathcal{A}$ controls corrupted parties and sees all communication. The output is $\mathcal{Z}$’s final bit.
Definition H.2 (Ideal Execution).
In the ideal execution $\mathsf{IDEAL}_{\mathcal{F},\mathcal{S},\mathcal{Z}}$, parties interact with ideal functionality $\mathcal{F}$. A simulator $\mathcal{S}$ simulates the adversary’s view. The output is $\mathcal{Z}$’s final bit.
Definition H.3 (UC-Realization).
Protocol $\Pi$ UC-realizes functionality $\mathcal{F}$ if there exists a PPT simulator $\mathcal{S}$ such that for all PPT environments $\mathcal{Z}$:
$|\Pr[\mathsf{REAL}_{\Pi,\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F},\mathcal{S},\mathcal{Z}}=1]|\leq\mathsf{negl}(\kappa)$
Theorem H.4 (UC Composition [Canetti01]).
If $\Pi$ UC-realizes $\mathcal{F}$, then any protocol $\rho$ that uses $\mathcal{F}$ as a subroutine can securely use $\Pi$ instead. Formally, $\rho^{\mathcal{F}}$ and $\rho^{\Pi}$ are indistinguishable to any environment.
H.2 Hybrid Functionalities
We work in the $(\mathcal{F}_{\mathsf{SPDZ}},\mathcal{F}_{\mathsf{edaBits}},\mathcal{F}_{\mathsf{Rand}},\mathcal{F}_{\mathsf{Broadcast}})$-hybrid model, where:
Functionality H.5 ($\mathcal{F}_{\mathsf{SPDZ}}$ – Authenticated Secret Sharing [DPSZ12]).
Parameterized by field $\mathbb{F}_{q}$, parties $\mathcal{P}=\{P_{1},\ldots,P_{n}\}$, and MAC key $\alpha\in\mathbb{F}_{q}$.
Share: On input $(\mathsf{SHARE},\mathsf{sid},x)$ from dealer $P_{i}$:
  1. 1.
    Sample random shares $x]_{1},\ldots,[x]_{n}$ with $\sum_{j}[x]_{j}=x$
  2. 2.
    Compute MAC shares $\alpha x]_{j}$ for each party
  3. 3.
    Send $([x]_{j},[\alpha x]_{j})$ to each $P_{j}$
  4. 4.
    Send $(\mathsf{SHARED},\mathsf{sid})$ to adversary $\mathcal{S}$
Add: On input $(\mathsf{ADD},\mathsf{sid},\mathsf{id}_{a},\mathsf{id}_{b},\mathsf{id}_{c})$:
  1. 1.
    Each party locally computes $c]_{j}=[a]_{j}+[b]_{j}$ and $\alpha c]_{j}=[\alpha a]_{j}+[\alpha b]_{j}$
Multiply: On input $(\mathsf{MULT},\mathsf{sid},\mathsf{id}_{a},\mathsf{id}_{b},\mathsf{id}_{c})$:
  1. 1.
    Using Beaver triples, compute shares of $c=a\cdot b$ with valid MACs
  2. 2.
    Notify adversary of completion
Open: On input $(\mathsf{OPEN},\mathsf{sid},\mathsf{id}_{x})$:
  1. 1.
    Collect $x]_{j}$ from all parties
  2. 2.
    Verify MACs: check $\sum_{j}[\alpha x]_{j}=\alpha\cdot\sum_{j}[x]_{j}$
  3. 3.
    If verification fails, output $\bot$ to all parties
  4. 4.
    Otherwise, output $x=\sum_{j}[x]_{j}$ to all parties
Security: $\mathcal{F}_{\mathsf{SPDZ}}$ provides information-theoretic hiding of shares and computational integrity (MAC forgery probability $\leq 1/q$ per check).
Functionality H.6 ($\mathcal{F}_{\mathsf{edaBits}}$ – Extended Doubly-Authenticated Bits [EKMOZ20]).
On input $(\mathsf{EDABIT},\mathsf{sid},m)$ requesting $m$ edaBits:
For $i=1,\ldots,m$:
  1. 1.
    Sample $r_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1,\ldots,2^{k}-1\}$ for specified bit-length $k$
  2. 2.
    Create arithmetic sharing: $\langle r_{i}\rangle_{q}$ over $\mathbb{F}_{q}$
  3. 3.
    Create binary sharing: $\langle r_{i}\rangle_{2}=(\langle r_{i}^{(0)}\rangle,\ldots,\langle r_{i}^{(k-1)}\rangle)$ over $\mathbb{F}_{2}$
  4. 4.
    Distribute shares to parties with appropriate MAC tags
Output $\{(\langle r_{i}\rangle_{q},\langle r_{i}\rangle_{2})\}_{i=1}^{m}$ to all parties.
Security: The correlation $(r,r)$ is hidden from any strict subset of parties. MAC tags ensure integrity.
Remark H.7 (edaBits Range and Modular Reduction).
The edaBits functionality samples $r_{i}\in\{0,\ldots,2^{k}-1\}$ where $k=\lceil\log_{2}q\rceil=23$ for ML-DSA. Since $2^{23}=8388608>q=8380417$, a direct reduction $r_{i}\bmod q$ introduces a small bias: values in $\{0,\ldots,2^{23}-q-1\}=\{0,\ldots,8191\}$ are slightly more likely.
Statistical distance: The bias is $(2^{23}-q)/2^{23}=8191/8388608<2^{-10}$. For $m=1536$ coefficients, the total statistical distance is at most $m\cdot 2^{-10}<2^{1}$ by union bound, which is not negligible.
Mitigation: We use rejection sampling within edaBits generation: sample $r\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,\ldots,2^{k}-1\}$; if $r\geq q$, resample. The expected number of samples is $2^{k}/q<1.001$, adding negligible overhead. With this modification, $r\bmod q$ is perfectly uniform over $\mathbb{F}_{q}$, and Lemma L.6 applies exactly.
Functionality H.8 ($\mathcal{F}_{\mathsf{Binary}}$ – Binary Circuit Evaluation).
On input $(\mathsf{EVAL},\mathsf{sid},C,\{\langle x_{i}\rangle_{2}\})$ where $C$ is a binary circuit:
  1. 1.
    Reconstruct inputs $x_{i}$ from shares
  2. 2.
    Evaluate $y=C(x_{1},\ldots,x_{m})$
  3. 3.
    Create fresh sharing $\langle y\rangle_{2}$
  4. 4.
    Distribute to parties
Implemented via Beaver triples: for AND gate on shared bits $\langle x\rangle,\langle y\rangle$:
  1. 1.
    Use preprocessed triple $(\langle a\rangle,\langle b\rangle,\langle c\rangle)$ with $c=a\land b$
  2. 2.
    Open $d=x\oplus a$ and $e=y\oplus b$
  3. 3.
    Compute $\langle xy\rangle=\langle c\rangle\oplus(e\land\langle a\rangle)\oplus(d\land\langle b\rangle)\oplus(d\land e)$
Appendix I Complete Ideal Functionality for Profile P2
We define the ideal functionality $\mathcal{F}_{r_{0}\text{-check}}$ that captures the security requirements of the r0-check subprotocol.
Functionality I.1 ($\mathcal{F}_{r_{0}\text{-check}}$ – Secure r0-Check).
Parameters:
  • $n$: number of parties
  • $q=8380417$: field modulus
  • $\gamma_{2}=(q-1)/32=261888$: rounding parameter
  • $\beta=196$: bound parameter
  • $\mathsf{bound}=\gamma_{2}-\beta=261692$: acceptance threshold
  • $m=k\cdot 256=1536$: number of coefficients (for ML-DSA-65)
State:
  • $\mathsf{received}[\cdot]$: stores received inputs
  • $\mathsf{result}\in\{0,1,\bot\}$: computation result
Input Phase:
On receiving $(\mathsf{INPUT},\mathsf{sid},i,[\mathbf{w}^{\prime}]_{i})$ from party $P_{i}$ where $\mathbf{w}^{\prime}]_{i}\in\mathbb{Z}_{q}^{m}$:
  1. 1.
    If $P_{i}$ is corrupted, forward $\mathbf{w}^{\prime}]_{i}$ to simulator $\mathcal{S}$
  2. 2.
    Store $\mathsf{received}[i]\leftarrow[\mathbf{w}^{\prime}]_{i}$
  3. 3.
    If $|\{j:\mathsf{received}[j]\neq\bot\}|=n$:
    1. (a)
      Reconstruct: $\mathbf{w}^{\prime}_{j}=\sum_{i=1}^{n}\mathsf{received}[i][j]\bmod q$ for each $j\in[m]$
    2. (b)
      For each $j\in[m]$: compute centered representative $\tilde{w}^{\prime}_{j}\in(-q/2,q/2]$
    3. (c)
      Compute pass/fail for each coefficient:
      $\mathsf{pass}_{j}=\begin{cases}1&\text{if }|\tilde{w}^{\prime}_{j}\bmod 2\gamma_{2}|_{[-\gamma_{2},\gamma_{2})}<\mathsf{bound}\\ 0&\text{otherwise}\end{cases}$
      where $|x|_{[-\gamma_{2},\gamma_{2})}$ denotes the centered representative of $x\bmod 2\gamma_{2}$
    4. (d)
      Compute final result: $\mathsf{result}=\bigwedge_{j=1}^{m}\mathsf{pass}_{j}$
  4. 4.
    Send $(\mathsf{COMPUTED},\mathsf{sid},\mathsf{result})$ to $\mathcal{S}$
Output Phase:
On receiving $(\mathsf{DELIVER},\mathsf{sid})$ from $\mathcal{S}$:
  1. 1.
    Send $(\mathsf{OUTPUT},\mathsf{sid},\mathsf{result})$ to all honest parties
On receiving $(\mathsf{ABORT},\mathsf{sid})$ from $\mathcal{S}$:
  1. 1.
    Send $(\mathsf{OUTPUT},\mathsf{sid},\bot)$ to all honest parties
Leakage: Only the single bit $\mathsf{result}$ is revealed to $\mathcal{S}$. The functionality does not reveal:
  • The reconstructed value $\mathbf{w}^{\prime}$
  • Individual coefficients or their pass/fail status
  • Honest parties’ input shares
Remark I.2 (Security with Abort).
$\mathcal{F}_{r_{0}\text{-check}}$ provides security-with-abort: the adversary can cause the protocol to abort after learning the result bit. This is standard for dishonest-majority MPC and matches the security of underlying primitives (SPDZ, edaBits).
Appendix J Real Protocol $\Pi_{r_{0}}$
We formally specify the real protocol $\Pi_{r_{0}}$ that realizes $\mathcal{F}_{r_{0}\text{-check}}$.
Protocol J.1 ($\Pi_{r_{0}}$ – MPC-Based r0-Check).
Notation:
  • $\langle x\rangle_{q}$: SPDZ sharing of $x$ over $\mathbb{F}_{q}$
  • $\langle x\rangle_{2}$: secret sharing of $x$ over $\mathbb{F}_{2}$
  • $x]_{i}$: party $P_{i}$’s share of $x$
Preprocessing (offline phase):
  1. 1.
    Generate $m$ edaBits $\{(\langle r_{j}\rangle_{q},\langle r_{j}\rangle_{2})\}_{j=1}^{m}$ via $\mathcal{F}_{\mathsf{edaBits}}$
  2. 2.
    Generate Beaver triples for binary AND gates (for comparison circuits)
  3. 3.
    Distribute MAC key shares for SPDZ
Online Protocol:
Round 1-2: Share Exchange and Aggregation
Each party $P_{i}$ holds input $\mathbf{w}^{\prime}]_{i}\in\mathbb{Z}_{q}^{m}$ (their share of $\mathbf{w}^{\prime}=\mathbf{w}-c\mathbf{s}_{2}$).
  1. 1.
    Each $P_{i}$ creates SPDZ sharing of their input: invoke $\mathcal{F}_{\mathsf{SPDZ}}.\mathsf{SHARE}$ for each $\mathbf{w}^{\prime}]_{i}[j]$
  2. 2.
    Parties invoke $\mathcal{F}_{\mathsf{SPDZ}}.\mathsf{ADD}$ to compute $\langle\mathbf{w}^{\prime}_{j}\rangle_{q}=\sum_{i}\langle[\mathbf{w}^{\prime}]_{i}[j]\rangle_{q}$ for each $j$
Round 3-4: Masked Reveal (Arithmetic to Public)
For each coefficient $j\in[m]$:
  1. 1.
    Parties compute $\langle\mathsf{masked}_{j}\rangle_{q}=\langle\mathbf{w}^{\prime}_{j}\rangle_{q}+\langle r_{j}\rangle_{q}$ using local addition
  2. 2.
    Each party $P_{i}$ computes commitment $\mathsf{Com}_{i,j}=H([\mathsf{masked}_{j}]_{i}\|\mathsf{sid}\|j\|i)$
  3. 3.
    Broadcast all commitments (Round 3)
  4. 4.
    After receiving all commitments, broadcast openings $\mathsf{masked}_{j}]_{i}$ (Round 4)
  5. 5.
    Verify commitments; abort if any mismatch
  6. 6.
    Reconstruct $\mathsf{masked}_{j}=\sum_{i}[\mathsf{masked}_{j}]_{i}\bmod q$
Round 5: Arithmetic-to-Binary Conversion
For each coefficient $j\in[m]$:
  1. 1.
    Convert public $\mathsf{masked}_{j}$ to binary: $\mathsf{masked}_{j}^{(b)}=(\mathsf{masked}_{j}^{(0)},\ldots,\mathsf{masked}_{j}^{(\lceil\log_{2}q\rceil-1)})$
  2. 2.
    Compute binary sharing of $\mathbf{w}^{\prime}_{j}$ via subtraction circuit:
    $\langle\mathbf{w}^{\prime}_{j}\rangle_{2}=\mathsf{masked}_{j}^{(b)}\boxminus\langle r_{j}\rangle_{2}$
[…]
  3. 3.
    The subtraction circuit has depth $O(\log q)$ and uses $O(\log q)$ AND gates
Round 6-7: Binary Comparison
For each coefficient $j\in[m]$:
  1. 1.
    Compute sign bit and absolute value from $\langle\mathbf{w}^{\prime}_{j}\rangle_{2}$
  2. 2.
    Evaluate comparison circuit: $\langle\mathsf{pass}_{j}\rangle_{2}=(|\langle\mathbf{w}^{\prime}_{j}\rangle_{2}|<\mathsf{bound})$
  3. 3.
    The comparison uses $O(\log\gamma_{2})$ AND gates
Round 8: AND Aggregation and Output
  1. 1.
    Compute $\langle\mathsf{result}\rangle_{2}=\bigwedge_{j=1}^{m}\langle\mathsf{pass}_{j}\rangle_{2}$ via AND-tree of depth $\lceil\log_{2}m\rceil$
  2. 2.
    Open $\mathsf{result}$: each party broadcasts $\mathsf{result}]_{i}$
  3. 3.
    Reconstruct $\mathsf{result}=\bigoplus_{i}[\mathsf{result}]_{i}$
  4. 4.
    Output $\mathsf{result}$ to all parties
Appendix K Simulator Construction
We construct a simulator $\mathcal{S}$ that interacts with $\mathcal{F}_{r_{0}\text{-check}}$ and simulates the view of adversary $\mathcal{A}$ controlling corrupted parties $C\subsetneq[n]$.
Simulator K.1 ($\mathcal{S}$ for $\Pi_{r_{0}}$).
$\mathcal{S}$ internally runs $\mathcal{A}$ and simulates the honest parties’ messages.
Setup:
  1. 1.
    $\mathcal{S}$ receives the set $C$ of corrupted parties from $\mathcal{A}$
  2. 2.
    Let $H=[n]\setminus C$ be the set of honest parties
Simulating Preprocessing:
  1. 1.
    For edaBits involving corrupted parties: $\mathcal{S}$ receives $\mathcal{A}$’s shares from $\mathcal{F}_{\mathsf{edaBits}}$ simulator
  2. 2.
    For honest parties: sample random shares consistent with the corrupted parties’ view
  3. 3.
    Key property: $\mathcal{S}$ does not know the actual edaBit values; only the corrupted shares
Simulating Rounds 1-2 (Share Exchange):
  1. 1.
    For each honest party $P_{i}\in H$:
    1. (a)
      Sample random SPDZ shares $\tilde{s}_{i,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{F}_{q}$ for $j\in[m]$
    2. (b)
      Compute MAC shares using the SPDZ simulator (consistent with corrupted MAC key shares)
    3. (c)
      Send simulated shares to $\mathcal{A}$
  2. 2.
    Receive corrupted parties’ inputs from $\mathcal{A}$
  3. 3.
    Forward corrupted inputs to $\mathcal{F}_{r_{0}\text{-check}}$
  4. 4.
    For honest parties: send dummy inputs to $\mathcal{F}_{r_{0}\text{-check}}$ (the functionality only checks if all inputs received)
Simulating Rounds 3-4 (Masked Reveal):
This is the critical simulation step. We must simulate $\mathsf{masked}_{j}=\mathbf{w}^{\prime}_{j}+r_{j}$ without knowing $\mathbf{w}^{\prime}_{j}$.
  1. 1.
    Round 3 (Commitments):
    1. (a)
      For each honest party $P_{i}\in H$ and each $j\in[m]$:
    2. (b)
      Sample $\tilde{m}_{i,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{F}_{q}$ uniformly at random
    3. (c)
      Compute commitment $\widetilde{\mathsf{Com}}_{i,j}=H(\tilde{m}_{i,j}\|\mathsf{sid}\|j\|i)$
    4. (d)
      Send $\widetilde{\mathsf{Com}}_{i,j}$ to $\mathcal{A}$
  2. 2.
    Receive corrupted parties’ commitments from $\mathcal{A}$
  3. 3.
    Round 4 (Openings):
    1. (a)
      For each honest party $P_{i}\in H$: send $\tilde{m}_{i,j}$ to $\mathcal{A}$
    2. (b)
      Receive corrupted parties’ openings from $\mathcal{A}$
    3. (c)
      Verify corrupted commitments; if any fail, simulate abort
    4. (d)
      Compute simulated $\widetilde{\mathsf{masked}}_{j}=\sum_{i\in H}\tilde{m}_{i,j}+\sum_{i\in C}[\mathsf{masked}_{j}]_{i}\bmod q$
Justification for Round 3-4 Simulation:
In the real protocol:
$\mathsf{masked}_{j}=\mathbf{w}^{\prime}_{j}+r_{j}=\sum_{i}[\mathbf{w}^{\prime}]_{i}[j]+\sum_{i}[r_{j}]_{i}\bmod q$
Since $r_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{F}_{q}$ from $\mathcal{F}_{\mathsf{edaBits}}$ and is unknown to $\mathcal{A}$ (only shares are known), by the one-time pad property:
$\mathsf{masked}_{j}=\mathbf{w}^{\prime}_{j}+r_{j}\equiv\mathsf{Uniform}(\mathbb{F}_{q})$
Thus sampling $\tilde{m}_{i,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{F}_{q}$ for honest parties produces an identical distribution.
Simulating Round 5 (A2B Conversion):
  1. 1.
    The public value $\widetilde{\mathsf{masked}}_{j}$ is converted to binary (same as real)
  2. 2.
    For binary subtraction circuit: use $\mathcal{F}_{\mathsf{Binary}}$ simulator
  3. 3.
    Honest parties’ binary shares are simulated as random (justified by Beaver AND security)
Simulating Rounds 6-7 (Comparison):
  1. 1.
    For each AND gate in the comparison circuit:
    1. (a)
      The gate opens $(d,e)=(x\oplus a,y\oplus b)$ where $(a,b)$ are Beaver triple components
    2. (b)
      Sample $(\tilde{d},\tilde{e})\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{2}$ uniformly
    3. (c)
      This is indistinguishable from real by Lemma L.10
  2. 2.
    Simulate all comparison circuits this way
Simulating Round 8 (Output):
  1. 1.
    Wait to receive $\mathsf{result}\in\{0,1\}$ from $\mathcal{F}_{r_{0}\text{-check}}$
  2. 2.
    Backward simulation of AND tree:
    1. (a)
      The AND tree computes $\mathsf{result}=\bigwedge_{j}\mathsf{pass}_{j}$
    2. (b)
      Starting from known $\mathsf{result}$, assign intermediate values:
      • If $\mathsf{result}=1$: set all $\mathsf{pass}_{j}=1$
      • If $\mathsf{result}=0$: set at least one $\mathsf{pass}_{j}=0$ (adversary cannot determine which)
    3. (c)
      For each AND gate computing $z=x\land y$:
      • Given target $z$, sample $(x,y)$ consistent with $z$
      • The Beaver triple values $(d,e)$ revealed during computation are uniform regardless of $(x,y)$
  3. 3.
    Simulate honest parties’ output shares consistent with $\mathsf{result}$
  4. 4.
    Send to $\mathcal{A}$
Handling Abort:
If $\mathcal{A}$ causes any verification to fail (commitment mismatch, MAC failure):
  1. 1.
    $\mathcal{S}$ sends $(\mathsf{ABORT},\mathsf{sid})$ to $\mathcal{F}_{r_{0}\text{-check}}$
  2. 2.
    Simulate abort to all honest parties
Appendix L Hybrid Argument
We prove indistinguishability through a sequence of hybrid experiments.
Definition L.1 (Hybrid Experiments).
We define hybrids $\mathsf{H}_{0},\ldots,\mathsf{H}_{6}$ where $\mathsf{H}_{0}$ is the real execution and $\mathsf{H}_{6}$ is the ideal execution.
L.1 Hybrid $\mathsf{H}_{0}$: Real Execution
This is the real protocol execution $\mathsf{REAL}_{\Pi_{r_{0}},\mathcal{A},\mathcal{Z}}$:
  • Honest parties execute $\Pi_{r_{0}}$ with their true inputs
  • All cryptographic operations use real keys and randomness
  • $\mathcal{A}$ controls corrupted parties and sees all communication
L.2 Hybrid $\mathsf{H}_{1}$: Replace SPDZ with Ideal
Replace the SPDZ protocol with ideal functionality $\mathcal{F}_{\mathsf{SPDZ}}$.
Lemma L.2.
$|\Pr[\mathsf{H}_{0}=1]-\Pr[\mathsf{H}_{1}=1]|\leq\epsilon_{\mathsf{SPDZ}}$
Proof L.3.
By the UC-security of SPDZ [DPSZ12], there exists a simulator $\mathcal{S}_{\mathsf{SPDZ}}$ such that no environment can distinguish real SPDZ execution from ideal $\mathcal{F}_{\mathsf{SPDZ}}$ execution with advantage better than $\epsilon_{\mathsf{SPDZ}}$.
The security parameter is $\kappa$ and the MAC key is in $\mathbb{F}_{q}$ where $q\approx 2^{23}$. Per SPDZ security analysis:
  • MAC forgery probability per verification: $1/q\approx 2^{-23}$
  • With $\ell$ verifications: $\epsilon_{\mathsf{SPDZ}}\leq\ell/q$
For our protocol with $m=1536$ coefficients and constant verifications per coefficient:
$\epsilon_{\mathsf{SPDZ}}\leq O(m)/q=O(1536)/8380417<2^{-12}$
L.3 Hybrid $\mathsf{H}_{2}$: Replace edaBits with Ideal
Replace the edaBits protocol with ideal functionality $\mathcal{F}_{\mathsf{edaBits}}$.
Lemma L.4.
$|\Pr[\mathsf{H}_{1}=1]-\Pr[\mathsf{H}_{2}=1]|\leq\epsilon_{\mathsf{edaBits}}$
Proof L.5.
By the UC-security of edaBits [EKMOZ20], the protocol UC-realizes $\mathcal{F}_{\mathsf{edaBits}}$ with negligible distinguishing advantage.
The edaBits protocol security relies on:
  • SPDZ security for arithmetic shares (already replaced with ideal)
  • Cut-and-choose for binary consistency (statistical security)
With security parameter $\kappa$ and cut-and-choose parameter $B$:
$\epsilon_{\mathsf{edaBits}}\leq 2^{-B}+\epsilon^{\prime}_{\mathsf{SPDZ}}$
where $\epsilon^{\prime}_{\mathsf{SPDZ}}$ is the SPDZ security bound for the internal operations used in edaBits generation.
For $B=64$ and ML-DSA-65 parameters ($m=nk=1536$ coefficients, $q=8380417$): the cut-and-choose term $2^{-64}$ is negligible, and $\epsilon^{\prime}_{\mathsf{SPDZ}}\leq m/q<2^{-12}$. Thus $\epsilon_{\mathsf{edaBits}}<2^{-64}+2^{-12}<2^{-11.9}$.
L.4 Hybrid $\mathsf{H}_{3}$: Simulate Masked Values
Replace honest parties’ contributions to $\mathsf{masked}_{j}$ with uniformly random values.
Lemma L.6 (One-Time Pad over $\mathbb{F}_{q}$).
For any $x\in\mathbb{F}_{q}$ and uniformly random $r\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{F}_{q}$ independent of $x$:
$x+r\bmod q\equiv\mathsf{Uniform}(\mathbb{F}_{q})$
Proof L.7.
For any target $t\in\mathbb{F}_{q}$:
$\Pr[x+r=t]=\Pr[r=t-x]=\frac{1}{q}$
since $r$ is uniform and independent of $x$. Thus $x+r$ is uniformly distributed.
Lemma L.8.
$\Pr[\mathsf{H}_{2}=1]=\Pr[\mathsf{H}_{3}=1]$ (perfect indistinguishability)
Proof L.9.
In $\mathsf{H}_{2}$, for each coefficient $j$:
$\mathsf{masked}_{j}=\mathbf{w}^{\prime}_{j}+r_{j}\bmod q$
where $r_{j}$ is the edaBit value, uniformly distributed in $\mathbb{F}_{q}$ and independent of all other values (by $\mathcal{F}_{\mathsf{edaBits}}$ security).
By Lemma L.6, $\mathsf{masked}_{j}$ is uniformly distributed in $\mathbb{F}_{q}$, regardless of $\mathbf{w}^{\prime}_{j}$.
In $\mathsf{H}_{3}$, we sample $\widetilde{\mathsf{masked}}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{F}_{q}$ directly.
Both distributions are identical (uniform over $\mathbb{F}_{q}$), so:
$\Pr[\mathsf{H}_{2}=1]=\Pr[\mathsf{H}_{3}=1]$
This is statistical (in fact, perfect) indistinguishability, not computational.
L.5 Hybrid $\mathsf{H}_{4}$: Replace Binary Circuits with Ideal
Replace Beaver-triple-based AND gates with ideal $\mathcal{F}_{\mathsf{Binary}}$.
Lemma L.10 (Beaver AND Security).
Let $(a,b,c)$ be a Beaver triple with $a,b\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}$ uniform and $c=a\land b$. For any inputs $x,y\in\{0,1\}$, the opened values $(d,e)=(x\oplus a,y\oplus b)$ are uniformly distributed in $\{0,1\}^{2}$ and independent of $(x,y)$.
Proof L.11.
Since $a\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}$ is uniform and independent of $x$:
$\Pr[d=0]=\Pr[x\oplus a=0]=\Pr[a=x]=\frac{1}{2}$
Similarly for $e$. Independence of $d$ and $e$ follows from independence of $a$ and $b$.
For any $(d^{*},e^{*})\in\{0,1\}^{2}$:
$\Pr[(d,e)=(d^{*},e^{*})]=\Pr[a=x\oplus d^{*}]\cdot\Pr[b=y\oplus e^{*}]=\frac{1}{4}$
Thus $(d,e)\equiv\mathsf{Uniform}(\{0,1\}^{2})$, independent of $(x,y)$.
Lemma L.12.
$\Pr[\mathsf{H}_{3}=1]=\Pr[\mathsf{H}_{4}=1]$ (perfect indistinguishability)
Proof L.13.
Each AND gate reveals $(d,e)$ which is uniform by Lemma L.10. Replacing real AND computation with $\mathcal{F}_{\mathsf{Binary}}$ (which also produces uniform $(d,e)$) yields identical distributions.
The A2B conversion and comparison circuits consist entirely of:
  • XOR gates (no communication, locally computable)
  • AND gates (each revealing uniform $(d,e)$)
Since all revealed values have identical distributions in both hybrids:
$\Pr[\mathsf{H}_{3}=1]=\Pr[\mathsf{H}_{4}=1]$
Remark L.14 (A2B Correctness and Borrow Propagation).
The A2B conversion computes $\langle\mathbf{w}^{\prime}_{j}\rangle_{2}=\mathsf{masked}_{j}^{(b)}\boxminus\langle r_{j}\rangle_{2}$ where $\mathsf{masked}_{j}=\mathbf{w}^{\prime}_{j}+r_{j}\bmod q$. Two potential issues arise:
(1) Negative intermediate result: If $\mathsf{masked}_{j}<r_{j}$ (as integers), the subtraction would wrap around. However, since $\mathbf{w}^{\prime}_{j}=\mathsf{masked}_{j}-r_{j}\bmod q$, the binary circuit must compute modular subtraction. Standard approaches:
  • Extended precision: Work in $k+1$ bits where $k=\lceil\log_{2}q\rceil$, allowing for the borrow.
  • Add-then-subtract: Compute $(\mathsf{masked}_{j}+q)-r_{j}$ when detecting borrow, using a multiplexer.
Both yield $\mathbf{w}^{\prime}_{j}\bmod q$ correctly.
(2) edaBits range mismatch: As noted in Remark H.7, with rejection sampling, $r_{j}$ is uniform in $\{0,\ldots,q-1\}$, matching the field $\mathbb{F}_{q}$.
Security implication: These are correctness considerations, not security concerns. The security proof (Lemma above) shows that the adversary’s view—the $(d,e)$ values from Beaver triples—is identically distributed regardless of the underlying $\mathbf{w}^{\prime}_{j}$ values. Correctness ensures the final result $\mathsf{result}$ is computed correctly; security ensures the adversary learns only $\mathsf{result}$.
L.6 Hybrid $\mathsf{H}_{5}$: Backward Simulate AND Tree
Simulate the AND tree computation given only the final result bit.
Lemma L.15.
$\Pr[\mathsf{H}_{4}=1]=\Pr[\mathsf{H}_{5}=1]$ (perfect indistinguishability)
Proof L.16.
In $\mathsf{H}_{4}$, the AND tree computes $\mathsf{result}=\bigwedge_{j=1}^{m}\mathsf{pass}_{j}$ level by level, revealing Beaver $(d,e)$ values at each gate.
In $\mathsf{H}_{5}$, given $\mathsf{result}$ from $\mathcal{F}_{r_{0}\text{-check}}$, we backward-simulate:
Case $\mathsf{result}=1$: All $\mathsf{pass}_{j}=1$. For each AND gate $z=x\land y$ with $z=1$: both $x=y=1$. The simulator sets all leaf values to 1 and propagates upward.
Case $\mathsf{result}=0$: At least one $\mathsf{pass}_{j}=0$. The simulator must assign leaf values such that the AND tree outputs 0. We use the following procedure:
  1. 1.
    Pick a uniformly random index $j^{*}\in[m]$ and set $\mathsf{pass}_{j^{*}}=0$.
  2. 2.
    For all $j\neq j^{*}$, set $\mathsf{pass}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}$ uniformly at random.
  3. 3.
    Propagate values through the AND tree (any path containing a 0 produces 0).
This simulation is not identical to the real distribution of leaf values. However, the adversary’s view consists only of Beaver $(d,e)$ pairs, not the leaf values themselves. The following argument shows indistinguishability:
In both cases, the revealed $(d,e)$ values are uniform by Lemma L.10, regardless of the underlying $(x,y)$. The simulator samples uniform $(d,e)\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{2}$ for each AND gate, matching the real distribution exactly.
Key observation: The adversary sees only $(d,e)$ values, never the underlying bits $x,y$ (which remain secret-shared). Since $(d,e)$ is uniform independent of $(x,y)$, backward simulation is perfect.
L.7 Hybrid $\mathsf{H}_{6}$: Ideal Execution
This is the ideal execution $\mathsf{IDEAL}_{\mathcal{F}_{r_{0}\text{-check}},\mathcal{S},\mathcal{Z}}$ with simulator $\mathcal{S}$ from Section K.
Lemma L.17.
$\Pr[\mathsf{H}_{5}=1]=\Pr[\mathsf{H}_{6}=1]$
Proof L.18.
$\mathsf{H}_{5}$ is exactly the execution of simulator $\mathcal{S}$ interacting with $\mathcal{F}_{r_{0}\text{-check}}$:
  • SPDZ replaced with $\mathcal{F}_{\mathsf{SPDZ}}$: done in $\mathsf{H}_{1}$
  • edaBits replaced with $\mathcal{F}_{\mathsf{edaBits}}$: done in $\mathsf{H}_{2}$
  • Masked values simulated as random: done in $\mathsf{H}_{3}$
  • Binary circuits replaced with $\mathcal{F}_{\mathsf{Binary}}$: done in $\mathsf{H}_{4}$
  • AND tree backward-simulated: done in $\mathsf{H}_{5}$
This matches the simulator construction exactly. Thus $\mathsf{H}_{5}\equiv\mathsf{H}_{6}$.
Appendix M Main Theorem
Theorem M.1 (UC Security of Profile P2).
Protocol $\Pi_{r_{0}}$ UC-realizes $\mathcal{F}_{r_{0}\text{-check}}$ in the $(\mathcal{F}_{\mathsf{SPDZ}},\mathcal{F}_{\mathsf{edaBits}},\mathcal{F}_{\mathsf{Rand}},\mathcal{F}_{\mathsf{Broadcast}})$-hybrid model against static malicious adversaries corrupting up to n1n-1 parties, achieving security-with-abort.
Concretely, for any PPT environment $\mathcal{Z}$ and adversary $\mathcal{A}$:
$|\Pr[\mathsf{REAL}_{\Pi_{r_{0}},\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}\text{-check}},\mathcal{S},\mathcal{Z}}=1]|\leq\epsilon$
where
$\epsilon=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$
for ML-DSA-65 parameters.
Proof M.2.
By the hybrid argument (Section L):
$|\Pr[\mathsf{H}_{0}=1]-\Pr[\mathsf{H}_{6}=1]|$
$\leq\sum_{i=0}^{5}|\Pr[\mathsf{H}_{i}=1]-\Pr[\mathsf{H}_{i+1}=1]|$
$=|\Pr[\mathsf{H}_{0}=1]-\Pr[\mathsf{H}_{1}=1]|+|\Pr[\mathsf{H}_{1}=1]-\Pr[\mathsf{H}_{2}=1]|$
$\quad+|\Pr[\mathsf{H}_{2}=1]-\Pr[\mathsf{H}_{3}=1]|+|\Pr[\mathsf{H}_{3}=1]-\Pr[\mathsf{H}_{4}=1]|$
$\quad+|\Pr[\mathsf{H}_{4}=1]-\Pr[\mathsf{H}_{5}=1]|+|\Pr[\mathsf{H}_{5}=1]-\Pr[\mathsf{H}_{6}=1]|$
$\leq\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}+0+0+0+0$
$=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}$
The zero terms arise from perfect (statistical) indistinguishability in Hybrids 3, 4, 5, and 6.
Concrete bound: For ML-DSA-65 with $q=8380417$ and $m=1536$ coefficients:
  • $\epsilon_{\mathsf{SPDZ}}\leq O(m)/q<2^{-12}$
  • $\epsilon_{\mathsf{edaBits}}<2^{-12}$ (with cut-and-choose parameter $B=64$)
Thus $\epsilon<2^{-11}$, which is negligible in the security parameter.
Corollary M.3 (Full Protocol P2 Security).
The complete Profile P2 threshold ML-DSA protocol achieves:
  1. 1.
    EUF-CMA security under Module-SIS (inherited from single-signer ML-DSA, Theorem 4.2)
  2. 2.
    Privacy against coalitions of up to n1n-1 parties (by UC security of $\mathcal{F}_{r_{0}\text{-check}}$: only the result bit is leaked)
  3. 3.
    Composition security: The protocol can be securely composed with other UC-secure protocols
Proof M.4.
EUF-CMA: The threshold signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ has identical distribution to single-signer ML-DSA (masks cancel, Lemma 3.2). The security proof of Theorem 4.2 applies directly.
Privacy: By Theorem M.1, the r0-check reveals only the result bit $\mathsf{result}\in\{0,1\}$. Combined with the mask hiding lemma (Lemma 4.6), an adversary’s view can be simulated from:
  • Public key $\mathsf{pk}$
  • Corrupted shares $\{\mathsf{sk}_{i}\}_{i\in C}$
  • Final signature $\sigma$
  • r0-check result bits (one per attempt)
Since $\mathsf{result}$ is derivable from $\sigma$ (a valid signature implies $\mathsf{result}=1$), no additional information is leaked.
Composition: Follows from the UC composition theorem (Theorem H.4).
Appendix N Communication and Round Complexity
Theorem N.1 (Complexity of $\Pi_{r_{0}}$).
Protocol $\Pi_{r_{0}}$ achieves:
  1. 1.
    Online rounds: 8 (as specified in Protocol J.1)
  2. 2.
    Online communication: $O(n\cdot m\cdot\log q)$ bits
  3. 3.
    Offline preprocessing: $O(m)$ edaBits and $O(m\cdot\log q)$ Beaver triples
Proof N.2.
Rounds:
  • Rounds 1-2: SPDZ share exchange and aggregation
  • Rounds 3-4: Commit-then-open for masked reveal
  • Round 5: A2B conversion (can be merged with Round 6)
  • Rounds 6-7: Comparison circuits
  • Round 8: AND aggregation and output
Communication: Each of $m=1536$ coefficients requires:
  • SPDZ sharing: $O(n)$ field elements
  • Masked reveal: $O(n)$ field elements
  • Binary operations: $O(\log q)$ bits per AND gate, $O(\log q)$ AND gates per coefficient
Total: $O(n\cdot m\cdot\log q)\approx n\cdot 1536\cdot 23\approx 35000\cdot n$ bits per party.
Preprocessing:
  • $m$ edaBits (one per coefficient)
  • $O(m\cdot\log q)$ Beaver triples for A2B and comparison circuits
Component Online Rounds Communication/Party
SPDZ aggregation 2 $O(m\log q)$ bits
Masked reveal 2 $O(m\log q)$ bits
A2B conversion 1 $O(m\log q)$ bits
Comparison 2 $O(m\log\gamma_{2})$ bits
AND tree 1 $O(\log m)$ bits
Total 8 $\approx 30$ KB
Table 10: Detailed round and communication breakdown for Profile P2 r0-check.
Appendix O Technical Lemmas
This section provides rigorous proofs of technical lemmas used throughout the paper.
O.1 Mask Properties
Lemma O.1 (Mask Cancellation – Formal Statement).
Let $S\subseteq[N]$ be a signing set with $|S|\geq 2$. For any domain separator $\mathsf{dom}$ and PRF $\mathsf{PRF}:\{0,1\}^{256}\times\{0,1\}^{*}\to R_{q}^{d}$, define:
$\mathbf{m}_{i}=\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{dom})$
Then $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}\in R_{q}^{d}$.
Proof O.2.
We prove this by showing each PRF term appears exactly once with coefficient +1+1 and once with coefficient 1-1.
For any pair $(i,j)\in S\times S$ with $i<j$, the term $\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})$ appears:
  1. 1.
    In $\mathbf{m}_{i}$ with coefficient +1+1 (from the sum over $k>i$, taking $k=j$)
  2. 2.
    In $\mathbf{m}_{j}$ with coefficient 1-1 (from the sum over $k<j$, taking $k=i$)
Thus:
$\sum_{i\in S}\mathbf{m}_{i}$ $=\sum_{i\in S}\left(\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{dom})\right)$
$=\sum_{\begin{subarray}{c}(i,j)\in S\times S\\ i<j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})-\sum_{\begin{subarray}{c}(i,j)\in S\times S\\ i<j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})$
$=\mathbf{0}$
Lemma O.3 (Mask Hiding – Formal Statement).
Let $C\subsetneq S$ be the set of corrupted parties in signing set $S$, with $|S\setminus C|\geq 2$. For any honest party $j\in S\setminus C$, the masked value:
$\mathbf{U}_{j}=\mathbf{y}_{j}+\lambda_{j}c\mathbf{s}_{1,j}+\mathbf{m}_{j}$
is computationally indistinguishable from uniform in $R_{q}^{\ell}$, given the adversary’s view consisting of:
  • Corrupted shares $\{\mathbf{s}_{1,i},\mathbf{s}_{2,i}\}_{i\in C}$
  • Corrupted seeds $\{\mathsf{seed}_{i,k}\}_{i\in C,k\in S}$
  • All protocol messages
Proof O.4.
Decompose $\mathbf{m}_{j}$ based on counterparty corruption status:
$\mathbf{m}_{j}$ $=\underbrace{\sum_{k\in C\cap S:k>j}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})-\sum_{k\in C\cap S:k<j}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{dom})}_{\mathbf{m}_{j}^{\text{known}}:\text{ computable by }\mathcal{A}}$
$\quad+\underbrace{\sum_{k\in(S\setminus C)\setminus\{j\}:k>j}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})-\sum_{k\in(S\setminus C)\setminus\{j\}:k<j}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{dom})}_{\mathbf{m}_{j}^{\text{unknown}}:\text{ unknown to }\mathcal{A}}$
Since $|S\setminus C|\geq 2$, there exists at least one honest party $k^{\prime}\in(S\setminus C)\setminus\{j\}$.
Claim: $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{dom})$ (assuming WLOG $j<k^{\prime}$) is computationally indistinguishable from uniform in $R_{q}^{\ell}$.
Proof of Claim:
  1. 1.
    The seed $\mathsf{seed}_{j,k^{\prime}}$ is shared only between honest parties $j$ and $k^{\prime}$
  2. 2.
    $\mathcal{A}$ does not know $\mathsf{seed}_{j,k^{\prime}}$ (parties $j,k^{\prime}\notin C$)
  3. 3.
    The domain separator $\mathsf{dom}$ includes a fresh nonce, so this is a fresh PRF evaluation
  4. 4.
    By PRF security, $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{dom})$ is indistinguishable from $\mathbf{r}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$
Conclusion:
$\mathbf{U}_{j}$ $=\underbrace{\mathbf{y}_{j}+\lambda_{j}c\mathbf{s}_{1,j}+\mathbf{m}_{j}^{\text{known}}}_{\text{fixed from }\mathcal{A}\text{'s view}}+\underbrace{\mathbf{m}_{j}^{\text{unknown}}}_{\text{contains uniform-looking term}}$
Since $\mathbf{m}_{j}^{\text{unknown}}$ contains at least one term indistinguishable from uniform, and uniform plus fixed equals uniform:
$\mathbf{U}_{j}\stackrel{{\scriptstyle c}}{{\approx}}\mathsf{Uniform}(R_{q}^{\ell})$
The distinguishing advantage is bounded by the PRF advantage:
$\mathsf{Adv}^{\text{mask-hiding}}\leq|S\setminus C|\cdot\epsilon_{\mathsf{PRF}}$
where we use a hybrid over all honest-honest PRF terms.
O.2 Irwin-Hall Distribution Analysis
We provide a self-contained analysis of the security implications of using Irwin-Hall distributed nonces.
Definition O.5 (Discrete Irwin-Hall Distribution).
Let $X_{1},\ldots,X_{n}$ be independent uniform random variables on the discrete interval $\{-B,-B+1,\ldots,B-1,B\}$. The sum $Y=\sum_{i=1}^{n}X_{i}$ follows the discrete Irwin-Hall distribution $\mathsf{IH}(n,B)$.
Lemma O.6 (Irwin-Hall Moments).
For $Y\sim\mathsf{IH}(n,B)$:
  1. 1.
    $\mathbb{E}[Y]=0$
  2. 2.
    $\mathsf{Var}(Y)=n\cdot\frac{(2B+1)^{2}-1}{12}=n\cdot\frac{B(B+1)}{3}$
  3. 3.
    For large $n$, $Y$ is approximately Gaussian by the Central Limit Theorem
Proof O.7.
Mean: Each $X_{i}$ has mean $\mathbb{E}[X_{i}]=0$ by symmetry. Thus $\mathbb{E}[Y]=\sum_{i}\mathbb{E}[X_{i}]=0$.
Variance: For uniform on $\{-B,\ldots,B\}$:
$\mathsf{Var}(X_{i})$ $=\mathbb{E}[X_{i}^{2}]=\frac{1}{2B+1}\sum_{k=-B}^{B}k^{2}=\frac{1}{2B+1}\cdot\frac{2B(B+1)(2B+1)}{6}=\frac{B(B+1)}{3}$
By independence, $\mathsf{Var}(Y)=n\cdot\mathsf{Var}(X_{1})=n\cdot\frac{B(B+1)}{3}$.
CLT: Standard application of Lindeberg-Lévy.
Definition O.8 (Rényi Divergence).
For distributions $P,Q$ over $\Omega$ with $P\ll Q$, the Rényi divergence of order $\alpha>1$ is:
$R_{\alpha}(P\|Q)=\left(\sum_{x\in\Omega}P(x)^{\alpha}Q(x)^{1-\alpha}\right)^{1/(\alpha-1)}$
Definition O.9 (Smooth Rényi Divergence [Raccoon2024]).
For $\epsilon\geq 0$:
$R_{\alpha}^{\epsilon}(P\|Q)=\min_{\begin{subarray}{c}P^{\prime}:\Delta(P^{\prime},P)\leq\epsilon\\ Q^{\prime}:\Delta(Q^{\prime},Q)\leq\epsilon\end{subarray}}R_{\alpha}(P^{\prime}\|Q^{\prime})$
where $\Delta$ denotes statistical distance.
Lemma O.10 (Security Reduction via Rényi Divergence [BLL+15, Raccoon2024]).
Let $P$ be the distribution of a cryptographic scheme’s output under distribution $D_{1}$, and $Q$ be the distribution under $D_{2}$. For any event $E$ (e.g., adversary succeeds):
$P(E)\leq(Q(E)+\epsilon)^{(\alpha-1)/\alpha}\cdot R_{\alpha}^{\epsilon}(P\|Q)+\epsilon$
Theorem O.11 (Irwin-Hall Security for Threshold ML-DSA).
Let $\Pi_{\mathsf{IH}}$ be threshold ML-DSA with Irwin-Hall nonces (sum of $|S|$ uniforms), and $\Pi_{\mathsf{U}}$ be the hypothetical scheme with uniform nonces. Then for any PPT adversary $\mathcal{A}$:
$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}(\mathcal{A})\leq R_{2}^{\epsilon}\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}(\mathcal{A})+q_{S}\cdot\epsilon$
where $q_{S}$ is the number of signing queries, and $R_{2}^{\epsilon},\epsilon$ are computed below.
Proof O.12.
We analyze the distribution of the signature response $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$.
In $\Pi_{\mathsf{U}}$ (hypothetical): $\mathbf{y}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1},\ldots,\gamma_{1}\}^{n\ell}$ uniform. Thus $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ is a shifted uniform.
In $\Pi_{\mathsf{IH}}$ (real): $\mathbf{y}=\sum_{i\in S}\mathbf{y}_{i}$ where each $\mathbf{y}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$.
For each coordinate, we compare:
  • $P$: Irwin-Hall distribution shifted by $(c\mathbf{s}_{1})_{j}$
  • $Q$: Uniform distribution shifted by $(c\mathbf{s}_{1})_{j}$
The key insight is that the shift is the same in both cases. We bound $R_{2}^{\epsilon}(P\|Q)$.
Applying Lemma 4.2 of [Raccoon2024]:
The Raccoon signature scheme [Raccoon2024] faces an analogous problem: threshold signers contribute nonce shares that sum to a non-uniform (Irwin-Hall) distribution, and security requires bounding the statistical distance from uniform. Lemma 4.2 of [Raccoon2024] provides a Rényi divergence bound for sums of discrete uniforms with bounded shifts. This lemma is directly applicable to our setting because:
  1. 1.
    Both schemes use additive secret sharing of nonces across threshold parties.
  2. 2.
    The shift arises from multiplying the challenge by the secret key share, bounded by $\|c\mathbf{s}_{1}\|_{\infty}\leq\beta$.
  3. 3.
    The discrete uniform ranges and modular arithmetic are structurally identical.
For $n=|S|$ uniforms on range $N=2\lfloor\gamma_{1}/|S|\rfloor$ with shift $c=\|(c\mathbf{s}_{1})_{j}\|_{\infty}\leq\beta$:
$R_{2}^{\epsilon}\leq 1+\frac{n^{2}c^{2}}{N^{2}}+O(1/n!)$
For ML-DSA-65 parameters:
  • $\gamma_{1}=2^{19}=524288$
  • $\beta=196$
  • $|S|=n$ (number of signers)
$R_{2}^{\epsilon}$ $\leq 1+\frac{n^{2}\cdot\beta^{2}}{(2\gamma_{1}/n)^{2}}=1+\frac{n^{4}\cdot\beta^{2}}{4\gamma_{1}^{2}}$
$=1+\frac{n^{4}\cdot 196^{2}}{4\cdot(524288)^{2}}=1+\frac{n^{4}\cdot 38416}{1.099\times 10^{12}}$
Concrete values:
$|S|$ $R_{2}^{\epsilon}-1$ Security loss
4 $8.9\times 10^{-8}$ Negligible
9 $2.3\times 10^{-5}$ Negligible
17 $2.9\times 10^{-3}$ $0.3\%$
25 $\times 10^{-2} <1.4<1.4\%$
33 $\times 10^{-2} <4.1<4.1\%$
The tail probability $\epsilon$ satisfies [Raccoon2024, Lemma 4.2]:
$\epsilon<\left(\frac{2e\cdot\beta}{N}\right)^{n}/\sqrt{2\pi n}$
For $|S|=17$: $\epsilon<(2e\cdot 196/61681)^{17}/\sqrt{34\pi}<10^{-30}$.
Conclusion: Over $q_{S}=2^{64}$ signing queries with $|S|\leq 33$:
$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}$ $\leq R_{2}^{\epsilon}\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+q_{S}\cdot\epsilon$
$\leq 1.05\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+2^{64}\cdot 10^{-30}$
$\leq 1.05\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+2^{-36}$
The multiplicative factor ($\leq 1.05$ for $|S|\leq 33$, or $\leq 1.003$ for $|S|\leq 17$) represents at most a 5% security loss, which is negligible for cryptographic purposes. The additive term ($2^{-36}$) is also negligible.
O.3 Lagrange Coefficient Bounds
Lemma O.13 (Lagrange Coefficient Characterization).
For evaluation points $S=\{i_{1},\ldots,i_{t}\}\subseteq[N]$ with $|S|=t$, the Lagrange coefficient for point $i_{k}$ is:
$\lambda_{i_{k}}=\prod_{j\in S\setminus\{i_{k}\}}\frac{j}{j-i_{k}}\bmod q$
Lemma O.14 (Coefficient Magnitude).
For consecutive evaluation points $S=\{1,2,\ldots,T\}$:
$\lambda_{i}=(-1)^{i-1}\cdot\binom{T}{i}\cdot\frac{T!}{i\cdot(T-i)!}=(-1)^{i-1}\cdot\binom{T}{i}$
The maximum magnitude is achieved at $i=\lceil T/2\rceil$:
$\max_{i}|\lambda_{i}|=\binom{T}{\lceil T/2\rceil}\approx\frac{2^{T}}{\sqrt{\pi T/2}}$
Proof O.15.
Standard combinatorial identity. The approximation follows from Stirling’s formula.
Corollary O.16 (Threshold Bound for Short Coefficients).
For ML-DSA-65 with $q=8380417\approx 2^{23}$:
  • For $T\leq 22$: $\max_{i}|\lambda_{i}|<q$, so Lagrange coefficients “fit” in $\mathbb{Z}_{q}$
  • For $T\geq 23$: $\max_{i}|\lambda_{i}|\geq q$, so coefficients wrap around mod $q$
This aligns with the Ball-Çakan-Malkin bound [BCM21]: short-coefficient LSSS cannot achieve $T>O(\log q)$.
O.4 Rejection Sampling Correctness
Lemma O.17 (Z-Bound Check Correctness).
Let $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ where:
  • $\mathbf{y}\in[-\gamma_{1},\gamma_{1}]^{n\ell}$
  • $c\in\mathcal{C}$ with $\|c\|_{1}=\tau$ (sparse polynomial)
  • $\mathbf{s}_{1}\in[-\eta,\eta]^{n\ell}$
Then $\|\mathbf{z}\|_{\infty}\leq\gamma_{1}+\tau\eta=\gamma_{1}+\beta$.
The check $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ ensures:
$\|\mathbf{y}\|_{\infty}\leq\|\mathbf{z}\|_{\infty}+\|c\mathbf{s}_{1}\|_{\infty}<(\gamma_{1}-\beta)+\beta=\gamma_{1}$
which is necessary for security (prevents leaking $\mathbf{s}_{1}$ via $\mathbf{z}-\mathbf{y}$).
Lemma O.18 (R0-Check Correctness).
The r0-check $\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\|_{\infty}<\gamma_{2}-\beta$ ensures that the hint $\mathbf{h}$ correctly recovers $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{w},2\gamma_{2})$ during verification.
Specifically, if the check passes, then:
$\mathsf{UseHint}(\mathbf{h},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d})=\mathbf{w}_{1}$
Proof O.19.
By the ML-DSA specification [FIPS204], the hint mechanism ensures that rounding errors in $\mathbf{A}\mathbf{z}-c\mathbf{t}$ are corrected when the low bits are bounded by $\gamma_{2}-\beta$. See [DKL18, Section 3.4] for detailed analysis.
O.5 Challenge Invertibility
Lemma O.20 (Challenge Invertibility in $R_{q}$).
Let $q=8380417$ and $R_{q}=\mathbb{Z}_{q}[X]/(X^{256}+1)$. For a uniformly random ML-DSA challenge $c\in\mathcal{C}$ (sparse polynomial with $\tau=49$ coefficients in $\{-1,+1\}$):
$\Pr[c\text{ is invertible in }R_{q}]\geq 1-256/q>1-2^{-15}$
Proof O.21.
Since $q\equiv 1\pmod{512}$, the polynomial $X^{256}+1$ splits completely over $\mathbb{Z}_{q}$:
$X^{256}+1=\prod_{i=0}^{255}(X-\omega^{2i+1})$
where $\omega$ is a primitive $512$-th root of unity modulo $q$.
A polynomial $c\in R_{q}$ is invertible iff $c(\omega^{2i+1})\neq 0$ for all $i\in[256]$.
For each evaluation point $\omega^{2i+1}$:
$c(\omega^{2i+1})=\sum_{j\in S}\sigma_{j}\cdot\omega^{(2i+1)j}$
where $S$ is the set of 49 nonzero positions and $\sigma_{j}\in\{-1,+1\}$.
By a union bound over 256 evaluation points, each with failure probability $\approx 1/q$:
$\Pr[c\text{ not invertible}]\leq 256/q<2^{-15}$
Monte Carlo verification with $10^{4}$ samples confirms 100% of random challenges are invertible.
Corollary O.22 (Security Implication).
Since virtually all challenges $c$ are invertible, leaking $(c,c\mathbf{s}_{2})$ for any single signing attempt allows recovery of $\mathbf{s}_{2}=c^{-1}(c\mathbf{s}_{2})$. Thus $c\mathbf{s}_{2}$ must be treated as secret material equivalent to $\mathbf{s}_{2}$.

### v3

@@
 Abstract
-We present masked Lagrange reconstruction, a technique that enables threshold ML-DSA (FIPS 204) with arbitrary thresholds $T$ while producing standard 3.3 KB signatures verifiable by unmodified FIPS 204 implementations.
-Concurrent approaches have limitations: Bienstock et al.
-(ePrint 2025/1163) achieve arbitrary $T$ but require honest-majority and 37–136 rounds; Celi et al.
-(ePrint 2026/013) achieve dishonest-majority but are limited to $T\leq 6$.
-Our technique addresses the barrier that Lagrange coefficients grow as $\Theta(q)$ for moderate $T$, making individual contributions too large for ML-DSA’s rejection sampling.
-Unlike ECDSA threshold schemes where pairwise masks suffice for correctness, ML-DSA requires solving three additional challenges absent in prior work: (1) rejection sampling on $\|\mathbf{z}\|_{\infty}$ must still pass after masking, (2) the r0-check exposes $c\mathbf{s}_{2}$ enabling key recovery if unprotected, and (3) the resulting Irwin-Hall nonce distribution must preserve EUF-CMA security.
-We solve all three.
-We instantiate this technique in three deployment profiles with full security proofs.
-Profile P1 (TEE-assisted) achieves 3-round signing with a trusted coordinator, with EUF-CMA security under Module-SIS (Theorem 4.2).
-Profile P2 (fully distributed) eliminates hardware trust via MPC in 8 rounds, achieving UC security against malicious adversaries corrupting up to n1n-1 parties (Theorem 6.2).
-Profile P3 (2PC-assisted) uses lightweight 2PC for the r0-check in 3–5 rounds, achieving UC security under a 1-of-2 CP honest assumption (Theorem 6.5) with the best empirical performance (249ms).
-Our scheme requires $|S|\geq T+1$ signers and achieves success rates of 23–32%, matching single-signer ML-DSA.
+[…]
+Our technique overcomes the barrier that Lagrange coefficients grow as $\Theta(q)$ for moderate $T$, making individual contributions too large for rejection sampling.
+We introduce pairwise-canceling masks that hide these coefficients while preserving correctness.
+Unlike ECDSA threshold schemes, ML-DSA presents three additional challenges: the $\|\mathbf{z}\|_{\infty}$ check must pass after masking, the r0-check must not leak $c\mathbf{s}_{2}$, and the resulting Irwin-Hall nonce distribution must preserve EUF-CMA security.
+We solve all three with complete UC proofs.
+We instantiate this technique in three deployment profiles.
+Profile P1 uses a TEE coordinator (3 rounds, 6ms).
+Profile P2 eliminates hardware trust via MPC (5 rounds, 22ms for small thresholds, up to 194ms for $T=16$).
+Profile P3+ uses lightweight 2PC with semi-asynchronous signing where signers precompute nonces offline and respond within a time window (2 logical rounds, 22ms), similar to FROST.
+Our Rust implementation scales from 2-of-3 to 32-of-45 thresholds.
+P1 and P3+ achieve sub-100ms latency for all configurations; P2 trades higher latency for eliminating hardware trust.
+Success rates of $\approx$21–45% are comparable to single-signer ML-DSA’s $\approx$20–25%.
 keywords:
@@
 Our scheme requires a two-honest requirement: signing sets must have size $|S|\geq T+1$ to ensure at least two honest parties for mask privacy.
-This addresses the threshold gap for $T\in[9,32]$ while remaining practical—configurations like $(3,5)$, $(11,15)$, and $(16,31)$ naturally satisfy $N\geq T+1$.
-The scheme achieves practical success rates comparable to single-signer ML-DSA (theoretical baseline $\approx 20\%$, observed $3232\%$ due to the more concentrated Irwin-Hall nonce distribution) and is fully verifier-compatible with FIPS 204: produced signatures are accepted by unmodified FIPS 204 verifiers.
+This addresses the threshold gap for $T\in[9,32]$ while remaining practical; configurations like $(3,5)$, $(11,15)$, and $(16,21)$ naturally satisfy $N\geq T+1$.
+The scheme achieves practical success rates comparable to single-signer ML-DSA (theoretical baseline $\approx 20\%$, observed $\approx$21–45% due to the more concentrated Irwin-Hall nonce distribution) and is fully verifier-compatible with FIPS 204: produced signatures are accepted by unmodified FIPS 204 verifiers.
 While our signer-side protocol is interactive and induces an Irwin-Hall nonce distribution rather than uniform, this affects only the signing process, not verification.
@@
 $\sum_{i\in S}\mathbf{U}_{i}=\sum_{i}\mathbf{y}_{i}+c\cdot\sum_{i}\lambda_{i}\mathbf{s}_{1,i}+\sum_{i}\mathbf{m}_{i}=\mathbf{y}+c\cdot\mathbf{s}_{1}+\mathbf{0}=\mathbf{z}$
-which correctly reconstructs the signature vector $\mathbf{z}$ and passes rejection sampling with probability matching single-signer ML-DSA (theoretical $\approx 20\%$, observed $3232\%$ for ML-DSA-65).
-We note that while the aggregated nonce $\mathbf{y}=\sum_{i}\mathbf{y}_{i}$ follows an Irwin-Hall distribution rather than uniform (see Definition O.5), this does not affect EUF-CMA security and only introduces a quantifiable statistical difference that we analyze in Section 3.
+which correctly reconstructs the signature vector $\mathbf{z}$ and passes rejection sampling with probability comparable to single-signer ML-DSA (theoretical $\approx 20\%$, observed $\approx$21–45% for ML-DSA-65).
+We note that while the aggregated nonce $\mathbf{y}=\sum_{i}\mathbf{y}_{i}$ follows an Irwin-Hall distribution rather than uniform (see Definition Q.5), this does not affect EUF-CMA security and only introduces a quantifiable statistical difference that we analyze in Section 3.
 Performance.
-Our implementation achieves practical signing times: approximately 1 second average for $(3,5)$-threshold and 4 seconds for $(16,31)$-threshold on commodity hardware (including retries due to rejection sampling), using NTT-accelerated polynomial arithmetic.
+Our optimized Rust implementation achieves practical signing times: 5.8ms (P1-TEE), 21.5ms (P2-MPC), and 22.1ms (P3+-Async) for $(3,5)$-threshold on commodity hardware (including retries).
+P1 and P3+ remain under 100ms for all configurations up to $(32,45)$; P2 trades higher latency ($\sim$130–194ms for $T\geq 16$) for eliminating hardware trust.
 The communication cost is approximately 12.3 KB per party per signing attempt.
@@
 A fully untrusted coordinator where the r0-check is computed via an MPC subprotocol using SPDZ [DPSZ12] and edaBits [EKMOZ20].
-We prove UC security against malicious adversaries corrupting up to n1n-1 parties in 8 online rounds (Theorem 6.2).
-Profile P3 (2PC-Assisted):
-A practical middle ground where two designated Computation Parties (CPs) jointly evaluate the r0-check via lightweight 2PC.
-Motivated by recent round-reduction efforts in threshold ECDSA [LLZD25], we pursue a similar goal for the lattice setting; while their Non-Interactive Multiplication technique does not directly apply to ML-DSA’s r0-check structure, we achieve analogous round improvements via 2PC.
-We prove UC security under the assumption that at least one CP is honest (Theorem 6.5)—weaker than TEE trust, more efficient than full MPC.
-Achieves the best empirical performance with 3–5 rounds and 249ms signing time.
+Using combiner-mediated commit-then-open and constant-depth comparison circuits, we achieve 5 online rounds (reduced from 8).
+We prove UC security against malicious adversaries corrupting up to n1n-1 parties (Theorem 6.2).
+Profile P3+ (Semi-Async 2PC):
+Designed for human-in-the-loop authorization.
+Two designated Computation Parties (CPs) jointly evaluate the r0-check via 2PC, but signers participate semi-asynchronously: they precompute nonces offline and respond within a time window (e.g., 5 minutes) rather than synchronizing for multiple rounds.
+This achieves 2 logical rounds (1 signer step + 1 server round).
+We prove UC security under the assumption that at least one CP is honest (Theorem 6.5).
+The “Async” column in Table 1 distinguishes our work from concurrent approaches, which require synchronous multi-round signer participation.
 1.3 Technical Overview
@@
 Table 1 summarizes threshold ML-DSA schemes, comparing arbitrary threshold support, dishonest-majority security, round complexity, and FIPS 204 compatibility.
-Scheme $T$ Size Rnd FIPS Dishon. UC
-del Pino-Niot [dPN25] $\leq 8$ 3.3 KB 3 $\times$
-Noise Flooding [BCDG21] $\infty$ 17 KB 3 $\times$ $\times$
-Ringtail [BMMSZ25] $\infty$ 13 KB 2 $\times$ $\times$
+Scheme $T$ Size Rnd FIPS Dishon. UC Async
+del Pino-Niot [dPN25] $\leq 8$ 3.3 KB 3 $\times$ $\times$
+Noise Flooding [BCDG21] $\infty$ 17 KB 3 $\times$ $\times$ $\times$
+Ringtail [BMMSZ25] $\infty$ 13 KB 2 $\times$ $\times$ $\times$
 Concurrent Work (2025–2026)
-Bienstock et al. [BDEPT25] $\infty$ 3.3 KB 37–136 $\times$
-Celi et al. [CDENP26] $\leq 6$ 3.3 KB 6 $\times$
-Trilithium [Trilithium25] 2 3.3 KB 3 $\times$
+Bienstock et al. [BDEPT25] $\infty$ 3.3 KB 37–136 $\times$ $\times$
+Celi et al. [CDENP26] $\leq 6$ 3.3 KB 6 $\times$ $\times$
+Trilithium [Trilithium25] 2 3.3 KB 3 $\times$ $\times$
 This work (P1) $\infty$ 3.3 KB 3 (TEE) $\times$
-This work (P2) $\infty$ 3.3 KB 8
-This work (P3) $\infty$ 3.3 KB 3–5 (1/2 CP)
-Table 1: Comparison of threshold ML-DSA schemes. “Dishon.” = dishonest-majority security; “(TEE)” and “(1/2 CP)” indicate trust assumptions.
+This work (P2) $\infty$ 3.3 KB 5 $\times$
+This work (P3+) $\infty$ 3.3 KB 2 (1/2 CP)
+Table 1: Comparison of threshold ML-DSA schemes.
+✓= supported, $\times$ = not supported.
+“Dishon.” = dishonest-majority security (tolerates n1n-1 corruptions); Bienstock et al.
+require honest majority.
+“(TEE)” and “(1/2 CP)” indicate additional trust assumptions beyond threshold.
+“Async” = semi-asynchronous signer participation.
+P3+ requires 1 signer step + 1 server round = 2 logical rounds.
+P1 UC security assumes ideal TEE (Theorem 4.10).
 Three concurrent works address threshold ML-DSA with FIPS 204 compatibility.
@@
 Trilithium [Trilithium25] is UC-secure but restricted to 2 parties in a server-phone architecture.
-Profile P2 combines arbitrary thresholds, dishonest-majority security, constant rounds (8 rounds), UC security, and FIPS 204 compatibility—a combination not achieved by prior work.
-Profile P3 also achieves UC security (Theorem 6.5) under the weaker 1-of-2 CP honest assumption, with the best performance (249ms, 3–5 rounds).
-We contribute: (1) Masked Lagrange reconstruction, adapting pairwise-canceling masks [CGGMP20] to lattices while addressing rejection sampling, r0-check key leakage, and Irwin-Hall nonce security; (2) Three deployment profiles (P1-TEE, P2-MPC, P3-2PC) with different trust/performance trade-offs; (3) Security proofs reducing to Module-SIS/LWE, including UC security for P2 and P3; (4) A reference implementation with benchmarks from $(3,5)$ to $(16,31)$ threshold configurations.
+Profile P2 combines arbitrary thresholds, dishonest-majority security, constant rounds (5 rounds), UC security, and FIPS 204 compatibility, a combination not achieved by prior work.
+Profile P3+ achieves UC security (Theorem 6.5) under the 1-of-2 CP honest assumption, with semi-asynchronous signer participation (2 logical rounds).
+We contribute: (1) Masked Lagrange reconstruction, adapting pairwise-canceling masks [CGGMP20] to lattices while addressing rejection sampling, r0-check key leakage, and Irwin-Hall nonce security; (2) Three deployment profiles (P1-TEE, P2-MPC, P3+-2PC) with different trust/performance trade-offs, including the novel semi-async model for P3+; (3) Security proofs reducing to Module-SIS/LWE, including complete UC security proofs for P2 and P3+; (4) A reference implementation with benchmarks from $(2,3)$ to $(32,45)$ threshold configurations.
 1.5 Paper Organization
@@
 3.2 Threat Model and Deployment Profiles
-We consider a $(T,N)$-threshold setting where an adversary may corrupt up to T1T-1 parties. Table 2 summarizes our two deployment profiles.
+We consider a $(T,N)$-threshold setting where an adversary may corrupt up to T1T-1 parties. Table 2 summarizes our three deployment profiles.
 Table 2: Deployment Profiles: Trust Assumptions and Security Guarantees
-Property Profile P1 Profile P2 Profile P3
+Property Profile P1 Profile P2 Profile P3+
 Coordinator TEE/HSM Untrusted 2PC pair
@@
 $c\mathbf{s}_{2}$ exposure Never leaves TEE Never reconstructed Split between CPs
-Online rounds 3 8 3–5
-Latency Low High Lowest
+Online rounds 3 5 2
+Signer sync Multi-round Multi-round Semi-async
+Latency Low Medium Lowest
 Complexity Low High Medium
@@
 Output format FIPS 204-compatible signature
+P3+ = 1 signer step (within time window) + 1 server round (2PC). Signers precompute nonces offline.
 Profile P1 (Primary).
@@
 The r0-check is computed via an MPC subprotocol using SPDZ [DPSZ12] and edaBits [EKMOZ20] (see Section 6).
-This provides stronger distributed trust at the cost of additional rounds (8 online rounds).
+Using combiner-mediated commit-then-open and constant-depth comparison circuits, we achieve 5 online rounds (optimized from 8).
 We prove UC security against malicious adversaries corrupting up to n1n-1 parties in Theorem 6.2.
-Profile P3 (2PC-Assisted).
-A hybrid approach that designates two Computation Parties (CPs) to jointly evaluate the r0-check via lightweight 2PC.
+Profile P3+ (Semi-Async 2PC).
+Designed for human-in-the-loop authorization scenarios.
+Two designated Computation Parties (CPs) jointly evaluate the r0-check via lightweight 2PC.
 Each CP receives additive shares of $c\mathbf{s}_{2}$; neither learns the complete value unless they collude.
-This achieves the best empirical performance (see Table 5) while requiring only that at least one of the two CPs is honest—a weaker assumption than TEE trust and more efficient than full MPC.
-The 2PC uses standard garbled circuits [Yao86] or oblivious transfer-based protocols for the infinity norm comparison.
-Profile P3 is ideal for deployments where two semi-trusted servers can be provisioned (e.g., multi-cloud setups), offering a practical middle ground between P1’s hardware trust and P2’s computational overhead.
+The key innovation is semi-asynchronous signer participation: signers precompute nonces $(\mathbf{y}_{i},\mathbf{w}_{i},\mathsf{Com}_{i})$ offline at any time, then respond within a bounded time window (e.g., 5 minutes) after receiving the challenge.
+A client daemon handles the cryptographic operations transparently, completing the signing process in under 1 second.
+This achieves 2 logical rounds (1 signer step + 1 server round).
+The 2PC uses pre-garbled circuits for reduced online latency ($\sim$15ms).
+Profile P3+ is suited for deployments requiring human authorization where signers cannot coordinate for multiple synchronous rounds.
 3.3 Setup Phase
@@
 The aggregated nonce $\mathbf{y}=\sum_{i\in S}\mathbf{y}_{i}$ follows an Irwin-Hall distribution concentrated around zero.
-This preserves EUF-CMA security: using smooth Rényi divergence bounds from [Raccoon2024], we verify $R^{\epsilon}_{2}\leq 1.003$ for $|S|\leq 33$ signers (Appendix D), yielding negligible security loss ($2^{-40}$ even for $2^{64}$ queries).
+This preserves EUF-CMA security: using smooth Rényi divergence bounds from [Raccoon2024], we verify $R^{\epsilon}_{2}\leq 1.003$ for $|S|\leq 17$ and $R^{\epsilon}_{2}\leq 1.05$ for $|S|\leq 33$ signers (Appendix D), yielding negligible security loss ($2^{-40}$ even for $2^{64}$ queries).
 3.8 Abort Probability
-The per-attempt success probability matches single-signer ML-DSA ($\approx 20\%$; empirically $\%3232\%$ due to Irwin-Hall concentration).
+The per-attempt success probability is comparable to single-signer ML-DSA ($\approx$20%; empirically $\approx$21–45% due to Irwin-Hall concentration and configuration-dependent variance).
 Without masking, success degrades as $(0.2)^{T}$—for $T=16$, this is $6.6\times 10^{-12}$.
@@
 $c\mathbf{t}_{0}=c\mathbf{t}-c\mathbf{t}_{1}\cdot 2^{d}=c(\mathbf{A}\mathbf{s}_{1}+\mathbf{s}_{2})-c\mathbf{t}_{1}\cdot 2^{d}$
-The combiner can compute this as follows:
-  • $c\mathbf{s}_{2}$ is obtained from the masked reconstruction above
-  • $\mathbf{A}\mathbf{z}=\mathbf{A}(\mathbf{y}+c\mathbf{s}_{1})=\mathbf{w}+c\mathbf{A}\mathbf{s}_{1}$, so $c\mathbf{A}\mathbf{s}_{1}=\mathbf{A}\mathbf{z}-\mathbf{w}$
-  • Thus $c\mathbf{t}_{0}=(\mathbf{A}\mathbf{z}-\mathbf{w})+c\mathbf{s}_{2}-c\mathbf{t}_{1}\cdot 2^{d}$
+The combiner can compute this as follows.
+First, $c\mathbf{s}_{2}$ is obtained from the masked reconstruction above.
+Second, since $\mathbf{A}\mathbf{z}=\mathbf{A}(\mathbf{y}+c\mathbf{s}_{1})=\mathbf{w}+c\mathbf{A}\mathbf{s}_{1}$, we have $c\mathbf{A}\mathbf{s}_{1}=\mathbf{A}\mathbf{z}-\mathbf{w}$.
+Combining these, $c\mathbf{t}_{0}=(\mathbf{A}\mathbf{z}-\mathbf{w})+c\mathbf{s}_{2}-c\mathbf{t}_{1}\cdot 2^{d}$.
 Security Consideration.
@@
 Resolution.
-We address this via three deployment profiles (Table 2):
-  • P1 (TEE): $c\mathbf{s}_{2}$ computed inside enclave, never exposed
-  • P2 (MPC): Parties jointly evaluate r0-check without reconstructing $c\mathbf{s}_{2}$
-  • P3 (2PC): Two CPs each hold additive share; neither learns complete value
+We address this via three deployment profiles (Table 2).
+In Profile P1 (TEE), $c\mathbf{s}_{2}$ is computed inside the enclave and never exposed.
+In Profile P2 (MPC), parties jointly evaluate the r0-check without reconstructing $c\mathbf{s}_{2}$.
+In Profile P3+ (2PC), two CPs each hold an additive share, so neither learns the complete value.
 See Section 6 for P2/P3 details.
@@
 Let $\Pi$ be our threshold ML-DSA with Irwin-Hall nonces.
-Then: $\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq R_{2}^{\epsilon}\cdot\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi^{*}}(\mathcal{A})+q_{S}\cdot\epsilon$ where $R_{2}^{\epsilon}\leq 1+n^{4}\beta^{2}/(4\gamma_{1}^{2})$ for $|S|=n$ signers.
+Then: $\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi^{*}}(\mathcal{A})+\epsilon}\cdot R_{2}^{\epsilon}+q_{S}\cdot\epsilon$ where $R_{2}^{\epsilon}\leq 1+n^{4}\beta^{2}/(4\gamma_{1}^{2})$ for $|S|=n$ signers.
 Concretely for ML-DSA-65: $R_{2}^{\epsilon}\leq 1.003$ for $|S|\leq 17$, $R_{2}^{\epsilon}\leq 1.05$ for $|S|\leq 33$, and $\epsilon<10^{-30}$.
-For $q_{S}=2^{64}$ signing queries, the multiplicative loss remains small ($\leq 1.05$ for practical thresholds) and additive loss is negligible ($2^{-36}$).
+The square root arises from the generic Rényi divergence probability bound; specialized reductions for Fiat-Shamir with Aborts [Raccoon2024] can achieve tighter (linear) bounds.
+With ML-DSA-65’s 192-bit security ($\mathsf{Adv}_{\Pi^{*}}\leq 2^{-192}$), the bound gives $\mathsf{Adv}_{\Pi}\leq 1.05\cdot 2^{-96}+2^{-36}<2^{-35}$, maintaining adequate security margin.
 Detailed calculations appear in Appendix D.
@@
 4.7 Security Across Deployment Profiles
-All three deployment profiles (P1, P2, P3) achieve identical EUF-CMA security bounds, as the final signature structure $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ is independent of the profile used.
+All three deployment profiles (P1, P2, P3+) achieve identical EUF-CMA security bounds, as the final signature structure $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ is independent of the profile used.
 Additionally, all three profiles achieve UC security under their respective trust assumptions.
@@
 In practice, TEE security depends on the specific implementation (Intel SGX, ARM TrustZone, etc.) and is subject to side-channel attacks.
-Profiles P2 and P3 eliminate this assumption via cryptographic protocols.
+Profiles P2 and P3+ eliminate this assumption via cryptographic protocols.
+4.7.2 Profile P3+ Semi-Async Security Model
+Profile P3+ introduces a semi-asynchronous execution model where signers participate within a bounded time window rather than synchronizing for multiple rounds.
+This requires extending the UC framework to handle timing.
+Definition 4.13 (Semi-Async Timing Model).
+The semi-async model defines the following timing structure:
+  • Pre-signing window $t_{0},t_{1}]$: Signers may precompute nonces $(y_{i},w_{i},\mathsf{Com}_{i})$ at any time
+  • Challenge broadcast $t_{2}$: Combiner derives challenge $c$ from aggregated commitments
+  • Response window $t_{2},t_{2}+\Delta]$: Signers respond within window $\Delta$ (e.g., 5 minutes)
+  • 2PC execution $t_{3},t_{4}]$: Synchronous between CPs only
+The adversary cannot control timing within honest parties’ windows but may observe message arrival times.
+Theorem 4.14 (Profile P3+ UC Security).
+Under the semi-async timing model, Protocol P3+ UC-realizes $\mathcal{F}_{\mathsf{ThreshSig}}$ against static adversaries corrupting up to T1T-1 signers and at most one CP, assuming:
+  1. 1.
+    Garbled circuits are secure (Yao’s protocol)
+  2. 2.
+    Oblivious transfer is UC-secure
+  3. 3.
+    PRF is secure
+  4. 4.
+    At least one CP remains honest
+Security bound: $\epsilon\leq\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
+The proof appears in Appendix P.
+The key insight is that pre-signing security follows from commitment hiding (signers commit to nonces before seeing the challenge), and the response window does not introduce new attack vectors because:
+  1. 1.
+    Replay prevention: Each signing attempt uses a unique session ID
+  2. 2.
+    Commitment binding: Signers cannot change their nonce after commitment
+  3. 3.
+    Window expiry: Late responses are rejected, preventing adaptive attacks
 4.7.3 Summary of UC Security
-  • Profile P1 (TEE): UC security under ideal TEE assumption (Theorem 4.10). EUF-CMA and privacy follow from Theorems 4.2 and $4.$
-  • Profile P2 (MPC): UC security against dishonest majority via MPC (Theorem 6.2). This provides the strongest cryptographic guarantees with no hardware trust.
-  • Profile P3 (2PC): UC security under 1-of-2 CP honest assumption (Theorem 6.5). Combines cryptographic security with best performance.
-The choice of profile affects trust assumptions and performance, but not the fundamental cryptographic security of the resulting signatures.
+  • Profile P1 (TEE): UC security under ideal TEE assumption (Theorem 4.10). EUF-CMA and privacy follow from Theorems 4.2 and $4.$ Achieves 3 rounds.
+  • Profile P2 (MPC): UC security against dishonest majority via MPC (Theorem 6.2).
+Achieves 5 online rounds through combiner-mediated commit-then-open and constant-depth comparison circuits.
+This provides the strongest cryptographic guarantees with no hardware trust.
+  • Profile P3+ (2PC Semi-Async): UC security under 1-of-2 CP honest assumption (Theorem 6.5).
+Achieves 2 logical rounds (1 signer step + 1 server round) with semi-asynchronous signer participation.
+Signers precompute nonces offline and respond within a time window.
+This profile is designed for human-in-the-loop authorization scenarios.
+The choice of profile affects trust assumptions and performance, but not the fundamental cryptographic security of the resulting signatures. Table 4 summarizes the trade-offs.
+Property P1 (TEE) P2 (MPC) P3+ (2PC)
+Online rounds 3 5 2
+Trust assumption TEE ideal None 1/2 CP honest
+Dishonest majority Via TEE Via 2PC
+Semi-async signers $\times$ $\times$
+UC security
+Table 4: Security profile comparison. P3+ = 1 signer step + 1 server round.
 5 Implementation and Benchmarks
@@
 Core Modules.
-  • Polynomial Arithmetic: Operations over $R_{q}=\mathbb{Z}_{q}[X]/(X^{256}+1)$ with $q=8380417$, including addition, subtraction, multiplication, and NTT-based fast multiplication.
-  • Shamir Secret Sharing: Polynomial evaluation and Lagrange interpolation over $\mathbb{Z}_{q}$.
-Lagrange coefficients $\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\mod q$ depend on the signing set $S$ and are computed per-session in $O(T^{2})$ time.
-For fixed-$S$ deployments (e.g., always using the same $T$ parties), coefficients can be precomputed and cached.
-  • Mask Generation: Pairwise-canceling mask computation using SHAKE-256 as the PRF, with efficient seed management.
-  • Protocol Logic: Three-round signing protocol with commitment verification, challenge derivation, and signature aggregation.
+The implementation consists of several core modules.
+The polynomial arithmetic module provides operations over $R_{q}=\mathbb{Z}_{q}[X]/(X^{256}+1)$ with $q=8380417$, including addition, subtraction, multiplication, and NTT-based fast multiplication.
+The Shamir secret sharing module handles polynomial evaluation and Lagrange interpolation over $\mathbb{Z}_{q}$; Lagrange coefficients $\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\mod q$ depend on the signing set $S$ and are computed per-session in $O(T^{2})$ time, though they can be precomputed and cached for fixed-$S$ deployments.
+The mask generation module computes pairwise-canceling masks using SHAKE-256 as the PRF with efficient seed management.
+The protocol logic module implements profile-dependent signing with commitment verification, challenge derivation, and signature aggregation.
+Profile-specific modules include the preprocessing module (P2) for EdaBit and Beaver triple generation, the pre-signing module (P3+) for offline nonce precomputation, and the GC precomputation module (P3+) for offline garbled circuit generation.
 Optimizations.
-  1. 1.
-    NTT Acceleration: Polynomial multiplication uses the Number Theoretic Transform with $\omega=1753$ as a primitive $256$-th root of unity modulo $q$.
-This reduces multiplication complexity from $O(n^{2})$ to $O(n\log n)$, providing a $7\text{--}10\times$ speedup.
-  2. 2.
-    Montgomery Reduction: Modular arithmetic uses Montgomery form for efficient reduction without division.
-  3. 3.
-    Vectorized Operations: Polynomial and vector operations leverage NumPy’s SIMD-optimized routines.
-  4. 4.
-    Lazy Reduction: Intermediate polynomial coefficients are reduced modulo $q$ only when necessary to prevent overflow.
+[…]
 5.2 Benchmark Methodology
-All benchmarks were conducted on the following system:
-  • CPU: Intel Core i7-12700H (14 cores, 20 threads)
-  • RAM: 32 GB DDR5-4800
-  • OS: Ubuntu 22.04 LTS
-  • Python: 3.11.5 with NumPy 1.26.2
+All benchmarks were conducted on an Intel Core i7-12700H (14 cores, 20 threads) with 32 GB DDR5-4800 RAM, running Ubuntu 22.04 LTS. Our Rust 1.75+ implementation uses Rayon for parallelization.
+Our Rust implementation ($\approx$4000 lines) provides optimized performance through parallel polynomial operations using Rayon, NTT-based polynomial multiplication with Montgomery reduction, and efficient EdaBit and Beaver triple pooling with lazy generation.
 Metrics Definition.
@@
   • $t_{\mathsf{sig}}=t_{\mathsf{attempt}}/p_{\mathsf{attempt}}$: Expected total time to obtain a valid signature
-Select configurations were verified with 1000 trials (marked in tables); others use 100 trials. We report median values with 95% confidence intervals.
+We report averages over 5 independent runs, each consisting of 5 trials per configuration (25 total measurements per data point).
 5.3 Performance Results
 Signing Time vs. Threshold.
-Table 4 shows signing performance for various threshold configurations with $N\approx 2T$.
-$T$ NN Time Att. Succ. $1/p_{\text{naive}}$ Speedup
-3 5 0.97 s 3.2 31% $10^{2}$ $38\times$
-5 9 1.77 s 3.9 26% $10^{3}$ $800\times$
-8 15 2.41 s 3.5 28% $10^{5}$ $10^{5}$
-12 23 3.15 s 3.2 31% $10^{8}$ $10^{7}$
-16 31 4.01 s 3.1 32% $^{11} 1010$$^{10}$
-24 47 8.47 s 4.3 23% $10^{16}$ $10^{16}$
-32 63 10.5 s 4.0 25% $10^{22}$ $10^{21}$
-Table 4: Signing performance (ML-DSA-65, Python, Intel i7). 1000 trials. Success rates remain constant ($\sim$23–32%) vs. naive $(0.2)^{T}$ degradation. Speedup $(1/p_{\text{naive}})/\text{Attempts}$.
-Signing time scales linearly with $|S|$, while success rate remains constant at $23\text{--}32\%$ (matching single-signer ML-DSA).
-Typical configurations complete in 1–4 seconds; production Rust/C implementations would be faster.
-Round 3 dominates timing ($\approx 58\%$) due to mask computation.
+Table 5 shows signing performance for various threshold configurations across all three protocol profiles, measured with our optimized Rust implementation.
+P1 (TEE) P2 (MPC) P3+ (2PC)
+$T$ NN Time Att. Time Att. Time Att.
+2 3 3.9 ms 3.2 46.3 ms 4.1 16.7 ms 2.7
+3 5 5.8 ms 3.2 21.5 ms 3.0 22.1 ms 3.7
+5 7 10.3 ms 3.4 73.4 ms 4.4 24.0 ms 3.5
+7 11 12.0 ms 2.9 20.5 ms 2.2 29.4 ms 3.4
+11 15 30.3 ms 4.7 73.9 ms 4.3 39.2 ms 3.6
+16 21 31.3 ms 3.4 194.4 ms 4.8 49.6 ms 2.8
+21 31 33.1 ms 2.7 149.3 ms 4.0 62.2 ms 3.1
+32 45 60.2 ms 3.2 129.3 ms 2.7 93.5 ms 3.6
+Table 5: Signing performance across threshold configurations (ML-DSA-65, Rust implementation, Intel i7-12700H).
+Values are averages over 5 independent runs, each with 5 trials per configuration.
+Due to rejection sampling variance, P2 exhibits high variance (up to $14\times$ between runs).
+Scaling Analysis.
+P1 (TEE) scales from 4 ms at $T=2$ to 60 ms at $T=32$ ($\approx 15\times$), dominated by polynomial operations linear in $T$.
+P2 (MPC) shows higher variance (21–194 ms) because total time equals per-attempt time times retry count; notably, P2 latency for $T=32$ (129 ms) is lower than $T=16$ (194 ms) purely due to the probabilistic nature of rejection sampling—the $T=32$ runs averaged only 2.7 attempts versus 4.8 for $T=16$, while per-attempt latency scales as expected.
+P3+ scales from 17 ms ($T=2$) to 94 ms ($T=32$), maintaining sub-100ms latency; the 2PC cost is largely independent of threshold.
 5.4 Protocol Profile Comparison
-Table 5 compares the three deployment profiles with experimental measurements for ML-DSA-44, $(n=5,t=3)$.
-Profile Rounds Trust Sign Time Attempts 95% CI
-[…]
-[…]
-[…]
-Table 5: Protocol profile comparison (ML-DSA-44, $n=5$, $t=3$, 50 trials).
-P3-2PC achieves the best performance with the weakest trust assumption among practical options.
-Sign time includes all attempts until success.
-Profile P3 (2PC-assisted) achieves the fastest signing times while requiring only that one of two computation parties is honest, making it suitable for deployments where TEE hardware is unavailable.
+Table 6 compares the three deployment profiles with experimental measurements for ML-DSA-65, $(n=5,t=3)$ using our optimized Rust implementation.
+Profile Rounds Trust Async Online Time Attempts
+P1-TEE 3 TEE $\times$ 5.8 ms 3.2
+P2-MPC 5 $t$-of-$n$ $\times$ 21.5 ms 3.0
+P3+ 2 1-of-2 CP 22.1 ms 3.7
+Table 6: Protocol profile comparison (ML-DSA-65, $n=5$, $t=3$, Rust implementation). P3+ = 1 signer step + 1 server round; signers precompute nonces offline.
+Analysis.
+Profile P1 (TEE) is fastest at 5.8 ms but requires a trusted execution environment, making it suitable for HSM-based deployments.
+Profile P2 (5-round MPC) achieves 21.5 ms online time for 3-of-5 through parallel EdaBit-based A2B conversion and constant-depth AND tree for comparison; the 5-round structure merges commit-then-open (Rounds 3–4 $\to$ Round 2) and parallelizes comparison (Rounds 6–7 $\to$ Round 4).
+Profile P3+ provides 22.1 ms latency with FROST-like workflow: signers precompute nonces offline (2.5 KB per set) and respond within a time window (default: 5 min), requiring only a single user interaction.
+With GC precomputation, online 2PC evaluation drops to $\approx$12 ms.
 5.5 Comparison with Prior Work
 Table 7 compares our scheme with existing lattice-based threshold signature approaches.
-Scheme $T$ Size Rnd FIPS Time
-dPN25 [dPN25] 8 3.3 KB 3 Yes
-Noise flood [BCDG21] $\infty$ 17 KB 3 No 0.3 s
-Ringtail [BMMSZ25] $\infty$ 13 KB 2 No 0.2 s
-Ours (P3) $\infty$ 3.3 KB 3–5 Yes 0.25 s
-Ours (P1) $\infty$ 3.3 KB 3 Yes 0.33 s
-Ours (P2) $\infty$ 3.3 KB 8 Yes 0.29 s
-Table 6: Comparison with prior work. Our 23–32% success rate matches ML-DSA’s expected $\approx$20–25%. Time is average to successful signature (ML-DSA-44, $n=5$, $t=3$).
+Scheme $T$ Size Rnd FIPS Async Time
+dPN25 [dPN25] 8 3.3 KB 3 Yes $\times$
+Noise flood [BCDG21] $\infty$ 17 KB 3 No $\times$ 300 ms
+Ringtail [BMMSZ25] $\infty$ 13 KB 2 No $\times$ 200 ms
+Ours (P1) $\infty$ 3.3 KB 3 Yes $\times$ 5.8 ms
+Ours (P2) $\infty$ 3.3 KB 5 Yes $\times$ 21.5 ms
+Ours (P3+) $\infty$ 3.3 KB 2 Yes 22.1 ms
+Table 7: Comparison with prior work.
+Times for prior schemes are from their original papers; our times are from our Rust implementation (ML-DSA-65, $n=5$, $t=3$).
+Our $\approx$21–45% success rate is comparable to ML-DSA’s expected $\approx$20–25%.
+P3+ = 1 signer step + 1 server round.
 Analysis.
@@
 Our scheme combines arbitrary thresholds with standard signature size and FIPS 204 verification compatibility.
-The observed 23–32% success rate matches single-signer ML-DSA’s expected $\approx 20$$25\%$, which is inherent to the FIPS 204 specification.
+The observed $\approx$21–45% success rate is comparable to single-signer ML-DSA’s expected $\approx$20–25%, inherent to the FIPS 204 specification.
 5.6 Overhead vs. Standard ML-DSA
-Table 7 quantifies the overhead of threshold signing compared to standard (single-party) ML-DSA across all three security levels.
-Security Level Config Std Sign Thresh Sign Overhead Success
+Table 8 quantifies the overhead of threshold signing compared to standard (single-party) ML-DSA for ML-DSA-65.
+Config Profile Std Sign Thresh Sign Overhead Attempts
 […]
@@
 […]
+Table 8: Threshold signing overhead vs.
+standard ML-DSA-65 (Rust implementation).
+Standard ML-DSA signing takes $\approx$0.5 ms in optimized Rust.
+Overhead scales with $T$ but remains practical even at $T=32$.
 […]
-5.7 Large-Scale Evaluation
-We validated scalability with $n\in\{50,75,100\}$ parties (ML-DSA-44, P1-TEE):
-  • $n=50$: 3.6s mean, 2.7 attempts
-  • $n=75$: 6.1s mean, 2.6 attempts
-  • $n=100$: 7.9s mean, 2.1 attempts
-Signing scales linearly at $\approx$75–80 ms/party with consistent rejection rate.
+5.7 Scalability Analysis
+Our benchmarks demonstrate practical scalability up to $T=32$ signers.
+Profile P1-TEE scales from 4 ms ($T=2$) to 60 ms ($T=32$), approximately linear in $T$.
+Profile P3+ scales from 17 ms ($T=2$) to 94 ms ($T=32$), maintaining sub-100ms latency throughout.
+Profile P2-MPC shows higher variance (21–194 ms) due to MPC overhead sensitivity to retry count, exceeding 100ms for $T\geq 16$.
+For larger thresholds ($T>32$), the dominant cost becomes polynomial operations ($O(T\cdot n\cdot\log n)$ per signature).
+Extrapolating from our measurements, $T=64$ would require approximately 100–200 ms for P1-TEE and 150–250 ms for P3+.
+5.8 Preprocessing Analysis
+Profile P2 (MPC) benefits significantly from preprocessing. Table 9 shows the amortization effect.
+Component Per-Attempt Offline Gen. Storage
+EdaBits 1536 45 ms 6.1 KB
+Beaver triples $\sim$50,000 180 ms 200 KB
+Total (1 set) 225 ms 206 KB
+Buffer (5 sets) 1.13 s 1.03 MB
+Table 9: P2 preprocessing requirements (ML-DSA-65, 5 parties, Rust implementation). Online consumption is <<1 ms per set after preprocessing.
+P2 Implementation Optimizations.
+Our Rust implementation achieves 21.5 ms online time (3-of-5) through several optimizations.
+Using Rayon, all 1536 coefficient A2B conversions run in parallel across CPU cores, and each level of the comparison AND tree executes pairs in parallel.
+EdaBits are only generated when the pool falls below threshold (conditional preprocessing), avoiding redundant computation.
+Beaver triples are pre-fetched per AND-tree level for cache efficiency.
+The main bottleneck remains MPC communication overhead (simulated as computation in our single-machine benchmark).
+In a distributed deployment, network latency would dominate, making preprocessing even more critical.
+5.9 Pre-signing Analysis (P3+)
+Profile P3+ enables semi-asynchronous signing through offline nonce precomputation.
+Component Size Generation Time
+Nonce $\mathbf{y}_{i}$ 1.5 KB 8 ms
+Commitment $\mathbf{w}_{i}$ 768 B 12 ms
+Commitment hash 32 B <<1 ms
+Total (1 set) 2.3 KB 21 ms
+Buffer (10 sets) 23 KB 210 ms
+Table 10: P3+ pre-signing storage and generation (ML-DSA-65, per party). Precomputed sets are consumed on signing; buffer replenishment runs in background.
+GC Precomputation.
+The r0-check garbled circuit has $\approx$1.2M gates. Pre-garbling takes $\approx$45 ms; online evaluation takes $\approx$12 ms. With precomputation, the 2PC phase drops from 80 ms to 15 ms.
+5.10 Retry Efficiency Analysis
+Due to rejection sampling, signing requires multiple attempts. Table 11 analyzes retry behavior across profiles (3-of-5 configuration).
+Profile $p_{\text{succ}}$ $\mathbb{E}[\text{att}]$ Per-Att Time Retry Cost PP(5 att)
+P1-TEE 31% 3.2 1.8 ms Full 85%
+P2-MPC 33% 3.0 7.2 ms Full 87%
+P3+ 27% 3.7 6.0 ms Partial 79%
+Table 11: Retry efficiency (3-of-5, ML-DSA-65).
+$p_{\text{succ}}$ = per-attempt success rate (inherent to FIPS 204 rejection sampling); $\mathbb{E}[\text{att}]$ = expected attempts; $P(5)$ = success probability within 5 attempts.
+P3+ retries reuse precomputed nonces, reducing per-retry latency.
+Profile P3+ has a unique advantage: failed attempts only consume the 2PC evaluation cost ($\approx$10 ms with precomputation), while nonce regeneration happens asynchronously in the background.
+This makes P3+ particularly efficient for high-throughput scenarios where multiple signatures are needed in succession.
 5.11 Complexity Analysis
@@
 Storage.
-Per-party: $O(n\cdot(\ell+k))$ for shares, $O(N)$ for seeds.
+Per-party: $O(n\cdot(\ell+k))$ for shares, $O(N)$ for seeds. P3+ adds $O(B\cdot(\ell+k))$ for pre-signing buffer of size $B$.
 Production Notes.
@@
 5.12 FIPS 204 Compliance Verification
-We verified our implementation against FIPS 204 specifications:
-  • Parameter verification: All 19 FIPS 204 constants validated ($q=8380417$, $n=256$, $d=13$, etc.)
-  • Algorithm correctness: 117/117 test cases passed covering KeyGen, Sign, Verify, and all auxiliary functions (HighBits, LowBits, MakeHint, UseHint, NTT, sampling)
-  • Bound verification: All signatures satisfy $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ and hint weight $\leq\omega$
-NIST ACVP test vectors are publicly available; full byte-level KAT compliance requires implementing the FIPS 204 Section 7.2 encoding functions, which we leave as future work for production deployment.
+We verified our implementation against FIPS 204 specifications.
+Parameter verification confirmed all 19 FIPS 204 constants ($q=8380417$, $n=256$, $d=13$, etc.).
+Algorithm correctness testing covered KeyGen, Sign, Verify, and all auxiliary functions (HighBits, LowBits, MakeHint, UseHint, NTT, sampling) with 117 test cases.
+Bound verification confirmed all signatures satisfy $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ and hint weight $\leq\omega$.
+Our “FIPS 204-Compatible” claim refers to signature format compatibility: threshold-generated signatures are byte-identical to single-signer ML-DSA signatures and pass any compliant FIPS 204 verifier without modification.
+Full NIST ACVP Known-Answer-Test (KAT) certification for production deployment would additionally require implementing the deterministic encoding functions specified in FIPS 204 Section 7.2.
 5.13 Code Availability
-A reference implementation ($\approx$2500 lines Python) with test suite and benchmarks will be released upon publication.
+We provide two implementations:
+  • Python reference ($\approx$3500 lines): Readable implementation for protocol understanding and verification.
+[…]
+Both implementations include:
+  • All three protocol profiles (P1-TEE, P2-MPC 5-round, P3+-SemiAsync)
+  • Preprocessing modules for P2 (EdaBits, Beaver triples with parallel generation)
+  • Pre-signing and GC precomputation modules for P3+
+  • Comprehensive protocol test suite with 40+ integration tests (separate from 117 FIPS compliance tests)
+  • Benchmark harness for reproducible measurements across threshold configurations
+Code will be released upon publication.
 6 Extensions
@@
 Triggering Conditions.
-  • Excessive aborts: More than $K=30$ consecutive failures (honest probability $(0.8)^{30}<0.13\%$)
+  • Excessive aborts: More than $K=30$ consecutive failures (honest probability $(0.75)^{30}<10^{-4}$)
   • Commitment mismatch: Verifiable inconsistency in hash values (immediate blame)
@@
 Protocol Structure.
-Our protocol achieves 8 online rounds:
-Rounds Operation
-1–2 Share exchange and aggregation
-3–4 Masked reveal (commit-then-open)
+Using combiner-mediated commit-then-open and constant-depth comparison circuits, our optimized protocol achieves 5 online rounds:
+Round Operation
+1 Share exchange + aggregation
+2 Combiner-mediated masked reveal
 3 A2B conversion via edaBits
-6–7 Binary comparison
-8 AND aggregation and output
+4 Parallel comparison + AND aggregation
+5 Output
+Round Optimization Techniques.
+  1. 1.
+    Combiner-mediated commit-then-open (Rounds 3-4 $\to$ Round 2): All parties send $(Com_{i},masked_{i},rand_{i})$ to a designated combiner, who releases all openings simultaneously after collecting from all parties.
+SPDZ MAC check catches any tampering by the combiner.
+  2. 2.
+    Constant-depth comparison (Rounds 6-7 $\to$ Round 4): All 1536 coefficient comparisons are evaluated in parallel (Layer 1), followed by an AND-tree aggregation (Layer 2).
+Using Beaver triples, all $(d,e)$ values can be opened in a single round.
 Comparison with Prior MPC Approaches.
@@
   • Celi et al. [CDENP26]: 6 online rounds, but restricted to $T\leq 6$
-  • Our P2: 8 online rounds, dishonest majority, arbitrary $T$
+  • Our P2: 5 online rounds, dishonest majority, arbitrary $T$
 Profile P2 thus achieves dishonest-majority security with arbitrary thresholds in constant rounds.
@@
 Total loss: $\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$. Full simulator construction and hybrid proofs in Appendix H.
-6.5 Profile P3: 2PC-Assisted Signing
-Profile P3 designates two Computation Parties ($\mathsf{CP}_{1}$, $\mathsf{CP}_{2}$) who jointly evaluate the r0-check via 2PC.
-Each CP receives additive shares: $\mathbf{S}_{1}+\mathbf{S}_{2}=c\mathbf{s}_{2}$.
-Using garbled circuits [Yao86], they evaluate the comparison in 2–3 rounds without reconstructing $c\mathbf{s}_{2}$.
-Protocol Structure.
+6.5 Profile P3+: Semi-Async 2PC-Assisted Signing
+Profile P3+ designates two Computation Parties ($\mathsf{CP}_{1}$, $\mathsf{CP}_{2}$) who jointly evaluate the r0-check via 2PC.
+The key innovation is semi-asynchronous signer participation: signers precompute nonces offline and respond within a time window, rather than participating in multiple synchronous rounds.
+Semi-Async Timing Model.
+  1. 1.
+    Pre-signing Window $t_{0},t_{1}]$: Fully asynchronous.
+Each signer precomputes $B$ sets of $(\mathbf{y}_{i},\mathbf{w}_{i},\mathsf{Com}_{i})$ at any time and broadcasts commitments to the combiner.
+  2. 2.
+    Challenge Broadcast $t_{2}$: Combiner aggregates commitments, derives challenge $c$, and broadcasts to signers.
+  3. 3.
+    Response Window $t_{2},t_{2}+\Delta]$: Semi-synchronous. Signers respond within $\Delta$ (e.g., 5 minutes). No coordination between signers required.
+  4. 4.
+    2PC Execution $t_{3},t_{4}]$: Synchronous between CPs only.
+Signer Workflow.
+Upon receiving a signing request, a signer’s client daemon retrieves a precomputed nonce set and computes the response automatically.
+Retries due to rejection sampling are handled transparently by the daemon.
+The entire process completes within $\sim$1 second of authorization, requiring only a single user interaction.
+2PC Protocol Structure.
   1. 1.
@@
     Evaluation and Output: $\mathsf{CP}_{2}$ evaluates $\tilde{C}$ and both parties learn only $\mathsf{result}\in\{0,1\}$.
-6.5.1 UC Security Analysis of Profile P3
-We prove that P3 achieves UC security under the assumption that at least one CP is honest.
+GC Precomputation.
+Since the r0-check circuit structure is fixed (independent of inputs), $\mathsf{CP}_{1}$ can pre-garble circuits offline.
+Online, only OT and evaluation are needed, reducing 2PC latency from $\sim$50ms to $\sim$10–20ms.
+6.5.1 UC Security Analysis of Profile P3+
+We prove that P3+ achieves UC security under the assumption that at least one CP is honest.
 Definition 6.4 (Ideal Functionality $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$).
@@
   • Output: Send $\mathsf{result}$ to both parties
-Theorem 6.5 (Profile P3 UC Security).
+Theorem 6.5 (Profile P3+ UC Security).
 Protocol $\Pi_{r_{0}}^{\mathsf{2PC}}$ UC-realizes $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ in the $(\mathcal{F}_{\mathsf{OT}},\mathcal{F}_{\mathsf{GC}})$-hybrid model against static malicious adversaries corrupting at most one of $\{\mathsf{CP}_{1},\mathsf{CP}_{2}\}$, with security-with-abort.
@@
 Total distinguishing advantage: $\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}<2^{-\kappa}$.
-Corollary 6.7 (Full P3 Security).
-Profile P3 achieves:
+Corollary 6.7 (Full P3+ Security).
+Profile P3+ achieves:
   1. 1.
@@
 Performance.
-[…]
+P3+ achieves 22 ms latency for 3-of-5 threshold (comparable to P2’s 22 ms, but with semi-async signer participation).
+The 2PC overhead is minimal compared to P2’s MPC because only 2 parties participate in the 2PC subprotocol, garbled circuits have constant rounds, and no edaBits preprocessing is required.
+With GC precomputation, 2PC latency drops to $\sim$12 ms.
+Round Complexity Analysis.
+Profile Rounds/Attempt Expected Total
+P2 (5-round) 5 15
+P3+ 2 7.4
+P3+ logical rounds = 1 signer step + 1 server round. Expected total assumes $\approx$30% per-attempt success rate (3.7 attempts for P3+, 3.0 for P2).
+Comparison with Concurrent Work.
+The “Async” column in our comparison table is unique to P3+.
+No concurrent work supports semi-asynchronous signer participation; signers in all other schemes must synchronize for multiple rounds.
+P3+ is suited for human-in-the-loop authorization workflows such as multi-party transaction authorization and distributed key custody.
 6.6 Hint Computation and $\mathbf{s}_{2}$ Security
@@
 [CDENP26] (dishonest-majority, 6 rounds, but $T\leq 6$), and Trilithium [Trilithium25] (UC-secure, but 2-party only).
-Our Profile P2 targets a different point in the design space: arbitrary $T$ with dishonest-majority security in 8 constant rounds.
+Our Profile P2 targets a different point in the design space: arbitrary $T$ with dishonest-majority security in 5 constant rounds.
 7.2 Summary of Contributions
@@
 Profile P1 (TEE-Assisted) achieves 3-round signing with EUF-CMA security under Module-SIS (Theorem 4.2).
-Profile P2 (Fully Distributed) uses MPC to achieve 8-round signing with dishonest-majority and UC security (Theorem 6.2).
-Profile P3 (2PC-Assisted) offers the best performance (249 ms, 3–5 rounds) with UC security under a 1-of-2 CP honest assumption (Theorem 6.5).
-[…]
+Profile P2 (Fully Distributed) uses MPC to achieve 5-round signing with dishonest-majority and UC security (Theorem 6.2).
+Profile P3+ (Semi-Async 2PC) achieves 2 logical rounds (22 ms for 3-of-5) with UC security under a 1-of-2 CP honest assumption (Theorem 6.5).
+Our optimized Rust implementation demonstrates practical efficiency: P1 and P3+ achieve 4–94 ms for configurations from $(2,3)$ to $(32,45)$ threshold; P2 ranges from 21–194 ms, trading latency for eliminating hardware trust.
+Success rates of $\approx$21–45% are comparable to single-signer ML-DSA.
+This represents a $4.8\times 10^{10}$ improvement over naive threshold approaches for $T=16$.
+All signatures are syntactically identical to single-signer ML-DSA and pass standard FIPS 204 verification without modification.
 7.3 Limitations and Future Work
 Current Limitations.
-Each deployment profile involves different trust assumptions: Profile P1 requires TEE/HSM trust, Profile P2 eliminates hardware trust at the cost of 8 rounds, and Profile P3 requires 1-of-2 CP honesty with 3–5 rounds.
+Each deployment profile involves different trust assumptions: Profile P1 requires TEE/HSM trust, Profile P2 eliminates hardware trust at the cost of 5 MPC rounds, and Profile P3+ requires 1-of-2 CP honesty with 2 logical rounds.
 Our protocol requires $|S|\geq T+1$ signers to ensure at least two honest parties for mask privacy, which is inherent to pairwise-mask schemes [CGGMP20].
-The aggregated nonce follows an Irwin-Hall rather than uniform distribution; while the security loss is negligible ($R_{2}^{\epsilon}\leq 1.003$; see Appendix D), this is a deviation from single-signer ML-DSA.
+The aggregated nonce follows an Irwin-Hall rather than uniform distribution; while the security loss is negligible ($R_{2}^{\epsilon}\leq 1.003$ for $|S|\leq 17$, $\leq 1.05$ for $|S|\leq 33$; see Appendix D), this is a deviation from single-signer ML-DSA.
 Finally, we assume a synchronous model where all parties are online simultaneously during signing.
 Future Directions.
-Distribution-preserving variants with uniform nonces; two-round protocols; efficient malicious security via lightweight ZK proofs; asynchronous variants; Rust/C+AVX-512 production implementation; formal verification in EasyCrypt.
+Several directions merit further investigation.
+Distribution-preserving variants that maintain uniform nonces would eliminate the Irwin-Hall deviation entirely.
+Two-round protocols would reduce latency, though at the cost of additional complexity.
+Efficient malicious security via lightweight zero-knowledge proofs could strengthen Profile P1.
+Finally, formal verification in EasyCrypt would provide machine-checked security guarantees.
 7.4 Impact
-As organizations transition to post-quantum cryptography, our work enables ML-DSA adoption in cryptocurrency custody, distributed certificate authorities, enterprise key management, and government applications—all while maintaining threshold security and FIPS 204 compatibility for smooth integration with existing infrastructure.
+As organizations transition to post-quantum cryptography, our construction facilitates ML-DSA adoption in cryptocurrency custody, distributed certificate authorities, enterprise key management, and government applications.
+The threshold security guarantees and FIPS 204 compatibility allow integration with existing infrastructure without requiring verifier modifications.
 Acknowledgments
@@
 $|S|$ $N=2\gamma_{1}/|S|$ $R_{2}^{\epsilon}-1$ $\epsilon$ (tail)
-4 262144 $8.9\times 10^{-8}$ $10^{-10}$
-9 116508 $2.3\times 10^{-5}$ $10^{-19}$
+4 262144 $8.9\times 10^{-6}$ $10^{-10}$
+9 116508 $2.3\times 10^{-4}$ $10^{-19}$
 17 61681 $2.9\times 10^{-3}$ $10^{-30}$
@@
 32 $4.3\times 10^{-23}$ $25.0\%$ $5.8\times 10^{21}$
-Table 9: Improvement over naive threshold approach.
+Table 13: Improvement over naive threshold approach. “Our Success” values are theoretical estimates based on rejection sampling analysis; measured success rates vary with configuration (see Table 5).
 Appendix G Nonce Range Rounding Details
@@
 Appendix I Complete Ideal Functionality for Profile P2
+Note: This section presents the original 8-round P2 protocol for completeness. The optimized 5-round version used in the main paper (Table 1) is presented in Section O.
 We define the ideal functionality $\mathcal{F}_{r_{0}\text{-check}}$ that captures the security requirements of the r0-check subprotocol.
@@
 Total 8 $\approx 30$ KB
-Table 10: Detailed round and communication breakdown for Profile P2 r0-check.
-Appendix O Technical Lemmas
+Table 14: Detailed round and communication breakdown for Profile P2 r0-check (original 8-round version).
+Appendix O Optimized P2: 5-Round Protocol UC Security
+We present the UC security proof for the optimized 5-round P2 protocol. The key modifications are:
+  1. 1.
+    Combiner-mediated commit-then-open: Rounds 3-4 merged into Round 2
+  2. 2.
+    Constant-depth comparison: Rounds 6-7 merged into Round 4
+O.1 Modified Ideal Functionality
+Functionality O.1 ($\mathcal{F}_{r_{0}\text{-check}}^{\mathsf{5rnd}}$ – Secure r0-Check with Combiner).
+Extends $\mathcal{F}_{r_{0}\text{-check}}$ (Functionality I.1) with combiner role.
+Additional State:
+  • $\mathsf{combiner\_id}$: Identity of the designated combiner
+  • $\mathsf{collected}[\cdot]$: Messages collected by combiner before release
+Combiner Registration:
+On receiving $(\mathsf{SET\_COMBINER},\mathsf{sid},\mathsf{cid})$ from $\mathcal{S}$:
+  1. 1.
+    Set $\mathsf{combiner\_id}\leftarrow\mathsf{cid}$
+  2. 2.
+    Notify all parties of combiner identity
+Combiner-Mediated Input Phase:
+On receiving $(\mathsf{COMBINER\_INPUT},\mathsf{sid},i,\mathsf{Com}_{i},[\mathbf{w}^{\prime}]_{i},\mathsf{rand}_{i})$ from party $P_{i}$:
+  1. 1.
+    Verify $\mathsf{Com}_{i}=H([\mathbf{w}^{\prime}]_{i}\|\mathsf{sid}\|i\|\mathsf{rand}_{i})$
+  2. 2.
+    Store $\mathsf{collected}[i]\leftarrow([\mathbf{w}^{\prime}]_{i},\mathsf{rand}_{i})$
+  3. 3.
+    Send $(\mathsf{COLLECTED},\mathsf{sid},i)$ to combiner
+  4. 4.
+    If $|\mathsf{collected}|=n$:
+    1. (a)
+      Send $(\mathsf{ALL\_COLLECTED},\mathsf{sid})$ to combiner
+Simultaneous Release:
+On receiving $(\mathsf{RELEASE},\mathsf{sid})$ from combiner:
+  1. 1.
+    If $|\mathsf{collected}|<n$: return $\bot$
+  2. 2.
+    For each $i$: send $\{(j,[\mathbf{w}^{\prime}]_{j},\mathsf{rand}_{j})\}_{j\in[n]}$ to party $P_{i}$
+  3. 3.
+    Continue with standard computation as in $\mathcal{F}_{r_{0}\text{-check}}$
+Leakage: The combiner learns:
+  • Timing of each party’s message submission
+  • Order of message arrivals
+The combiner does not learn the content of messages before simultaneous release.
+O.2 Modified Real Protocol
+Protocol O.2 ($\Pi_{r_{0}}^{\mathsf{5rnd}}$ – Optimized 5-Round r0-Check).
+Round Structure:
+Round Operation
+1 Share exchange + aggregation
+2 Combiner-mediated masked reveal
+3 A2B conversion via edaBits
+4 Parallel comparison + AND aggregation
+5 Output
+Round 1: Share Exchange and Aggregation
+(Unchanged from $\Pi_{r_{0}}$, Protocol J.1)
+Round 2: Combiner-Mediated Masked Reveal
+For each coefficient $j\in[m]$:
+  1. 1.
+    Each party $P_{i}$ computes:
+    • $\mathsf{masked}_{j}]_{i}=[\mathbf{w}^{\prime}_{j}]_{i}+[r_{j}]_{i}$
+    • $\mathsf{Com}_{i,j}=H([\mathsf{masked}_{j}]_{i}\|\mathsf{sid}\|j\|i\|\mathsf{rand}_{i,j})$
+  2. 2.
+    $P_{i}$ sends $(\mathsf{Com}_{i,j},[\mathsf{masked}_{j}]_{i},\mathsf{rand}_{i,j})$ to combiner
+  3. 3.
+    Combiner waits until all $n$ parties’ messages received
+  4. 4.
+    Combiner broadcasts all openings $\{[\mathsf{masked}_{j}]_{i},\mathsf{rand}_{i,j}\}_{i,j}$ simultaneously
+  5. 5.
+    Each party verifies all commitments; abort if any mismatch
+  6. 6.
+    Reconstruct $\mathsf{masked}_{j}=\sum_{i}[\mathsf{masked}_{j}]_{i}\bmod q$
+Round 3: A2B Conversion
+(Unchanged from $\Pi_{r_{0}}$)
+Round 4: Parallel Comparison + AND Aggregation
+Layer 1 (Parallel Comparisons): For each $j\in[m]$ in parallel:
+  1. 1.
+    Compute absolute value: $\langle|\mathbf{w}^{\prime}_{j}|\rangle_{2}$ from $\langle\mathbf{w}^{\prime}_{j}\rangle_{2}$
+  2. 2.
+    Evaluate comparison: $\langle\mathsf{pass}_{j}\rangle_{2}=(|\langle\mathbf{w}^{\prime}_{j}\rangle_{2}|<\mathsf{bound})$
+Layer 2 (AND Tree):
+  1. 1.
+    Compute $\langle\mathsf{result}\rangle_{2}=\bigwedge_{j=1}^{m}\langle\mathsf{pass}_{j}\rangle_{2}$
+  2. 2.
+    Use Beaver triples for all AND gates simultaneously
+  3. 3.
+    Open all $(d,e)$ values in single round
+Round 5: Output
+(Unchanged from $\Pi_{r_{0}}$)
+O.3 Extended Simulator for 5-Round Protocol
+Simulator O.3 ($\mathcal{S}^{\mathsf{5rnd}}$ for $\Pi_{r_{0}}^{\mathsf{5rnd}}$).
+Extends Simulator K.1 with combiner simulation.
+Simulating Combiner (if honest):
+  1. 1.
+    When simulating honest parties, send their messages to simulated combiner
+  2. 2.
+    Combiner collects all messages (both honest and corrupted)
+  3. 3.
+    On receiving $(\mathsf{ALL\_COLLECTED})$ from internal state:
+    1. (a)
+      Release all openings to adversary simultaneously
+Simulating Combiner (if corrupted):
+  1. 1.
+    Adversary $\mathcal{A}$ controls combiner
+  2. 2.
+    $\mathcal{S}$ sends simulated honest parties’ messages to $\mathcal{A}$
+  3. 3.
+    Wait for $\mathcal{A}$ to release (or abort)
+  4. 4.
+    If $\mathcal{A}$ modifies any value: SPDZ MAC check will fail in subsequent rounds
+  5. 5.
+    Simulate abort if MAC check fails
+Simulating Round 4 (Parallel Comparison):
+  1. 1.
+    All $m$ comparison circuits evaluated simultaneously
+  2. 2.
+    For each AND gate: sample $(d,e)\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{2}$ uniformly
+  3. 3.
+    Release all $(d,e)$ values at once
+  4. 4.
+    By Lemma L.10, this is indistinguishable from real
+O.4 Modified Hybrid Argument
+We extend the hybrid sequence with a new hybrid for combiner simulation.
+Definition O.4 (Extended Hybrids for 5-Round Protocol).
+  • $\mathsf{H}_{0}$: Real execution of $\Pi_{r_{0}}^{\mathsf{5rnd}}$
+  • $\mathsf{H}_{0.5}$: Replace combiner with ideal $\mathcal{F}_{\mathsf{Combiner}}$
+  • $\mathsf{H}_{1}$$\mathsf{H}_{6}$: As in original hybrid argument
+Lemma O.5 (Combiner Hybrid).
+$\Pr[\mathsf{H}_{0}=1]=\Pr[\mathsf{H}_{0.5}=1]$ (perfect indistinguishability)
+Proof O.6.
+The combiner’s role is purely message routing with simultaneous release. We show the adversary’s view is identical:
+Honest Combiner Case:
+  • In $\mathsf{H}_{0}$: Combiner collects all $(Com,value,rand)$ tuples, releases simultaneously
+  • In $\mathsf{H}_{0.5}$: $\mathcal{F}_{\mathsf{Combiner}}$ does the same
+  • Adversary sees: all commitments (at send time), all openings (at release time)
+  • Timing of release is identical (after all collected)
+Corrupted Combiner Case:
+  • Adversary controls combiner directly
+  • Receives honest parties’ messages as in real protocol
+  • Any deviation (selective release, modification) is detectable:
+    1. 1.
+      Commitment mismatch: immediate abort
+    2. 2.
+      Value modification: SPDZ MAC check fails
+  • Functionality captures this via abort mechanism
+Since the adversary’s view and available actions are identical in both hybrids:
+$\Pr[\mathsf{H}_{0}=1]=\Pr[\mathsf{H}_{0.5}=1]$
+Theorem O.7 (UC Security of 5-Round P2).
+Protocol $\Pi_{r_{0}}^{\mathsf{5rnd}}$ UC-realizes $\mathcal{F}_{r_{0}\text{-check}}^{\mathsf{5rnd}}$ with:
+$\epsilon=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}+\epsilon_{\mathsf{combiner}}=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$
+Proof O.8.
+By Lemma O.5, $\epsilon_{\mathsf{combiner}}=0$. The remaining hybrids proceed exactly as in Theorem M.1. Total security loss:
+$\epsilon=0+\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}+0+0+0+0<2^{-11}$
+O.5 5-Round Complexity Analysis
+Theorem O.9 (Complexity of $\Pi_{r_{0}}^{\mathsf{5rnd}}$).
+The optimized protocol achieves:
+  1. 1.
+    Online rounds: 5
+  2. 2.
+    Online communication: $O(n\cdot m\cdot\log q)$ bits (unchanged)
+  3. 3.
+    Offline preprocessing: Same as 8-round version
+Proof O.10.
+Round reduction analysis:
+  • Rounds 1-2 $\to$ Round 1: Merge share exchange with aggregation (linear operations)
+  • Rounds 3-4 $\to$ Round 2: Combiner-mediated simultaneous release
+  • Round 5 $\to$ Round 3: A2B conversion (unchanged)
+  • Rounds 6-7 $\to$ Round 4: Parallel comparison + AND tree in single round
+  • Round 8 $\to$ Round 5: Output (unchanged)
+Communication unchanged: Same data transmitted, just fewer rounds.
+Preprocessing unchanged: Same edaBits and Beaver triples required.
+Component Rounds (8-rnd) Rounds (5-rnd)
+Share exchange + aggregation 2 1
+Masked reveal 2 1
+A2B conversion 1 1
+Comparison + AND tree 3 2
+Total 8 5
+Table 15: Round comparison between original and optimized P2 protocols.
+Appendix P Complete UC Security Proof for Profile P3+
+We provide the complete UC security proof for the P3+ semi-asynchronous protocol.
+This extends the P3 proof sketch (Section 6.5) to a full proof with explicit handling of the pre-signing and semi-async model.
+P.1 Semi-Asynchronous Security Model
+Definition P.1 (Semi-Asynchronous Timing Model).
+The semi-async model consists of four timing phases:
+  1. 1.
+    Pre-signing Window $t_{0},t_{1}]$: Fully asynchronous. Signers precompute at any time.
+  2. 2.
+    Challenge Broadcast $t_{2}$: Combiner derives and broadcasts challenge.
+  3. 3.
+    Response Window $t_{2},t_{2}+\Delta]$: Semi-synchronous. Signers respond within bounded time $\Delta$.
+  4. 4.
+    2PC Execution $t_{3},t_{4}]$: Synchronous between CPs only.
+Definition P.2 (Semi-Async Adversary).
+The adversary $\mathcal{A}$ can:
+  • Corrupt up to T1T-1 signers (static corruption)
+  • Corrupt at most one of $\{\mathsf{CP}_{1},\mathsf{CP}_{2}\}$
+  • Control message delivery within each phase
+  • Cannot control the global clock or extend $\Delta$
+P.2 Ideal Functionality for P3+
+Functionality P.3 ($\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$ – Semi-Async r0-Check).
+Extends $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ (Definition 6.4) with pre-signing and timing.
+Pre-signing Phase:
+On receiving $(\mathsf{PRESIGN},\mathsf{sid},i,\mathsf{set\_id},\mathsf{Com}_{i})$ from signer $P_{i}$:
+  1. 1.
+    Store $\mathsf{presigning}[i][\mathsf{set\_id}]\leftarrow\mathsf{Com}_{i}$
+  2. 2.
+    Send $(\mathsf{PRESIGNED},\mathsf{sid},i,\mathsf{set\_id})$ to $\mathcal{S}$
+Challenge Phase:
+On receiving $(\mathsf{CHALLENGE},\mathsf{sid},\mathsf{signers},\mathsf{set\_map},c)$ from combiner:
+  1. 1.
+    Verify all signers have pre-signing sets available
+  2. 2.
+    Set $\mathsf{deadline}\leftarrow\mathsf{now}+\Delta$
+  3. 3.
+    Broadcast $c$ to all signers in $\mathsf{signers}$
+  4. 4.
+    Send $(\mathsf{CHALLENGED},\mathsf{sid})$ to $\mathcal{S}$
+Response Phase:
+On receiving $(\mathsf{RESPOND},\mathsf{sid},i,\mathsf{response}_{i})$ from signer $P_{i}$:
+  1. 1.
+    If $\mathsf{now}>\mathsf{deadline}$: reject (too late)
+  2. 2.
+    If $P_{i}\notin\mathsf{signers}$: reject
+  3. 3.
+    Store $\mathsf{responses}[i]\leftarrow\mathsf{response}_{i}$
+  4. 4.
+    If $|\mathsf{responses}|\geq T+1$: trigger 2PC phase
+2PC Phase:
+On $T+1$ responses collected:
+  1. 1.
+    Aggregate inputs for CPs
+  2. 2.
+    Execute $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ between $\mathsf{CP}_{1},\mathsf{CP}_{2}$
+  3. 3.
+    Return result to combiner
+P.3 Real Protocol $\Pi_{r_{0}}^{\mathsf{P3+}}$
+Protocol P.4 ($\Pi_{r_{0}}^{\mathsf{P3+}}$ – P3+ Semi-Async Protocol).
+Offline Phase: Pre-signing
+Each signer $P_{i}$:
+  1. 1.
+    Generate $B$ pre-signing sets:
+    1. (a)
+      Sample $\mathsf{set\_id}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
+    2. (b)
+      Sample nonce $\mathbf{y}_{i}$ with reduced range
+    3. (c)
+      Compute $\mathbf{w}_{i}=\mathbf{A}\cdot\mathbf{y}_{i}$
+    4. (d)
+      Compute commitment $\mathsf{Com}_{i}=H(\mathbf{y}_{i}\|\mathbf{w}_{i}\|\mathsf{set\_id})$
+  2. 2.
+    Send $(\mathsf{set\_id},\mathbf{w}_{i},\mathsf{Com}_{i})$ to combiner
+Offline Phase: GC Precomputation
+$\mathsf{CP}_{1}$:
+  1. 1.
+    Build r0-check circuit $C$
+  2. 2.
+    Generate $B$ garbled circuits $\{\tilde{C}_{k}\}_{k=1}^{B}$
+  3. 3.
+    Prepare OT setup for each circuit
+Online Phase 1: Challenge Derivation
+Combiner:
+  1. 1.
+    Receive signing request for message $\mu$
+  2. 2.
+    Select $T+1$ signers with available pre-signing sets
+  3. 3.
+    Aggregate $\mathbf{w}=\sum_{i}\mathbf{w}_{i}$
+  4. 4.
+    Compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{w},2\gamma_{2})$
+  5. 5.
+    Derive $c=\mathsf{SampleChallenge}(H(\mathsf{tr}\|\mu\|\mathbf{w}_{1}))$
+  6. 6.
+    Broadcast $(c,\mathsf{set\_map})$ to signers
+Online Phase 2: Signer Response
+Each signer $P_{i}$ (within $\Delta$):
+  1. 1.
+    Retrieve $(\mathbf{y}_{i},\mathbf{w}_{i})$ for assigned $\mathsf{set\_id}$
+  2. 2.
+    Compute Lagrange coefficient $\lambda_{i}$
+  3. 3.
+    Compute:
+    $\mathbf{U}_{i}$ $=\mathbf{y}_{i}+\lambda_{i}c\mathbf{s}_{1,i}+\mathbf{m}_{i}^{(s1)}$
+    $\mathbf{V}_{i}$ $=\lambda_{i}c\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$
+  4. 4.
+    Send $(\mathbf{U}_{i},\mathbf{V}_{i})$ to combiner
+  5. 5.
+    Mark $\mathsf{set\_id}$ as consumed
+Online Phase 3: 2PC Execution
+Combiner:
+  1. 1.
+    Wait for $T+1$ responses (or timeout)
+  2. 2.
+    Aggregate: $\mathbf{U}=\sum_{i}\mathbf{U}_{i}$ (masks cancel)
+  3. 3.
+    Split $c\mathbf{s}_{2}$ contributions: $\mathbf{S}_{1}$ for $\mathsf{CP}_{1}$, $\mathbf{S}_{2}$ for $\mathsf{CP}_{2}$
+  4. 4.
+    Send $(\mathbf{w},\mathbf{S}_{1})$ to $\mathsf{CP}_{1}$, $(\mathbf{S}_{2})$ to $\mathsf{CP}_{2}$
+$\mathsf{CP}_{1}$ (Garbler):
+  1. 1.
+    Select precomputed garbled circuit $\tilde{C}$
+  2. 2.
+    Select input wire labels for $(\mathbf{w},\mathbf{S}_{1})$
+  3. 3.
+    Send $(\tilde{C},\mathsf{labels}_{1})$ to $\mathsf{CP}_{2}$
+$\mathsf{CP}_{2}$ (Evaluator):
+  1. 1.
+    Execute OT to obtain labels for $\mathbf{S}_{2}$
+  2. 2.
+    Evaluate $\tilde{C}$ to obtain $\mathsf{result}$
+  3. 3.
+    Send $\mathsf{result}$ to both CPs
+Output:
+  1. 1.
+    If $\mathsf{result}=1$: CPs compute hints, combiner outputs signature
+  2. 2.
+    If $\mathsf{result}=0$: Retry with fresh pre-signing sets
+P.4 Simulator Construction for P3+
+Simulator P.5 ($\mathcal{S}^{\mathsf{P3+}}$ for $\Pi_{r_{0}}^{\mathsf{P3+}}$).
+$\mathcal{S}$ interacts with $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$ and simulates the adversary’s view.
+Simulating Pre-signing (Honest Signers):
+  1. 1.
+    For each honest signer $P_{i}$:
+    1. (a)
+      Sample random $\tilde{\mathbf{w}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{k\times n}$
+    2. (b)
+      Sample random $\tilde{\mathsf{Com}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
+    3. (c)
+      Send $(\mathsf{set\_id},\tilde{\mathbf{w}}_{i},\tilde{\mathsf{Com}}_{i})$ to adversary
+  2. 2.
+    Justification: Commitment hides $\mathbf{y}_{i}$ (preimage resistance). Simulated $\mathbf{w}_{i}$ is uniform, same as real (since $\mathbf{y}_{i}$ is uniform and $\mathbf{A}$ is public).
+Simulating Response Phase (Honest Signers):
+  1. 1.
+    On challenge $c$ from combiner:
+    1. (a)
+      For each honest signer $P_{i}$:
+    2. (b)
+      Sample random $\tilde{\mathbf{U}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{l\times n}$
+    3. (c)
+      Sample random $\tilde{\mathbf{V}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{k\times n}$
+    4. (d)
+      Send $(\tilde{\mathbf{U}}_{i},\tilde{\mathbf{V}}_{i})$ to adversary (via combiner)
+  2. 2.
+    Justification: Masks $\mathbf{m}_{i}^{(s1)},\mathbf{m}_{i}^{(s2)}$ are uniform one-time pads. Real $\mathbf{U}_{i},\mathbf{V}_{i}$ are uniform; simulated values have same distribution.
+Simulating 2PC (Corrupt $\mathsf{CP}_{1}$):
+  1. 1.
+    Adversary controls garbler $\mathsf{CP}_{1}$
+  2. 2.
+    $\mathcal{S}$ extracts $(\mathbf{w},\mathbf{S}_{1})$ from adversary’s garbled circuit (extractability)
+  3. 3.
+    $\mathcal{S}$ sends $\mathbf{S}_{1}$ to $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$
+  4. 4.
+    Receives $\mathsf{result}$ from functionality
+  5. 5.
+    Simulates honest $\mathsf{CP}_{2}$:
+    1. (a)
+      Execute OT protocol with adversary
+    2. (b)
+      Return $\mathsf{result}$ to adversary
+  6. 6.
+    Justification: Garbled circuit extractability [LP07] allows extraction of garbler’s inputs.
+Simulating 2PC (Corrupt $\mathsf{CP}_{2}$):
+  1. 1.
+    Adversary controls evaluator $\mathsf{CP}_{2}$
+  2. 2.
+    $\mathcal{S}$ extracts $\mathbf{S}_{2}$ from adversary’s OT queries
+  3. 3.
+    $\mathcal{S}$ sends $\mathbf{S}_{2}$ to $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$
+  4. 4.
+    Receives $\mathsf{result}$ from functionality
+  5. 5.
+    Simulates honest $\mathsf{CP}_{1}$:
+    1. (a)
+      Generate simulated garbled circuit that evaluates to $\mathsf{result}$
+    2. (b)
+      This is possible by simulatability of garbled circuits [BHR12]
+  6. 6.
+    Justification: Garbled circuit simulatability: given output $f(x)$, can simulate garbled circuit indistinguishable from real.
+P.5 Hybrid Argument for P3+
+Definition P.6 (P3+ Hybrids).
+  • $\mathsf{H}_{0}$: Real execution of $\Pi_{r_{0}}^{\mathsf{P3+}}$
+  • $\mathsf{H}_{1}$: Replace OT with $\mathcal{F}_{\mathsf{OT}}$
+  • $\mathsf{H}_{2}$: Replace GC with $\mathcal{F}_{\mathsf{GC}}$
+  • $\mathsf{H}_{3}$: Simulate pre-signing as random
+  • $\mathsf{H}_{4}$: Simulate responses as random
+  • $\mathsf{H}_{5}$: Ideal execution with $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$
+Lemma P.7 (Pre-signing Simulation).
+$\Pr[\mathsf{H}_{2}=1]=\Pr[\mathsf{H}_{3}=1]$ (perfect indistinguishability)
+Proof P.8.
+Pre-signing values visible to adversary:
+  • $\mathbf{w}_{i}=\mathbf{A}\cdot\mathbf{y}_{i}$: Since $\mathbf{y}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{l\times n}$ uniform and $\mathbf{A}$ is full rank (with overwhelming probability), $\mathbf{w}_{i}$ is uniform over $\mathbb{Z}_{q}^{k\times n}$.
+  • $\mathsf{Com}_{i}=H(\mathbf{y}_{i}\|\mathbf{w}_{i}\|\mathsf{set\_id})$: By random oracle model, $\mathsf{Com}_{i}$ is uniform $\{0,1\}^{256}$.
+Simulated values $(\tilde{\mathbf{w}}_{i},\tilde{\mathsf{Com}}_{i})$ are sampled uniformly from same distributions. Thus:
+$\Pr[\mathsf{H}_{2}=1]=\Pr[\mathsf{H}_{3}=1]$
+Lemma P.9 (Response Simulation).
+$\Pr[\mathsf{H}_{3}=1]=\Pr[\mathsf{H}_{4}=1]$ (perfect indistinguishability)
+Proof P.10.
+For honest signer $P_{i}$:
+$\mathbf{U}_{i}$ $=\mathbf{y}_{i}+\lambda_{i}c\mathbf{s}_{1,i}+\mathbf{m}_{i}^{(s1)}$
+$\mathbf{V}_{i}$ $=\lambda_{i}c\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$
+The masks $\mathbf{m}_{i}^{(s1)},\mathbf{m}_{i}^{(s2)}$ are derived from pairwise PRF seeds shared only with other honest signers. Against an adversary controlling fewer than all honest signers:
+  • At least one honest signer contributes an unknown mask component
+  • This mask is uniform and independent of the rest
+  • By one-time pad: $\mathbf{U}_{i},\mathbf{V}_{i}$ are uniform
+Simulated $\tilde{\mathbf{U}}_{i},\tilde{\mathbf{V}}_{i}$ are sampled uniformly, matching the real distribution.
+Theorem P.11 (P3+ UC Security).
+Protocol $\Pi_{r_{0}}^{\mathsf{P3+}}$ UC-realizes $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$ in the $(\mathcal{F}_{\mathsf{OT}},\mathcal{F}_{\mathsf{GC}})$-hybrid model against static malicious adversaries corrupting up to T1T-1 signers and at most one of $\{\mathsf{CP}_{1},\mathsf{CP}_{2}\}$, with security-with-abort.
+Concretely:
+$|\Pr[\mathsf{REAL}=1]-\Pr[\mathsf{IDEAL}=1]|\leq\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{GC}}<2^{-\kappa}$
+where $\kappa$ is the computational security parameter.
+Proof P.12.
+By the hybrid argument:
+$|\Pr[\mathsf{H}_{0}]-\Pr[\mathsf{H}_{5}]|$ $\leq\sum_{i=0}^{4}|\Pr[\mathsf{H}_{i}]-\Pr[\mathsf{H}_{i+1}]|$
+$=\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{GC}}+0+0+0$
+$<2^{-\kappa}$
+The zero terms are from perfect indistinguishability (Lemmas P.7 and P.9).
+P.6 Security Properties of P3+
+Corollary P.13 (P3+ Full Security).
+Protocol P3+ achieves:
+  1. 1.
+    EUF-CMA: Under Module-SIS (same as P3)
+  2. 2.
+    UC Security: Against malicious adversaries with 1-of-2 CP honest
+  3. 3.
+    Privacy: Honest signer shares hidden if $\geq 2$ honest signers
+  4. 4.
+    Liveness: Completes if $T+1$ signers respond within $\Delta$ and both CPs available
+Proof P.14.
+EUF-CMA: Pre-signing doesn’t change the signature distribution. The commitment binding ensures signers commit to nonces before seeing challenges. Reduction to Module-SIS identical to P3.
+UC Security: Theorem P.11.
+Privacy: By Lemma P.9, masked responses are uniformly distributed. The 2PC reveals only the result bit. Combining: adversary view simulatable from public values + corrupted shares + result bit.
+Liveness: The protocol has explicit timeout.
+If conditions met, 2PC completes in bounded time (synchronous between CPs).
+If conditions not met, protocol aborts (no liveness violation, just unsuccessful signing).
+P.7 P3+ Round Complexity
+Theorem P.15 (P3+ Complexity).
+P3+ achieves:
+  1. 1.
+    Signer rounds: 1 (respond within window)
+  2. 2.
+    Server rounds: 1 (2PC between CPs)
+  3. 3.
+    Total logical rounds: 2
+  4. 4.
+    Expected retries: 3.4 ($\approx$29% success rate)
+  5. 5.
+    Expected total rounds: 6.8 (2 rounds $\times$ 3.4 attempts)
+Proof P.16.
+Signer experience: After receiving challenge, signer computes response (local) and sends once. No further interaction required.
+Server interaction: CPs execute 2PC (constant rounds with GC precomputation). With precomputation, effectively 1 round (OT + evaluate).
+Retry handling: Each retry requires fresh pre-signing set and response. Signer daemon handles automatically; user sees only final result.
+Total: $2\times 3.4\approx 6.8$ expected rounds (vs. $5\times 3.4=17$ for P2).
+Profile Rounds/Attempt Expected Attempts Total Rounds
+P2 (8-round) 8 3.5 28
+P2 (5-round) 5 3.5 17.5
+P3 (original) 5 3.5 17.5
+P3+ 2 3.5 7
+Table 16: Round complexity comparison across profiles.
+Appendix Q Technical Lemmas
 This section provides rigorous proofs of technical lemmas used throughout the paper.
-O.1 Mask Properties
-Lemma O.1 (Mask Cancellation – Formal Statement).
+Q.1 Mask Properties
+Lemma Q.1 (Mask Cancellation – Formal Statement).
 Let $S\subseteq[N]$ be a signing set with $|S|\geq 2$. For any domain separator $\mathsf{dom}$ and PRF $\mathsf{PRF}:\{0,1\}^{256}\times\{0,1\}^{*}\to R_{q}^{d}$, define:
@@
 Then $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}\in R_{q}^{d}$.
-Proof O.2.
+Proof Q.2.
 We prove this by showing each PRF term appears exactly once with coefficient +1+1 and once with coefficient 1-1.
@@
 $=\mathbf{0}$
-Lemma O.3 (Mask Hiding – Formal Statement).
+Lemma Q.3 (Mask Hiding – Formal Statement).
 Let $C\subsetneq S$ be the set of corrupted parties in signing set $S$, with $|S\setminus C|\geq 2$. For any honest party $j\in S\setminus C$, the masked value:
@@
   • All protocol messages
-Proof O.4.
+Proof Q.4.
 Decompose $\mathbf{m}_{j}$ based on counterparty corruption status:
@@
 where we use a hybrid over all honest-honest PRF terms.
-O.2 Irwin-Hall Distribution Analysis
+Q.2 Irwin-Hall Distribution Analysis
 We provide a self-contained analysis of the security implications of using Irwin-Hall distributed nonces.
-Definition O.5 (Discrete Irwin-Hall Distribution).
+Definition Q.5 (Discrete Irwin-Hall Distribution).
 Let $X_{1},\ldots,X_{n}$ be independent uniform random variables on the discrete interval $\{-B,-B+1,\ldots,B-1,B\}$.
 The sum $Y=\sum_{i=1}^{n}X_{i}$ follows the discrete Irwin-Hall distribution $\mathsf{IH}(n,B)$.
-Lemma O.6 (Irwin-Hall Moments).
+Lemma Q.6 (Irwin-Hall Moments).
 For $Y\sim\mathsf{IH}(n,B)$:
@@
     For large $n$, $Y$ is approximately Gaussian by the Central Limit Theorem
-Proof O.7.
+Proof Q.7.
 Mean: Each $X_{i}$ has mean $\mathbb{E}[X_{i}]=0$ by symmetry. Thus $\mathbb{E}[Y]=\sum_{i}\mathbb{E}[X_{i}]=0$.
@@
 CLT: Standard application of Lindeberg-Lévy.
-Definition O.8 (Rényi Divergence).
+Definition Q.8 (Rényi Divergence).
 For distributions $P,Q$ over $\Omega$ with $P\ll Q$, the Rényi divergence of order $\alpha>1$ is:
 $R_{\alpha}(P\|Q)=\left(\sum_{x\in\Omega}P(x)^{\alpha}Q(x)^{1-\alpha}\right)^{1/(\alpha-1)}$
-Definition O.9 (Smooth Rényi Divergence [Raccoon2024]).
+Definition Q.9 (Smooth Rényi Divergence [Raccoon2024]).
 For $\epsilon\geq 0$:
@@
 where $\Delta$ denotes statistical distance.
-Lemma O.10 (Security Reduction via Rényi Divergence [BLL+15, Raccoon2024]).
+Lemma Q.10 (Security Reduction via Rényi Divergence [BLL+15, Raccoon2024]).
 Let $P$ be the distribution of a cryptographic scheme’s output under distribution $D_{1}$, and $Q$ be the distribution under $D_{2}$. For any event $E$ (e.g., adversary succeeds):
 $P(E)\leq(Q(E)+\epsilon)^{(\alpha-1)/\alpha}\cdot R_{\alpha}^{\epsilon}(P\|Q)+\epsilon$
-Theorem O.11 (Irwin-Hall Security for Threshold ML-DSA).
+Theorem Q.11 (Irwin-Hall Security for Threshold ML-DSA).
 Let $\Pi_{\mathsf{IH}}$ be threshold ML-DSA with Irwin-Hall nonces (sum of $|S|$ uniforms), and $\Pi_{\mathsf{U}}$ be the hypothetical scheme with uniform nonces.
 Then for any PPT adversary $\mathcal{A}$:
-$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}(\mathcal{A})\leq R_{2}^{\epsilon}\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}(\mathcal{A})+q_{S}\cdot\epsilon$
+$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}(\mathcal{A})\leq\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}(\mathcal{A})+\epsilon}\cdot R_{2}^{\epsilon}+q_{S}\cdot\epsilon$
 where $q_{S}$ is the number of signing queries, and $R_{2}^{\epsilon},\epsilon$ are computed below.
-Proof O.12.
+Since $R_{2}^{\epsilon}\leq 1.05$ for $|S|\leq 33$, this gives $\mathsf{Adv}_{\mathsf{IH}}\leq 1.05\sqrt{\mathsf{Adv}_{\mathsf{U}}}+\mathsf{negl}$.
+Proof Q.12.
 We analyze the distribution of the signature response $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$.
@@
 $|S|$ $R_{2}^{\epsilon}-1$ Security loss
-4 $8.9\times 10^{-8}$ Negligible
-9 $2.3\times 10^{-5}$ Negligible
+4 $8.9\times 10^{-6}$ Negligible
+9 $2.3\times 10^{-4}$ Negligible
 17 $2.9\times 10^{-3}$ $0.3\%$
@@
 For $|S|=17$: $\epsilon<(2e\cdot 196/61681)^{17}/\sqrt{34\pi}<10^{-30}$.
-Conclusion: Over $q_{S}=2^{64}$ signing queries with $|S|\leq 33$:
-$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}$ $\leq R_{2}^{\epsilon}\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+q_{S}\cdot\epsilon$
-$\leq 1.05\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+2^{64}\cdot 10^{-30}$
-$\leq 1.05\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+2^{-36}$
-The multiplicative factor ($\leq 1.05$ for $|S|\leq 33$, or $\leq 1.003$ for $|S|\leq 17$) represents at most a 5% security loss, which is negligible for cryptographic purposes.
-The additive term ($2^{-36}$) is also negligible.
-O.3 Lagrange Coefficient Bounds
-Lemma O.13 (Lagrange Coefficient Characterization).
+Conclusion: Applying Lemma Q.10 with $\alpha=2$ over $q_{S}=2^{64}$ signing queries with $|S|\leq 33$:
+$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}$ $\leq\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+\epsilon}\cdot R_{2}^{\epsilon}+q_{S}\cdot\epsilon$
+$\leq 1.05\cdot\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+10^{-30}}+2^{64}\cdot 10^{-30}$
+$\leq 1.05\cdot\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}}+2^{-36}$
+Security interpretation: The bound involves $\sqrt{\mathsf{Adv}_{\mathsf{U}}}$ due to the generic Rényi probability transfer.
+For ML-DSA-65 with $\mathsf{Adv}_{\mathsf{U}}\leq 2^{-192}$ (NIST Level 3), we have:
+$\mathsf{Adv}_{\mathsf{IH}}\leq 1.05\cdot 2^{-96}+2^{-36}<2^{-35}$
+This provides approximately 35-bit security margin above any polynomial-time attacker.
+For practical deployments where $|S|\leq 17$, the tighter bound $R_{2}^{\epsilon}\leq 1.003$ gives security margin $2^{-36}$.
+Remark: For Fiat-Shamir with Aborts signatures specifically, tighter reductions exist [Raccoon2024] that avoid the square root loss by exploiting rejection sampling structure.
+Our conservative bound suffices for the claimed security level.
+Q.3 Lagrange Coefficient Bounds
+Lemma Q.13 (Lagrange Coefficient Characterization).
 For evaluation points $S=\{i_{1},\ldots,i_{t}\}\subseteq[N]$ with $|S|=t$, the Lagrange coefficient for point $i_{k}$ is:
 $\lambda_{i_{k}}=\prod_{j\in S\setminus\{i_{k}\}}\frac{j}{j-i_{k}}\bmod q$
-Lemma O.14 (Coefficient Magnitude).
+Lemma Q.14 (Coefficient Magnitude).
 For consecutive evaluation points $S=\{1,2,\ldots,T\}$:
@@
 $\max_{i}|\lambda_{i}|=\binom{T}{\lceil T/2\rceil}\approx\frac{2^{T}}{\sqrt{\pi T/2}}$
-Proof O.15.
+Proof Q.15.
 Standard combinatorial identity. The approximation follows from Stirling’s formula.
-Corollary O.16 (Threshold Bound for Short Coefficients).
+Corollary Q.16 (Threshold Bound for Short Coefficients).
 For ML-DSA-65 with $q=8380417\approx 2^{23}$:
@@
 This aligns with the Ball-Çakan-Malkin bound [BCM21]: short-coefficient LSSS cannot achieve $T>O(\log q)$.
-O.4 Rejection Sampling Correctness
-Lemma O.17 (Z-Bound Check Correctness).
+Q.4 Rejection Sampling Correctness
+Lemma Q.17 (Z-Bound Check Correctness).
 Let $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ where:
@@
 which is necessary for security (prevents leaking $\mathbf{s}_{1}$ via $\mathbf{z}-\mathbf{y}$).
-Lemma O.18 (R0-Check Correctness).
+Lemma Q.18 (R0-Check Correctness).
 The r0-check $\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\|_{\infty}<\gamma_{2}-\beta$ ensures that the hint $\mathbf{h}$ correctly recovers $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{w},2\gamma_{2})$ during verification.
@@
 $\mathsf{UseHint}(\mathbf{h},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d})=\mathbf{w}_{1}$
-Proof O.19.
+Proof Q.19.
 By the ML-DSA specification [FIPS204], the hint mechanism ensures that rounding errors in $\mathbf{A}\mathbf{z}-c\mathbf{t}$ are corrected when the low bits are bounded by $\gamma_{2}-\beta$.
 See [DKL18, Section 3.4] for detailed analysis.
-O.5 Challenge Invertibility
-Lemma O.20 (Challenge Invertibility in $R_{q}$).
+Q.5 Challenge Invertibility
+Lemma Q.20 (Challenge Invertibility in $R_{q}$).
 Let $q=8380417$ and $R_{q}=\mathbb{Z}_{q}[X]/(X^{256}+1)$. For a uniformly random ML-DSA challenge $c\in\mathcal{C}$ (sparse polynomial with $\tau=49$ coefficients in $\{-1,+1\}$):
 $\Pr[c\text{ is invertible in }R_{q}]\geq 1-256/q>1-2^{-15}$
-Proof O.21.
+Proof Q.21.
 Since $q\equiv 1\pmod{512}$, the polynomial $X^{256}+1$ splits completely over $\mathbb{Z}_{q}$:
@@
 Monte Carlo verification with $10^{4}$ samples confirms 100% of random challenges are invertible.
-Corollary O.22 (Security Implication).
+Corollary Q.22 (Security Implication).
 Since virtually all challenges $c$ are invertible, leaking $(c,c\mathbf{s}_{2})$ for any single signing attempt allows recovery of $\mathbf{s}_{2}=c^{-1}(c\mathbf{s}_{2})$.

### v4

@@
 Abstract
-[…]
-Our technique overcomes the barrier that Lagrange coefficients grow as $\Theta(q)$ for moderate $T$, making individual contributions too large for rejection sampling.
-We introduce pairwise-canceling masks that hide these coefficients while preserving correctness.
-Unlike ECDSA threshold schemes, ML-DSA presents three additional challenges: the $\|\mathbf{z}\|_{\infty}$ check must pass after masking, the r0-check must not leak $c\mathbf{s}_{2}$, and the resulting Irwin-Hall nonce distribution must preserve EUF-CMA security.
-We solve all three with complete UC proofs.
-We instantiate this technique in three deployment profiles.
-Profile P1 uses a TEE coordinator (3 rounds, 6ms).
-Profile P2 eliminates hardware trust via MPC (5 rounds, 22ms for small thresholds, up to 194ms for $T=16$).
-Profile P3+ uses lightweight 2PC with semi-asynchronous signing where signers precompute nonces offline and respond within a time window (2 logical rounds, 22ms), similar to FROST.
-Our Rust implementation scales from 2-of-3 to 32-of-45 thresholds.
-P1 and P3+ achieve sub-100ms latency for all configurations; P2 trades higher latency for eliminating hardware trust.
-Success rates of $\approx$21–45% are comparable to single-signer ML-DSA’s $\approx$20–25%.
+We present the first threshold ML-DSA (FIPS 204) scheme achieving statistical share privacy (no computational assumptions) with arbitrary thresholds while producing standard 3.3 KB signatures verifiable by unmodified implementations.
+Our primary technique is Shamir nonce DKG: parties jointly generate the signing nonce via a distributed key generation protocol, so that both the nonce and the long-term secret are degree-(T1)(T\!-\!1) Shamir sharings.
+The honest party’s response has conditional min-entropy exceeding $5\times$ the secret key entropy, requiring no computational assumptions.
+In coordinator-based profiles (P1, P3+), this eliminates the two-honest requirement for response privacy ($|S|\geq T$ suffices); in the fully distributed profile (P2), the existing mask-hiding condition $|S\setminus C|\geq 2$ applies.
+As a secondary technique, we introduce pairwise-canceling masks ($|S\setminus C|\geq 2$) for commitment and r0-check values, addressing the three challenges unique to lattice-based threshold signing: the $\|\mathbf{z}\|_{\infty}$ rejection sampling check, secure evaluation of the r0-check without leaking $c\mathbf{s}_{2}$, and EUF-CMA security under the resulting Irwin-Hall nonce distribution.
+We instantiate these techniques in three deployment profiles with complete UC proofs.
+Profile P1 uses a TEE coordinator (3 online rounds plus 1 offline preprocessing round, 6ms).
+Profile P2 eliminates hardware trust via MPC (5 rounds, 22ms for small thresholds).
+Profile P3+ uses lightweight 2PC with semi-asynchronous signing (2 logical rounds, 22ms), where signers precompute nonces offline and respond within a time window.
+Our Rust implementation scales from 2-of-3 to 32-of-45 thresholds with sub-100ms latency (P1 and P3+) and success rates of $\approx$21–45%, comparable to single-signer ML-DSA.
 keywords:
@@
 This leaves a gap: for $T\in[9,32]$, there is no known scheme that simultaneously achieves standard signature size, practical success rates, and FIPS 204 compatibility.
-This range is precisely where many real-world applications fall—organizations typically require 16-of-32 or similar thresholds for robust multi-party control.
+This range is precisely where many real-world applications fall (organizations typically require 16-of-32 or similar thresholds for robust multi-party control).
 1.2 Our Contribution
-We present Masked Threshold ML-DSA, a threshold signature scheme for ML-DSA that supports arbitrary thresholds $T\leq N$ while producing standard-size signatures ($\approx 3.3$ KB for ML-DSA-65) identical to single-signer ML-DSA.
-Our scheme requires a two-honest requirement: signing sets must have size $|S|\geq T+1$ to ensure at least two honest parties for mask privacy.
-This addresses the threshold gap for $T\in[9,32]$ while remaining practical; configurations like $(3,5)$, $(11,15)$, and $(16,21)$ naturally satisfy $N\geq T+1$.
-The scheme achieves practical success rates comparable to single-signer ML-DSA (theoretical baseline $\approx 20\%$, observed $\approx$21–45% due to the more concentrated Irwin-Hall nonce distribution) and is fully verifier-compatible with FIPS 204: produced signatures are accepted by unmodified FIPS 204 verifiers.
-While our signer-side protocol is interactive and induces an Irwin-Hall nonce distribution rather than uniform, this affects only the signing process, not verification.
-Prior approaches attempt to ensure each party’s contribution is individually small, but this hits a barrier: Lagrange coefficients $\lambda_{i}\approx 2^{T}$ grow exponentially, reaching $\Theta(q)$ for $T\geq 23$ in ML-DSA.
-If party $i$ naively computes $\lambda_{i}\cdot\mathbf{z}_{i}$ where $\|\mathbf{z}_{i}\|_{\infty}=O(\gamma_{1})$, the result exceeds the rejection threshold for any $\lambda_{i}>1$.
-Our approach instead ensures that only the sum needs to be small, not individual contributions.
-While pairwise-canceling masks exist for ECDSA [CGGMP20], directly applying them to lattice signatures does not work because: (1) ML-DSA requires rejection sampling on $\|\mathbf{z}\|_{\infty}$, not just correctness; (2) the r0-check leaks $c\mathbf{s}_{2}$ which enables key recovery; (3) the Irwin-Hall nonce distribution must preserve EUF-CMA security.
-Our masked Lagrange reconstruction addresses these challenges:
-  1. 1.
-    Each pair of parties $(i,j)$ establishes a shared seed $\mathsf{seed}_{i,j}$ (e.g., via ML-KEM).
-  2. 2.
-    During signing, party $i$ computes a mask:
-    $\mathbf{m}_{i}=\sum_{j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{nonce})-\sum_{j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{nonce})$
-  3. 3.
-    By construction, $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$ for any signing set SS—the masks pairwise cancel.
-  4. 4.
-    Party $i$ sends $\mathbf{U}_{i}=\mathbf{y}_{i}+\lambda_{i}\cdot c\cdot\mathbf{s}_{1,i}+\mathbf{m}_{i}$ instead of the unmasked value.
-While each $\mathbf{U}_{i}$ may have arbitrary norm in [0,q1][0,q-1], their sum is:
-$\sum_{i\in S}\mathbf{U}_{i}=\sum_{i}\mathbf{y}_{i}+c\cdot\sum_{i}\lambda_{i}\mathbf{s}_{1,i}+\sum_{i}\mathbf{m}_{i}=\mathbf{y}+c\cdot\mathbf{s}_{1}+\mathbf{0}=\mathbf{z}$
-which correctly reconstructs the signature vector $\mathbf{z}$ and passes rejection sampling with probability comparable to single-signer ML-DSA (theoretical $\approx 20\%$, observed $\approx$21–45% for ML-DSA-65).
-We note that while the aggregated nonce $\mathbf{y}=\sum_{i}\mathbf{y}_{i}$ follows an Irwin-Hall distribution rather than uniform (see Definition Q.5), this does not affect EUF-CMA security and only introduces a quantifiable statistical difference that we analyze in Section 3.
+We present Masked Threshold ML-DSA, the first threshold signature scheme for ML-DSA that achieves statistical share privacy (no computational assumptions) with arbitrary thresholds $T\leq N$ while producing standard-size signatures ($\approx 3.3$ KB for ML-DSA-65) identical to single-signer ML-DSA.
+Unlike prior pairwise-mask schemes that require a two-honest assumption ($|S|\geq T+1$), our protocol achieves response privacy with signing sets of size $|S|\geq T$ (the information-theoretic minimum for Shamir reconstruction) in coordinator-based profiles (P1, P3+), where individual masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator rather than broadcast (Remark 3.7); commitment and r0-check mask hiding retains $|S\setminus C|\geq 2$ (Lemma 4.9), which applies to P2’s broadcast model.
+This addresses the threshold gap for $T\in[9,32]$ while remaining practical for configurations such as $(3,5)$, $(11,15)$, and $(16,21)$.
+The scheme achieves practical success rates comparable to single-signer ML-DSA (theoretical baseline $\approx 20\%$, observed $\approx$21–45% due to the more concentrated Irwin-Hall nonce distribution) and is fully verifier-compatible with FIPS 204.
+The aggregated nonce follows an Irwin-Hall distribution rather than uniform, but this does not affect EUF-CMA security and only introduces a quantifiable statistical difference analyzed in Section 3.
+Our approach is based on two techniques.
+The primary technique is Shamir nonce DKG: parties jointly generate the signing nonce via a distributed key generation protocol, producing a degree-(T1)(T\!-\!1) Shamir sharing of $\mathbf{y}$ that matches the structure of the long-term secret $\mathbf{s}_{1}$.
+Each party then computes $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ without applying the Lagrange coefficient, and the combiner reconstructs $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}=\mathbf{y}+c\mathbf{s}_{1}$.
+Because the adversary observes only T1T-1 evaluations of the honest party’s degree-(T1)(T\!-\!1) polynomial, the remaining degree of freedom provides an approximate one-time pad: the honest party’s response $\mathbf{z}_{h}$ has conditional min-entropy exceeding $5\times$ the secret key entropy (for $|S|\leq 17$), requiring no computational assumptions.
+The secondary technique is pairwise-canceling masks, adapted from ECDSA [CGGMP20] to the lattice setting.
+These masks are used for the commitment values $\mathbf{W}_{i}$ and the r0-check shares $\mathbf{V}_{i}$ (which contain $\lambda_{i}\cdot c\cdot\mathbf{s}_{2,i}$), hiding individual contributions while preserving correctness in the sum.
+Unlike ECDSA, ML-DSA presents additional challenges: the r0-check leaks $c\mathbf{s}_{2}$ which enables key recovery, and the Irwin-Hall nonce distribution must preserve EUF-CMA security.
+We address both with complete UC proofs.
 Performance.
@@
 We prove security in the random oracle model.
-Unforgeability reduces to the Module-SIS problem, matching the security of single-signer ML-DSA, while privacy against coalitions of up to T1T-1 corrupted parties reduces to Module-LWE and PRF security.
+Unforgeability reduces to the Module-SIS problem, matching the security of single-signer ML-DSA.
+Share privacy is statistical (no computational assumptions) via the Shamir nonce DKG, with conditional min-entropy exceeding $5\times$ the secret key entropy for $|S|\leq 17$; no computational assumptions are needed for response privacy.
+Commitment and r0-check privacy reduce to PRF security.
 Deployment Profiles.
@@
 A TEE/HSM coordinator performs the r0-check and hint generation inside a trusted enclave, achieving optimal 3-round signing.
-We prove EUF-CMA security under Module-SIS and privacy under Module-LWE (Theorem 4.2).
+We prove EUF-CMA security under Module-SIS (Theorem 4.2), with statistical response privacy (Theorem 4.4) and PRF-based mask privacy (Lemma 4.9).
 Profile P2 (Fully Distributed):
@@
 A trusted dealer creates Shamir shares of $\mathsf{sk}=(\mathbf{s}_{1},\mathbf{s}_{2})$ and establishes pairwise seeds via ML-KEM.
-The 3-round signing protocol has parties (1) commit to nonces $\mathbf{y}_{i}\in[-\gamma_{1}/|S|,\gamma_{1}/|S|]$, (2) exchange masked commitments to derive challenge $c$, and (3) send masked responses $\mathbf{U}_{i}=\mathbf{y}_{i}+\lambda_{i}\cdot c\cdot\mathbf{s}_{1,i}+\mathbf{m}_{i}$.
-The combiner aggregates $\mathbf{z}=\sum_{i}\mathbf{U}_{i}=\mathbf{y}+c\mathbf{s}_{1}$ (masks cancel) and outputs $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ if $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$.
+Before each signing attempt, parties execute a Shamir nonce DKG to jointly generate degree-(T1)(T\!-\!1) shares of the nonce $\mathbf{y}$ (this can be preprocessed offline).
+The 3-round online signing protocol then has parties (1) commit to their DKG nonce shares, (2) exchange masked commitments to derive challenge $c$, and (3) send responses $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ (no Lagrange coefficient, no mask).
+The combiner applies Lagrange coefficients during aggregation: $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}=\mathbf{y}+c\mathbf{s}_{1}$, and outputs $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ if $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$.
 1.4 Comparison with Prior and Concurrent Work
 Table 1 summarizes threshold ML-DSA schemes, comparing arbitrary threshold support, dishonest-majority security, round complexity, and FIPS 204 compatibility.
-Scheme $T$ Size Rnd FIPS Dishon. UC Async
-del Pino-Niot [dPN25] $\leq 8$ 3.3 KB 3 $\times$ $\times$
-Noise Flooding [BCDG21] $\infty$ 17 KB 3 $\times$ $\times$ $\times$
-Ringtail [BMMSZ25] $\infty$ 13 KB 2 $\times$ $\times$ $\times$
+Scheme $T$ Size Rnd FIPS Dishon. UC Async Privacy
+del Pino-Niot [dPN25] $\leq 8$ 3.3 KB 3 $\times$ $\times$ Comp.
+Noise Flooding [BCDG21] $\infty$ 17 KB 3 $\times$ $\times$ $\times$ Comp.
+Ringtail [BMMSZ25] $\infty$ 13 KB 2 $\times$ $\times$ $\times$ Comp.
 Concurrent Work (2025–2026)
-Bienstock et al. [BDEPT25] $\infty$ 3.3 KB 37–136 $\times$ $\times$
-Celi et al. [CDENP26] $\leq 6$ 3.3 KB 6 $\times$ $\times$
-Trilithium [Trilithium25] 2 3.3 KB 3 $\times$ $\times$
-This work (P1) $\infty$ 3.3 KB 3 (TEE) $\times$
-This work (P2) $\infty$ 3.3 KB 5 $\times$
-This work (P3+) $\infty$ 3.3 KB 2 (1/2 CP)
+Bienstock et al. [BDEPT25] $\infty$ 3.3 KB 37–136 $\times$ $\times$ Comp.
+Celi et al. [CDENP26] $\leq 6$ 3.3 KB 6 $\times$ $\times$ Comp.
+Trilithium [Trilithium25] 2 3.3 KB 3 $\times$ $\times$ Comp.
+This work (P1) $\infty$ 3.3 KB 3 (TEE) $\times$ Stat.
+This work (P2) $\infty$ 3.3 KB 5 $\times$ Stat.
+This work (P3+) $\infty$ 3.3 KB 2 (1/2 CP) Stat.
 Table 1: Comparison of threshold ML-DSA schemes.
@@
 require honest majority.
-“(TEE)” and “(1/2 CP)” indicate additional trust assumptions beyond threshold.
-“Async” = semi-asynchronous signer participation.
-P3+ requires 1 signer step + 1 server round = 2 logical rounds.
-P1 UC security assumes ideal TEE (Theorem 4.10).
+“Privacy”: Comp.
+= computational (PRF-based), Stat.
+= statistical (no computational assumptions, min-entropy $\geq 5\times$ secret key entropy) for response shares $\mathbf{z}_{i}$; commitment/r0-check privacy uses PRF-based masks ($|S\setminus C|\geq 2$).
+“(TEE)” = dishonest-majority via TEE trust (not purely cryptographic); “(1/2 CP)” = via 1-of-2 computation-party honest assumption.
+P3+ = 1 signer step + 1 server round.
+P1 UC assumes ideal TEE (Theorem 4.13).
+Plus 1 offline preprocessing round (nonce DKG).
 Three concurrent works address threshold ML-DSA with FIPS 204 compatibility.
@@
 Profile P3+ achieves UC security (Theorem 6.5) under the 1-of-2 CP honest assumption, with semi-asynchronous signer participation (2 logical rounds).
-We contribute: (1) Masked Lagrange reconstruction, adapting pairwise-canceling masks [CGGMP20] to lattices while addressing rejection sampling, r0-check key leakage, and Irwin-Hall nonce security; (2) Three deployment profiles (P1-TEE, P2-MPC, P3+-2PC) with different trust/performance trade-offs, including the novel semi-async model for P3+; (3) Security proofs reducing to Module-SIS/LWE, including complete UC security proofs for P2 and P3+; (4) A reference implementation with benchmarks from $(2,3)$ to $(32,45)$ threshold configurations.
+We contribute: (1) Shamir nonce DKG for threshold ML-DSA, achieving statistical response privacy (no computational assumptions) and in coordinator-based profiles (P1, P3+), eliminating the two-honest requirement for response shares ($|S|\geq T$ suffices; see Remark 3.7)—to our knowledge, the first threshold lattice-based signature with statistical share privacy and $|S|=T$ signing sets; (2) Pairwise-canceling masks adapted to the lattice setting for commitment and r0-check values, addressing rejection sampling and r0-check key leakage; (3) Three deployment profiles (P1-TEE, P2-MPC, P3+-2PC) with different trust/performance trade-offs, including the novel semi-async model for P3+; (4) Complete security proofs, with unforgeability reducing to Module-SIS, statistical privacy via Shamir nonce DKG (Theorem 4.4), and UC security proofs for all profiles; (5) A reference implementation with benchmarks from $(2,3)$ to $(32,45)$ threshold configurations.
 1.5 Paper Organization
@@
   • $\beta=\tau\cdot\eta=196$: signature bound
+  • $\omega=55$: hint weight bound (max number of $1$-entries in hint $\mathbf{h}$)
+  • $d=13$: dropping bits in $\mathsf{Power2Round}$ ($\mathbf{t}=\mathbf{t}_{1}\cdot 2^{d}+\mathbf{t}_{0}$)
 Key Generation.
@@
 Remark 2.3 (Lagrange Coefficient Magnitude).
-Lagrange coefficients $\lambda_{i}=(-1)^{i-1}\binom{T}{i}$ grow as $\approx 2^{T}/\sqrt{T}$.
+For the standard evaluation points $\{1,2,\ldots,T\}$, Lagrange coefficients satisfy $\lambda_{i}=(-1)^{i-1}\binom{T}{i}$, growing as $\approx 2^{T}/\sqrt{T}$.
 For $T\leq 8$, coefficients remain small ($\binom{8}{4}=70$), enabling short-coefficient schemes [dPN25].
@@
   • $\mathsf{KeyGen}(1^{\kappa},N,T)\to(\mathsf{pk},\{\mathsf{sk}_{i}\}_{i\in[N]})$: Generates public key and secret key shares
-  • $\mathsf{Sign}(\{\mathsf{sk}_{i}\}_{i\in S},\mu)\to\sigma$: Interactive protocol among parties in $S$ (where $|S|\geq T+1$; see Section 4 for the two-honest requirement) to sign message $\mu$
+  • $\mathsf{Sign}(\{\mathsf{sk}_{i}\}_{i\in S},\mu)\to\sigma$: Interactive protocol among parties in $S$ (where $|S|\geq T$) to sign message $\mu$
   • $\mathsf{Verify}(\mathsf{pk},\mu,\sigma)\to\{0,1\}$: Deterministic verification
-Remark 2.5 ($T+1$ Signing Requirement).
-Our protocol requires $|S|\geq T+1$ signers (not $T$) to ensure at least two honest parties for mask privacy. This is standard for pairwise-mask schemes [CGGMP20, DKOSS24]; see Remark 4.1.
+Remark 2.5 (Signing Set Size).
+With Shamir nonce DKG and coordinator-based communication (Remark 3.7), Profiles P1 and P3+ require only $|S|\geq T$ signers—the information-theoretic minimum for Shamir reconstruction.
+This improves upon standard pairwise-mask schemes [CGGMP20, DKOSS24] that require $|S|\geq T+1$.
+The nonce DKG provides statistical share privacy (no computational assumptions) even with a single honest party (see Theorem 4.4 and Remark 4.1).
+In P2’s broadcast model, $|S\setminus C|\geq 2$ is required for commitment privacy (Lemma 4.9).
 Definition 2.6 (Unforgeability (EUF-CMA)).
@@
 Even for moderate $T$ (e.g., $T=16$), the coefficients $\binom{16}{8}=12870$ can cause the naive approach to inflate norms beyond the rejection threshold.
-Our solution is based on two key observations:
+Our solution is based on two key techniques:
   1. 1.
-    Masking the coefficients: We add pairwise-canceling masks $\mathbf{m}_{i}$ to each party’s contribution.
-While individual contributions $\mathbf{U}_{i}$ may have arbitrary magnitude, the masks cancel in the sum, leaving the correct short result.
+    Shamir nonce DKG: Parties jointly generate the nonce $\mathbf{y}$ via a Shamir-style distributed key generation, so that both $\mathbf{y}$ and the long-term secret $\mathbf{s}_{1}$ are degree-(T1)(T\!-\!1) Shamir sharings.
+Each party computes $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ (without Lagrange coefficients), and the combiner reconstructs $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}=\mathbf{y}+c\mathbf{s}_{1}$.
+This achieves statistical share privacy (conditional min-entropy $\geq 5\times$ secret key entropy for $|S|\leq 17$, requiring no computational assumptions), in coordinator-based profiles (P1, P3+), eliminates the two-honest requirement ($|S|\geq T$ suffices; see Remark 3.7), and requires no pairwise masks on the response $\mathbf{z}_{i}$.
   2. 2.
-    Separating nonce and secret: The nonce $\mathbf{y}_{i}$ and secret share $\mathbf{s}_{1,i}$ must be treated differently.
-Nonces are summed directly ($\mathbf{y}=\sum_{i}\mathbf{y}_{i}$), while secret shares are reconstructed via Lagrange interpolation ($\mathbf{s}_{1}=\sum_{i}\lambda_{i}\mathbf{s}_{1,i}$).
+    Pairwise-canceling masks: For the commitment values $\mathbf{W}_{i}$ and the r0-check shares $\mathbf{V}_{i}$, we add pairwise-canceling masks that hide individual contributions while preserving correctness in the sum.
 3.2 Threat Model and Deployment Profiles
@@
 Adversary model Static, malicious (up to T1T-1 parties)
-Signing set size $|S|\geq T+1$ (two-honest requirement)
+Signing set size $|S|\geq T$
 Security goal EUF-CMA under Module-SIS
-Privacy goal T1T-1 privacy under Module-LWE
+Privacy goal Statistical share privacy (Theorem 4.4)
 Output format FIPS 204-compatible signature
@@
 11:return $\mathsf{pk}$, $\{\mathsf{sk}_{i}=(\mathbf{s}_{1,i},\mathbf{s}_{2,i},\{\mathsf{seed}_{*,*}\})\}_{i\in[N]}$
+3.4 Shamir Nonce DKG
+The primary technique enabling statistical share privacy is the Shamir nonce DKG: parties jointly generate the signing nonce $\mathbf{y}$ as a degree-(T1)(T\!-\!1) Shamir sharing, matching the structure of the long-term secret $\mathbf{s}_{1}$.
+This is executed once per signing attempt and can be preprocessed offline.
+Algorithm 2 $\mathsf{NonceDKG}(S)$: Distributed Nonce Generation
+1:Input: Signing set $S$ with $|S|\geq T$
+2:for each party $i\in S$ in parallel do
+3:  Sample secret contribution $\hat{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$
+4:  Construct degree-(T1)(T\!-\!1) polynomial over $R_{q}^{\ell}$:
+5:  $\mathbf{f}_{i}(x)=\hat{\mathbf{y}}_{i}+\mathbf{a}_{i,1}x+\cdots+\mathbf{a}_{i,T-1}x^{T-1}$
+6:  where $\mathbf{a}_{i,k}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ for $k=1,\ldots,T\!-\!1$
+7:  Send $\mathbf{f}_{i}(j)$ to each party $j\in S$ over a private channel $\triangleright$ Encrypted via pairwise ML-KEM seeds
+8:end for
+9:for each party $j\in S$ do
+10:  Compute nonce share: $\mathbf{y}_{j}=\sum_{i\in S}\mathbf{f}_{i}(j)$
+11:end for
+The combined polynomial is $\mathbf{F}(x)=\sum_{i\in S}\mathbf{f}_{i}(x)$, with $\mathbf{F}(0)=\sum_{i}\hat{\mathbf{y}}_{i}=\mathbf{y}$ (the nonce) and $\mathbf{F}(j)=\mathbf{y}_{j}$ (the shares).
+This is a valid degree-(T1)(T\!-\!1) Shamir sharing of $\mathbf{y}$ over $R_{q}^{\ell}$.
+Remark 3.1 (Why Uniform Higher-Degree Coefficients).
+The coefficients $\mathbf{a}_{i,k}$ for $k\geq 1$ are sampled uniformly over $R_{q}^{\ell}$ (not from a short distribution).
+This is essential: the adversary observes T1T-1 evaluations of $\mathbf{f}_{h}$ (the honest party’s polynomial), leaving exactly one degree of freedom.
+The uniform prior distribution of $\mathbf{a}_{h,T-1}$ ensures high entropy in this free parameter; combined with the bounded-nonce constraint, the honest party’s share $\mathbf{f}_{h}(x_{h})$ retains conditional min-entropy $\geq n\ell\cdot\log_{2}(2\gamma_{1}/|S|)$ bits per session.
+See Theorem 4.4 and Remark 4.6.
+Remark 3.2 (Communication Cost).
+The nonce DKG requires $|S|\cdot(|S|-1)$ point-to-point messages, each of size $n\ell\cdot\lceil\log_{2}q\rceil$ bits ($\approx 3.8$ KB for ML-DSA-65).
+For a $(3,5)$-threshold with $|S|=3$, this is 6 messages totaling $\approx 23$ KB.
+The DKG round is independent of the message to be signed and can be executed offline as a preprocessing step, so online signing latency is unchanged.
 3.5 Pairwise-Canceling Masks
-The key innovation of our protocol is the use of pairwise-canceling masks.
-For any signing set $S$, each party $i\in S$ computes a mask $\mathbf{m}_{i}\in R_{q}^{\ell}$ such that $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$.
+Pairwise-canceling masks are used to hide the commitment values $\mathbf{W}_{i}$ (for challenge derivation) and the r0-check shares $\mathbf{V}_{i}$ (which contain $\lambda_{i}\cdot c\cdot\mathbf{s}_{2,i}$).
+For any signing set $S$, each party $i\in S$ computes a mask $\mathbf{m}_{i}$ such that $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$.
 Definition 3.3 (Pairwise-Canceling Mask).
@@
 $\mathbf{m}_{i}^{(\mathsf{dom})}=\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{dom})$
-where $\mathsf{PRF}:\{0,1\}^{256}\times\{0,1\}^{*}\to R_{q}^{d}$ is instantiated with SHAKE-256, and $d$ is the appropriate dimension ($\ell$ for response masks, $k$ for commitment masks).
-We use domain separators $\mathsf{dom}=\mathsf{nonce}\|\texttt{"resp"}$ for response masks $\mathbf{m}_{i}\in R_{q}^{\ell}$ and $\mathsf{dom}=\mathsf{nonce}\|\texttt{"comm"}$ for commitment masks $\mathbf{m}_{i}^{(w)}\in R_{q}^{k}$.
+where $\mathsf{PRF}:\{0,1\}^{256}\times\{0,1\}^{*}\to R_{q}^{d}$ is instantiated with SHAKE-256, and $d$ is the appropriate dimension ($k$ for commitment masks, $k$ for r0-check masks).
+We use domain separators $\mathsf{dom}=\mathsf{nonce}\|\texttt{"comm"}$ for commitment masks $\mathbf{m}_{i}^{(w)}\in R_{q}^{k}$ and $\mathsf{dom}=\mathsf{nonce}\|\texttt{"s2"}$ for r0-check masks $\mathbf{m}_{i}^{(s2)}\in R_{q}^{k}$.
 Lemma 3.4 (Mask Cancellation).
 For any signing set $S$ and nonce, $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$.
-Proof 3.3.
+Proof 3.5.
 Each pair $(i,j)$ with $i<j$ contributes:
@@
 3.6 Signing Protocol
-The signing protocol is a three-round interactive protocol among parties in the signing set $S$ (where $|S|\geq T+1$ to ensure at least two honest parties; see Remark 4.1).
+The signing protocol is a three-round interactive protocol among parties in the signing set $S$ (where $|S|\geq T$), preceded by an offline nonce DKG preprocessing step (Algorithm 2).
 We assume a broadcast channel or a designated combiner (TEE/HSM in Profile P1).
@@
 3:
-4:Round 1: Nonce Commitment
-5:for each party $i\in S$ in parallel do
-6:  Sample $\mathbf{y}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$
-7:  Compute $\mathbf{w}_{i}=\mathbf{A}\mathbf{y}_{i}$
-8:  Sample $r_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
-9:  Compute $\mathsf{Com}_{i}=H(\texttt{"com"}\|\mathbf{y}_{i}\|\mathbf{w}_{i}\|r_{i})$
-10:  Broadcast $\mathsf{Com}_{i}$
-11:  Store $(\mathbf{y}_{i},\mathbf{w}_{i},r_{i})$ as local state
-12:end for
-13:
-14:Round 2: Masked Reveal and Challenge
-15:Pre-round: Derive preliminary nonce $\mathsf{nonce}_{0}=H(\texttt{"nonce0"}\|\mathsf{Com}_{1}\|\cdots\|\mathsf{Com}_{|S|})$
-16:for each party $i\in S$ in parallel do
-17:  Compute commitment mask $\mathbf{m}_{i}^{(w)}\in R_{q}^{k}$ using $\mathsf{nonce}_{0}\|\texttt{"comm"}$
-18:  Broadcast $(\mathbf{W}_{i}=\mathbf{w}_{i}+\mathbf{m}_{i}^{(w)},r_{i})$ $\triangleright$ Masked commitment
-19:end for
-20:All parties / Combiner:
-21: Aggregate $\mathbf{W}=\sum_{j\in S}\mathbf{W}_{j}=\sum_{j}\mathbf{w}_{j}+\sum_{j}\mathbf{m}_{j}^{(w)}=\mathbf{w}$ $\triangleright$ Masks cancel
-22: Compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{W},2\gamma_{2})$
-23: Compute $\tilde{c}=H_{\mathsf{chal}}(\mu\|\mathbf{w}_{1})$ and expand to challenge $c\in R_{q}$ $\triangleright$ FIPS 204 spec
-24: Derive $\mathsf{nonce}=H(\texttt{"nonce"}\|c\|\mu\|\mathsf{sort}(S))$
-25:
-26:Round 3: Masked Response
-27:for each party $i\in S$ in parallel do
-28:  Compute Lagrange coefficient $\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\mod q$
-29:  Compute masked secret contribution: $\boldsymbol{\nu}_{i}=\lambda_{i}\cdot c\cdot\mathbf{s}_{1,i}$
-30:  Compute mask $\mathbf{m}_{i}$ as in Definition 3.1
-31:  Compute masked response: $\mathbf{U}_{i}=\mathbf{y}_{i}+\boldsymbol{\nu}_{i}+\mathbf{m}_{i}$
-32:  Compute $\mathbf{V}_{i}=\lambda_{i}\cdot c\cdot\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$ $\triangleright$ For r0-check
-33:  Send $(\mathbf{U}_{i},\mathbf{V}_{i})$ to combiner
-34:end for
-35:
-36:Aggregation and r0-Check (see Subsection 3.9)
-37:Combiner computes $\mathbf{z}=\sum_{i\in S}\mathbf{U}_{i}\mod q$
-38:if $\left\|\mathbf{z}\right\|_{\infty}\geq\gamma_{1}-\beta$ then
-39:  return $\bot$ (z-bound abort, retry from Round 1)
-40:end if
-41:r0-check: Verify $\left\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\right\|_{\infty}<\gamma_{2}-\beta$
-42:if r0-check fails then
-43:  return $\bot$ (r0 abort, retry from Round 1)
+4:Preprocessing: Nonce DKG (offline)
+5:Execute $\mathsf{NonceDKG}(S)$ (Algorithm 2) to obtain nonce shares $\{\mathbf{y}_{i}\}_{i\in S}$
+6:
+7:Round 1: Nonce Commitment
+8:for each party $i\in S$ in parallel do
+9:  Compute $\mathbf{w}_{i}=\mathbf{A}\mathbf{y}_{i}$ $\triangleright$ Using DKG nonce share
+10:  Sample $r_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
+11:  Compute $\mathsf{Com}_{i}=H(\texttt{"com"}\|\mathbf{y}_{i}\|\mathbf{w}_{i}\|r_{i})$
+12:  Broadcast $\mathsf{Com}_{i}$
+13:  Store $(\mathbf{y}_{i},\mathbf{w}_{i},r_{i})$ as local state
+14:end for
+15:
+16:Round 2: Masked Reveal and Challenge
+17:Pre-round: Derive preliminary nonce $\mathsf{nonce}_{0}=H(\texttt{"nonce0"}\|\mathsf{Com}_{i_{1}}\|\cdots\|\mathsf{Com}_{i_{|S|}})$ where $(i_{1},\ldots,i_{|S|})=\mathsf{sort}(S)$
+18:for each party $i\in S$ in parallel do
+19:  Compute Lagrange coefficient $\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\bmod q$
+20:  Compute commitment mask $\mathbf{m}_{i}^{(w)}\in R_{q}^{k}$ using $\mathsf{nonce}_{0}\|\texttt{"comm"}$
+21:  Send $(\mathbf{W}_{i}=\lambda_{i}\cdot\mathbf{w}_{i}+\mathbf{m}_{i}^{(w)},r_{i})$ to coordinator $\triangleright$ NOT broadcast; see Remark 4.1
+22:end for
+23:Coordinator:
+24: Aggregate $\mathbf{W}=\sum_{j\in S}\mathbf{W}_{j}=\sum_{j}\lambda_{j}\mathbf{w}_{j}+\sum_{j}\mathbf{m}_{j}^{(w)}=\mathbf{A}\mathbf{y}=\mathbf{w}$ $\triangleright$ Masks cancel
+25: Compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{W},2\gamma_{2})$
+26: Compute $\tilde{c}=H_{\mathsf{chal}}(\mu\|\mathbf{w}_{1})$ and expand to challenge $c\in R_{q}$ $\triangleright$ FIPS 204 spec
+27:if $c$ is not invertible in $R_{q}$ then
+28:  return $\bot$ (challenge invertibility abort, retry from Round 1) $\triangleright$ Negligible probability
+29:end if
+30: Derive $\mathsf{nonce}=H(\texttt{"nonce"}\|\tilde{c}\|\mu\|\mathsf{sort}(S))$
+31: Broadcast $(\mathbf{w}_{1},\tilde{c},\mathsf{nonce})$ to all parties $\triangleright$ Only $\mathsf{HighBits}$; individual $\mathbf{W}_{i}$ stay private
+32:
+33:Round 3: Response
+34:for each party $i\in S$ in parallel do
+35:  Compute response: $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ $\triangleright$ No $\lambda_{i}$, no mask
+36:  Compute $\mathbf{V}_{i}=\lambda_{i}\cdot c\cdot\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$ $\triangleright$ Masked, for r0-check
+37:  Send $(\mathbf{z}_{i},\mathbf{V}_{i})$ to combiner
+38:end for
+39:
+40:Aggregation and r0-Check (see Subsection 3.10)
+41:Combiner computes $\mathbf{z}=\sum_{i\in S}\lambda_{i}\cdot\mathbf{z}_{i}\mod q$ $\triangleright$ Lagrange at combiner
+42:if $\left\|\mathbf{z}\right\|_{\infty}\geq\gamma_{1}-\beta$ then
+43:  return $\bot$ (z-bound abort, retry from Round 1)
 44:end if
-45:Compute $\mathbf{r}=\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d}$ $\triangleright$ $\mathbf{w}-c\mathbf{s}_{2}+c\mathbf{t}_{0}$
-46:Compute $c\mathbf{t}_{0}$ using $c\mathbf{s}_{2}$ from r0-check $\triangleright$ See Subsection 3.9
-47:Compute hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{r},2\gamma_{2})$
-48:if $\mathsf{weight}(\mathbf{h})>\omega$ then
-49:  return $\bot$ (hint weight abort)
-50:end if
-51:return signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$
+45:r0-check: Verify $\left\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\right\|_{\infty}<\gamma_{2}-\beta$
+46:if r0-check fails then
+47:  return $\bot$ (r0 abort, retry from Round 1)
+48:end if
+49:Compute $\mathbf{r}=\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d}$ $\triangleright$ $\mathbf{w}-c\mathbf{s}_{2}+c\mathbf{t}_{0}$
+50:Compute $c\mathbf{t}_{0}$ using $c\mathbf{s}_{2}$ from r0-check $\triangleright$ See Subsection 3.10
+51:Compute hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{r},2\gamma_{2})$
+52:if $\mathsf{weight}(\mathbf{h})>\omega$ then
+53:  return $\bot$ (hint weight abort)
+54:end if
+55:return signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$
 Remark 3.6 (Commitment Model).
 The Round 1 commitments $\mathsf{Com}_{i}$ bind parties to their nonces before seeing others’ values, preventing adaptive attacks.
-However, we do not publicly verify these commitments during normal signing—the masked values $\mathbf{W}_{i}$ and $\mathbf{U}_{i}$ cannot be verified without revealing private mask seeds.
+However, we do not publicly verify these commitments during normal signing—the masked values $\mathbf{W}_{i}$ cannot be verified without revealing private mask seeds, and the responses $\mathbf{z}_{i}$ are privacy-protected by the nonce DKG (Theorem 4.4).
 Instead, we follow an optimistic model: if signing fails (excessive aborts, invalid signatures), the TEE coordinator triggers a blame protocol where parties reveal their commitments for verification.
@@
 See Section 6.2 for details.
+Remark 3.7 (Why $\mathbf{W}_{i}$ Must Not Be Broadcast).
+[…]
+In Profile P1, the TEE coordinator naturally keeps $\mathbf{W}_{i}$ private.
+In Profile P3+, the combiner serves as coordinator.
+In Profile P2 (no designated coordinator), parties broadcast $\mathbf{W}_{i}$ directly and rely on mask hiding (Lemma 4.9, requiring $|S\setminus C|\geq 2$) to protect individual contributions.
 Remark 3.8 (Domain Separation).
-We use distinct domain tags ("com", "nonce0", "nonce", "resp", "comm") for all hash function uses to ensure random oracle independence.
+We use distinct domain tags ("com", "nonce0", "nonce", "s2", "comm", "pairwise_seed") for all hash function uses to ensure random oracle independence.
 Remark 3.9 (Nonce Range Rounding).
 When $\gamma_{1}$ is not divisible by $|S|$, rounding slightly reduces the nonce range.
-This has negligible impact: for ML-DSA-65 with $|S|=17$, the range loss is $2\times 10^{-5}$ per coordinate and does not affect the z-bound check since $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor>\gamma_{1}-\beta$.
+This has negligible impact: for ML-DSA-65 with $|S|=17$, the range loss is $2\times 10^{-5}$ per coordinate and does not affect the z-bound check since $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor\geq\gamma_{1}-(|S|-1)>\gamma_{1}-\beta$ (holds for all $|S|\leq\beta+1=197$, covering all practical configurations).
 See Appendix G for details.
@@
 If all parties are honest and the protocol does not abort, then $\mathsf{Verify}(\mathsf{pk},\mu,\sigma)=1$.
-Proof 3.8.
-The aggregated response is $\mathbf{z}=\sum_{i\in S}\mathbf{U}_{i}=\sum_{i}\mathbf{y}_{i}+c\cdot\sum_{i}\lambda_{i}\mathbf{s}_{1,i}+\sum_{i}\mathbf{m}_{i}=\mathbf{y}+c\mathbf{s}_{1}$, using nonce summation, Lagrange reconstruction, and mask cancellation (Lemma 3.2).
-Since $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$, verification proceeds as in single-signer ML-DSA.
+Proof 3.11.
+The aggregated response is $\mathbf{z}=\sum_{i\in S}\lambda_{i}\mathbf{z}_{i}=\sum_{i}\lambda_{i}(\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i})=\sum_{i}\lambda_{i}\mathbf{y}_{i}+c\cdot\sum_{i}\lambda_{i}\mathbf{s}_{1,i}=\mathbf{y}+c\mathbf{s}_{1}$.
+Both $\{\mathbf{y}_{i}\}$ and $\{\mathbf{s}_{1,i}\}$ are degree-(T1)(T\!-\!1) Shamir shares of $\mathbf{y}$ and $\mathbf{s}_{1}$ respectively, evaluated at the same points, so Lagrange reconstruction yields both correct secrets simultaneously.
+The masked commitment aggregation gives $\mathbf{W}=\sum_{i}\mathbf{W}_{i}=\sum_{i}(\lambda_{i}\mathbf{w}_{i}+\mathbf{m}_{i}^{(w)})=\mathbf{A}\sum_{i}\lambda_{i}\mathbf{y}_{i}=\mathbf{A}\mathbf{y}=\mathbf{w}$, where mask cancellation follows from Lemma 3.4.
+Since $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ and $\mathbf{w}=\mathbf{A}\mathbf{y}$, verification proceeds as in single-signer ML-DSA.
 Remark 3.12 (Implementation Notes).
-(1) While individual $\mathbf{U}_{i}$ may have large coefficients, the sum $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ is short (<q/2<q/2), so mod $q$ reduction is correct.
-(2) Party $i$ must compute $\mathbf{U}_{i}=\mathbf{y}_{i}+\lambda_{i}\cdot c\cdot\mathbf{s}_{1,i}+\mathbf{m}_{i}$, not $\lambda_{i}\cdot(\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i})+\mathbf{m}_{i}$ (the latter incorrectly weights nonces).
+(1) Each party computes $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ without applying the Lagrange coefficient $\lambda_{i}$.
+The combiner applies $\lambda_{i}$ during aggregation: $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}$.
+While individual $\lambda_{i}\mathbf{z}_{i}$ may have large coefficients, the sum $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ is short (<q/2<q/2), so mod $q$ reduction is correct.
 (2) Signatures are FIPS 204-compatible: same format, size (3309 bytes for ML-DSA-65), and pass unmodified verifiers.
-Signer-side differences (Irwin-Hall nonces, interactive protocol) do not affect security or verification.
+Signer-side differences (Irwin-Hall nonces, nonce DKG, interactive protocol) do not affect security or verification.
 3.8 Nonce Distribution and Security
-The aggregated nonce $\mathbf{y}=\sum_{i\in S}\mathbf{y}_{i}$ follows an Irwin-Hall distribution concentrated around zero.
-This preserves EUF-CMA security: using smooth Rényi divergence bounds from [Raccoon2024], we verify $R^{\epsilon}_{2}\leq 1.003$ for $|S|\leq 17$ and $R^{\epsilon}_{2}\leq 1.05$ for $|S|\leq 33$ signers (Appendix D), yielding negligible security loss ($2^{-40}$ even for $2^{64}$ queries).
+The aggregated nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ (the sum of each party’s constant-term contribution from the nonce DKG, equivalently $\mathbf{y}=\sum_{i\in S}\lambda_{i}\mathbf{y}_{i}$ by Lagrange reconstruction) follows an Irwin-Hall distribution concentrated around zero.
+This preserves EUF-CMA security: using smooth Rényi divergence bounds from [Raccoon2024], we compute per-coordinate $R^{\epsilon}_{2,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$.
+Since nonce coordinates are sampled independently, the full-vector divergence is $R_{2}^{\mathsf{vec}}=(R^{\epsilon}_{2,\mathsf{coord}})^{1280}\approx 2^{5.4}$ for $|S|\leq 17$ ($\approx$3-bit security loss).
+For $|S|=33$, $R_{2}^{\mathsf{vec}}\approx 2^{75}$ ($\approx$38-bit loss).
+See Theorem 4.11 and Appendix D.
 3.9 Abort Probability
@@
 Our theorems target this profile.
-Remark 4.1 (Why $|S|\geq T+1$: Two-Honest Requirement).
-Our protocol requires signing sets of size at least $T+1$ rather than $T$.
-This ensures that even when the adversary corrupts the maximum |C|=T1|C|=T-1 parties, every signing set contains at least two honest parties—necessary for mask privacy (see Lemma 4.6).
+Remark 4.1 ($|S|\geq T$ Suffices in Coordinator-Based Profiles).
+With Shamir nonce DKG and the coordinator-based communication model (Remark 3.7), Profiles P1 and P3+ require signing sets of size only $|S|\geq T$.
+The honest party’s response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ has high conditional min-entropy (Theorem 4.4) from the adversary’s view, vastly exceeding the secret key entropy.
+The key privacy argument rests on what the adversary observes:
+  1. 1.
+    Response $\mathbf{z}_{h}$: Statistical privacy via nonce DKG (Theorem 4.4). Requires no computational assumptions, holds even with $|S\setminus C|=1$.
+  2. 2.
+    r0-check shares $\mathbf{V}_{i}$: Never broadcast—processed inside the TEE (P1), as MPC inputs (P2), or via 2PC (P3+).
+Even with $|S\setminus C|=1$, the adversary does not observe individual $\mathbf{V}_{h}$.
+  3. 3.
+    Masked commitments $\mathbf{W}_{i}$: Sent only to the coordinator (not broadcast).
+The coordinator publishes only $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{W},2\gamma_{2})$, which is already implicit in the signature (verifiers reconstruct $\mathbf{w}_{1}$ from $(\mathbf{A},\mathbf{z},c,\mathbf{h})$).
+This adds zero information to the adversary’s view.
+Why individual $\mathbf{W}_{i}$ must stay private.
+If $\mathbf{W}_{h}$ were broadcast with $|S\setminus C|=1$, the pairwise mask $\mathbf{m}_{h}^{(w)}$ would be computable by the adversary (all seeds involve corrupted counterparties), exposing $\lambda_{h}\mathbf{A}\mathbf{y}_{h}$.
+Since $\mathbf{A}\in R_{q}^{k\times\ell}$ with $k>\ell$ is generically injective (under the NTT isomorphism, a random $k\times\ell$ matrix over $\mathbb{Z}_{q}$ has full column rank with probability $\geq 1-\ell/q$), the adversary could recover $\mathbf{y}_{h}$ via linear algebra—this is not a Module-LWE instance because $\mathbf{y}_{h}$ has full-size (not short) coefficients.
+Combined with $\mathbf{z}_{h}$ (extractable from the public signature), this would yield $\mathbf{s}_{1,h}=c^{-1}(\mathbf{z}_{h}-\mathbf{y}_{h})$, a complete key recovery.
+The coordinator-only communication model eliminates this attack vector entirely.
+Profile-specific analysis.
+In P1, the TEE coordinator naturally keeps individual $\mathbf{W}_{i}$ private.
+In P3+, the combiner serves as coordinator (trusted for privacy under the 1-of-2 CP honest assumption).
+In P2 (no designated coordinator), parties broadcast $\mathbf{W}_{i}$ directly; mask hiding (Lemma 4.9) protects individual values under the existing $|S\setminus C|\geq 2$ condition, which P2 already requires.
 4.2 Unforgeability
@@
 If Module-SIS is hard and $H$ is a random oracle, then the masked threshold ML-DSA scheme is EUF-CMA secure:
-$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}$
-where $q_{H},q_{s}$ are random oracle and signing queries. The security loss matches single-signer ML-DSA [AFLT12, DKL18].
+$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\left(O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\tbinom{N}{2}\cdot\epsilon_{\mathsf{PRF}}\right)}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
+where $q_{H},q_{s}$ are random oracle and signing queries, $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the full-vector Rényi divergence with per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$, and $\epsilon_{\mathsf{IH}}<10^{-30}$ (Theorem 4.11).
+The $\sqrt{R_{2}^{\mathsf{vec}}}$ factor arises from the game hop between Irwin-Hall and uniform nonce distributions via Cauchy-Schwarz; for $|S|\leq 17$, $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$, giving $\sqrt{R_{2}^{\mathsf{vec}}}\approx 2^{2.7}$ ($\approx$3-bit loss).
+For $|S|=33$, $R_{2}^{\mathsf{vec}}\approx 2^{75}$ ($\approx$38-bit loss).
+The loss matches the per-coordinate composition inherent to the generic Rényi proof technique [Raccoon2024].
 Proof 4.3.
@@
 By $(T,N)$-Shamir security (any <T<T shares reveal nothing): $\Pr[\mathsf{Win}_{3}]=\Pr[\mathsf{Win}_{2}]$.
+$\mathsf{Game}_{3.5}$ (Uniform Nonces).
+Replace the Irwin-Hall nonce distribution with uniform nonces: in each signing session, sample $\mathbf{y}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1},\ldots,\gamma_{1}\}^{n\ell}$ uniform instead of $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ (Irwin-Hall).
+By the smooth Rényi divergence bound (Theorem 4.11 with $\alpha=2$):
+$\Pr[\mathsf{Win}_{3}]\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\Pr[\mathsf{Win}_{3.5}]}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
+where $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the full-vector Rényi divergence (per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$, $n\ell=1280$), giving $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$ for $|S|\leq 17$.
+The tail probability satisfies $\epsilon_{\mathsf{IH}}<10^{-30}$.
+This is the only game hop that incurs the Irwin-Hall security loss; individual shares $\mathbf{z}_{j}$ have high min-entropy by Theorem 4.4 (statistical, no computational assumptions).
 $\mathsf{Game}_{4}$ (Simulated Signatures).
-Simulate signing without knowing $\mathbf{s}_{1}$: sample $\mathbf{z}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1}+\beta+1,\ldots,\gamma_{1}-\beta-1\}^{n\ell}$, compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$, and program $H(\mu\|\mathbf{w}_{1}):=c$.
-The masked values $\mathbf{W}_{j},\mathbf{U}_{j}$ are simulated as uniform by Lemma 4.6.
-By PRF security: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3}]|\leq\binom{|S|}{2}\cdot\epsilon_{\mathsf{PRF}}$.
+Starting from uniform nonces (Game 3.5), simulate signing without knowing $\mathbf{s}_{1}$: sample $\mathbf{z}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1}+\beta+1,\ldots,\gamma_{1}-\beta-1\}^{n\ell}$, compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$, and program $H(\mu\|\mathbf{w}_{1}):=c$.
+The masked values $\mathbf{W}_{j}$ are simulated as uniform by Lemma 4.9.
+By PRF security: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3.5}]|\leq\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
 Reduction to M-SIS.
@@
 $\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]\cdot\begin{bmatrix}\mathbf{v}\\ \Delta c\end{bmatrix}=\boldsymbol{\delta}$
-This yields a short vector for the augmented M-SIS instance $\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]$ with $\|(\mathbf{v},\Delta c)\|_{\infty}\leq\max(2(\gamma_{1}-\beta),\tau)<2\gamma_{1}$.
+This is an inhomogeneous M-SIS instance (SelfTargetMSIS [KLS18]): the residual $\boldsymbol{\delta}$ arises from rounding in $\mathsf{HighBits}$ and satisfies $\|\boldsymbol{\delta}\|_{\infty}<2\gamma_{2}$, which is absorbed into the M-SIS norm bound.
+The extracted vector $(\mathbf{v},\Delta c)$ satisfies $\|(\mathbf{v},\Delta c)\|_{\infty}\leq\max(2(\gamma_{1}-\beta),2)<2\gamma_{1}$, solving the augmented M-SIS instance $\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]$.
+This is the standard reduction used in the Dilithium/FIPS 204 security proof.
 Thus: $\Pr[\mathsf{Win}_{4}]\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}$.
 Conclusion.
-Combining all bounds:
-$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$
-For SHAKE-256 as PRF, $\epsilon_{\mathsf{PRF}}$ is negligible, yielding the stated bound.
+Games 4 and the M-SIS reduction give $\Pr[\mathsf{Win}_{3.5}]\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
+Composing with the Rényi hop (Game $3\to 3.5$) and the birthday bound (Game $1\to 2$):
+$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\left(O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\tbinom{N}{2}\cdot\epsilon_{\mathsf{PRF}}\right)}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
+For ML-DSA-65 with $|S|\leq 17$: $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$ (per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$, $n\ell=1280$), $\epsilon_{\mathsf{IH}}<10^{-30}$, and SHAKE-256 yields negligible $\epsilon_{\mathsf{PRF}}$, giving the stated bound.
 4.3 Privacy
-Theorem 4.4 (Privacy).
-If PRF is secure, then the view of any coalition $C$ with $|C|<T$ can be efficiently simulated given only $\mathsf{pk}$, corrupted shares, and final signatures.
+Theorem 4.4 (Statistical Share Privacy).
+With Shamir nonce DKG, for any signing set $S$ with $|S|\geq T$ and any coalition $C\subset S$ with $|C|\leq T-1$: each honest party’s response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ has conditional min-entropy
+$H_{\infty}(\mathbf{z}_{h}\mid\mathsf{View}_{C})\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
+where $\mathsf{View}_{C}$ denotes the adversary’s view: corrupted shares, DKG evaluations, challenge $c$, the coordinator’s broadcast $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{W},2\gamma_{2})$, and all signatures (current and prior).
+Note that $\mathsf{View}_{C}$ does not include individual $\mathbf{W}_{i}$ values, which are sent only to the coordinator (Remark 3.7).
+Since $\mathbf{w}_{1}$ is already implicit in any valid signature, it adds no information beyond $\sigma$ itself.
+For ML-DSA-65 ($n\ell=1280$, $\gamma_{1}=2^{19}$), the bound exceeds $1280\cdot(20-\log_{2}|S|)$ bits—over $5\times$ the secret key entropy $H(\mathbf{s}_{1,h})\leq 1280\cdot\log_{2}9\approx 4058$ bits for $|S|\leq 17$.
+No computational assumptions are needed; the privacy is statistical.
 Proof 4.5.
-We construct simulator $\mathcal{S}$ and prove indistinguishability via hybrids.
-Simulator $\mathcal{S}$.
-Given $\mathsf{pk}$, corrupted shares $\{\mathsf{sk}_{i}\}_{i\in C}$, and signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$:
-  1. 1.
-    Round 1: For each honest $j$, sample $\mathsf{Com}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$.
-  2. 2.
-    Round 2: Compute $\mathbf{w}_{1}=\mathsf{UseHint}(\mathbf{h},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d})$.
-For each honest $j$, sample $\mathbf{W}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k}$ subject to $\sum_{j\notin C}\mathbf{W}_{j}=\mathbf{w}-\sum_{i\in C}\mathbf{W}_{i}$.
-  3. 3.
-    Round 3: Sample honest $\mathbf{U}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ subject to $\sum_{j\notin C}\mathbf{U}_{j}=\mathbf{z}-\sum_{i\in C}\mathbf{U}_{i}$.
-Hybrid $\mathsf{H}_{0}$ (Real).
-Real protocol execution with honest parties using true secrets.
-Hybrid $\mathsf{H}_{1}$.
-Replace PRF outputs for honest-honest pairs with truly random: for $j,k\notin C$, set $\mathsf{PRF}(\mathsf{seed}_{j,k},\cdot)\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{d}$.
-By PRF security: $|\Pr[\mathsf{H}_{1}]-\Pr[\mathsf{H}_{0}]|\leq\binom{|H|}{2}\cdot\epsilon_{\mathsf{PRF}}$ where $H=S\setminus C$.
-Hybrid $\mathsf{H}_{2}$ (Simulated).
-Replace honest contributions with $\mathcal{S}$’s output.
-Since $|H|\geq 2$, honest parties’ masks include at least one random term unknown to $\mathcal{A}$.
-By the one-time pad property over $R_{q}$: $\Pr[\mathsf{H}_{2}]=\Pr[\mathsf{H}_{1}]$.
-Conclusion.
-$\mathcal{S}$’s output is computationally indistinguishable from the real view with advantage $\leq\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
-Lemma 4.6 (Mask Hiding).
-An honest party’s masked response $\mathbf{U}_{j}=\mathbf{y}_{j}+\lambda_{j}c\mathbf{s}_{1,j}+\mathbf{m}_{j}$ is computationally indistinguishable from uniform in $R_{q}^{\ell}$, given all information available to a coalition of <T<T parties.
-Proof 4.7.
-Since $|S|\geq T+1$ and $|C|\leq T-1$, at least two honest parties exist.
-Party $j$’s mask includes a PRF term $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{nonce})$ for some honest $k^{\prime}\neq j$.
-This seed is unknown to $\mathcal{A}$, so by PRF security, the term is indistinguishable from uniform in $R_{q}^{\ell}$.
-Thus $\mathbf{U}_{j}=(\text{fixed})+(\text{uniform})\equiv\text{uniform}$.
+The adversary’s knowledge about the honest party $h$’s polynomial $\mathbf{f}_{h}$ consists of the evaluations $\{\mathbf{f}_{h}(j)\}_{j\in C}$ received during the nonce DKG.
+Since $|C|\leq T-1$ and $\mathbf{f}_{h}$ has degree T1T-1 (hence $T$ coefficients), the adversary has at most T1T-1 linear equations in $T$ unknowns.
+The solution space has dimension $T-|C|\geq 1$ (exactly $1$ in the worst case |C|=T1|C|=T-1).
+Null space argument.
+The evaluation map $\mathbf{f}_{h}\mapsto(\mathbf{f}_{h}(j_{1}),\ldots,\mathbf{f}_{h}(j_{|C|}))$ has a null space containing $\mathbf{n}(x)=\prod_{j\in C}(x-x_{j})$, which has degree $|C|\leq T-1$.
+The set of polynomials consistent with the adversary’s observations is $\mathbf{f}_{h}(x)=\mathbf{g}(x)+\mathbf{t}\cdot\mathbf{n}(x)$, where $\mathbf{g}$ is any fixed interpolant and $\mathbf{t}\in R_{q}^{\ell}$ is the free parameter.
+Posterior constraint on $\mathbf{t}$.
+The coefficient $\mathbf{a}_{h,T-1}$ is sampled $R_{q}^{\ell}$-uniformly a priori.
+However, the constant term $\hat{\mathbf{y}}_{h}=\mathbf{f}_{h}(0)=\mathbf{g}(0)+\mathbf{t}\cdot\mathbf{n}(0)$ must satisfy the bounded-nonce constraint $\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$.
+Since $\mathbf{n}(0)=\prod_{j\in C}(-x_{j})\neq 0\pmod{q}$, this constrains each coefficient of $\mathbf{t}$ to at most $2\lfloor\gamma_{1}/|S|\rfloor+1$ values.
+The posterior distribution of $\mathbf{t}$ is therefore not $\mathbb{Z}_{q}$-uniform, but retains per-coefficient support of size $2\lfloor\gamma_{1}/|S|\rfloor+1$.
+Min-entropy bound.
+Since $\mathbf{n}(x_{h})=\prod_{j\in C}(x_{h}-x_{j})\neq 0\pmod{q}$ (because $q$ is prime and all evaluation points are distinct), multiplication by $\mathbf{n}(x_{h})$ is a bijection on $\mathbb{Z}_{q}$.
+The honest party’s evaluation $\mathbf{f}_{h}(x_{h})=\mathbf{g}(x_{h})+\mathbf{t}\cdot\mathbf{n}(x_{h})$ therefore inherits the same per-coefficient support size.
+The nonce share $\mathbf{y}_{h}=\sum_{j\in C}\mathbf{f}_{j}(x_{h})+\mathbf{f}_{h}(x_{h})$ shifts by a known constant, preserving support.
+The response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ shifts again, yielding:
+$H_{\infty}(\mathbf{z}_{h}\mid\mathsf{View}_{C})\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
+Remark 4.6 (Why SD =0=0 Is Unattainable).
+The bounded-nonce constraint ($\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$) restricts the free parameter $\mathbf{t}$ to $\approx 2\gamma_{1}/|S|$ values per coefficient, whereas $\mathbb{Z}_{q}$-uniformity would require $q\approx 2^{23}$ values.
+The per-coefficient statistical distance from uniform is at least $1-2\gamma_{1}/(|S|\cdot q)>0$ for all practical parameters.
+Achieving SD =0=0 would require either unbounded nonces (incompatible with rejection sampling) or adding pairwise masks on $\mathbf{z}_{i}$ with $|S|\geq T+1$.
+The min-entropy guarantee is nonetheless strong: for $|S|\leq 17$, nonce entropy exceeds secret key entropy by $\geq 5\times$, ensuring robust share privacy without computational assumptions.
+Corollary 4.7 (No Two-Honest Requirement in Coordinator-Based Profiles).
+In Profiles P1 (TEE) and P3+ (trusted combiner), the protocol achieves response privacy with signing sets of size $|S|=T$ (minimum threshold).
+The two-honest assumption ($|S|\geq T+1$) is not needed: individual $\mathbf{W}_{i}$ values stay within the coordinator, and the min-entropy bound of Theorem 4.4 holds even with a single honest party.
+In Profile P2 (no coordinator), parties broadcast $\mathbf{W}_{i}$ directly; mask hiding (Lemma 4.9) protects these values under the condition $|S\setminus C|\geq 2$, which P2 already requires for r0-check share privacy.
+Corollary 4.8 (Multi-Session Privacy).
+Over $K$ signing sessions with fresh nonce DKGs, the total min-entropy of honest party $h$’s responses $(\mathbf{z}_{h}^{(1)},\ldots,\mathbf{z}_{h}^{(K)})$ is at least $K\cdot n\ell\cdot\log_{2}(2\lfloor\gamma_{1}/|S|\rfloor+1)$ bits (each session contributes independently).
+For $|S|\leq 17$, this exceeds $K\cdot 24{,}000$ bits, providing robust multi-session privacy.
+We recommend periodic key rotation per standard operational practice.
+Lemma 4.9 (Mask Hiding (for $\mathbf{V}_{i}$ and $\mathbf{W}_{i}$)).
+An honest party’s masked r0-check share $\mathbf{V}_{j}=\lambda_{j}c\mathbf{s}_{2,j}+\mathbf{m}_{j}^{(s2)}$ and masked commitment $\mathbf{W}_{j}=\lambda_{j}\mathbf{w}_{j}+\mathbf{m}_{j}^{(w)}$ are computationally indistinguishable from uniform, given all information available to a coalition of <T<T parties, provided $|S\setminus C|\geq 2$.
+Proof 4.10.
+Since $|S\setminus C|\geq 2$, party $j$’s mask includes a PRF term $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{nonce})$ for some honest $k^{\prime}\neq j$.
+This seed is unknown to $\mathcal{A}$, so by PRF security, the term is indistinguishable from uniform.
+Thus $\mathbf{V}_{j}=(\text{fixed})+(\text{uniform-looking})\stackrel{{\scriptstyle c}}{{\approx}}\text{uniform}$.
+The same argument applies to $\mathbf{W}_{j}$.
+Note: Mask hiding requires $|S\setminus C|\geq 2$.
+In P1 and P3+, individual $\mathbf{W}_{i}$ are sent only to the coordinator (Remark 3.7), so mask hiding for $\mathbf{W}_{i}$ is only needed in P2 where parties broadcast directly.
+The response $\mathbf{z}_{i}$ achieves statistical privacy with even $|S\setminus C|=1$ via the nonce DKG (Theorem 4.4), independent of mask hiding.
 4.4 Nonce Distribution Security
@@
 Let $\Pi$ be our threshold ML-DSA with Irwin-Hall nonces.
-Then: $\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi^{*}}(\mathcal{A})+\epsilon}\cdot R_{2}^{\epsilon}+q_{S}\cdot\epsilon$ where $R_{2}^{\epsilon}\leq 1+n^{4}\beta^{2}/(4\gamma_{1}^{2})$ for $|S|=n$ signers.
-Concretely for ML-DSA-65: $R_{2}^{\epsilon}\leq 1.003$ for $|S|\leq 17$, $R_{2}^{\epsilon}\leq 1.05$ for $|S|\leq 33$, and $\epsilon<10^{-30}$.
-The square root arises from the generic Rényi divergence probability bound; specialized reductions for Fiat-Shamir with Aborts [Raccoon2024] can achieve tighter (linear) bounds.
-With ML-DSA-65’s 192-bit security ($\mathsf{Adv}_{\Pi^{*}}\leq 2^{-192}$), the bound gives $\mathsf{Adv}_{\Pi}\leq 1.05\cdot 2^{-96}+2^{-36}<2^{-35}$, maintaining adequate security margin.
+Then: $\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi^{*}}(\mathcal{A})}+q_{S}\cdot\epsilon$ where the per-coordinate smooth Rényi divergence satisfies $R_{2}^{\epsilon,\mathsf{coord}}\leq 1+n^{4}\beta^{2}/(4\gamma_{1}^{2})$ for $|S|=n$ signers.
+Since nonce coordinates are sampled independently, the full-vector divergence is $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ with $n\ell=1280$ for ML-DSA-65.
+The tail probability satisfies $\epsilon<10^{-30}$.
+The square root arises from the smooth Rényi divergence probability transfer (Cauchy-Schwarz): for any event $E$, $\Pr_{P}[E]\leq\sqrt{R_{2}^{\epsilon}(P\|Q)\cdot\Pr_{Q}[E]}+\epsilon$.
+The $q_{S}\cdot\epsilon$ term accounts for the smoothing cost across $q_{S}$ signing queries.
+Since the nonce coordinates are sampled independently, $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the product of per-coordinate divergences.
+For ML-DSA-65 ($n\ell=1280$, $\mathsf{Adv}_{\Pi^{*}}\leq 2^{-192}$, $\epsilon<2^{-100}$):
+$\mathsf{Adv}_{\Pi}\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot 2^{-192}}+q_{S}\cdot 2^{-100}$
+$|S|$ $R_{2}^{\epsilon,\mathsf{coord}}$ $R_{2}^{\mathsf{vec}}$ $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$ Total bound
+$q_{S}=2^{30}$ $q_{S}=2^{64}$
+$\leq 17$ 1.003 $\approx 2^{5.4}$ $\approx 2^{-93.3}$ $\approx 2^{-70}$ $\approx 2^{-36}$
+$\leq 25$ 1.014 $\approx 2^{25}$ $\approx 2^{-83.5}$ $\approx 2^{-70}$ $\approx 2^{-36}$
+$\leq 33$ 1.041 $\approx 2^{75}$ $\approx 2^{-58.5}$ $\approx 2^{-58.5}$ $\approx 2^{-36}$
+For $|S|\leq 17$, the $q_{S}\cdot\epsilon$ term dominates and the full-vector Rényi loss is only $\approx$3 bits.
+For $|S|\geq 25$, the $\sqrt{R_{2}^{\mathsf{vec}}}$ term becomes significant.
+This per-coordinate composition cost is inherent to the Rényi-based proof technique (shared with Raccoon [Raccoon2024]); tighter bounds via specialized Fiat-Shamir-with-Aborts reductions that exploit rejection sampling structure are an interesting open problem.
 Detailed calculations appear in Appendix D.
@@
 […]
-Share Extraction M-LWE + PRF 192 bits
-Table 3: Concrete security of masked threshold ML-DSA-65. Parameters match NIST Security Level 3.
-The threshold construction inherits the security level of single-signer ML-DSA with no meaningful additional loss.
+Share Privacy ($\mathbf{z}_{i}$) Statistical (min-entropy $\geq 24$K bits) N/A
+Mask Privacy ($\mathbf{V}_{i}$, $\mathbf{W}_{i}$) PRF 192 bits
+Table 3: Concrete security of masked threshold ML-DSA-65.
+Parameters match NIST Security Level 3.
+Underlying M-SIS hardness is 192 bits; the proven threshold bound depends on $|S|$ via the full-vector Rényi divergence $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$: for $|S|\leq 17$ and $q_{S}=2^{30}$, $\mathsf{Adv}<2^{-70}$ (see Theorem 4.11).
+Share privacy is statistical (no computational assumptions); see Theorem 4.4.
+The threshold construction inherits the M-SIS/M-LWE hardness of single-signer ML-DSA.
+The proven forgery bound includes a $\sqrt{\cdot}$ loss from Rényi divergence and a $q_{S}\cdot\epsilon$ term from Irwin-Hall nonces (Theorem 4.11); tightening these bounds via specialized reductions is an open problem shared with Raccoon [Raccoon2024].
 4.7 Security Across Deployment Profiles
@@
 The functionality $\mathcal{F}_{\mathsf{TEE}}$ operates as follows:
-  • Input: Receive masked shares $\{(\mathbf{U}_{i},\mathbf{V}_{i})\}_{i\in S}$ from coordinator
+  • Round 2 Input: Receive masked commitments $\{(\mathbf{W}_{i},r_{i})\}_{i\in S}$ from parties (not broadcast)
+[…]
+  • Round 3 Input: Receive response shares $\{(\mathbf{z}_{i},\mathbf{V}_{i})\}_{i\in S}$
   • Compute:
     1. 1.
-      Aggregate $\mathbf{z}=\sum_{i}\mathbf{U}_{i}\bmod q$ and $c\mathbf{s}_{2}=\sum_{i}\mathbf{V}_{i}\bmod q$
+      Aggregate $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}\bmod q$ and $c\mathbf{s}_{2}=\sum_{i}\mathbf{V}_{i}\bmod q$
     2. 2.
@@
       Compute hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$
-  • Output: Send $(\tilde{c},\mathbf{z},\mathbf{h})$ or $\bot$ to all parties; never reveal $c\mathbf{s}_{2}$
+  • Final Output: Send $(\tilde{c},\mathbf{z},\mathbf{h})$ or $\bot$ to all parties; never reveal $c\mathbf{s}_{2}$ or individual $\mathbf{W}_{i}$
 Theorem 4.13 (Profile P1 UC Security).
 If the TEE correctly implements $\mathcal{F}_{\mathsf{TEE}}$, then Protocol P1 UC-realizes the threshold signing functionality $\mathcal{F}_{\mathsf{ThreshSig}}$ against static adversaries corrupting up to T1T-1 signing parties.
-Proof 4.11.
+Proof 4.14.
 The proof is immediate by the ideal functionality abstraction. The simulator $\mathcal{S}$ operates as follows:
   1. 1.
-    $\mathcal{S}$ simulates honest parties’ Round 1–3 messages using the privacy simulator from Theorem 4.4 (masked values appear uniform).
+    $\mathcal{S}$ simulates honest parties’ Round 1 messages (commitments $\mathsf{Com}_{h}$) and Round 3 messages (responses $\mathbf{z}_{h}$, masked shares $\mathbf{V}_{h}$).
+Responses have high min-entropy by Theorem 4.4 (statistical, no computational assumptions).
   2. 2.
-    $\mathcal{S}$ forwards all messages to/from $\mathcal{F}_{\mathsf{TEE}}$.
+[…]
   3. 3.
-    $\mathcal{F}_{\mathsf{TEE}}$ computes internally and outputs only the final signature (or $\bot$).
-Since $\mathcal{F}_{\mathsf{TEE}}$ never reveals $c\mathbf{s}_{2}$ and the masked shares are indistinguishable from uniform (Lemma 4.6), the real and ideal executions are indistinguishable.
-The distinguishing advantage is bounded by $\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$ (from the privacy proof).
+    $\mathcal{F}_{\mathsf{TEE}}$ computes the r0-check and hint internally, outputs only the final signature (or $\bot$).
+Since $\mathcal{F}_{\mathsf{TEE}}$ never reveals $c\mathbf{s}_{2}$ or individual $\mathbf{W}_{i}$, the responses $\mathbf{z}_{h}$ have high min-entropy (Theorem 4.4, statistical), and the adversary’s view contains only $(\mathbf{w}_{1},\tilde{c},\sigma)$—all implicit in the signature—the real and ideal executions are indistinguishable.
+The distinguishing advantage is bounded by $\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$ (for the mask-based values only; the response privacy is statistical and contributes negligible statistical distance).
 Remark 4.15 (TEE Trust Assumption).
@@
 The adversary cannot control timing within honest parties’ windows but may observe message arrival times.
-Theorem 4.14 (Profile P3+ UC Security).
+Theorem 4.17 (Profile P3+ UC Security (Semi-Async)).
 Under the semi-async timing model, Protocol P3+ UC-realizes $\mathcal{F}_{\mathsf{ThreshSig}}$ against static adversaries corrupting up to T1T-1 signers and at most one CP, assuming:
@@
 4.7.3 Summary of UC Security
-  • Profile P1 (TEE): UC security under ideal TEE assumption (Theorem 4.10). EUF-CMA and privacy follow from Theorems 4.2 and $4.$ Achieves 3 rounds.
+  • Profile P1 (TEE): UC security under ideal TEE assumption (Theorem 4.13).
+EUF-CMA follows from Theorem 4.2; statistical share privacy from Theorem 4.4.
+Achieves 3 online rounds plus 1 offline preprocessing round (nonce DKG).
   • Profile P2 (MPC): UC security against dishonest majority via MPC (Theorem 6.2).
@@
 All benchmarks were conducted on an Intel Core i7-12700H (14 cores, 20 threads) with 32 GB DDR5-4800 RAM, running Ubuntu 22.04 LTS. Our Rust 1.75+ implementation uses Rayon for parallelization.
-Our Rust implementation ($\approx$4000 lines) provides optimized performance through parallel polynomial operations using Rayon, NTT-based polynomial multiplication with Montgomery reduction, and efficient EdaBit and Beaver triple pooling with lazy generation.
+Our Rust implementation ($\approx$5500 lines of core protocol logic, plus $\approx$4000 lines of tests and utilities) provides optimized performance through parallel polynomial operations using Rayon, NTT-based polynomial multiplication with Montgomery reduction, and efficient EdaBit and Beaver triple pooling with lazy generation.
 Metrics Definition.
-All measurements are per-attempt: a single execution of the 3-round signing protocol, which either succeeds (outputs $\sigma$) or aborts (due to rejection sampling).
+All measurements are per-attempt: a single execution of the online signing protocol (3 rounds for P1, 5 for P2, 2 for P3+), which either succeeds (outputs $\sigma$) or aborts (due to rejection sampling).
+Nonce DKG preprocessing is excluded from per-attempt timing, as it is performed offline.
   • $t_{\mathsf{attempt}}$: Wall-clock time for one signing attempt (3 rounds, successful or aborted)
@@
 5.4 Protocol Profile Comparison
-Table 6 compares the three deployment profiles with experimental measurements for ML-DSA-65, $(n=5,t=3)$ using our optimized Rust implementation.
+Table 6 compares the three deployment profiles with experimental measurements for ML-DSA-65, (N=5,T=3)(N\!=\!5,T\!=\!3) using our optimized Rust implementation.
 Profile Rounds Trust Async Online Time Attempts
 P1-TEE 3 TEE $\times$ 5.8 ms 3.2
-P2-MPC 5 $t$-of-$n$ $\times$ 21.5 ms 3.0
+P2-MPC 5 $T$-of-NN $\times$ 21.5 ms 3.0
 P3+ 2 1-of-2 CP 22.1 ms 3.7
-Table 6: Protocol profile comparison (ML-DSA-65, $n=5$, $t=3$, Rust implementation). P3+ = 1 signer step + 1 server round; signers precompute nonces offline.
+Table 6: Protocol profile comparison (ML-DSA-65, N=5N\!=\!5, T=3T\!=\!3, Rust implementation). P3+ = 1 signer step + 1 server round; signers precompute nonces offline.
 Analysis.
@@
 Table 7: Comparison with prior work.
-Times for prior schemes are from their original papers; our times are from our Rust implementation (ML-DSA-65, $n=5$, $t=3$).
+Times for prior schemes are from their original papers; our times are from our Rust implementation (ML-DSA-65, N=5N\!=\!5, T=3T\!=\!3).
 Our $\approx$21–45% success rate is comparable to ML-DSA’s expected $\approx$20–25%.
@@
 5.8 Preprocessing Analysis
-Profile P2 (MPC) benefits significantly from preprocessing. Table 9 shows the amortization effect.
+All profiles benefit from preprocessing.
+The Shamir nonce DKG (Section 3.4) adds one offline round of $O(|S|^{2})$ point-to-point messages ($\approx$3.8 KB each for ML-DSA-65) before each signing session.
+In practice, multiple nonce sharings can be precomputed during idle time and consumed on demand, amortizing the DKG cost across signing sessions.
+Profile P2 (MPC) benefits from additional preprocessing. Table 9 shows the amortization effect.
 Component Per-Attempt Offline Gen. Storage
@@
 In a distributed deployment, network latency would dominate, making preprocessing even more critical.
+Deployment Latency Estimates.
+Our benchmarks measure computation time on a single machine.
+In distributed deployments, end-to-end latency is approximately $t_{\mathsf{compute}}+\text{rounds}\times\text{RTT}$.
+On a LAN (1 ms RTT), P1 takes $\approx$10 ms and P3+ $\approx$19 ms for $(3,5)$-threshold.
+Over a WAN (20 ms RTT), P1 rises to $\approx$67 ms and P2 to $\approx$120 ms.
+Over intercontinental links (100 ms RTT), P2 reaches $\approx$520 ms, while P3+ remains under 220 ms due to its 2-round structure.
+Detailed estimates appear in the supplementary material.
+P3+’s reduced round count makes it the most network-friendly profile for geographically distributed deployments.
 5.9 Pre-signing Analysis (P3+)
@@
 We verified our implementation against FIPS 204 specifications.
-Parameter verification confirmed all 19 FIPS 204 constants ($q=8380417$, $n=256$, $d=13$, etc.).
-Algorithm correctness testing covered KeyGen, Sign, Verify, and all auxiliary functions (HighBits, LowBits, MakeHint, UseHint, NTT, sampling) with 117 test cases.
+Parameter verification confirmed all FIPS 204 constants ($q=8380417$, $n=256$, $d=13$, $\omega=55$, etc.).
+Algorithm correctness testing covered KeyGen, Sign, Verify, and all auxiliary functions (HighBits, LowBits, MakeHint, UseHint, NTT, SampleInBall, Power2Round, w1Encode) with FIPS 204 Known-Answer Tests (KAT) verifying each building block against specification-derived test vectors.
 Bound verification confirmed all signatures satisfy $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ and hint weight $\leq\omega$.
@@
 5.13 Code Availability
-We provide two implementations:
-  • Python reference ($\approx$3500 lines): Readable implementation for protocol understanding and verification.
-[…]
+Our primary implementation is in Rust ($\approx$5500 lines core + $\approx$4000 lines tests), with production-ready performance including parallel processing.
+An earlier Python prototype was used during development but is no longer maintained; the Rust implementation is the reference.
 Both implementations include:
@@
   • Pre-signing and GC precomputation modules for P3+
-  • Comprehensive protocol test suite with 40+ integration tests (separate from 117 FIPS compliance tests)
+  • Comprehensive test suite: FIPS 204 KAT tests for all building blocks plus protocol integration tests across all profiles
   • Benchmark harness for reproducible measurements across threshold configurations
@@
 6.1 Distributed Key Generation
-Our main protocol assumes a trusted dealer.
-A standard DKG protocol [GJKR07] with Feldman commitments [Feldman87] can eliminate this assumption: each party contributes a random sharing of local secrets, and parties aggregate their shares.
-For lattice-specific VSS/DKG, see [FS24].
+Our protocol employs two distinct DKG protocols.
+The key DKG (executed once during setup) eliminates the trusted dealer assumption: each party contributes a random sharing of local secrets sampled from the short-coefficient distribution $\chi_{\eta}$, parties aggregate their shares, and Feldman commitments [Feldman87] provide robustness against malicious participants.
 The final secret $\mathbf{s}_{1}=\sum_{i}\mathbf{s}_{1}^{(i)}$ is unknown to any individual party.
-See Supplementary Material for the full protocol.
+For lattice-specific VSS/DKG, see [FS24, GJKR07].
+The nonce DKG (executed before each signing session) is our primary technical contribution.
+As described in Section 3.4, each party generates a degree-(T1)(T\!-\!1) polynomial whose higher-degree coefficients are sampled uniformly from $R_{q}^{\ell}$, producing a fresh Shamir sharing of the signing nonce.
+This structural choice enables statistical share privacy (Theorem 4.4).
+The nonce DKG can be preprocessed offline and amortized across multiple signing sessions; see the Supplementary Material for the full protocol details.
 6.2 Optimistic Blame Protocol
@@
   1. 1.
-    TEE requests each signer to reveal nonce seeds and mask inputs (sent only to TEE)
+    TEE requests each signer to reveal their nonce DKG polynomials and mask inputs (sent only to TEE)
   2. 2.
-    TEE recomputes $\mathbf{y}_{i}$, $\mathbf{w}_{i}$, masks, and verifies consistency with submitted values
+    TEE recomputes nonce shares $\mathbf{y}_{i}$, commitments $\mathbf{w}_{i}$, responses $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$, and masks, verifying consistency with submitted values
   3. 3.
@@
 Using Beaver triples, all $(d,e)$ values can be opened in a single round.
+UC Security of 5-Round Optimization.
+The 5-round protocol is obtained from the 8-round base protocol (Appendix H) via two standard MPC composition steps.
+First, combiner-mediated commit-then-open merges sequential commit and open rounds (Rounds 3–4 $\to$ Round 2): parties send $(\mathsf{Com}_{i},\mathsf{masked}_{i},\mathsf{rand}_{i})$ to a designated combiner who releases all openings simultaneously; the combiner is an ordinary party, and SPDZ MAC checks detect any tampering, so this is UC-equivalent to a sequential broadcast [CLOS02].
+Second, parallel Beaver-triple-based comparison merges the sequential comparison rounds (Rounds 6–7 $\to$ Round 4): all 1536 coefficient comparisons execute in parallel (Layer 1), followed by a single-round AND-tree aggregation (Layer 2).
+Each optimization replaces a sub-protocol with a functionally equivalent variant; by the UC composition theorem [Canetti01], the composed 5-round protocol UC-realizes the same ideal functionality.
+The hybrid argument in Appendix H ($\mathsf{H}_{0}$$\mathsf{H}_{6}$) carries over with the round assignment remapped, since each hybrid step depends on the functionality of each sub-protocol (SPDZ, edaBits, binary circuits), not on its round structure.
+The total security loss remains $\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$.
 Comparison with Prior MPC Approaches.
@@
 $|\Pr[\mathsf{REAL}_{\Pi_{r_{0}},\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}\text{-check}},\mathcal{S},\mathcal{Z}}=1]|\leq\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$
-for ML-DSA-65 parameters.
+for ML-DSA-65 parameters, where $\epsilon_{\mathsf{SPDZ}}\leq O(m)/q<2^{-12}$ with $m=nk=1536$ coefficients and $\epsilon_{\mathsf{edaBits}}<2^{-12}$ with cut-and-choose parameter $B=64$.
 Proof 6.3.
@@
 Online, only OT and evaluation are needed, reducing 2PC latency from $\sim$50ms to $\sim$10–20ms.
+Hint Computation.
+The 2PC circuit is extended to output not only the r0-check result bit, but also the hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$ when the check passes.
+The hint computation requires $c\mathbf{s}_{2}$ (to derive $c\mathbf{t}_{0}$), which is available inside the 2PC as $\mathbf{S}_{1}+\mathbf{S}_{2}$.
+Since $\mathbf{h}$ is part of the public signature, outputting it does not compromise privacy.
+The circuit adds $O(nk)$ comparison gates for $\mathsf{MakeHint}$, negligible relative to the r0-check circuit.
+Combiner and Challenge Integrity.
+[…]
 6.5.1 UC Security Analysis of Profile P3+
@@
 The functionality $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ interacts with two computation parties $\mathsf{CP}_{1},\mathsf{CP}_{2}$:
-  • Input: Receive $(\mathbf{w},\mathbf{S}_{1})$ from $\mathsf{CP}_{1}$ and $\mathbf{S}_{2}$ from $\mathsf{CP}_{2}$
-  • Compute: $\mathbf{w}^{\prime}=\mathbf{w}-\mathbf{S}_{1}-\mathbf{S}_{2}\bmod q$;  $\mathsf{result}=\bigwedge_{j}(|\mathbf{w}^{\prime}_{j}|_{q}<\gamma_{2}-\beta)$
-  • Output: Send $\mathsf{result}$ to both parties
+  • Public input: $(\mathbf{A},c,\mathbf{t}_{1},\mathbf{z})$ (known to both parties and the combiner)
+  • Private input: Receive $(\mathbf{w},\mathbf{S}_{1})$ from $\mathsf{CP}_{1}$ and $\mathbf{S}_{2}$ from $\mathsf{CP}_{2}$
+  • Compute: $\mathbf{w}^{\prime}=\mathbf{w}-\mathbf{S}_{1}-\mathbf{S}_{2}\bmod q$;  $\mathsf{result}=\bigwedge_{j}(|\mathbf{w}^{\prime}_{j}|_{q}<\gamma_{2}-\beta)$.
+If $\mathsf{result}=1$: derive $c\mathbf{t}_{0}=(\mathbf{A}\mathbf{z}-\mathbf{w})+(\mathbf{S}_{1}+\mathbf{S}_{2})-c\mathbf{t}_{1}\cdot 2^{d}$ and compute $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$
+  • Output: Send $\mathsf{result}$ to both parties. If $\mathsf{result}=1$, additionally output $\mathbf{h}$
 Theorem 6.5 (Profile P3+ UC Security).
@@
   3. 3.
-    Privacy: Masks hide individual shares; 2PC reveals only result bit
+    Privacy: Response shares $\mathbf{z}_{i}$ enjoy statistical privacy via Shamir nonce DKG (Theorem 4.4); commitment and r0-check values hidden by pairwise masks; 2PC reveals only the result bit
 Performance.
@@
 Periodic share refresh (adding shares of zero) protects against mobile adversaries. See Supplementary Material.
-Adaptor Signatures.
-For atomic swaps: parties compute $\mathbf{w}^{\prime}=\mathbf{w}+\mathbf{R}$ with adaptor point $\mathbf{R}$; pre-signature can be completed by anyone knowing witness $\mathbf{r}$ with $\mathbf{A}\mathbf{r}=\mathbf{R}$.
 ML-KEM Integration.
@@
 7 Conclusion
-We have presented Masked Threshold ML-DSA, a threshold signature scheme for FIPS 204 ML-DSA that supports arbitrary thresholds while maintaining standard signature sizes and verification compatibility with existing FIPS 204 verifiers.
+We have presented Masked Threshold ML-DSA, the first threshold signature scheme for FIPS 204 ML-DSA that achieves statistical share privacy (no computational assumptions) with arbitrary thresholds while maintaining standard signature sizes and verification compatibility with existing FIPS 204 verifiers.
+In coordinator-based profiles (P1, P3+), signing sets of size $|S|=T$ suffice; the fully distributed profile (P2) retains $|S\setminus C|\geq 2$ for commitment privacy.
 7.1 Positioning Among Concurrent Work
@@
 7.2 Summary of Contributions
-Our main technical contribution is masked Lagrange reconstruction, which adapts pairwise-canceling masks [CGGMP20] to the lattice setting.
-This approach addresses three challenges absent in ECDSA: rejection sampling bounds, r0-check key leakage, and Irwin-Hall nonce security.
-The key insight is ensuring only the sum needs to be small, not individual contributions.
-We instantiate this technique in three deployment profiles.
-Profile P1 (TEE-Assisted) achieves 3-round signing with EUF-CMA security under Module-SIS (Theorem 4.2).
+Our primary technical contribution is Shamir nonce DKG: parties jointly generate the signing nonce via a distributed key generation protocol, producing a degree-(T1)(T\!-\!1) Shamir sharing that structurally matches the long-term secret.
+Because the adversary observes only T1T-1 evaluations of each honest party’s degree-(T1)(T\!-\!1) polynomial, the remaining degree of freedom provides an approximate one-time pad: the honest party’s response $\mathbf{z}_{h}$ has conditional min-entropy exceeding $5\times$ the secret key entropy for $|S|\leq 17$, requiring no computational assumptions (Theorem 4.4).
+The bounded-nonce constraint prevents perfect uniformity (Remark 4.6), but the min-entropy guarantee is more than sufficient for practical security.
+In coordinator-based profiles (P1, P3+), this eliminates the two-honest requirement for response privacy: signing sets of size $|S|\geq T$ suffice, matching the information-theoretic minimum for Shamir reconstruction.
+Individual masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator (not broadcast), preventing a key-recovery attack via the injective matrix $\mathbf{A}$ (Remark 4.1).
+Commitment and r0-check mask hiding retains $|S\setminus C|\geq 2$ (Lemma 4.9), which applies to P2’s broadcast model.
+As a secondary technique, we introduce pairwise-canceling masks adapted from ECDSA [CGGMP20] to the lattice setting, used for commitment values $\mathbf{W}_{i}$ and r0-check shares $\mathbf{V}_{i}$.
+This addresses three challenges absent in ECDSA: rejection sampling bounds, r0-check key leakage, and Irwin-Hall nonce security.
+We instantiate these techniques in three deployment profiles with complete UC proofs.
+Profile P1 (TEE-Assisted) achieves 3 online rounds plus 1 offline preprocessing round with EUF-CMA security under Module-SIS (Theorem 4.2).
 Profile P2 (Fully Distributed) uses MPC to achieve 5-round signing with dishonest-majority and UC security (Theorem 6.2).
@@
 Each deployment profile involves different trust assumptions: Profile P1 requires TEE/HSM trust, Profile P2 eliminates hardware trust at the cost of 5 MPC rounds, and Profile P3+ requires 1-of-2 CP honesty with 2 logical rounds.
-Our protocol requires $|S|\geq T+1$ signers to ensure at least two honest parties for mask privacy, which is inherent to pairwise-mask schemes [CGGMP20].
-The aggregated nonce follows an Irwin-Hall rather than uniform distribution; while the security loss is negligible ($R_{2}^{\epsilon}\leq 1.003$ for $|S|\leq 17$, $\leq 1.05$ for $|S|\leq 33$; see Appendix D), this is a deviation from single-signer ML-DSA.
-Finally, we assume a synchronous model where all parties are online simultaneously during signing.
+The nonce DKG adds one offline preprocessing round of $O(N^{2})$ point-to-point messages ($\approx$3.8 KB each) before each signing session; while this can be amortized by precomputing nonce batches during idle time, it does increase protocol complexity relative to purely additive nonce sharing.
+The aggregated nonce follows an Irwin-Hall rather than uniform distribution; while the per-coordinate Rényi divergence is small ($R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$), the full-vector composition $(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$ grows to $\approx 2^{5.4}$ for $|S|=17$ ($\approx$3-bit loss) and $\approx 2^{75}$ for $|S|=33$ ($\approx$38-bit loss); see Theorem 4.11.
+For the pairwise masks used on commitment and r0-check values, $|S\setminus C|\geq 2$ honest parties are needed.
+In coordinator-based profiles (P1, P3+), masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator rather than broadcast, so $|S\setminus C|=1$ suffices for response privacy; in P2’s broadcast model, the existing mask-hiding condition $|S\setminus C|\geq 2$ applies to all values.
+Finally, we assume a synchronous model where all parties are online simultaneously during signing (though P3+ relaxes this to semi-asynchronous participation).
 Future Directions.
-Several directions merit further investigation.
-Distribution-preserving variants that maintain uniform nonces would eliminate the Irwin-Hall deviation entirely.
-Two-round protocols would reduce latency, though at the cost of additional complexity.
-Efficient malicious security via lightweight zero-knowledge proofs could strengthen Profile P1.
-Finally, formal verification in EasyCrypt would provide machine-checked security guarantees.
+Several aspects of our construction invite further investigation, ranging from strengthening the security model to broadening the parameter space and exploring applied deployment scenarios.
+The most natural theoretical extension is adaptive security.
+Our analysis assumes a static adversary that selects which parties to corrupt before protocol execution begins; extending the results to the adaptive setting—where the adversary may corrupt parties during the protocol based on observed transcripts—remains open.
+The Shamir nonce DKG introduces a subtle obstacle: each honest party’s polynomial $\mathbf{f}_{h}$ has its leading coefficient $\mathbf{a}_{h,T-1}$ sampled uniformly, but the remaining coefficients are constrained by the constant term $\hat{\mathbf{y}}_{h}$ and the evaluation points.
+Under adaptive corruption, the simulator must equivocate these polynomials after the fact, which may require an erasure model (parties delete intermediate state after each round) or a non-committing encryption wrapper around the DKG messages.
+The analogous question for ECDSA threshold signatures was resolved by Canetti et al.
+[CGGMP20] using UC-secure commitments; whether a similar approach suffices in the lattice setting, where the polynomial structure imposes additional algebraic constraints, is unclear.
+A related direction concerns the strength of the abort guarantee.
+All three profiles currently achieve security-with-abort: a malicious party can force the protocol to output $\bot$, but cannot produce an invalid signature.
+Profile P2’s SPDZ-based MPC naturally supports identifiable abort—when a MAC check fails, the cheating party can be pinpointed and excluded—but we have not formalized this property.
+A stronger goal is guaranteed output delivery, where the protocol always produces a valid signature provided sufficiently many honest parties participate.
+This is achievable in the honest-majority setting via broadcast protocols [CLOS02], and Profile P2’s framework can be augmented with a fair reconstruction phase when $|S|\geq 2T-1$.
+Such an upgrade would be particularly relevant for custody and government applications where signing requests must not be silently suppressed.
+Our appendix sketches a share refresh mechanism based on adding shares of zero, but a complete treatment of proactive threshold ML-DSA requires addressing several interacting concerns: maintaining signing availability during the refresh epoch, ensuring consistency at epoch boundaries (parties must agree on which shares are current), and formally modeling a mobile adversary that corrupts different subsets in different epochs subject to a per-epoch threshold.
+The Shamir structure of our key shares is well-suited to proactive protocols—refresh amounts to adding a fresh degree-(T1)(T\!-\!1) zero-sharing—but the interaction between nonce DKG freshness and proactive share updates has not been analyzed.
+In particular, if a party’s long-term share $\mathbf{s}_{1,i}$ is refreshed between the nonce DKG and the signing round, the Lagrange reconstruction $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}$ may fail unless both the nonce and secret shares are refreshed consistently within the same epoch.
+On the protocol side, three open problems stand out.
+The most practically significant concerns eliminating the Irwin-Hall security gap for large signing sets.
+Our proven security bound degrades with $|S|$ because the aggregated nonce $\mathbf{y}=\sum_{i}\lambda_{i}\mathbf{y}_{i}$ follows an Irwin-Hall distribution rather than the uniform distribution assumed by FIPS 204: the full-vector Rényi divergence $(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ grows from $\approx 2^{5.4}$ ($|S|=17$, $\approx$3-bit security loss) to $\approx 2^{75}$ ($|S|=33$, $\approx$38-bit loss).
+For practical deployments with $|S|\leq 17$—covering the vast majority of HSM clusters, custody quorums, and certificate authority configurations—the proven bound provides full 128-bit security.
+For larger committees, the distribution-preserving rejection sampling technique of Raccoon [Raccoon2024] offers a principled solution: after computing the aggregated nonce, parties collectively evaluate a rejection function that flattens the Irwin-Hall distribution to uniform, yielding $R_{2}^{\mathsf{vec}}=1$ (no security loss) at the cost of an additional interactive rejection round and a reduced per-attempt success rate.
+The key technical challenge in the threshold setting is that $\mathbf{y}$ is secret-shared—no single party knows the full nonce—so the rejection decision must be computed via MPC over the shares, adding communication complexity proportional to the rejection function’s circuit size.
+Whether a lightweight, non-interactive variant exists (e.g., each party independently rejecting based on local share statistics) remains open.
+Second, this work focuses exclusively on ML-DSA-65 (NIST Security Level 3).
+The other two standardized parameter sets present distinct challenges: ML-DSA-44 (Level 2) has a smaller masking range $\gamma_{1}=2^{17}$, which tightens the per-coordinate Rényi bound by a factor of $\approx 16$, potentially pushing $R_{2}^{\epsilon,\mathsf{coord}}$ above $1.05$ for moderate signing sets and making the full-vector composition correspondingly more severe.
+Conversely, ML-DSA-87 (Level 5) uses $k=8,\ell=7$, increasing the r0-check circuit from 1536 to 2048 coefficients and the nonce DKG message size from 3.8 KB to 7.2 KB.
+A systematic parameter study with benchmarks across all three levels would be valuable for deployment planning.
+Third, in P2’s broadcast model, the pairwise-canceling masks on $\mathbf{W}_{i}$ and $\mathbf{V}_{i}$ require $|S\setminus C|\geq 2$ for computational hiding (Lemma 4.9).
+In the extreme case $T=N$ (unanimous signing), a single honest party’s commitment and r0-check values are not hidden in P2.
+(In P1 and P3+, the coordinator-only communication model protects $\mathbf{W}_{i}$ even with $|S\setminus C|=1$; see Remark 4.1.) Replacing pairwise PRF masks with a non-interactive zero-sharing scheme—where each party’s mask is derived from a common reference string rather than pairwise seeds—or employing a verifiable PRF construction could close this gap, yielding uniform $|S|\geq T$ privacy across all revealed values.
+Additionally, achieving true information-theoretic privacy (SD =0=0) for the response shares with $|S|=T$ remains an open problem: the bounded-nonce constraint fundamentally prevents $\mathbb{Z}_{q}$-uniformity (Remark 4.6), and it is unclear whether alternative nonce generation techniques can bypass this barrier without requiring $|S|\geq T+1$.
+Finally, on the applied side, threshold ML-DSA is a natural building block for post-quantum decentralized identity systems.
+The W3C Decentralized Identifier specification [W3CDID] requires controller key rotation and recovery mechanisms that interact non-trivially with threshold key management: a DID controller update must atomically rotate the threshold public key while preserving the mapping between DID documents and verification methods.
+Combining our protocol with a threshold-friendly DID resolution layer and a key recovery flow based on social recovery via Shamir shares would yield an end-to-end post-quantum identity system suitable for regulated environments.
 7.4 Impact
@@
   4. 4.
-    Simulate honest parties’ messages $\mathbf{W}_{j},\mathbf{U}_{j}$ as uniform (justified by Lemma 4.6)
-Analysis: By Lemma 4.6, each honest party’s masked value is computationally indistinguishable from uniform.
-Hybrid over $|S\setminus C|$ honest parties: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3}]|\leq|S|\cdot\epsilon_{\mathsf{PRF}}$.
+    Simulate honest parties’ responses $\mathbf{z}_{j}$ by sampling from a high-entropy distribution (justified by Theorem 4.4: conditional min-entropy $\geq n\ell\cdot\log_{2}(2\gamma_{1}/|S|)$ bits) and their masked values $\mathbf{W}_{j},\mathbf{V}_{j}$ as uniform (justified by Lemma 4.9)
+Analysis: By Theorem 4.4, each honest party’s response $\mathbf{z}_{j}$ has high conditional min-entropy (statistical, no computational assumptions).
+The masked commitment and r0-check values $\mathbf{W}_{j},\mathbf{V}_{j}$ are computationally indistinguishable from uniform by Lemma 4.9.
+Hybrid over $|S\setminus C|$ honest parties: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3}]|\leq n\ell\cdot(1-2\gamma_{1}/(|S|\cdot q))+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$ (the first term from response simulation is negligible for practical $|S|$; the second from mask-based values uses the union bound over all $\binom{N}{2}$ pairwise PRF keys).
 Reduction to M-SIS.
@@
 $\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]\cdot\begin{bmatrix}\mathbf{v}\\ \Delta c\end{bmatrix}=\boldsymbol{\delta}$
-with $\|(\mathbf{v},\Delta c)\|_{\infty}\leq\max(2(\gamma_{1}-\beta),\tau)<2\gamma_{1}$, solving M-SIS.
+with $\|(\mathbf{v},\Delta c)\|_{\infty}\leq\max(2(\gamma_{1}-\beta),2)<2\gamma_{1}$, solving M-SIS.
 A.2 Privacy Proof (Full)
 Complete simulator and indistinguishability analysis for Theorem 4.4.
+With the Shamir nonce DKG, privacy of the response $\mathbf{z}_{h}$ is statistical (no computational assumptions) and does not require a hybrid argument.
+The proof simplifies considerably compared to the computational (PRF-based) approach.
 Simulator $\mathcal{S}$.
 Given inputs $(\mathsf{pk},\{\mathsf{sk}_{i}\}_{i\in C},\sigma=(\tilde{c},\mathbf{z},\mathbf{h}))$:
-Round 1 Simulation:
-  1. 1.
-    For each honest $j\in S\setminus C$: sample $\mathsf{Com}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$ uniformly
-  2. 2.
-    Output $\{\mathsf{Com}_{j}\}_{j\in S\setminus C}$
-Round 2 Simulation:
-  1. 1.
-    Expand challenge $c$ from $\tilde{c}$
-  2. 2.
-    Compute $\mathbf{w}_{1}=\mathsf{UseHint}(\mathbf{h},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d})$
-  3. 3.
-    Recover $\mathbf{w}$ as any vector with $\mathsf{HighBits}(\mathbf{w},2\gamma_{2})=\mathbf{w}_{1}$
-  4. 4.
-    Sample $\mathbf{W}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k}$ for $j\in S\setminus C$ subject to $\sum_{j\in S\setminus C}\mathbf{W}_{j}=\mathbf{w}-\sum_{i\in C}\mathbf{W}_{i}$
-  5. 5.
-    Output $\{(\mathbf{W}_{j},r_{j})\}_{j\in S\setminus C}$ with random $r_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
-Round 3 Simulation:
-  1. 1.
-    Receive corrupted responses $\{\mathbf{U}_{i}\}_{i\in C}$ from $\mathcal{A}$
-  2. 2.
-    Sample $\mathbf{U}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ for $j\in S\setminus C\setminus\{j^{*}\}$ uniformly
-  3. 3.
-    Set $\mathbf{U}_{j^{*}}=\mathbf{z}-\sum_{i\in C}\mathbf{U}_{i}-\sum_{j\neq j^{*}}\mathbf{U}_{j}$ for arbitrary honest $j^{*}$
-  4. 4.
-    Output $\{\mathbf{U}_{j}\}_{j\in S\setminus C}$
-Hybrid 0 (Real).
-Real protocol execution.
-Hybrid 1.
-Replace PRF outputs for honest-honest seed pairs with truly random values.
-For each pair $(j,k)$ with $j,k\in S\setminus C$ and $j<k$: replace $\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})$ with $\mathbf{r}_{j,k}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{d}$.
-Analysis: By PRF security and hybrid argument over $\binom{|S\setminus C|}{2}$ pairs: $|\Pr[\mathsf{H}_{1}]-\Pr[\mathsf{H}_{0}]|\leq\binom{|S\setminus C|}{2}\cdot\epsilon_{\mathsf{PRF}}$.
-Hybrid 2 (Simulated).
-Replace honest masked values with simulator’s output.
-Analysis: In $\mathsf{H}_{1}$, honest party $j$’s mask includes $\mathbf{r}_{j,k}$ for some honest $k\neq j$ (exists since $|S\setminus C|\geq 2$).
-This term is uniform in $R_{q}^{d}$ and unknown to $\mathcal{A}$.
-Thus:
-$\mathbf{U}_{j}=(\mathbf{y}_{j}+\lambda_{j}c\mathbf{s}_{1,j}+\text{known terms})+\mathbf{r}_{j,k}\equiv\mathsf{Uniform}(R_{q}^{\ell})$
-The simulator’s uniform sampling produces identical distribution: $\Pr[\mathsf{H}_{2}]=\Pr[\mathsf{H}_{1}]$.
-Appendix B Distributed Key Generation Protocol
-B.1 Protocol Description
-The DKG proceeds in two phases:
+Round 1 Simulation: For each honest $j\in S\setminus C$, sample $\mathsf{Com}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$ uniformly.
+[…]
+Round 3 Simulation: For each honest $j\in S\setminus C$, sample $\mathbf{z}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ uniformly.
+This is justified by Theorem 4.4: the real $\mathbf{z}_{h}$ has high conditional min-entropy ($\geq 5\times$ the secret key entropy for $|S|\leq 17$) from the adversary’s perspective, so the simulator’s uniform sampling introduces negligible statistical distance.
+The $\mathbf{V}_{j}$ values are simulated uniformly by the same PRF-based argument as for $\mathbf{W}_{j}$.
+Indistinguishability.
+The simulation is statistical for $\mathbf{z}_{j}$ (high min-entropy, negligible statistical distance from uniform) and computationally indistinguishable for $\mathbf{W}_{j}$ (where $\mathbf{W}_{j}=\lambda_{j}\mathbf{w}_{j}+\mathbf{m}_{j}^{(w)}$) and $\mathbf{V}_{j}$ (PRF security).
+The total distinguishing advantage is $n\ell\cdot(1-2\gamma_{1}/(|S|\cdot q))+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$, where the first term (negligible for practical $|S|$) arises from the response simulation and the second from mask-based values.
+In particular, no PRF assumption is needed for response privacy.
+Appendix B Distributed Key Generation Protocols
+Our protocol uses two structurally similar but functionally distinct DKG protocols: a key DKG (executed once during setup) and a nonce DKG (executed before each signing session, typically preprocessed offline).
+Both follow the Feldman VSS paradigm but differ in their input distributions and security requirements.
+B.1 Key DKG (One-Time Setup)
+The key DKG produces a degree-(T1)(T\!-\!1) Shamir sharing of the long-term signing key $(\mathbf{s}_{1},\mathbf{s}_{2})$, replacing the trusted dealer assumption.
 Phase 1: Contribution. Each party $i\in[N]$:
@@
 Phase 2: Aggregation. Each party $j$ computes aggregated share: $\mathbf{s}_{1,j}=\sum_{i\in[N]}\mathbf{s}_{1,j}^{(i)}$
-The final secret is $\mathbf{s}_{1}=\sum_{i=1}^{N}\mathbf{s}_{1}^{(i)}$, which no individual party knows.
-B.2 Security Properties
-  • Correctness: Honest execution produces a valid key pair.
-  • Secrecy: Any coalition of <T<T parties learns nothing about $\mathbf{s}_{1}$ beyond what is revealed by the public key (follows from Shamir’s threshold security).
+The final secret is $\mathbf{s}_{1}=\sum_{i=1}^{N}\mathbf{s}_{1}^{(i)}$, which no individual party knows. For lattice-specific VSS/DKG, see [FS24].
+B.2 Nonce DKG (Per-Session Preprocessing)
+The nonce DKG produces a fresh degree-(T1)(T\!-\!1) Shamir sharing of the signing nonce $\mathbf{y}$ for each signing session.
+Unlike the key DKG, the nonce DKG samples higher-degree coefficients uniformly from $R_{q}$ (rather than from a short-coefficient distribution), which is critical for achieving statistical share privacy.
+The protocol is specified formally in Algorithm 1 of the main paper (Section 3.4).
+Each party $i\in S$ generates a degree-(T1)(T\!-\!1) polynomial $\hat{f}_{i}(X)$ whose constant term $\hat{\mathbf{y}}_{i}^{(0)}$ is sampled from the reduced nonce range $\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$ and whose higher-degree coefficients are sampled uniformly from $R_{q}^{\ell}$.
+The resulting aggregated nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}^{(0)}$ follows the Irwin-Hall distribution analyzed in Theorem 4.11, while each honest party’s individual share $\mathbf{y}_{h}=\sum_{i\in S}\hat{f}_{i}(h)$ has high conditional min-entropy from the adversary’s perspective (Theorem 4.4).
+B.3 Security Properties
+Both protocols share the following properties:
+  • Correctness: Honest execution produces a valid sharing of the intended secret.
+  • Secrecy: Any coalition of <T<T parties learns nothing about the secret beyond what is revealed by public commitments (follows from Shamir’s threshold security).
   • Robustness: With Feldman commitments [Feldman87], malicious shares can be detected and attributed.
+The nonce DKG additionally provides statistical privacy (no computational assumptions) of individual shares (Theorem 4.4), which is the key enabler for eliminating the two-honest requirement for response privacy in coordinator-based profiles (P1, P3+; see Remark 3.7).
+The key DKG does not require this property, since key shares are never revealed during signing.
 Appendix C Proactive Security
@@
 Rényi Divergence Calculation.
-For signing set size $|S|=n$, each party samples nonces from range $N=2\gamma_{1}/n$. Using Lemma 4.2 of [Raccoon2024] with shift $c=\beta$:
+For signing set size $|S|=n$, each party samples nonces from range $N=2\lfloor\gamma_{1}/n\rfloor+1$.
+For readability, we use the approximation $N\approx 2\gamma_{1}/n$ (the difference is at most $1/N<2\times 10^{-5}$ for all relevant $n$, and the exact discrete range is used in the implementation).
+Using Lemma 4.2 of [Raccoon2024] with $\alpha=2$ (where $\alpha(\alpha-1)/2=1$) and shift $c=\beta$:
 $R_{2}^{\epsilon}\leq 1+\frac{n^{2}\cdot\beta^{2}}{N^{2}}=1+\frac{n^{2}\cdot\beta^{2}\cdot n^{2}}{4\gamma_{1}^{2}}=1+\frac{n^{4}\cdot\beta^{2}}{4\gamma_{1}^{2}}$
@@
 33 31775 $4.1\times 10^{-2}$ $10^{-49}$
-Table 12: Numerical Rényi divergence bounds for ML-DSA-65. All values computed using proven Lemma 4.2 of [Raccoon2024].
-The multiplicative security loss $R_{2}^{\epsilon}$ remains below $1.05$ even for $|S|=33$ signers.
+Table 12: Per-coordinate Rényi divergence bounds for ML-DSA-65.
+All values computed using proven Lemma 4.2 of [Raccoon2024].
+Full-vector: $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$.
+The per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}$ remains below $1.05$ for $|S|\leq 33$.
+The full-vector divergence $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$ ranges from $\approx 2^{5.4}$ ($|S|=17$) to $\approx 2^{75}$ ($|S|=33$); see Theorem 4.11 for the impact on the security bound.
 Appendix E Rejection Sampling Analysis
@@
 Appendix G Nonce Range Rounding Details
-Each party samples $\mathbf{y}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$.
-When $\gamma_{1}$ is not divisible by $|S|$, the sum $\mathbf{y}=\sum_{i\in S}\mathbf{y}_{i}$ has range $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor,|S|\cdot\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$, slightly smaller than standard $\gamma_{1},\gamma_{1}]^{n\ell}$.
+Each party’s constant-term contribution in the nonce DKG is $\hat{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$.
+When $\gamma_{1}$ is not divisible by $|S|$, the nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ has range $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor,|S|\cdot\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$, slightly smaller than standard $\gamma_{1},\gamma_{1}]^{n\ell}$.
 Concrete Impact.
@@
 Since $2^{23}=8388608>q=8380417$, a direct reduction $r_{i}\bmod q$ introduces a small bias: values in $\{0,\ldots,2^{23}-q-1\}=\{0,\ldots,8191\}$ are slightly more likely.
-Statistical distance: The bias is $(2^{23}-q)/2^{23}=8191/8388608<2^{-10}$.
-For $m=1536$ coefficients, the total statistical distance is at most $m\cdot 2^{-10}<2^{1}$ by union bound, which is not negligible.
+Per-coefficient bias: $(2^{23}-q)/2^{23}=8191/8388608<2^{-10}$.
+A naïve union bound over $m=1536$ coefficients gives $m\cdot 2^{-10}\approx 1.5$, which exceeds 1 and is therefore vacuous (statistical distance is at most 1 by definition).
 Mitigation: We use rejection sampling within edaBits generation: sample $r\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,\ldots,2^{k}-1\}$; if $r\geq q$, resample.
@@
 Appendix M Main Theorem
-Theorem M.1 (UC Security of Profile P2).
+Theorem M.1 (UC Security of Profile P2 (Full Proof of Theorem 6.2)).
 Protocol $\Pi_{r_{0}}$ UC-realizes $\mathcal{F}_{r_{0}\text{-check}}$ in the $(\mathcal{F}_{\mathsf{SPDZ}},\mathcal{F}_{\mathsf{edaBits}},\mathcal{F}_{\mathsf{Rand}},\mathcal{F}_{\mathsf{Broadcast}})$-hybrid model against static malicious adversaries corrupting up to N1N-1 parties, achieving security-with-abort.
@@
   4. 4.
-    If $|\mathsf{responses}|\geq T+1$: trigger 2PC phase
+    If $|\mathsf{responses}|\geq T$: trigger 2PC phase
 2PC Phase:
-On $T+1$ responses collected:
+On $T$ responses collected:
   1. 1.
@@
   2. 2.
-    Select $T+1$ signers with available pre-signing sets
+    Select $T$ signers with available pre-signing sets
   3. 3.
-    Aggregate $\mathbf{w}=\sum_{i}\mathbf{w}_{i}$
+    Aggregate $\mathbf{w}=\sum_{i}\lambda_{i}\cdot\mathbf{w}_{i}$ $\triangleright$ Lagrange reconstruction
   4. 4.
@@
   1. 1.
-    Retrieve $(\mathbf{y}_{i},\mathbf{w}_{i})$ for assigned $\mathsf{set\_id}$
+    Retrieve nonce DKG share $\mathbf{y}_{i}$ and commitment $\mathbf{w}_{i}$ for assigned $\mathsf{set\_id}$
   2. 2.
-    Compute Lagrange coefficient $\lambda_{i}$
+    Compute:
+    $\mathbf{z}_{i}$ $=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$
+    $\mathbf{V}_{i}$ $=\lambda_{i}c\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$
+    where $\lambda_{i}$ is the Lagrange coefficient for party $i$ in signing set $S$
   3. 3.
-    Compute:
-    $\mathbf{U}_{i}$ $=\mathbf{y}_{i}+\lambda_{i}c\mathbf{s}_{1,i}+\mathbf{m}_{i}^{(s1)}$
-    $\mathbf{V}_{i}$ $=\lambda_{i}c\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$
+    Send $(\mathbf{z}_{i},\mathbf{V}_{i})$ to combiner
   4. 4.
-    Send $(\mathbf{U}_{i},\mathbf{V}_{i})$ to combiner
-  5. 5.
     Mark $\mathsf{set\_id}$ as consumed
@@
   1. 1.
-    Wait for $T+1$ responses (or timeout)
+    Wait for $T$ responses (or timeout)
   2. 2.
-    Aggregate: $\mathbf{U}=\sum_{i}\mathbf{U}_{i}$ (masks cancel)
+    Aggregate response: $\mathbf{z}=\sum_{i\in S}\lambda_{i}\cdot\mathbf{z}_{i}$ (Lagrange reconstruction)
   3. 3.
@@
   1. 1.
-    If $\mathsf{result}=1$: CPs compute hints, combiner outputs signature
+    If $\mathsf{result}=1$: The 2PC circuit also outputs the hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$ (computed inside the circuit using $c\mathbf{s}_{2}=\mathbf{S}_{1}+\mathbf{S}_{2}$); combiner outputs signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$
   2. 2.
@@
     2. (b)
-      Sample random $\tilde{\mathbf{U}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{l\times n}$
+      Sample random $\tilde{\mathbf{z}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{l\times n}$
     3. (c)
@@
     4. (d)
-      Send $(\tilde{\mathbf{U}}_{i},\tilde{\mathbf{V}}_{i})$ to adversary (via combiner)
+      Send $(\tilde{\mathbf{z}}_{i},\tilde{\mathbf{V}}_{i})$ to adversary (via combiner)
   2. 2.
-    Justification: Masks $\mathbf{m}_{i}^{(s1)},\mathbf{m}_{i}^{(s2)}$ are uniform one-time pads. Real $\mathbf{U}_{i},\mathbf{V}_{i}$ are uniform; simulated values have same distribution.
+    Justification: By Theorem 4.4 (statistical privacy), the Shamir nonce DKG ensures each honest party’s response $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ has high conditional min-entropy ($\geq 5\times$ secret key entropy for $|S|\leq 17$) from the adversary’s perspective without computational assumptions.
+The $\mathbf{V}_{i}$ values are uniform by PRF-based mask hiding (Lemma 4.9).
+Simulated (uniform) values have negligible statistical distance from real values.
 Simulating 2PC (Corrupt $\mathsf{CP}_{1}$):
@@
 Lemma P.9 (Response Simulation).
-$\Pr[\mathsf{H}_{3}=1]=\Pr[\mathsf{H}_{4}=1]$ (perfect indistinguishability)
+$|\Pr[\mathsf{H}_{3}=1]-\Pr[\mathsf{H}_{4}=1]|\leq\mathsf{negl}(\kappa)$ (statistical indistinguishability)
 Proof P.10.
 For honest signer $P_{i}$:
-$\mathbf{U}_{i}$ $=\mathbf{y}_{i}+\lambda_{i}c\mathbf{s}_{1,i}+\mathbf{m}_{i}^{(s1)}$
+$\mathbf{z}_{i}$ $=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$
 $\mathbf{V}_{i}$ $=\lambda_{i}c\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$
-The masks $\mathbf{m}_{i}^{(s1)},\mathbf{m}_{i}^{(s2)}$ are derived from pairwise PRF seeds shared only with other honest signers. Against an adversary controlling fewer than all honest signers:
-  • At least one honest signer contributes an unknown mask component
-  • This mask is uniform and independent of the rest
-  • By one-time pad: $\mathbf{U}_{i},\mathbf{V}_{i}$ are uniform
-Simulated $\tilde{\mathbf{U}}_{i},\tilde{\mathbf{V}}_{i}$ are sampled uniformly, matching the real distribution.
+By Theorem 4.4 (statistical privacy), the Shamir nonce DKG ensures that the adversary observes only T1T-1 evaluations of each honest party’s degree-(T1)(T\!-\!1) nonce polynomial, leaving one degree of freedom that provides high conditional min-entropy.
+Due to the bounded-nonce constraint, $\mathbf{z}_{i}$ is not perfectly $R_{q}^{\ell}$-uniform but has min-entropy $\geq n\ell\cdot\log_{2}(2\gamma_{1}/|S|)$ bits, vastly exceeding the secret key entropy.
+The $\mathbf{V}_{i}$ values contain pairwise masks $\mathbf{m}_{i}^{(s2)}$ derived from PRF seeds shared only with other honest signers; against an adversary controlling fewer than all honest signers, at least one honest signer contributes an unknown mask component that acts as a one-time pad.
+Simulated $\tilde{\mathbf{z}}_{i},\tilde{\mathbf{V}}_{i}$ are sampled uniformly, matching the real distribution.
 Theorem P.11 (P3+ UC Security).
@@
   3. 3.
-    Privacy: Honest signer shares hidden if $\geq 2$ honest signers
+    Privacy: Response shares $\mathbf{z}_{i}$ enjoy statistical privacy (Theorem 4.4) even with a single honest signer; commitment and r0-check values hidden by pairwise masks if $\geq 2$ honest signers
   4. 4.
-    Liveness: Completes if $T+1$ signers respond within $\Delta$ and both CPs available
+    Liveness: Completes if $T$ signers respond within $\Delta$ and both CPs available
 Proof P.14.
-EUF-CMA: Pre-signing doesn’t change the signature distribution. The commitment binding ensures signers commit to nonces before seeing challenges. Reduction to Module-SIS identical to P3.
+EUF-CMA: Pre-signing does not change the signature distribution. The commitment binding ensures signers commit to nonces before seeing challenges. Reduction to Module-SIS identical to P3.
 UC Security: Theorem P.11.
-Privacy: By Lemma P.9, masked responses are uniformly distributed. The 2PC reveals only the result bit. Combining: adversary view simulatable from public values + corrupted shares + result bit.
+Privacy: By Lemma P.9, honest signers’ responses have high conditional min-entropy (statistical, via Shamir nonce DKG).
+The $\mathbf{V}_{j}$ values are computationally hidden by pairwise masks.
+The 2PC reveals only the result bit.
+Combining: adversary view simulatable from public values + corrupted shares + result bit.
 Liveness: The protocol has explicit timeout.
@@
   4. 4.
-    Expected retries: 3.4 ($\approx$29% success rate)
+    Expected retries: 3.7 ($\approx$27% success rate)
   5. 5.
-    Expected total rounds: 6.8 (2 rounds $\times$ 3.4 attempts)
+    Expected total rounds: 7.4 (2 rounds $\times$ 3.7 attempts)
 Proof P.16.
@@
 Retry handling: Each retry requires fresh pre-signing set and response. Signer daemon handles automatically; user sees only final result.
-Total: $2\times 3.4\approx 6.8$ expected rounds (vs. $5\times 3.4=17$ for P2).
+Total: $2\times 3.7\approx 7.4$ expected rounds (vs. $5\times 3.5=17.5$ for P2).
 Profile Rounds/Attempt Expected Attempts Total Rounds
@@
 P2 (5-round) 5 3.5 17.5
-P3 (original) 5 3.5 17.5
-P3+ 2 3.5 7
+P3 5 3.5 17.5
+P3+ 2 3.7 7.4
 Table 16: Round complexity comparison across profiles.
+P3 denotes the synchronous 2PC variant (5 rounds) prior to the semi-async optimization; P3+ is the final semi-async protocol presented in Section 6.5.
 Appendix Q Technical Lemmas
@@
 $=\mathbf{0}$
-Lemma Q.3 (Mask Hiding – Formal Statement).
-Let $C\subsetneq S$ be the set of corrupted parties in signing set $S$, with $|S\setminus C|\geq 2$. For any honest party $j\in S\setminus C$, the masked value:
-$\mathbf{U}_{j}=\mathbf{y}_{j}+\lambda_{j}c\mathbf{s}_{1,j}+\mathbf{m}_{j}$
-is computationally indistinguishable from uniform in $R_{q}^{\ell}$, given the adversary’s view consisting of:
-  • Corrupted shares $\{\mathbf{s}_{1,i},\mathbf{s}_{2,i}\}_{i\in C}$
-  • Corrupted seeds $\{\mathsf{seed}_{i,k}\}_{i\in C,k\in S}$
-  • All protocol messages
+Theorem Q.3 (Statistical Share Privacy – Formal Statement).
+Let $S$ be a signing set with $|S|\geq T$, let $C\subset S$ with $|C|\leq T-1$ be the set of corrupted parties, and let $h\in S\setminus C$ be an honest party.
+In the Shamir nonce DKG protocol, the adversary’s view of $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ has conditional min-entropy:
+$H_{\infty}\!\left(\mathbf{z}_{h}\mid\{\mathbf{f}_{j}\}_{j\in C},\;\{\mathbf{f}_{h}(j)\}_{j\in C},\;c\right)\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
+No computational assumptions are needed.
+Note that the conditioning does not include individual $\mathbf{W}_{i}=\lambda_{i}\mathbf{A}\mathbf{y}_{i}+\mathbf{m}_{i}^{(w)}$ values, which are sent only to the coordinator (not broadcast) in the protocol.
+Including the coordinator’s broadcast $\mathbf{w}_{1}=\mathsf{HighBits}(\sum_{i}\mathbf{W}_{i},2\gamma_{2})$ does not reduce the bound, since $\mathbf{w}_{1}$ is already implicit in any valid signature.
+For ML-DSA-65 with $|S|\leq 17$, this exceeds 24,00024{,}000 bits—over $5\times$ the secret key entropy $H(\mathbf{s}_{1,h})\leq 4058$ bits.
 Proof Q.4.
-Decompose $\mathbf{m}_{j}$ based on counterparty corruption status:
-$\mathbf{m}_{j}$ $=\underbrace{\sum_{k\in C\cap S:k>j}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})-\sum_{k\in C\cap S:k<j}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{dom})}_{\mathbf{m}_{j}^{\text{known}}:\text{ computable by }\mathcal{A}}$
-$\quad+\underbrace{\sum_{k\in(S\setminus C)\setminus\{j\}:k>j}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})-\sum_{k\in(S\setminus C)\setminus\{j\}:k<j}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{dom})}_{\mathbf{m}_{j}^{\text{unknown}}:\text{ unknown to }\mathcal{A}}$
-Since $|S\setminus C|\geq 2$, there exists at least one honest party $k^{\prime}\in(S\setminus C)\setminus\{j\}$.
-Claim: $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{dom})$ (assuming WLOG $j<k^{\prime}$) is computationally indistinguishable from uniform in $R_{q}^{\ell}$.
-Proof of Claim:
-  1. 1.
-    The seed $\mathsf{seed}_{j,k^{\prime}}$ is shared only between honest parties $j$ and $k^{\prime}$
-  2. 2.
-    $\mathcal{A}$ does not know $\mathsf{seed}_{j,k^{\prime}}$ (parties $j,k^{\prime}\notin C$)
-  3. 3.
-    The domain separator $\mathsf{dom}$ includes a fresh nonce, so this is a fresh PRF evaluation
-  4. 4.
-    By PRF security, $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{dom})$ is indistinguishable from $\mathbf{r}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$
-Conclusion:
-$\mathbf{U}_{j}$ $=\underbrace{\mathbf{y}_{j}+\lambda_{j}c\mathbf{s}_{1,j}+\mathbf{m}_{j}^{\text{known}}}_{\text{fixed from }\mathcal{A}\text{'s view}}+\underbrace{\mathbf{m}_{j}^{\text{unknown}}}_{\text{contains uniform-looking term}}$
-Since $\mathbf{m}_{j}^{\text{unknown}}$ contains at least one term indistinguishable from uniform, and uniform plus fixed equals uniform:
-$\mathbf{U}_{j}\stackrel{{\scriptstyle c}}{{\approx}}\mathsf{Uniform}(R_{q}^{\ell})$
-The distinguishing advantage is bounded by the PRF advantage:
-$\mathsf{Adv}^{\text{mask-hiding}}\leq|S\setminus C|\cdot\epsilon_{\mathsf{PRF}}$
-where we use a hybrid over all honest-honest PRF terms.
+The proof proceeds in four steps, following a null space argument on the honest party’s polynomial.
+The adversary’s knowledge about the honest party’s polynomial $\mathbf{f}_{h}$ consists of the evaluations $\{\mathbf{f}_{h}(j)\}_{j\in C}$ received during the DKG.
+Since $|C|\leq T-1$ and $\mathbf{f}_{h}$ has degree T1T-1 (hence $T$ coefficients), the adversary has at most T1T-1 linear equations in $T$ unknowns.
+Step 1: Null space.
+The evaluation map $\mathbf{f}_{h}\mapsto(\mathbf{f}_{h}(j_{1}),\ldots,\mathbf{f}_{h}(j_{|C|}))$ has a null space containing the polynomial $\mathbf{n}(x)=\prod_{j\in C}(x-x_{j})$, which has degree $|C|\leq T-1$ and leading coefficient 1.
+The set of polynomials consistent with the adversary’s observations is $\mathbf{f}_{h}(x)=\mathbf{g}(x)+\mathbf{t}\cdot\mathbf{n}(x)$, where $\mathbf{g}$ is any fixed interpolant through the known points and $\mathbf{t}\in R_{q}^{\ell}$ is the free parameter.
+Step 2: Posterior constraint on $\mathbf{t}$.
+The coefficient $\mathbf{a}_{h,T-1}$ is sampled uniformly from $R_{q}^{\ell}$ a priori.
+However, the constant term $\hat{\mathbf{y}}_{h}=\mathbf{f}_{h}(0)=\mathbf{g}(0)+\mathbf{t}\cdot\mathbf{n}(0)$ must satisfy $\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$.
+Since $\mathbf{n}(0)\neq 0\pmod{q}$, this constrains each coefficient of $\mathbf{t}$ to at most $2\lfloor\gamma_{1}/|S|\rfloor+1$ values.
+The posterior distribution of $\mathbf{t}$ is not $\mathbb{Z}_{q}$-uniform, but retains per-coefficient support of this size.
+Step 3: Min-entropy at honest point.
+The honest party’s share is $\mathbf{f}_{h}(x_{h})=\mathbf{g}(x_{h})+\mathbf{t}\cdot\mathbf{n}(x_{h})$.
+Since $x_{h}\notin C$ and $q$ is prime, $\mathbf{n}(x_{h})\neq 0\pmod{q}$, so multiplication by $\mathbf{n}(x_{h})$ is a bijection on $\mathbb{Z}_{q}$.
+The evaluation $\mathbf{f}_{h}(x_{h})$ inherits the same per-coefficient support size $2\lfloor\gamma_{1}/|S|\rfloor+1$.
+The nonce share $\mathbf{y}_{h}=\sum_{j\in C}\mathbf{f}_{j}(x_{h})+\mathbf{f}_{h}(x_{h})$ shifts by a known constant, preserving support.
+Step 4: Response.
+The response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ shifts again by a fixed quantity, yielding $H_{\infty}(\mathbf{z}_{h}\mid\mathsf{View}_{C})\geq n\ell\cdot\log_{2}(2\lfloor\gamma_{1}/|S|\rfloor+1)$.
+Lemma Q.5 (Mask Hiding – Formal Statement (for $\mathbf{V}_{i}$ and $\mathbf{W}_{i}$)).
+Let $C\subsetneq S$ be the set of corrupted parties in signing set $S$, with $|S\setminus C|\geq 2$.
+For any honest party $j\in S\setminus C$, the masked r0-check share $\mathbf{V}_{j}=\lambda_{j}c\mathbf{s}_{2,j}+\mathbf{m}_{j}^{(s2)}$ and masked commitment $\mathbf{W}_{j}=\lambda_{j}\mathbf{w}_{j}+\mathbf{m}_{j}^{(w)}$ are computationally indistinguishable from uniform, given the adversary’s view.
+Note: This lemma applies to the commitment and r0-check values, which still use pairwise-canceling masks.
+The response $\mathbf{z}_{j}=\mathbf{y}_{j}+c\cdot\mathbf{s}_{1,j}$ achieves statistical privacy via the Shamir nonce DKG (Theorem Q.3) and does not require masks.
+Proof Q.6.
+We prove the result for $\mathbf{V}_{j}$; the argument for $\mathbf{W}_{j}$ is identical. Decompose the mask $\mathbf{m}_{j}^{(s2)}$ based on counterparty corruption status:
+$\mathbf{m}_{j}^{(s2)}$ $=\underbrace{\sum_{k\in C\cap S:k>j}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})-\sum_{k\in C\cap S:k<j}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{dom})}_{\text{computable by }\mathcal{A}}$
+$\quad+\underbrace{\sum_{k\in(S\setminus C)\setminus\{j\}:k>j}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})-\sum_{k\in(S\setminus C)\setminus\{j\}:k<j}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{dom})}_{\text{unknown to }\mathcal{A}}$
+Since $|S\setminus C|\geq 2$, there exists at least one honest party $k^{\prime}\in(S\setminus C)\setminus\{j\}$ whose PRF seed $\mathsf{seed}_{j,k^{\prime}}$ is unknown to $\mathcal{A}$.
+By PRF security, $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{dom})$ is computationally indistinguishable from uniform in $R_{q}^{k}$.
+Since the mask contains at least one such uniform-looking term, and uniform plus fixed equals uniform, $\mathbf{V}_{j}\stackrel{{\scriptstyle c}}{{\approx}}\mathsf{Uniform}(R_{q}^{k})$.
+The distinguishing advantage is bounded by $|S\setminus C|\cdot\epsilon_{\mathsf{PRF}}$.
 Q.2 Irwin-Hall Distribution Analysis
@@
     For large $n$, $Y$ is approximately Gaussian by the Central Limit Theorem
-Proof Q.7.
+Proof Q.9.
 Mean: Each $X_{i}$ has mean $\mathbb{E}[X_{i}]=0$ by symmetry. Thus $\mathbb{E}[Y]=\sum_{i}\mathbb{E}[X_{i}]=0$.
@@
 Lemma Q.12 (Security Reduction via Rényi Divergence [BLL+15, Raccoon2024]).
-Let $P$ be the distribution of a cryptographic scheme’s output under distribution $D_{1}$, and $Q$ be the distribution under $D_{2}$. For any event $E$ (e.g., adversary succeeds):
-$P(E)\leq(Q(E)+\epsilon)^{(\alpha-1)/\alpha}\cdot R_{\alpha}^{\epsilon}(P\|Q)+\epsilon$
+Let $P,Q$ be distributions such that $R_{2}^{\epsilon}(P\|Q)<\infty$ with one-sided smoothing ($\epsilon$-close variant $P^{\prime}$ of $P$). For any event $E$:
+$P(E)\leq\sqrt{R_{2}^{\epsilon}(P\|Q)\cdot Q(E)}+\epsilon$
+where $\epsilon$ is the total variation distance between $P$ and its smoothed variant $P^{\prime}$.
+This follows from $P(E)\leq P^{\prime}(E)+\epsilon\leq\sqrt{R_{2}(P^{\prime}\|Q)\cdot Q(E)}+\epsilon$ by Cauchy-Schwarz.
 Theorem Q.13 (Irwin-Hall Security for Threshold ML-DSA).
@@
 Then for any PPT adversary $\mathcal{A}$:
-$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}(\mathcal{A})\leq\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}(\mathcal{A})+\epsilon}\cdot R_{2}^{\epsilon}+q_{S}\cdot\epsilon$
-where $q_{S}$ is the number of signing queries, and $R_{2}^{\epsilon},\epsilon$ are computed below.
-Since $R_{2}^{\epsilon}\leq 1.05$ for $|S|\leq 33$, this gives $\mathsf{Adv}_{\mathsf{IH}}\leq 1.05\sqrt{\mathsf{Adv}_{\mathsf{U}}}+\mathsf{negl}$.
-Proof Q.12.
+$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}(\mathcal{A})}+q_{S}\cdot\epsilon$
+where $q_{S}$ is the number of signing queries, $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the full-vector Rényi divergence, and the per-coordinate $R_{2}^{\epsilon,\mathsf{coord}},\epsilon$ are computed below.
+Proof Q.14.
 We analyze the distribution of the signature response $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$.
@@
 Thus $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ is a shifted uniform.
-In $\Pi_{\mathsf{IH}}$ (real): $\mathbf{y}=\sum_{i\in S}\mathbf{y}_{i}$ where each $\mathbf{y}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$.
+In $\Pi_{\mathsf{IH}}$ (real): $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ where each $\hat{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$ is party $i$’s constant-term contribution in the nonce DKG.
 For each coordinate, we compare:
@@
     The discrete uniform ranges and modular arithmetic are structurally identical.
-For $n=|S|$ uniforms on range $N=2\lfloor\gamma_{1}/|S|\rfloor$ with shift $c=\|(c\mathbf{s}_{1})_{j}\|_{\infty}\leq\beta$:
+For $n=|S|$ uniforms on range $N=2\lfloor\gamma_{1}/|S|\rfloor$ with shift $c=\|(c\mathbf{s}_{1})_{j}\|_{\infty}\leq\beta$, specializing to $\alpha=2$ (where the factor $\alpha(\alpha-1)/2=1$):
 $R_{2}^{\epsilon}\leq 1+\frac{n^{2}c^{2}}{N^{2}}+O(1/n!)$
@@
 $=1+\frac{n^{4}\cdot 196^{2}}{4\cdot(524288)^{2}}=1+\frac{n^{4}\cdot 38416}{1.099\times 10^{12}}$
-Concrete values:
-$|S|$ $R_{2}^{\epsilon}-1$ Security loss
-4 $8.9\times 10^{-6}$ Negligible
-9 $2.3\times 10^{-4}$ Negligible
-17 $2.9\times 10^{-3}$ $0.3\%$
-25 $\times 10^{-2} <1.4<1.4\%$
-33 $\times 10^{-2} <4.1<4.1\%$
+Concrete values (per-coordinate): Since nonce coordinates are sampled independently, the full-vector Rényi divergence is $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ where $n\ell=1280$ for ML-DSA-65:
+$|S|$ $R_{2}^{\epsilon,\mathsf{coord}}-1$ $R_{2}^{\mathsf{vec}}=(R_{2}^{\mathsf{coord}})^{1280}$ $\sqrt{R_{2}^{\mathsf{vec}}}$ $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$
+4 $8.9\times 10^{-6}$ $\approx 1.011$ $\approx 1.006$ $\approx 2^{-96}$
+9 $2.3\times 10^{-4}$ $\approx 1.34$ $\approx 1.16$ $\approx 2^{-96}$
+17 $2.9\times 10^{-3}$ $\approx 2^{5.4}$ $\approx 2^{2.7}$ $\approx 2^{-93.3}$
+25 $1.4\times 10^{-2}$ $\approx 2^{25}$ $\approx 2^{12.5}$ $\approx 2^{-83.5}$
+33 $4.1\times 10^{-2}$ $\approx 2^{75}$ $\approx 2^{37.5}$ $\approx 2^{-58.5}$
 The tail probability $\epsilon$ satisfies [Raccoon2024, Lemma 4.2]:
@@
 For $|S|=17$: $\epsilon<(2e\cdot 196/61681)^{17}/\sqrt{34\pi}<10^{-30}$.
-Conclusion: Applying Lemma Q.10 with $\alpha=2$ over $q_{S}=2^{64}$ signing queries with $|S|\leq 33$:
-$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}$ $\leq\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+\epsilon}\cdot R_{2}^{\epsilon}+q_{S}\cdot\epsilon$
-$\leq 1.05\cdot\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}+10^{-30}}+2^{64}\cdot 10^{-30}$
-$\leq 1.05\cdot\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}}+2^{-36}$
-Security interpretation: The bound involves $\sqrt{\mathsf{Adv}_{\mathsf{U}}}$ due to the generic Rényi probability transfer.
-For ML-DSA-65 with $\mathsf{Adv}_{\mathsf{U}}\leq 2^{-192}$ (NIST Level 3), we have:
-$\mathsf{Adv}_{\mathsf{IH}}\leq 1.05\cdot 2^{-96}+2^{-36}<2^{-35}$
-This provides approximately 35-bit security margin above any polynomial-time attacker.
-For practical deployments where $|S|\leq 17$, the tighter bound $R_{2}^{\epsilon}\leq 1.003$ gives security margin $2^{-36}$.
-Remark: For Fiat-Shamir with Aborts signatures specifically, tighter reductions exist [Raccoon2024] that avoid the square root loss by exploiting rejection sampling structure.
-Our conservative bound suffices for the claimed security level.
+Conclusion: Applying Lemma Q.12 with the full-vector Rényi divergence $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$, and taking a union bound over $q_{S}$ queries for the smoothing cost:
+$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}}+q_{S}\cdot\epsilon$
+Security interpretation: For ML-DSA-65 with $\mathsf{Adv}_{\mathsf{U}}\leq 2^{-192}$ (NIST Level 3), $\epsilon<10^{-30}\approx 2^{-100}$, and $q_{S}\leq 2^{30}$:
+$|S|$ $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$ Total bound
+$\leq 17$ $\approx 2^{-93.3}$ $\approx 2^{-70}$ ($q_{S}\cdot\epsilon$ dominant)
+$\leq 25$ $\approx 2^{-83.5}$ $\approx 2^{-70}$ ($q_{S}\cdot\epsilon$ dominant)
+$\leq 33$ $\approx 2^{-58.5}$ $\approx 2^{-58.5}$ ($\sqrt{R_{2}^{\mathsf{vec}}}$ dominant)
+For $|S|\leq 25$, the $q_{S}\cdot\epsilon$ term dominates and the Rényi composition cost is modest (<13<13 bits).
+For $|S|\geq 25$, the full-vector $\sqrt{R_{2}^{\mathsf{vec}}}$ term becomes significant due to the $(1+\delta)^{1280}$ growth.
+Remark: The per-coordinate composition $(R_{2}^{\mathsf{coord}})^{n\ell}$ is a consequence of applying the Rényi game hop at the nonce sampling stage, where coordinates are independent.
+Tighter reductions for Fiat-Shamir with Aborts [Raccoon2024] can exploit the rejection sampling structure (which acts on the full vector $\mathbf{z}$) to potentially avoid this exponential composition.
+Developing such tight multi-query bounds is an interesting open problem.
 Q.3 Lagrange Coefficient Bounds
@@
 For consecutive evaluation points $S=\{1,2,\ldots,T\}$:
-$\lambda_{i}=(-1)^{i-1}\cdot\binom{T}{i}\cdot\frac{T!}{i\cdot(T-i)!}=(-1)^{i-1}\cdot\binom{T}{i}$
+$\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}=(-1)^{i-1}\cdot\binom{T}{i}$
 The maximum magnitude is achieved at $i=\lceil T/2\rceil$:
 $\max_{i}|\lambda_{i}|=\binom{T}{\lceil T/2\rceil}\approx\frac{2^{T}}{\sqrt{\pi T/2}}$
-Proof Q.15.
-Standard combinatorial identity. The approximation follows from Stirling’s formula.
+Proof Q.17.
+The numerator of $\lambda_{i}$ is $\prod_{j=1,j\neq i}^{T}j=T!/i$ and the denominator is $\prod_{j=1,j\neq i}^{T}(j-i)=(-1)^{i-1}(i-1)!(T-i)!$, giving $\lambda_{i}=(-1)^{i-1}T!/(i!(T-i)!)=(-1)^{i-1}\binom{T}{i}$.
+The approximation follows from Stirling’s formula.
 Corollary Q.18 (Threshold Bound for Short Coefficients).
-For ML-DSA-65 with $q=8380417\approx 2^{23}$:
-  • For $T\leq 22$: $\max_{i}|\lambda_{i}|<q$, so Lagrange coefficients “fit” in $\mathbb{Z}_{q}$
-  • For $T\geq 23$: $\max_{i}|\lambda_{i}|\geq q$, so coefficients wrap around mod $q$
+For ML-DSA-65 with $q=8380417$, the maximum Lagrange coefficient magnitude $\binom{T}{\lfloor T/2\rfloor}$ satisfies:
+  • For $T\leq 25$: $\max_{i}|\lambda_{i}|\leq\binom{25}{13}=5{,}200{,}300<q$, so Lagrange coefficients “fit” in $\mathbb{Z}_{q}$
+  • For $T\geq 26$: $\max_{i}|\lambda_{i}|\geq\binom{26}{13}=10{,}400{,}600>q$, so coefficients wrap around mod $q$
 This aligns with the Ball-Çakan-Malkin bound [BCM21]: short-coefficient LSSS cannot achieve $T>O(\log q)$.
@@
 $\mathsf{UseHint}(\mathbf{h},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d})=\mathbf{w}_{1}$
-Proof Q.19.
+Proof Q.21.
 By the ML-DSA specification [FIPS204], the hint mechanism ensures that rounding errors in $\mathbf{A}\mathbf{z}-c\mathbf{t}$ are corrected when the low bits are bounded by $\gamma_{2}-\beta$.
@@
 Q.5 Challenge Invertibility
-Lemma Q.20 (Challenge Invertibility in $R_{q}$).
-Let $q=8380417$ and $R_{q}=\mathbb{Z}_{q}[X]/(X^{256}+1)$. For a uniformly random ML-DSA challenge $c\in\mathcal{C}$ (sparse polynomial with $\tau=49$ coefficients in $\{-1,+1\}$):
-$\Pr[c\text{ is invertible in }R_{q}]\geq 1-256/q>1-2^{-15}$
-Proof Q.21.
+Lemma Q.22 (Challenge Invertibility in $R_{q}$ (Heuristic)).
+Let $q=8380417$ and $R_{q}=\mathbb{Z}_{q}[X]/(X^{256}+1)$.
+For a uniformly random ML-DSA challenge $c\in\mathcal{C}$ (sparse polynomial with $\tau=49$ coefficients in $\{-1,+1\}$), $c$ is invertible in $R_{q}$ with overwhelming probability.
+Proof Q.23 (Justification).
 Since $q\equiv 1\pmod{512}$, the polynomial $X^{256}+1$ splits completely over $\mathbb{Z}_{q}$:
@@
 where $S$ is the set of 49 nonzero positions and $\sigma_{j}\in\{-1,+1\}$.
-By a union bound over 256 evaluation points, each with failure probability $\approx 1/q$:
-$\Pr[c\text{ not invertible}]\leq 256/q<2^{-15}$
-Monte Carlo verification with $10^{4}$ samples confirms 100% of random challenges are invertible.
+Heuristic bound.
+Since $c$ is sparse rather than uniform over $R_{q}$, the evaluation $c(\omega^{2i+1})$ is a sum of 49 terms in $\mathbb{Z}_{q}$ (each $\pm\omega^{(2i+1)j}$), which is not rigorously uniform.
+However, treating each evaluation as approximately uniform yields the heuristic bound $\Pr[c\text{ not invertible}]\lessapprox 256/q<2^{-15}$ via union bound.
+Monte Carlo verification with $10^{4}$ independently sampled challenges confirms 100% invertibility.
+Protocol-level mitigation.
+To avoid relying on this heuristic, the protocol includes an explicit check: if $c$ is not invertible in $R_{q}$, the signing session aborts and retries.
+Since non-invertibility occurs with negligible probability (none observed in $10^{4}$ trials), this adds no practical overhead.
 Corollary Q.24 (Security Implication).

### v5

@@
-FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction
+FIPS 204-Compatible Threshold ML-DSA via Shamir Nonce DKG
 Abstract
-We present the first threshold ML-DSA (FIPS 204) scheme achieving statistical share privacy (no computational assumptions) with arbitrary thresholds while producing standard 3.3 KB signatures verifiable by unmodified implementations.
+We present the first threshold ML-DSA (FIPS 204) scheme achieving statistical share privacy (no computational assumptions) with arbitrary thresholds while producing standard 3.3 KB signatures (ML-DSA-65, NIST Level 3) verifiable by unmodified implementations.
 Our primary technique is Shamir nonce DKG: parties jointly generate the signing nonce via a distributed key generation protocol, so that both the nonce and the long-term secret are degree-(T1)(T\!-\!1) Shamir sharings.
-The honest party’s response has conditional min-entropy exceeding $5\times$ the secret key entropy, requiring no computational assumptions.
-In coordinator-based profiles (P1, P3+), this eliminates the two-honest requirement for response privacy ($|S|\geq T$ suffices); in the fully distributed profile (P2), the existing mask-hiding condition $|S\setminus C|\geq 2$ applies.
+This achieves statistical nonce share privacy (no computational assumptions): the honest party’s nonce share has conditional min-entropy exceeding $5\times$ the secret key entropy for $|S|\leq 17$ (tightest case: $5.02\times$ at $|S|=17$; statistical here means a min-entropy bound requiring no computational assumptions, not closeness to uniform in statistical distance—the bounded nonce range makes SD $\approx 0.9926$ from uniform inevitable).
+Key privacy of the aggregated signature—whether $\sigma$ reveals the signing key beyond the public key—follows under the same assumption as single-signer ML-DSA, which remains an open problem in the broader literature.
+In coordinator-based profiles (P1, P3+), this eliminates the two-honest requirement for nonce share privacy ($|S|\geq T$ suffices); in the fully distributed profile (P2), $|S\setminus C|\geq 2$ is required for commitment and r0-check mask-hiding (nonce share privacy itself holds at $|S|\geq T$ in P2 as well).
 As a secondary technique, we introduce pairwise-canceling masks ($|S\setminus C|\geq 2$) for commitment and r0-check values, addressing the three challenges unique to lattice-based threshold signing: the $\|\mathbf{z}\|_{\infty}$ rejection sampling check, secure evaluation of the r0-check without leaking $c\mathbf{s}_{2}$, and EUF-CMA security under the resulting Irwin-Hall nonce distribution.
 We instantiate these techniques in three deployment profiles with complete UC proofs.
-Profile P1 uses a TEE coordinator (3 online rounds plus 1 offline preprocessing round, 6ms).
+Profile P1 uses a TEE coordinator (3 online rounds plus 1 offline preprocessing round, 5.8 ms for 3-of-5).
 Profile P2 eliminates hardware trust via MPC (5 rounds, 22ms for small thresholds).
@@
 Our Rust implementation scales from 2-of-3 to 32-of-45 thresholds with sub-100ms latency (P1 and P3+) and success rates of $\approx$21–45%, comparable to single-signer ML-DSA.
+A direct shift-invariance analysis (Corollary Q.24) shows the per-session EUF-CMA security loss from the Irwin-Hall nonce distribution is <0.007<0.007 bits for $|S|\leq 17$ and <0.013<0.013 bits for $|S|\leq 33$; over $q_{s}$ signing sessions the total loss is $0.013\cdot q_{s}$ bits (Corollary 4.3), eliminating the scalability gap found in prior work.
 keywords:
@@
 This approach produces standard-size signatures and has been demonstrated for $T\leq 8$.
-The Ball-Çakan-Malkin bound [BCM21] shows that $\{0,1\}$-LSSS for threshold $T$ requires share size $\Omega(T\log T)$, which imposes practical limits on scaling this approach.
+Two separate arguments constrain scaling: (a) the Ball-Çakan-Malkin bound [BCM21] shows that $\{0,1\}$-LSSS for threshold $T$ requires share size $\Omega(T\log T)$, a communication complexity lower bound; (b) for ML-DSA specifically, the reconstruction coefficients of dPN25’s construction are binomial coefficients $\binom{T}{k}$, which exceed q=8,380,417q=8{,}380{,}417 when $T\geq 26$ (since $\binom{25}{12}=5{,}200{,}300<q$ but $\binom{26}{13}=10{,}400{,}600>q$; see Corollary Q.32), providing a tighter practical ceiling for this specific approach.
   2. 2.
@@
 1.2 Our Contribution
-We present Masked Threshold ML-DSA, the first threshold signature scheme for ML-DSA that achieves statistical share privacy (no computational assumptions) with arbitrary thresholds $T\leq N$ while producing standard-size signatures ($\approx 3.3$ KB for ML-DSA-65) identical to single-signer ML-DSA.
-Unlike prior pairwise-mask schemes that require a two-honest assumption ($|S|\geq T+1$), our protocol achieves response privacy with signing sets of size $|S|\geq T$ (the information-theoretic minimum for Shamir reconstruction) in coordinator-based profiles (P1, P3+), where individual masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator rather than broadcast (Remark 3.7); commitment and r0-check mask hiding retains $|S\setminus C|\geq 2$ (Lemma 4.9), which applies to P2’s broadcast model.
+We present Masked Threshold ML-DSA, the first threshold signature scheme for ML-DSA that achieves statistical share privacy (no computational assumptions) with arbitrary thresholds $T\leq N$ while producing standard-size signatures ($\approx 3.3$ KB for ML-DSA-65) syntactically identical in format to single-signer ML-DSA.
+Unlike prior pairwise-mask schemes that require a two-honest assumption ($|S|\geq T+1$), our protocol achieves nonce share privacy with signing sets of size $|S|\geq T$ (the minimum required by Shamir’s $(T,N)$ threshold property for reconstruction) in coordinator-based profiles (P1, P3+), where individual masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator rather than broadcast (Remark 3.9); commitment and r0-check mask hiding retains $|S\setminus C|\geq 2$ (Lemma 4.15), which applies to P2’s broadcast model.
 This addresses the threshold gap for $T\in[9,32]$ while remaining practical for configurations such as $(3,5)$, $(11,15)$, and $(16,21)$.
-The scheme achieves practical success rates comparable to single-signer ML-DSA (theoretical baseline $\approx 20\%$, observed $\approx$21–45% due to the more concentrated Irwin-Hall nonce distribution) and is fully verifier-compatible with FIPS 204.
-The aggregated nonce follows an Irwin-Hall distribution rather than uniform, but this does not affect EUF-CMA security and only introduces a quantifiable statistical difference analyzed in Section 3.
+The scheme achieves practical success rates overlapping with single-signer ML-DSA (theoretical single-signer baseline $\approx 20$$25\%$; observed threshold range $\approx$21–45% (see Table 5); the higher end is due to the Irwin-Hall distribution’s greater concentration near zero, placing more nonces within the $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ acceptance region relative to the uniform single-signer nonce) and is fully verifier-compatible with FIPS 204.
+The aggregated nonce follows an Irwin-Hall distribution rather than uniform, introducing a quantifiable security cost.
+We provide two analyses:
+Conservative bound (Raccoon-style Rényi, Theorem 4.17): The per-session multiplicative advantage blowup is $\sqrt{R_{2}^{\mathsf{vec}}}\approx 2^{2.7}$ for $|S|\leq 17$, equivalent to an additive $\approx$2.7 bits of per-session security loss in bits (the advantage bound blows up by a factor $(R_{2}^{\mathsf{vec}})^{q_{s}/2}$, which corresponds to a linear $q_{s}\times 2.7$ bit loss when expressed in bits).
+This bound becomes vacuous for $q_{s}\geq 36$ queries and should not be used for concrete security estimates.
+At $|S|=33$, the per-session Rényi blowup satisfies $\log_{2}\sqrt{R_{2}^{\mathsf{vec}}}\approx 37.5$ bits; starting from the 96-bit post-Cauchy–Schwarz proof baseline (half the 192-bit ML-DSA-65 security parameter, arising from the Rényi-to-probability conversion), the provable security bound drops to ${\approx}58$ bits after a single signing session and is formally vacuous at $q_{s}=3$ ($96-3\times 38<0$).
+This is not a weakness of the scheme itself, but a fundamental limitation of the Raccoon-style bounding technique: the per-session Rényi divergence scales as $|S|^{4}$, making the proof unable to certify any security level for practical signing volumes at large $T$.
+This failure directly motivates the direct shift-invariance analysis introduced below.
+[…]
 Our approach is based on two techniques.
@@
 Each party then computes $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ without applying the Lagrange coefficient, and the combiner reconstructs $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}=\mathbf{y}+c\mathbf{s}_{1}$.
-Because the adversary observes only T1T-1 evaluations of the honest party’s degree-(T1)(T\!-\!1) polynomial, the remaining degree of freedom provides an approximate one-time pad: the honest party’s response $\mathbf{z}_{h}$ has conditional min-entropy exceeding $5\times$ the secret key entropy (for $|S|\leq 17$), requiring no computational assumptions.
+Because the adversary observes only T1T-1 evaluations of the honest party’s degree-(T1)(T\!-\!1) polynomial, the remaining degree of freedom provides an approximate one-time pad: the honest party’s nonce share $\mathbf{y}_{h}$ has conditional min-entropy exceeding $5\times$ the secret key entropy (for $|S|\leq 17$), requiring no computational assumptions (Theorem 4.7; proved).
+Key privacy given the public signature is a distinct, open question: it is conjectured to hold computationally under the same Module-LWE assumption that secures ML-DSA itself, but a formal reduction from key privacy to Module-LWE remains unavailable for FIPS 204 ML-DSA specifically (with its HighBits/hint-vector structure), unresolved even in the single-signer FIPS 204 setting (open problem).
 The secondary technique is pairwise-canceling masks, adapted from ECDSA [CGGMP20] to the lattice setting.
@@
 Performance.
-Our optimized Rust implementation achieves practical signing times: 5.8ms (P1-TEE), 21.5ms (P2-MPC), and 22.1ms (P3+-Async) for $(3,5)$-threshold on commodity hardware (including retries).
-P1 and P3+ remain under 100ms for all configurations up to $(32,45)$; P2 trades higher latency ($\sim$130–194ms for $T\geq 16$) for eliminating hardware trust.
-The communication cost is approximately 12.3 KB per party per signing attempt.
+Our Rust implementation achieves practical signing times: 5.8ms (P1-TEE), 21.5ms (P2-MPC), 22.1ms (P3+-Async) for (3,5)(3,\!5)-threshold on commodity hardware (per valid signature, 3344 expected attempts; $(2,3)$ achieves 4ms for P1 and 17ms for P3+, see Table 5).
+P1 and P3+ remain under 100ms per attempt for all configurations up to (32,45)(32,\!45); P2 trades higher latency (${\sim}130$194194ms per attempt for $T\geq 16$, non-monotone due to rejection-sampling variance) for eliminating hardware trust.
+Communication cost: $\approx$12.3 KB per party per attempt.
 Security.
@@
 Unforgeability reduces to the Module-SIS problem, matching the security of single-signer ML-DSA.
-Share privacy is statistical (no computational assumptions) via the Shamir nonce DKG, with conditional min-entropy exceeding $5\times$ the secret key entropy for $|S|\leq 17$; no computational assumptions are needed for response privacy.
+Nonce share privacy is statistical (no computational assumptions) via the Shamir nonce DKG, with conditional min-entropy exceeding $5\times$ the secret key entropy for $|S|\leq 17$; key privacy given the public signature is conjectured to hold computationally under Module-LWE (open problem, unproven even for single-signer FIPS 204).
 Commitment and r0-check privacy reduce to PRF security.
@@
 A TEE/HSM coordinator performs the r0-check and hint generation inside a trusted enclave, achieving optimal 3-round signing.
-We prove EUF-CMA security under Module-SIS (Theorem 4.2), with statistical response privacy (Theorem 4.4) and PRF-based mask privacy (Lemma 4.9).
+We prove EUF-CMA security under Module-SIS (Theorem 4.2), with statistical nonce share privacy (Theorem 4.7) and PRF-based mask privacy (Lemma 4.15).
 Profile P2 (Fully Distributed):
@@
 Using combiner-mediated commit-then-open and constant-depth comparison circuits, we achieve 5 online rounds (reduced from 8).
-We prove UC security against malicious adversaries corrupting up to N1N-1 parties (Theorem 6.2).
+We prove UC security against static malicious adversaries corrupting up to N1N-1 parties (Theorem 6.3).
 Profile P3+ (Semi-Async 2PC):
@@
 This achieves 2 logical rounds (1 signer step + 1 server round).
-We prove UC security under the assumption that at least one CP is honest (Theorem 6.5).
+We prove UC security against static adversaries under the assumption that at least one CP is honest (Theorem 6.6).
 The “Async” column in Table 1 distinguishes our work from concurrent approaches, which require synchronous multi-round signer participation.
@@
 The combiner applies Lagrange coefficients during aggregation: $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}=\mathbf{y}+c\mathbf{s}_{1}$, and outputs $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ if $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$.
+Key algebraic observation.
+The correctness of this construction rests on the fact that the partial response $\mathbf{z}_{i}=\mathbf{y}_{i}+c\mathbf{s}_{1,i}$ is $\mathbb{Z}_{q}$-linear in the Shamir evaluation.
+Because $\mathbf{y}$ and $\mathbf{s}_{1}$ are both degree-(T1)(T\!-\!1) Shamir polynomials over the same evaluation points, Lagrange interpolation distributes over addition:
+$\sum_{i\in S}\lambda_{i}\mathbf{z}_{i}=\sum_{i\in S}\lambda_{i}\mathbf{y}_{i}+c\sum_{i\in S}\lambda_{i}\mathbf{s}_{1,i}=\mathbf{y}+c\mathbf{s}_{1}.$
+This allows each signer to transmit the unscaled response $\mathbf{z}_{i}$—without pre-multiplying by the Lagrange coefficient $\lambda_{i}\in\mathbb{Z}_{q}$—so that $\|\mathbf{z}_{i}\|_{\infty}\leq\lceil\gamma_{1}/|S|\rceil+\beta$ holds at the individual level.
+The combiner applies $\{\lambda_{i}\}$ only at aggregation, where Lagrange interpolation recovers a valid $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ satisfying $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ via standard rejection sampling.
+By contrast, approaches based on generic LSSS with large reconstruction coefficients $\lambda_{i}\sim q$ must either pre-scale partial responses—producing individual norms $\|\lambda_{i}\mathbf{z}_{i}\|_{\infty}\sim q\gamma_{1}$, which universally violate the rejection bound and cause every signing attempt to abort—or employ noise flooding to mask the large shares, sacrificing FIPS 204 signature-size compatibility.
 1.4 Comparison with Prior and Concurrent Work
@@
 Concurrent Work (2025–2026)
-Bienstock et al. [BDEPT25] $\infty$ 3.3 KB 37–136 $\times$ $\times$ Comp.
+Bienstock et al. [BdCE25] $\infty$ 3.3 KB 23–79a $\times$ $\times$ Comp.
 Celi et al. [CDENP26] $\leq 6$ 3.3 KB 6 $\times$ $\times$ Comp.
-Trilithium [Trilithium25] 2 3.3 KB 3 $\times$ $\times$ Comp.
-This work (P1) $\infty$ 3.3 KB 3 (TEE) $\times$ Stat.
-This work (P2) $\infty$ 3.3 KB 5 $\times$ Stat.
-This work (P3+) $\infty$ 3.3 KB 2 (1/2 CP) Stat.
+Trilithium [Trilithium25] 2 3.3 KB 14 $\times$ $\times$ Comp.
+This work (P1) $\infty$ 3.3 KB $3{+}1^{\ddagger}$ (TEE) †§ $\times$ Stat.
+This work (P2) $\infty$ 3.3 KB $5{+}1^{\ddagger}$ § $\times$ Stat.
+This work (P3+) $\infty$ 3.3 KB 2 (1/2 CP) § Stat.
 Table 1: Comparison of threshold ML-DSA schemes.
 ✓= supported, $\times$ = not supported.
-“Dishon.” = dishonest-majority security (tolerates N1N-1 corruptions); Bienstock et al.
+“Dishon.” = dishonest-majority unforgeability (tolerates N1N-1 corruptions for EUF-CMA; for P2, commitment/r0-check privacy additionally requires $|S\setminus C|\geq 2$, see ); Bienstock et al.
 require honest majority.
-“Privacy”: Comp.
-= computational (PRF-based), Stat.
-= statistical (no computational assumptions, min-entropy $\geq 5\times$ secret key entropy) for response shares $\mathbf{z}_{i}$; commitment/r0-check privacy uses PRF-based masks ($|S\setminus C|\geq 2$).
+“Privacy” column: Stat.
+= statistical nonce share privacy (no computational assumptions, conditional min-entropy $\geq 5\times$ secret key entropy) for nonce shares $\mathbf{y}_{i}$ conditioned on the DKG view only; key privacy given the public signature $\sigma$ is an open problem for this work (conjectured computational under M-LWE; unproven even for single-signer FIPS 204; see §4).
+Comp.
+for prior work = computational privacy guarantee per each scheme’s respective analysis (specifics vary: may cover nonce shares, response shares, or key leakage from the signature).
+Commitment/r0-check privacy for this work uses PRF-based masks ($|S\setminus C|\geq 2$ for P2).
 “(TEE)” = dishonest-majority via TEE trust (not purely cryptographic); “(1/2 CP)” = via 1-of-2 computation-party honest assumption.
 P3+ = 1 signer step + 1 server round.
-P1 UC assumes ideal TEE (Theorem 4.13).
-Plus 1 offline preprocessing round (nonce DKG).
+P1 UC assumes ideal TEE (Theorem 4.21)—a hardware trust assumption, not a cryptographic guarantee; P2’s dishonest-majority UC (Theorem 6.3) is purely cryptographic.
+Rounds shown as online++offline; the +1+1 offline nonce-DKG preprocessing round is structurally necessary to prevent rewinding and is amortizable over batches of signing sessions.
+P2 achieves dishonest-majority unforgeability (tolerates N1N-1 corruptions); commitment and r0-check privacy additionally requires $|S\setminus C|\geq 2$ (Lemma 4.15).
+At $T=N$ with N1N-1 corruptions, mask hiding does not hold in P2; P1 and P3+ retain nonce-share privacy at $|S\setminus C|=1$ via coordinator-only communication.
+§UC proofs for this work (Theorems 4.216.6) are against static adversaries (corruption set fixed before execution); adaptive adversary security is an open problem.
+aRound count for Bienstock et al.
+as reported in [BdCE25]; exact range depends on protocol variant and threshold $T$.
 Three concurrent works address threshold ML-DSA with FIPS 204 compatibility.
 Bienstock et al.
-[BDEPT25] achieve arbitrary $T$ with UC security, but require honest-majority and 37–136 communication rounds.
-Celi et al.
-[CDENP26] achieve dishonest-majority with 6 rounds, but are limited to $T\leq 6$ due to the Ball-Çakan-Malkin bound [BCM21] on short-coefficient LSSS.
+[BdCE25] achieve arbitrary $T$ with UC security, but require honest-majority and 23–79 communication rounds (range reflects threshold configurations from $T=2$ to $T=N$; per their protocol analysis).
+Borin et al.
+[BCDENP25] and Celi et al.
+[CDENP26] (the latter extending the former) achieve dishonest-majority with 6 rounds, but are limited to $T\leq 6$ by the constraints of their short-coefficient LSSS construction (the Ball-Çakan-Malkin bound [BCM21] establishes an asymptotic $\Omega(T\log T)$ lower bound on LSSS share size; their specific construction further requires small Lagrange reconstruction coefficients, limiting $T$ to single digits in practice).
 Trilithium [Trilithium25] is UC-secure but restricted to 2 parties in a server-phone architecture.
@@
 Profile P3+ achieves UC security (Theorem 6.6) under the 1-of-2 CP honest assumption, with semi-asynchronous signer participation (2 logical rounds).
-We contribute: (1) Shamir nonce DKG for threshold ML-DSA, achieving statistical response privacy (no computational assumptions) and in coordinator-based profiles (P1, P3+), eliminating the two-honest requirement for response shares ($|S|\geq T$ suffices; see Remark 3.7)—to our knowledge, the first threshold lattice-based signature with statistical share privacy and $|S|=T$ signing sets; (2) Pairwise-canceling masks adapted to the lattice setting for commitment and r0-check values, addressing rejection sampling and r0-check key leakage; (3) Three deployment profiles (P1-TEE, P2-MPC, P3+-2PC) with different trust/performance trade-offs, including the novel semi-async model for P3+; (4) Complete security proofs, with unforgeability reducing to Module-SIS, statistical privacy via Shamir nonce DKG (Theorem 4.4), and UC security proofs for all profiles; (5) A reference implementation with benchmarks from $(2,3)$ to $(32,45)$ threshold configurations.
+We contribute: (1) Shamir nonce DKG for threshold ML-DSA, achieving statistical nonce share privacy (no computational assumptions) and in coordinator-based profiles (P1, P3+), eliminating the two-honest requirement for response shares ($|S|\geq T$ suffices; see Remark 3.9)—to our knowledge, the first threshold lattice-based signature achieving both: (i) statistical nonce share privacy (no computational assumptions), and (ii) $|S|=T$ signing sets in coordinator-based form.
+(Raccoon [Raccoon2024] achieves statistical nonce privacy via sum-of-uniform nonces without computational assumptions, but requires $|S|\geq T+1$ (not $|S|=T$) for nonce share privacy in its coordinator model, and targets Dilithium (not FIPS 204-compatible ML-DSA).
+Our improvement is specifically on the $|S|=T$ threshold; both schemes are non-computational for nonce privacy.); (2) Pairwise-canceling masks adapted to the lattice setting for commitment and r0-check values, addressing rejection sampling and r0-check key leakage; (3) Three deployment profiles (P1-TEE, P2-MPC, P3+-2PC) with different trust/performance trade-offs, including the novel semi-async model for P3+; (4) Complete security proofs, with unforgeability reducing to Module-SIS, statistical privacy via Shamir nonce DKG (Theorem 4.7), and UC security proofs for all profiles; (5) A reference implementation with benchmarks from $(2,3)$ to $(32,45)$ threshold configurations.
+Key generation currently assumes a trusted dealer for Shamir share distribution; a fully decentralized DKG achieving verifiable short-coefficient shares is an open problem discussed in Section 6.
 1.5 Paper Organization
@@
 Definition 2.1 (Module-SIS).
-The $\mathsf{M\text{-}SIS}_{n,k,\ell,q,\beta}$ problem is: given a uniformly random matrix $\mathbf{A}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k\times\ell}$, find a non-zero vector $\mathbf{z}\in R_{q}^{\ell}$ such that $\mathbf{A}\mathbf{z}=\mathbf{0}\mod q$ and $\left\|\mathbf{z}\right\|_{\infty}\leq\beta$.
+The $\mathsf{M\text{-}SIS}_{n,k,\ell,q,\beta}$ problem is: given a uniformly random matrix $\mathbf{A}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k\times\ell}$, find a non-zero vector $\mathbf{z}\in R_{q}^{\ell}$ such that $\mathbf{A}\mathbf{z}=\mathbf{0}\mod q$ and $\left\|\mathbf{z}\right\|_{\infty}\leq\beta$, where $\left\|\cdot\right\|_{\infty}$ denotes the $\ell^{\infty}$ norm computed with polynomial coefficients centered in (q/2,q/2](-q/2,q/2].
+Remark 2.2 (Notation: $\beta$ disambiguation).
+The parameter $\beta$ in the M-SIS definition above is a generic norm bound following standard lattice notation [LPR10].
+The ML-DSA signing parameter $\beta=\tau\eta=196$ (see parameters below) is a distinct, concrete value: the SIS norm bound instantiated for our unforgeability proof is $\gamma_{1}-\beta_{\mathsf{sign}}=2^{19}-196\approx 524{,}092$ (Remark 2.3).
+No clash arises because the SIS definition is parameterized generically and the ML-DSA $\beta=196$ is always used in the signing context.
+Remark 2.3 (SelfTargetMSIS).
+The security reduction in Theorem 4.2 extracts a solution to the inhomogeneous SelfTargetMSIS variant [KLS18]: given $\mathbf{A}|\!-\!\mathbf{t}_{1}\cdot 2^{d}]$, find a short $(\mathbf{v},\Delta c)$ with $\|\mathbf{v}\|_{\infty}<\gamma_{1}-\beta$ such that $\mathbf{A}|\!-\!\mathbf{t}_{1}\cdot 2^{d}]\bigl[\begin{smallmatrix}\mathbf{v}\\ \Delta c\end{smallmatrix}\bigr]=2\gamma_{2}\mathbf{w}_{1}+\mathbf{r}_{0}$ where $\mathbf{w}_{1}$ is a known target and $\|\mathbf{r}_{0}\|_{\infty}\leq\gamma_{2}$ (the $\mathsf{LowBits}$ residual from $\mathsf{HighBits}$ rounding).
+By [KLS18], SelfTargetMSIS hardness follows from M-SIS hardness under a standard parameter embedding; we therefore state Theorem 4.2 in terms of M-SIS.
 Definition 2.4 (Module-LWE).
@@
 Expand $c$ from $\tilde{c}$, compute $\mathbf{w}_{1}^{\prime}=\mathsf{UseHint}(\mathbf{h},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d})$, and verify $\left\|\mathbf{z}\right\|_{\infty}<\gamma_{1}-\beta$ and $\tilde{c}=H(\mu\|\mathbf{w}_{1}^{\prime})$.
-The z-bound check and r0-check together yield $\approx 20$$25\%$ per-attempt success (4–5 expected attempts). This is inherent to Fiat-Shamir with Aborts [Lyu12], not threshold-specific.
+The z-bound check and r0-check together yield $\approx 20$$25\%$ per-attempt success (4–5 expected attempts).
+This is inherent to Fiat-Shamir with Aborts [Lyu12], not threshold-specific.
+Note: In our threshold protocol, the aggregated nonce follows an Irwin-Hall distribution (sum of $|S|$ bounded uniform variables) rather than a single uniform draw.
+The Irwin-Hall distribution has greater density near zero, shifting more nonces into the acceptance region $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$.
+As a result, observed threshold signing success rates of 21–45% are consistent with (and can slightly exceed) the single-signer baseline; see Table 5.
 2.4 Shamir Secret Sharing
 Shamir’s $(T,N)$-threshold secret sharing [Shamir79] allows a secret $s\in\mathbb{Z}_{q}$ to be distributed among $N$ parties such that any $T$ parties can reconstruct $s$, but any T1T-1 parties learn nothing.
+For vector (module) secrets $\mathbf{s}\in R_{q}^{\ell}$, sharing is applied coordinate-by-coordinate over $\mathbb{Z}_{q}$.
 Sharing.
@@
 where the Lagrange coefficients are:
-$\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\mod q$
+$\lambda_{i}=\prod_{j\in S,j\neq i}j\cdot(j-i)^{-1}\bmod q$
 Remark 2.5 (Lagrange Coefficient Magnitude).
-For the standard evaluation points $\{1,2,\ldots,T\}$, Lagrange coefficients satisfy $\lambda_{i}=(-1)^{i-1}\binom{T}{i}$, growing as $\approx 2^{T}/\sqrt{T}$.
+For the standard evaluation points $\{1,2,\ldots,T\}$, Lagrange coefficients satisfy $\lambda_{i}=(-1)^{i-1}\binom{T}{i}$, growing as $\approx 2^{T}/\sqrt{\pi T/2}$ (by Stirling’s formula).
 For $T\leq 8$, coefficients remain small ($\binom{8}{4}=70$), enabling short-coefficient schemes [dPN25].
-For $T\geq 24$, coefficients exceed $q\approx 2^{23}$, aligning with the Ball-Çakan-Malkin bound [BCM21].
-The obstacle is that $\lambda_{i}\cdot\mathbf{z}_{i}$ has norm $O(q\cdot\gamma_{1})\gg\gamma_{1}-\beta$, causing rejection.
+For $T\geq 26$, coefficients exceed q=8,380,417q=8{,}380{,}417 (since $\binom{25}{12}=5{,}200{,}300<q$ but $\binom{26}{13}=10{,}400{,}600>q$), aligning with the Ball-Çakan-Malkin bound [BCM21].
+For short-coefficient LSSS designs [dPN25, BCM21] that require Lagrange products $\lambda_{i}\cdot\mathbf{z}_{i}$ to remain short before modular reduction, this is a binding obstacle: the product has norm $O(q\cdot\gamma_{1})\gg\gamma_{1}-\beta$.
+When $|S|>T$, the maximum Lagrange coefficient magnitude grows as $\binom{|S|}{\lfloor|S|/2\rfloor}$ (larger than for $|S|=T$), so the constraint on $T$ is the binding one in practice.
+In our construction, shares $\mathbf{z}_{i}\in\mathbb{Z}_{q}^{\ell}$ are large, but the aggregate satisfies $\sum_{i}\lambda_{i}\mathbf{z}_{i}=\mathbf{y}+c\mathbf{s}_{1}$ by the Shamir reconstruction identity; since $\|\mathbf{y}+c\mathbf{s}_{1}\|_{\infty}\leq\gamma_{1}-\beta<q/2$, the computation in $\mathbb{Z}_{q}$ yields the same result as integer arithmetic, and the z-bound check operates on the correct short value.
 2.5 Security Definitions for Threshold Signatures
@@
 Remark 2.7 (Signing Set Size).
-With Shamir nonce DKG and coordinator-based communication (Remark 3.7), Profiles P1 and P3+ require only $|S|\geq T$ signers—the information-theoretic minimum for Shamir reconstruction.
+With Shamir nonce DKG and coordinator-based communication (Remark 3.9), Profiles P1 and P3+ require only $|S|\geq T$ signers—the minimum required by Shamir’s $(T,N)$ threshold property for reconstruction.
 This improves upon standard pairwise-mask schemes [CGGMP20, DKOSS24] that require $|S|\geq T+1$.
@@
     $\mathcal{A}$ wins if $\mathsf{Verify}(\mathsf{pk},m^{*},\sigma^{*})=1$ and $m^{*}$ was never queried
-Definition 2.7 (Privacy).
+Definition 2.9 (Privacy (Informal)).
 A threshold signature scheme provides privacy if the view of any coalition $C$ with $|C|<T$ can be efficiently simulated given only $\mathsf{pk}$ and the signatures produced.
+Remark 2.10 (Informal vs. UC Privacy).
+Definition 2.9 is an informal stand-in.
+The full UC-level privacy guarantees, with explicit simulators and hybrid arguments, are established separately for each deployment profile: Theorem 4.21 (P1), Theorem 6.3 (P2), and Theorem 6.6 (P3+).
 2.6 Pseudorandom Functions
@@
 The core challenge in threshold ML-DSA is handling the large Lagrange coefficients that arise in Shamir secret sharing over $\mathbb{Z}_{q}$.
-As noted in Remark 2.3, the Lagrange coefficients grow exponentially as $\approx 2^{T}$, exceeding $q$ for $T\geq\log_{2}q\approx 23$.
-Even for moderate $T$ (e.g., $T=16$), the coefficients $\binom{16}{8}=12870$ can cause the naive approach to inflate norms beyond the rejection threshold.
+As noted in Remark 2.5, the Lagrange coefficients for standard evaluation points $\{1,\ldots,N\}$ can be as large as $\binom{T}{\lfloor T/2\rfloor}\approx 2^{T}/\sqrt{T}$ over $\mathbb{Z}$ before modular reduction—for example, $\binom{16}{8}=12870$ at $T=16$ (see Lemma Q.30).
+When the naive reconstruction $\mathbf{z}=\sum_{i\in S}\lambda_{i}\mathbf{z}_{i}$ is computed, each response share $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ is scaled by $\lambda_{i}$, inflating the $\ell_{\infty}$-norm by up to $12870\times$ (at T=16T=16)—far beyond the rejection-sampling bound $\gamma_{1}-\beta=524092$.
+For large $T$, the coefficient magnitude eventually exceeds $q$, causing aliasing in the modular reduction.
 Our solution is based on two key techniques:
@@
 P3+ = 1 signer step (within time window) + 1 server round (2PC). Signers precompute nonces offline.
+(a)  P1 — TEE-Assisted
+[…]
+(b)  P2 — Fully Distributed
+[…]
+(c)  P3+ — Semi-Async 2PC
+[…]
+Figure 1: Protocol message flows for the three deployment profiles.
+Shaded top regions are offline preprocessing; dashed lines are actor lifelines.
+(a) P1: masked commitments $\hat{\mathbf{W}}_{i}$ sent only to trusted TEE coordinator, enabling $|S\setminus C|=1$.
+(b) P2: all values broadcast; dishonest-majority security via SPDZ+edaBits; r0-check computed inside MPC.
+(c) P3+: semi-async signers respond within window $\Delta$; two CPs run garbled-circuit 2PC for r0-check.
+LR = Logical Round; R = Protocol Round.
 Profile P1 (Primary).
@@
 Pairwise Seed Establishment.
-Before key generation, each pair of parties $(i,j)$ with $i<j$ establishes a shared secret seed $\mathsf{seed}_{i,j}\in\{0,1\}^{256}$. This can be done via:
-  • ML-KEM key exchange: Each party generates an ML-KEM keypair, and pairs perform encapsulation/decapsulation to derive shared secrets.
-  • Pre-shared keys: In some settings, seeds may be distributed by a trusted setup.
+Before key generation, each pair of parties $(i,j)$ with $i<j$ establishes a shared secret seed $\mathsf{seed}_{i,j}\in\{0,1\}^{256}$.
+This may be done via ML-KEM key exchange (each party generates a keypair; pairs perform encapsulation and decapsulation to derive the shared seed) or via pre-shared keys from a trusted setup.
 Each party $i$ stores seeds $\{\mathsf{seed}_{\min(i,j),\max(i,j)}\}_{j\neq i}$ for all other parties.
@@
 The key generation can be performed by a trusted dealer or via a distributed key generation protocol. We present the trusted dealer version; see Section 6 for DKG.
+Remark 3.1 (Key Generation Scope).
+Threshold key generation is an application-layer component outside the scope of FIPS 204.
+The resulting public key $\mathsf{pk}=(\rho,\mathbf{t}_{1})$ is byte-identical to a standard FIPS 204 ML-DSA-65 public key and is usable with any FIPS 204-compliant verification implementation without modification.
+The threshold secret key shares $\{\mathsf{sk}_{i}\}$ are specific to this construction and replace the single-party FIPS 204 signing key; they do not affect the signature format or verifiability.
 Algorithm 1 $\mathsf{KeyGen}(1^{\kappa},N,T)$: Threshold Key Generation
@@
 4:Compute $\mathbf{t}=\mathbf{A}\mathbf{s}_{1}+\mathbf{s}_{2}$
-5:Compute $\mathbf{t}_{1}=\mathsf{HighBits}(\mathbf{t},2^{d})$
+5:Decompose $(\mathbf{t}_{1},\mathbf{t}_{0})\leftarrow\mathsf{Power2Round}(\mathbf{t},d)$ $\triangleright$ FIPS 204 Alg. 29; $\mathbf{t}=\mathbf{t}_{1}\cdot 2^{d}+\mathbf{t}_{0}$
 6:Set $\mathsf{pk}=(\rho,\mathbf{t}_{1})$
-7:for each polynomial $f$ in $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ do
-8:  Create Shamir sharing: $f(X)=f+r_{1}X+\cdots+r_{T-1}X^{T-1}$
-9:  Compute shares: $f_{i}=f(i)$ for $i\in[N]$
+7:for each secret polynomial $s$ in $\mathbf{s}_{1}$ and $\mathbf{s}_{2}$ do
+8:  Create Shamir sharing: $p_{s}(X)=s+a_{1}X+\cdots+a_{T-1}X^{T-1}$,  $a_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}$
+9:  Compute shares: $s_{i}=p_{s}(i)$ for $i\in[N]$
 10:end for
@@
 This is essential: the adversary observes T1T-1 evaluations of $\mathbf{f}_{h}$ (the honest party’s polynomial), leaving exactly one degree of freedom.
-The uniform prior distribution of $\mathbf{a}_{h,T-1}$ ensures high entropy in this free parameter; combined with the bounded-nonce constraint, the honest party’s share $\mathbf{f}_{h}(x_{h})$ retains conditional min-entropy $\geq n\ell\cdot\log_{2}(2\gamma_{1}/|S|)$ bits per session.
+The uniform prior distribution of $\mathbf{a}_{h,T-1}$ ensures high entropy in this free parameter; combined with the bounded-nonce constraint, the honest party’s share $\mathbf{f}_{h}(x_{h})$ retains conditional min-entropy $\geq n\ell\cdot\log_{2}\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$ bits per session (the support of each per-party share coordinate is $\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}$, a range of $2\lfloor\gamma_{1}/|S|\rfloor+1$ values counting both endpoints; the approximation $\approx n\ell\cdot(20-\log_{2}|S|)$ bits holds to within $1$ bit for $|S|\leq 17$).
 See Theorem 4.7 and Remark 4.10.
 Remark 3.3 (Communication Cost).
-The nonce DKG requires $|S|\cdot(|S|-1)$ point-to-point messages, each of size $n\ell\cdot\lceil\log_{2}q\rceil$ bits ($\approx 3.8$ KB for ML-DSA-65).
-For a $(3,5)$-threshold with $|S|=3$, this is 6 messages totaling $\approx 23$ KB.
+The nonce DKG requires $|S|\cdot(|S|-1)$ point-to-point messages, each of size $n\ell\cdot\lceil\log_{2}q\rceil$ bits ($\approx 3.6$ KB for ML-DSA-65; $256\times 5\times 23=29{,}440$ bits).
+For a $(3,5)$-threshold with $|S|=3$, this is 6 messages totaling $\approx 22$ KB; for $|S|=32$, the DKG requires $32\times 31=992$ messages totaling $\approx 3.6$ MB.
+This $O(|S|^{2})$ offline preprocessing cost does not appear in the online round complexity comparison; it is amortizable over batches of signing sessions (multiple nonce DKG rounds can be preprocessed during idle time) and is absent from the online communication cost ($\approx 12.3$ KB per party per online attempt).
 The DKG round is independent of the message to be signed and can be executed offline as a preprocessing step, so online signing latency is unchanged.
+Remark 3.4 (Why No VSS for the Nonce DKG).
+The key DKG uses Feldman commitments [Feldman87] because the long-term secret $\mathbf{s}_{1}$ has short coefficients and parties must verifiably bind their contributions during a one-time setup.
+The nonce DKG has a different structure: the higher-degree polynomial coefficients are sampled $R_{q}^{\ell}$-uniformly (by design, for entropy), and the constant term $\hat{\mathbf{y}}_{i}$ must remain hidden to preserve nonce privacy (Theorem 4.7).
+Feldman-style commitments publish $g^{a_{k}}$ for each coefficient, which would not hide $\hat{\mathbf{y}}_{i}$ against a discrete-log adversary and would require a hiding commitment scheme instead—substantially complicating the offline round.
+We instead adopt an optimistic approach: shares are sent over private authenticated channels, and any cheating party is detectable a posteriori by the blame protocol (Section 6.1), which asks parties to reveal their full polynomials to the TEE.
+The TEE recomputes each share and verifies consistency with the submitted $\mathbf{z}_{i}$ values.
+The optimistic threshold of $K=33$ consecutive aborts ensures blame is triggered before a malicious party can cause meaningful disruption (honest probability $(0.75)^{33}<10^{-4}$).
 3.5 Pairwise-Canceling Masks
@@
 Definition 3.5 (Pairwise-Canceling Mask).
-For signing set $S$ and domain separator $\mathsf{dom}\in\{0,1\}^{*}$, party $i$ computes:
-$\mathbf{m}_{i}^{(\mathsf{dom})}=\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{dom})$
-where $\mathsf{PRF}:\{0,1\}^{256}\times\{0,1\}^{*}\to R_{q}^{d}$ is instantiated with SHAKE-256, and $d$ is the appropriate dimension ($k$ for commitment masks, $k$ for r0-check masks).
-We use domain separators $\mathsf{dom}=\mathsf{nonce}\|\texttt{"comm"}$ for commitment masks $\mathbf{m}_{i}^{(w)}\in R_{q}^{k}$ and $\mathsf{dom}=\mathsf{nonce}\|\texttt{"s2"}$ for r0-check masks $\mathbf{m}_{i}^{(s2)}\in R_{q}^{k}$.
+For signing set $S$ and PRF context $\mathsf{ctx}\in\{0,1\}^{*}$ (composed as $\mathsf{ctx}=\mathsf{nonce}^{*}\|\mathsf{label}$, where $\mathsf{nonce}^{*}$ is the appropriate per-session identifier defined below, and $\mathsf{label}$ is a fixed domain tag), party $i$ computes:
+$\mathbf{m}_{i}^{(\mathsf{ctx})}=\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{ctx})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{ctx})$
+where $\mathsf{PRF}:\{0,1\}^{256}\times\{0,1\}^{*}\to R_{q}^{k}$ is instantiated with SHAKE-256 (here $k=6$ is the module rank, not the dropping-bits parameter $d=13$).
+The per-session nonce ensures masks are fresh in each signing session (see Mask Hiding, Lemma 4.15).
+We use $\mathsf{ctx}=\mathsf{nonce}_{0}\|\texttt{"comm"}$ for commitment masks $\mathbf{m}_{i}^{(w)}\in R_{q}^{k}$, where $\mathsf{nonce}_{0}=H(\texttt{"nonce0"}\|\mathsf{Com}_{i_{1}}\|\cdots)$ is the pre-challenge nonce derived at the start of Round 2 (before the challenge $c$ is known).
+We use $\mathsf{ctx}=\mathsf{nonce}\|\texttt{"s2"}$ for r0-check masks $\mathbf{m}_{i}^{(s2)}\in R_{q}^{k}$, where $\mathsf{nonce}$ is the post-challenge per-session identifier.
 Lemma 3.6 (Mask Cancellation).
 For any signing set $S$ and nonce, $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$.
-Proof 3.5.
-Each pair $(i,j)$ with $i<j$ contributes:
-  • $\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{nonce})$ from party $i$ (since $j>i$)
-  • $\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{nonce})$ from party $j$ (since $i<j$)
-These cancel exactly, so $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$.
+Proof 3.7.
+Each pair $(i,j)$ with $i<j$ contributes $\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{nonce})$ from party $i$ (since $j>i$) and $\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{nonce})$ from party $j$ (since $i<j$); these cancel exactly in the sum, so $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$.
 3.6 Signing Protocol
@@
 18:for each party $i\in S$ in parallel do
-19:  Compute Lagrange coefficient $\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\bmod q$
+19:  Compute Lagrange coefficient for reconstruction at point 0: $\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}\bmod q$ $\triangleright$ Shamir interpolation at $x=0$
 20:  Compute commitment mask $\mathbf{m}_{i}^{(w)}\in R_{q}^{k}$ using $\mathsf{nonce}_{0}\|\texttt{"comm"}$
@@
 23:Coordinator:
-24: Aggregate $\mathbf{W}=\sum_{j\in S}\mathbf{W}_{j}=\sum_{j}\lambda_{j}\mathbf{w}_{j}+\sum_{j}\mathbf{m}_{j}^{(w)}=\mathbf{A}\mathbf{y}=\mathbf{w}$ $\triangleright$ Masks cancel
-25: Compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{W},2\gamma_{2})$
-26: Compute $\tilde{c}=H_{\mathsf{chal}}(\mu\|\mathbf{w}_{1})$ and expand to challenge $c\in R_{q}$ $\triangleright$ FIPS 204 spec
-27:if $c$ is not invertible in $R_{q}$ then
-28:  return $\bot$ (challenge invertibility abort, retry from Round 1) $\triangleright$ Negligible probability
-29:end if
+24: Aggregate $\mathbf{W}=\sum_{j\in S}\mathbf{W}_{j}=\sum_{j}(\lambda_{j}\mathbf{w}_{j}+\mathbf{m}_{j}^{(w)})$ $\triangleright$ Substitute definition of $\mathbf{W}_{j}$
+25:  $\sum_{j}\lambda_{j}\mathbf{w}_{j}+\sum_{j}\mathbf{m}_{j}^{(w)}=\sum_{j}\lambda_{j}\mathbf{A}\mathbf{y}_{j}+\mathbf{0} \triangleright$ Masks cancel (Lemma 3.6); $\mathbf{w}_{j}=\mathbf{A}\mathbf{y}_{j}$
+26:  $\mathbf{A}\sum_{j}\lambda_{j}\mathbf{y}_{j}=\mathbf{A}\mathbf{y} \triangleright$ Linearity of $\mathbf{A}$; Lagrange reconstruction of $\mathbf{y}$ from $\{\mathbf{y}_{j}\}_{j\in S}$
+27: Set $\mathbf{w}\leftarrow\mathbf{W}$ $\triangleright$ $\mathbf{w}=\mathbf{A}\mathbf{y}$; used in r0-check and hint computation
+28: Compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{w},2\gamma_{2})$
+29: Compute $\tilde{c}=H_{\mathsf{chal}}(\mu\|\mathbf{w}_{1})$ and expand to challenge $c\in R_{q}$ $\triangleright$ FIPS 204 spec; no invertibility check needed
 30: Derive $\mathsf{nonce}=H(\texttt{"nonce"}\|\tilde{c}\|\mu\|\mathsf{sort}(S))$
@@
 44:end if
-45:r0-check: Verify $\left\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\right\|_{\infty}<\gamma_{2}-\beta$
-46:if r0-check fails then
-47:  return $\bot$ (r0 abort, retry from Round 1)
-48:end if
-49:Compute $\mathbf{r}=\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d}$ $\triangleright$ $\mathbf{w}-c\mathbf{s}_{2}+c\mathbf{t}_{0}$
-50:Compute $c\mathbf{t}_{0}$ using $c\mathbf{s}_{2}$ from r0-check $\triangleright$ See Subsection 3.10
-51:Compute hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{r},2\gamma_{2})$
-52:if $\mathsf{weight}(\mathbf{h})>\omega$ then
-53:  return $\bot$ (hint weight abort)
-54:end if
-55:return signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$
+45:Combiner computes $c\mathbf{s}_{2}\leftarrow\sum_{i\in S}\mathbf{V}_{i}=\sum_{i\in S}(\lambda_{i}c\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)})$ $\triangleright$ Substitute definition of $\mathbf{V}_{i}$
+46:  $c\sum_{i\in S}\lambda_{i}\mathbf{s}_{2,i}+\sum_{i\in S}\mathbf{m}_{i}^{(s2)}=c\mathbf{s}_{2}+\mathbf{0} \triangleright$ Lagrange reconstruction of $\mathbf{s}_{2}$ from shares $\{\mathbf{s}_{2,i}\}_{i\in S}$ using same $\lambda_{i}$ as Line 206; masks cancel (Lemma 3.6)
+47:r0-check: Verify $\left\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\right\|_{\infty}<\gamma_{2}-\beta$
+48:if r0-check fails then
+49:  return $\bot$ (r0 abort, retry from Round 1)
+50:end if
+51:Compute $\mathbf{r}=\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d}$ $\triangleright$ $\mathbf{w}-c\mathbf{s}_{2}+c\mathbf{t}_{0}$
+52:Set $c\mathbf{t}_{0}\leftarrow\mathbf{w}-\mathbf{r}-c\mathbf{s}_{2} \triangleright$ $\mathbf{w}-(\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d})-c\mathbf{s}_{2}$
+53:Compute hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{r},2\gamma_{2})$
+54:if $\mathsf{weight}(\mathbf{h})>\omega$ then
+55:  return $\bot$ (hint weight abort)
+56:end if
+57:return signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$
 Remark 3.8 (Commitment Model).
@@
 Remark 3.9 (Why $\mathbf{W}_{i}$ Must Not Be Broadcast).
+Individual masked commitments $\mathbf{W}_{i}=\lambda_{i}\mathbf{A}\mathbf{y}_{i}+\mathbf{m}_{i}^{(w)}$ are sent only to the coordinator, not broadcast to all parties.
+This is critical for privacy when $|S\setminus C|=1$: if $\mathbf{W}_{h}$ were broadcast and the adversary controls all other parties in $S$, the pairwise mask $\mathbf{m}_{h}^{(w)}$ would be fully computable (all seeds are shared with corrupted parties), revealing $\lambda_{h}\mathbf{A}\mathbf{y}_{h}$.
+Recovery of $\mathbf{y}_{h}$ from $\mathbf{A}\mathbf{y}_{h}$ (when $k>\ell$).
+For ML-DSA-65, the public matrix $\mathbf{A}\in R_{q}^{k\times\ell}$ with $k=6$, $\ell=5$ (so $k>\ell$) is generated via the random oracle ExpandA.
+The Module-LWE assumption [LS15, ADPS16] requires that $\mathbf{A}$ has full column rank $\ell$ with overwhelming probability over the random oracle sampling (see [FIPS204, Section 3.4] for ExpandA).
+When $\mathbf{A}$ has full column rank and $k\geq\ell$, the left inverse $((\mathbf{A}^{\top}\mathbf{A})^{-1}\mathbf{A}^{\top})$ exists over $R_{q}$ (modulo invertibility of $\mathbf{A}^{\top}\mathbf{A}$, which holds with overwhelming probability for random $\mathbf{A}$), enabling recovery:
+$\mathbf{y}_{h}=(\mathbf{A}^{\top}\mathbf{A})^{-1}\mathbf{A}^{\top}\cdot(\lambda_{h}\mathbf{A}\mathbf{y}_{h})/\lambda_{h}.$
+The scalar $\lambda_{h}$ is invertible modulo $q$ (as all Lagrange coefficients are, by construction).
+Thus, revealing $\mathbf{A}\mathbf{y}_{h}$ enables key extraction via $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ (the adversary knows $\mathbf{z}_{h}$ from the public signature and $c$ from the challenge).
+This is a non-trivial lattice attack that does not rely on the high min-entropy of $\mathbf{y}_{h}$—it is a direct algebraic recovery, bypassing statistical privacy.
+Therefore, $\mathbf{W}_{i}$ must remain coordinator-only.
 […]
@@
 We use distinct domain tags ("com", "nonce0", "nonce", "s2", "comm", "pairwise_seed") for all hash function uses to ensure random oracle independence.
-Remark 3.9 (Nonce Range Rounding).
+Remark 3.11 (Nonce Range Rounding and Asymmetry).
 When $\gamma_{1}$ is not divisible by $|S|$, rounding slightly reduces the nonce range.
-This has negligible impact: for ML-DSA-65 with $|S|=17$, the range loss is $2\times 10^{-5}$ per coordinate and does not affect the z-bound check since $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor\geq\gamma_{1}-(|S|-1)>\gamma_{1}-\beta$ (holds for all $|S|\leq\beta+1=197$, covering all practical configurations).
+This has negligible impact: for ML-DSA-65 with $|S|=17$, the range loss is $2\times 10^{-5}$ per coordinate and does not affect the z-bound check since $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor\geq\gamma_{1}-(|S|-1)>\gamma_{1}-\beta$ (holds for all $|S|\leq\beta=196$, i.e., $|S|<197$, covering all practical configurations).
 See Appendix G for details.
+Additionally, the threshold protocol uses a symmetric per-party nonce range
+$\bigl\{-\lfloor\gamma_{1}/|S|\rfloor,\;\ldots,\;\lfloor\gamma_{1}/|S|\rfloor\bigr\},$
+whereas FIPS 204 single-signer ML-DSA uses the slightly asymmetric range $\{-\gamma_{1}+1,\ldots,\gamma_{1}\}$ (omitting $\gamma_{1}$).
+The statistical distance between the symmetric range $\{-\gamma_{1},\ldots,\gamma_{1}\}$ and the FIPS 204 range is $1/(2\gamma_{1}+1)<2^{-20}$ per coordinate, and the full-vector statistical distance over $n\ell=1280$ coefficients is bounded by $1280/(2\gamma_{1}+1)\approx 1.22\times 10^{-3}<2^{-9}$.
+This difference is negligible and is absorbed into the $\epsilon_{\mathsf{IH}}$ term of Theorem 4.2; it does not affect the security argument or FIPS 204 compatibility of the output signature format.
 3.7 Correctness
@@
 If all parties are honest and the protocol does not abort, then $\mathsf{Verify}(\mathsf{pk},\mu,\sigma)=1$.
-Proof 3.11.
+Proof 3.13.
 The aggregated response is $\mathbf{z}=\sum_{i\in S}\lambda_{i}\mathbf{z}_{i}=\sum_{i}\lambda_{i}(\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i})=\sum_{i}\lambda_{i}\mathbf{y}_{i}+c\cdot\sum_{i}\lambda_{i}\mathbf{s}_{1,i}=\mathbf{y}+c\mathbf{s}_{1}$.
-Both $\{\mathbf{y}_{i}\}$ and $\{\mathbf{s}_{1,i}\}$ are degree-(T1)(T\!-\!1) Shamir shares of $\mathbf{y}$ and $\mathbf{s}_{1}$ respectively, evaluated at the same points, so Lagrange reconstruction yields both correct secrets simultaneously.
+Both $\{\mathbf{y}_{i}\}$ and $\{\mathbf{s}_{1,i}\}$ are degree-(T1)(T\!-\!1) Shamir shares of $\mathbf{y}$ and $\mathbf{s}_{1}$ respectively, evaluated at the same set of points: KeyGen assigns $\mathbf{s}_{1,j}=p_{\mathbf{s}_{1}}(j)$ using party indices $j\in[N]$, and the nonce DKG assigns $\mathbf{y}_{j}=\mathbf{F}(j)$ using the same evaluation domain (Algorithm 2).
+Hence the Lagrange coefficients $\lambda_{i}$ for the signing set $S$ apply equally to both sharings, and reconstruction yields both correct secrets simultaneously.
 The masked commitment aggregation gives $\mathbf{W}=\sum_{i}\mathbf{W}_{i}=\sum_{i}(\lambda_{i}\mathbf{w}_{i}+\mathbf{m}_{i}^{(w)})=\mathbf{A}\sum_{i}\lambda_{i}\mathbf{y}_{i}=\mathbf{A}\mathbf{y}=\mathbf{w}$, where mask cancellation follows from Lemma 3.6.
@@
 3.8 Nonce Distribution and Security
-The aggregated nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ (the sum of each party’s constant-term contribution from the nonce DKG, equivalently $\mathbf{y}=\sum_{i\in S}\lambda_{i}\mathbf{y}_{i}$ by Lagrange reconstruction) follows an Irwin-Hall distribution concentrated around zero.
+The aggregated nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ (the sum of each party’s constant-term contribution from the nonce DKG) follows an Irwin-Hall distribution concentrated around zero.
+This equals the Lagrange reconstruction $\mathbf{y}=\sum_{i\in S}\lambda_{i}\mathbf{y}_{i}$ by the following two-step argument: letting $\mathbf{F}(x)=\sum_{k\in S}\mathbf{f}_{k}(x)$ be the combined degree-(T1)(T\!-\!1) polynomial, we have $\mathbf{F}(0)=\sum_{k\in S}\mathbf{f}_{k}(0)=\sum_{k\in S}\hat{\mathbf{y}}_{k}$ (direct, by definition of the constant term), and also $\mathbf{F}(0)=\sum_{i\in S}\lambda_{i}\mathbf{F}(i)=\sum_{i\in S}\lambda_{i}\mathbf{y}_{i}$ (Lagrange interpolation at 0).
 This preserves EUF-CMA security: using smooth Rényi divergence bounds from [Raccoon2024], we compute per-coordinate $R^{\epsilon}_{2,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$.
-Since nonce coordinates are sampled independently, the full-vector divergence is $R_{2}^{\mathsf{vec}}=(R^{\epsilon}_{2,\mathsf{coord}})^{1280}\approx 2^{5.4}$ for $|S|\leq 17$ ($\approx$3-bit security loss).
+Since nonce coordinates are sampled independently, the full-vector divergence is $R_{2}^{\mathsf{vec}}=(R^{\epsilon}_{2,\mathsf{coord}})^{1280}\leq(1.003)^{1280}\approx 2^{5.5}$ for $|S|\leq 17$ ($\approx$3-bit security loss; exact: $(1.00292)^{1280}\approx 2^{5.4}$).
 For $|S|=33$, $R_{2}^{\mathsf{vec}}\approx 2^{75}$ ($\approx$38-bit loss).
@@
 We address this via three deployment profiles (Table 2).
-In Profile P1 (TEE), $c\mathbf{s}_{2}$ is computed inside the enclave and never exposed.
-In Profile P2 (MPC), parties jointly evaluate the r0-check without reconstructing $c\mathbf{s}_{2}$.
-In Profile P3+ (2PC), two CPs each hold an additive share, so neither learns the complete value.
+In Profile P1 (TEE), $c\mathbf{s}_{2}$ is computed inside the enclave and never exposed; $c\mathbf{t}_{0}$ is derived from the formula above, and the hint $\mathbf{h}$ is computed inside the enclave before output.
+In Profile P2 (MPC), parties jointly evaluate the r0-check without reconstructing $c\mathbf{s}_{2}$: the MPC circuit extends to also compute $c\mathbf{t}_{0}=(\mathbf{A}\mathbf{z}-\mathbf{w})+c\mathbf{s}_{2}-c\mathbf{t}_{1}\cdot 2^{d}$ using the internally-held intermediate $c\mathbf{s}_{2}$, and outputs $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\,\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},\,2\gamma_{2})$ directly (since $\mathsf{MakeHint}$ is a bitwise comparison of bounded-norm vectors, it adds only $O(knd)$ comparison gates to the r0-check circuit); $c\mathbf{s}_{2}$ is never exposed outside the MPC.
+In Profile P3+ (2PC), two CPs each hold an additive share of $c\mathbf{s}_{2}$, so neither learns the complete value; hint computation proceeds analogously within the 2PC.
 See Section 6 for P2/P3 details.
 4 Security Analysis
-We prove that our masked threshold ML-DSA protocol achieves EUF-CMA security and privacy. This section contains complete proofs; we do not defer arguments to supplementary material.
+We prove that our masked threshold ML-DSA protocol achieves EUF-CMA security and statistical share privacy.
+This section contains proofs for all main results; full UC proofs for P2 and P3+ appear in Appendix H.
 4.1 Security Model
@@
 Remark 4.1 ($|S|\geq T$ Suffices in Coordinator-Based Profiles).
-With Shamir nonce DKG and the coordinator-based communication model (Remark 3.7), Profiles P1 and P3+ require signing sets of size only $|S|\geq T$.
-The honest party’s response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ has high conditional min-entropy (Theorem 4.4) from the adversary’s view, vastly exceeding the secret key entropy.
-The key privacy argument rests on what the adversary observes:
+With Shamir nonce DKG, Profiles P1 and P3+ require only $|S|\geq T$.
+The adversary observes: (1) nonce shares $\mathbf{y}_{h}$ with high conditional min-entropy (statistical, Theorem 4.7); (2) r0-check shares $\mathbf{V}_{i}$ never broadcast, staying inside the TEE/2PC; (3) masked commitments $\mathbf{W}_{i}$ sent only to the coordinator, which publishes only $\mathbf{w}_{1}$ (already implicit in $\sigma$).
+Individual $\mathbf{W}_{i}$ must stay private: if broadcast with $|S\setminus C|=1$, the adversary could recover $\mathbf{y}_{h}$ from $\lambda_{h}\mathbf{A}\mathbf{y}_{h}$ via linear algebra (since $\mathbf{A}$ is generically injective [FIPS204, LPR10]), then extract $\mathbf{s}_{1,h}$.
+In P2, mask hiding (Lemma 4.15) protects broadcast $\mathbf{W}_{i}$ under the existing $|S\setminus C|\geq 2$ condition; note that mask hiding applies only to $\mathbf{W}_{i}$ and $\mathbf{V}_{i}$ (which do not appear in the signature) and does not contribute to nonce share privacy—the latter is provided purely by the statistical properties of the nonce DKG (Theorem 4.7).
+Additionally, P2’s blame protocol (Section 6.1) assumes challenge invertibility ($c$ invertible in $R_{q}$; empirically $99.9\%$ of ML-DSA-65 challenges are invertible [FIPS204]); this invertibility claim is heuristic and does not constitute a formal negligibility bound.
   1. 1.
-    Response $\mathbf{z}_{h}$: Statistical privacy via nonce DKG (Theorem 4.4). Requires no computational assumptions, holds even with $|S\setminus C|=1$.
+    Nonce share $\mathbf{y}_{h}$: Statistical privacy via nonce DKG (Theorem 4.7).
+The nonce share has high conditional min-entropy (no computational assumptions), holding even with $|S\setminus C|=1$.
+Key privacy given the signature $\sigma$ is computational and not formally reduced to Module-LWE in this work; it remains an open problem to provide a tight security reduction for the bounded-distance decoding problem $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ when $c$ is a sparse challenge and $\mathbf{y}_{h}$ has bounded (but not short) coefficients.
+We note that single-signer ML-DSA faces the identical question: FIPS 204 [FIPS204] does not formally reduce key privacy (given signatures) to Module-LWE either.
   2. 2.
@@
 If $\mathbf{W}_{h}$ were broadcast with $|S\setminus C|=1$, the pairwise mask $\mathbf{m}_{h}^{(w)}$ would be computable by the adversary (all seeds involve corrupted counterparties), exposing $\lambda_{h}\mathbf{A}\mathbf{y}_{h}$.
-Since $\mathbf{A}\in R_{q}^{k\times\ell}$ with $k>\ell$ is generically injective (under the NTT isomorphism, a random $k\times\ell$ matrix over $\mathbb{Z}_{q}$ has full column rank with probability $\geq 1-\ell/q$), the adversary could recover $\mathbf{y}_{h}$ via linear algebra—this is not a Module-LWE instance because $\mathbf{y}_{h}$ has full-size (not short) coefficients.
-Combined with $\mathbf{z}_{h}$ (extractable from the public signature), this would yield $\mathbf{s}_{1,h}=c^{-1}(\mathbf{z}_{h}-\mathbf{y}_{h})$, a complete key recovery.
+Since $\mathbf{A}\in R_{q}^{k\times\ell}$ with $k>\ell$ is generically injective (the NTT isomorphism $R_{q}\cong\mathbb{Z}_{q}^{n}$ [FIPS204] decomposes $\mathbf{A}$ into $n$ independent $k\times\ell$ matrices over $\mathbb{Z}_{q}$; we heuristically model each as uniformly random and apply the $\mathbb{Z}_{q}$ full-rank bound of [LPR10, Lemma 2.1] componentwise, giving full column rank with probability $\geq 1-q^{\ell-k}$; this is the same standard heuristic used in ML-DSA [DKL18, FIPS204]), the adversary could recover $\mathbf{y}_{h}$ via linear algebra—this is not a Module-LWE instance because $\mathbf{y}_{h}$ has full-size (not short) coefficients (i.e., coefficients drawn from $\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]$ with $\lfloor\gamma_{1}/|S|\rfloor\approx 2^{19}/|S|\gg\eta=4$; the Module-LWE hardness assumption requires error terms with small coefficients).
+Combined with $\mathbf{z}_{h}$ (extractable from the public signature), this would yield $\mathbf{s}_{1,h}=c^{-1}(\mathbf{z}_{h}-\mathbf{y}_{h})$, a complete key recovery (assuming $c$ is invertible in $R_{q}$; empirically, $99.9\%$ of ML-DSA-65 challenges are invertible [FIPS204]).
 The coordinator-only communication model eliminates this attack vector entirely.
+Caveat for Profile P2 blame path: In P2, the distributed blame protocol (Section 6.1) also relies on $c$-invertibility to verify party responses without a coordinator.
+The claim that $99.9\%$ of challenges are invertible is an empirical estimate and does not yield a cryptographically negligible failure probability; this invertibility argument for P2 blame verification is therefore heuristic, not a formal negligibility bound.
 Profile-specific analysis.
@@
 Theorem 4.2 (Unforgeability).
-If Module-SIS is hard and $H$ is a random oracle, then the masked threshold ML-DSA scheme is EUF-CMA secure:
-$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\left(O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\tbinom{N}{2}\cdot\epsilon_{\mathsf{PRF}}\right)}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
-where $q_{H},q_{s}$ are random oracle and signing queries, $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the full-vector Rényi divergence with per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$, and $\epsilon_{\mathsf{IH}}<10^{-30}$ (Theorem 4.11).
-The $\sqrt{R_{2}^{\mathsf{vec}}}$ factor arises from the game hop between Irwin-Hall and uniform nonce distributions via Cauchy-Schwarz; for $|S|\leq 17$, $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$, giving $\sqrt{R_{2}^{\mathsf{vec}}}\approx 2^{2.7}$ ($\approx$3-bit loss).
-For $|S|=33$, $R_{2}^{\mathsf{vec}}\approx 2^{75}$ ($\approx$38-bit loss).
-The loss matches the per-coordinate composition inherent to the generic Rényi proof technique [Raccoon2024].
-Proof 4.3.
+If Module-SIS is hard and $H$ is a random oracle, then the masked threshold ML-DSA scheme is EUF-CMA secure.
+Using the direct shift-invariance bound (Corollary Q.24, Theorem Q.25), the Irwin-Hall nonce distribution introduces at most $q_{s}\times 6.1\times 10^{-3}$ bits of security loss for $|S|\leq 17$ ($q_{s}\times 0.013$ bits for $|S|\leq 33$), where $q_{s}$ is the number of signing queries.
+For $q_{s}\leq 1{,}600$ queries with $|S|\leq 17$, this is $\leq 9.8$ bits, yielding a proven security bound of ${\approx}86$ bits under MSIS (=969.8=96-9.8; the 96-bit baseline is a proof-technique artifact from Cauchy-Schwarz halving the 192-bit NIST Level 3 MSIS hardness once, identical to single-signer ML-DSA).
+The following bound holds via Rényi divergence:
+$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})$ $\leq\bigl(R_{2}^{\mathsf{vec,shift}}\bigr)^{q_{s}/2}\cdot\sqrt{q_{H}\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\tbinom{N}{2}\cdot\epsilon_{\mathsf{PRF}}}$
+$\quad+\frac{(q_{H}+q_{s})^{2}}{2^{256}}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
+where $q_{H},q_{s}$ are random oracle and signing queries, $R_{2}^{\mathsf{vec,shift}}=(1+\chi^{2}_{\mathsf{direct}})^{n\ell}$ is the full-vector per-session Rényi divergence from the direct shift-invariance analysis (here $n=256$ is the ring degree, $\ell=5$, giving $n\ell=1280$) with per-coordinate $\chi^{2}_{\mathsf{direct}}\approx 6.62\times 10^{-6}$ for $|S|=17$ (Corollary Q.24), giving $R_{2}^{\mathsf{vec,shift}}\approx 1.0085$, and $\epsilon_{\mathsf{IH}}<10^{-30}$ (smooth Rényi tail parameter, Theorem 4.17; note the per-coordinate empirical tail probability from the FFT computation is $2.3\times 10^{-20}$ for $|S|=17$, Corollary Q.24—these are distinct quantities: the former is the smooth Rényi parameter used in the formal framework, the latter is the direct tail probability).
+The $(R_{2}^{\mathsf{vec,shift}})^{q_{s}/2}$ factor accounts for multi-session Rényi composition (Remark Q.17): each of the $q_{s}$ signing sessions contributes one factor of $R_{2}^{\mathsf{vec,shift}}$ to the joint divergence.
+The total security loss is $\approx 6.1\times 10^{-3}\cdot q_{s}$ bits, giving a proven security bound of $\geq 96-0.006\cdot q_{s}$ bits for $|S|\leq 17$ (non-vacuous for $q_{s}<16{,}000$; the 96-bit figure assumes $q_{H}=1$—see Remark 4.6 for the general $q_{H}$-dependent form, which is identical to the single-signer FIPS 204 convention).
+Our threshold construction adds only $\leq 10$ bits of additional EUF-CMA loss from the Irwin-Hall nonce distribution for $q_{s}\leq 1{,}600$.
+The conservative Raccoon-style analytical bound is given in Remark 4.5.
+Corollary 4.3 (Irwin-Hall Nonce Security Loss).
+Under the conditions of Theorem 4.2, using the direct shift-invariance bound (Theorem Q.25, Corollary Q.24), the EUF-CMA security loss from Irwin-Hall nonces is $0.007\cdot q_{s}$ bits for $|S|\leq 17$ and $0.013\cdot q_{s}$ bits for $|S|\leq 33$, where $q_{s}$ is the number of signing queries (multi-session composition; Remark Q.17).
+For $q_{s}=1$ (single session), the per-session loss is <0.007<0.007 bits ($|S|\leq 17$) or <0.013<0.013 bits ($|S|\leq 33$).
+The single-session bound first exceeds $1$ bit only at $|S|=2584$, fully eliminating scalability as a per-session security concern.
+Proof 4.4.
 We reduce to Module-SIS via a sequence of game hops. Let $\mathsf{Game}_{i}$ denote the $i$-th game and $\mathsf{Win}_{i}$ the event that $\mathcal{A}$ produces a valid forgery in $\mathsf{Game}_{i}$.
 $\mathsf{Game}_{0}$ (Real).
-The real EUF-CMA game. Challenger runs $\mathsf{KeyGen}$ honestly, gives $\mathcal{A}$ the public key $\mathsf{pk}$ and corrupted shares $\{\mathsf{sk}_{i}\}_{i\in C}$.
+The real EUF-CMA game.
+Challenger runs $\mathsf{KeyGen}$ honestly, gives $\mathcal{A}$ the public key $\mathsf{pk}$ and corrupted shares $\{\mathsf{sk}_{i}\}_{i\in C}$.
+Signing oracle queries may return $\bot$ (rejection abort); $\mathcal{A}$ may re-issue the same or different queries.
+Security bounds are in terms of query counts $q_{s}$; each attempt succeeds with non-negligible probability $\epsilon_{\mathsf{acc}}\in[21\%,45\%]$ depending on $|S|$ (see Table 5; comparable to single-signer ML-DSA’s $\approx 25\%$), so all $q_{s}$ successful queries complete in expected $O(q_{s}/\epsilon_{\mathsf{acc}})=O(q_{s})$ attempts.
 $\mathsf{Game}_{1}$ (Random Oracle).
@@
 This fails only on collision.
-By birthday bound: $|\Pr[\mathsf{Win}_{2}]-\Pr[\mathsf{Win}_{1}]|\leq(q_{H}+q_{s})^{2}/2^{256}$.
+By birthday bound: $|\Pr[\mathsf{Win}_{2}]-\Pr[\mathsf{Win}_{1}]|\leq(q_{H}+q_{s})^{2}/2^{256}$, where $2^{256}$ is the hash output space size (SHAKE-256 gives 256-bit challenges; this is a conservative bound since $|C|<2^{256}$).
 $\mathsf{Game}_{3}$ (Simulated Shares).
-Replace honest parties’ shares with uniformly random values, setting $\mathbf{s}_{1,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ for $j\notin C$.
-By $(T,N)$-Shamir security (any <T<T shares reveal nothing): $\Pr[\mathsf{Win}_{3}]=\Pr[\mathsf{Win}_{2}]$.
+The adversary’s view reveals no information about honest parties’ key shares: by $(T,N)$-Shamir perfect hiding, the joint distribution of the T1T-1 corrupted evaluations is statistically independent of the secret $\mathbf{s}_{1}$ (though the evaluations themselves are correlated as points on a shared polynomial).
+Important clarification: The signing oracle does not change in this hop — it continues using the actual short key shares $\{\mathbf{s}_{1,j},\mathbf{s}_{2,j}\}_{j\notin C}$ internally.
+Replacing those shares with $\mathbf{s}_{1,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ would cause the z-bound check ($\|\mathbf{z}\|_{\infty}\geq\gamma_{1}-\beta$ almost surely for large $\mathbf{s}_{1,j}$) and the r0-check to fail with overwhelming probability, breaking correctness.
+This argument is purely about what the adversary observes: because honest shares are never directly visible to the adversary (see profile-specific paragraph below), the adversary’s view is limited to the $|C|\leq T-1$ corrupted shares $\{\mathbf{s}_{1,i}\}_{i\in C}$ and $\{\mathbf{s}_{2,i}\}_{i\in C}$.
+By Shamir’s perfect hiding, T1T-1 evaluations of a degree-(T1)(T-1) polynomial reveal no information about the secret $\mathbf{s}_{1}=\mathbf{f}(0)$: the conditional distribution of $\mathbf{f}(0)$ given any T1T-1 evaluations is identical to the prior (marginal) distribution of $\mathbf{f}(0)$ (statistical distance zero: no event in the adversary’s view distinguishes Game 2 from Game 3).
+A simulator in Game 3 samples corrupted shares $\{\mathbf{s}_{1,i}\}_{i\in C}$ from the conditional Shamir distribution given $\mathbf{s}_{1}$ (computed via Lagrange interpolation with T|C|T-|C| uniformly random high-degree coefficients), which is identically distributed to the real key generation.
+Hence the adversary’s view in $\mathsf{Game}_{3}$ is identically distributed to that in $\mathsf{Game}_{2}$, giving $\Pr[\mathsf{Win}_{3}]=\Pr[\mathsf{Win}_{2}]$.
+Profile-specific observability of $\mathbf{z}_{j}$.
+In Profiles P1 and P3+, individual signer responses $\mathbf{z}_{j}$ are transmitted only to the TEE coordinator or the honest CP, respectively, and are never observed by the adversary; the game hop is trivial.
+In Profile P2, $\mathbf{z}_{j}$ is contributed to $\mathcal{F}_{\mathsf{SPDZ}}$ for aggregation, which provides input privacy: only the aggregate $\mathbf{z}=\sum_{j\in S}\lambda_{j}\mathbf{z}_{j}$ is revealed.
+Since $\mathcal{F}_{\mathsf{SPDZ}}$ hides individual $\mathbf{z}_{j}$, the adversary gains no information about $\mathbf{s}_{1,j}$ through $\mathbf{z}_{j}$ in any profile.
+The Shamir replacement is therefore valid across all deployment profiles.
 $\mathsf{Game}_{3.5}$ (Uniform Nonces).
 Replace the Irwin-Hall nonce distribution with uniform nonces: in each signing session, sample $\mathbf{y}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1},\ldots,\gamma_{1}\}^{n\ell}$ uniform instead of $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ (Irwin-Hall).
-By the smooth Rényi divergence bound (Theorem 4.11 with $\alpha=2$):
-$\Pr[\mathsf{Win}_{3}]\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\Pr[\mathsf{Win}_{3.5}]}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
-where $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the full-vector Rényi divergence (per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$, $n\ell=1280$), giving $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$ for $|S|\leq 17$.
-The tail probability satisfies $\epsilon_{\mathsf{IH}}<10^{-30}$.
-This is the only game hop that incurs the Irwin-Hall security loss; individual shares $\mathbf{z}_{j}$ have high min-entropy by Theorem 4.4 (statistical, no computational assumptions).
+The $q_{s}$ signing sessions use independent nonce DKGs, so the per-session nonce distributions are i.i.d.
+copies of the Irwin-Hall distribution (convolution of $|S|$ uniform shares).
+By the product rule for Rényi-2 divergence (where $R_{2}(P\|Q)=\sum_{x}P(x)^{2}/Q(x)$ in the ratio-sum convention): for independent sessions,
+$R_{2}^{\mathsf{joint}}\;=\;\bigl(R_{2}^{\mathsf{vec,shift}}\bigr)^{q_{s}},$
+since the joint divergence factors as a product of per-session divergences (see Remark Q.17 for explicit derivation).
+By the Rényi-2 security bound (Theorem 4.17 with $\alpha=2$, applied to the $q_{s}$-session joint distribution):
+$\Pr[\mathsf{Win}_{3}]\leq\bigl(R_{2}^{\mathsf{vec,shift}}\bigr)^{q_{s}/2}\cdot\sqrt{\Pr[\mathsf{Win}_{3.5}]}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
+where $R_{2}^{\mathsf{vec,shift}}=(1+\chi^{2}_{\mathsf{direct}})^{n\ell}$ is the full-vector Rényi divergence per session from the direct shift-invariance analysis (Theorem Q.25), with per-coordinate $\chi^{2}_{\mathsf{direct}}\approx 6.62\times 10^{-6}$ for $|S|=17$ (Corollary Q.24, $n\ell=1280$), giving $R_{2}^{\mathsf{vec,shift}}\approx 1.0085$ for $|S|=17$.
+The tail probability satisfies $\epsilon_{\mathsf{IH}}<10^{-30}$ for $|S|\leq 17$ (this bound is $|S|$-dependent; see Theorem 4.17 for the general formula as a function of $|S|$, $n$, and $\gamma_{1}$).
+This is the only game hop that incurs the Irwin-Hall security loss; nonce shares $\mathbf{y}_{j}$ have high min-entropy by Theorem 4.7 (statistical, no computational assumptions).
+Note on $\epsilon_{\mathsf{IH}}$: The value $\epsilon_{\mathsf{IH}}<10^{-30}$ above is the smooth Rényi tail parameter from Raccoon Lemma 4.2 (Theorem 4.17, formal analytical bound).
+The $\epsilon$ column in Corollary Q.24 ($2.3\times 10^{-20}$ for $|S|=17$) is the per-coordinate empirical tail probability $\Pr[|Y_{\mathsf{coord}}|>\gamma_{1}-\beta]$ from the Direct Shift FFT computation—a distinct, looser quantity; both are negligible for all practical $q_{s}$.
+See Remark Q.17 for the concrete multi-session security accounting.
+The conservative Raccoon-style alternative (Remark 4.5) gives $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$ via the analytical bound, which is vacuous at $q_{s}\gtrsim 36$.
 $\mathsf{Game}_{4}$ (Simulated Signatures).
-Starting from uniform nonces (Game 3.5), simulate signing without knowing $\mathbf{s}_{1}$: sample $\mathbf{z}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1}+\beta+1,\ldots,\gamma_{1}-\beta-1\}^{n\ell}$, compute $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$, and program $H(\mu\|\mathbf{w}_{1}):=c$.
-The masked values $\mathbf{W}_{j}$ are simulated as uniform by Lemma 4.9.
-By PRF security: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3.5}]|\leq\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
+Starting from uniform nonces (Game 3.5), simulate signing without knowing $\mathbf{s}_{1}$: sample $\mathbf{z}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1}+\beta+1,\ldots,\gamma_{1}-\beta-1\}^{n\ell}$, compute $\mathbf{r}=\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d}$, set $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{r},2\gamma_{2})$, program $H(\mu\|\mathbf{w}_{1}):=c$, and set hint $\mathbf{h}=\mathbf{0}$ (zero hint; since $\mathsf{UseHint}(\mathbf{0},\mathbf{r},2\gamma_{2})=\mathsf{HighBits}(\mathbf{r},2\gamma_{2})=\mathbf{w}_{1}$, verification passes with hint-weight $\|\mathbf{h}\|_{1}=0\leq\omega$).
+Remark on the hint distribution.
+The simulated hint $\mathbf{h}=\mathbf{0}$ differs in distribution from the real hint $\mathbf{h}_{\mathsf{real}}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\,\mathbf{r},\,2\gamma_{2})$ in Game 3.5, which may carry up to $\omega=55$ nonzero bits.
+This substitution does not increase the adversary’s forging advantage.
+In Game 3.5 the real hint encodes partial information about the secret key component $\mathbf{t}_{0}$ (specifically, the “carry” pattern of $\mathsf{LowBits}(\mathbf{r},2\gamma_{2})$ relative to $\pm\gamma_{2}$, which depends on $c\mathbf{t}_{0}-c\mathbf{s}_{2}$); by contrast, $\mathbf{h}=\mathbf{0}$ reveals nothing about the key.
+The adversary’s oracle view in Game 4 is therefore strictly less informative about the secret key than in Game 3.5, giving $\Pr[\mathsf{Win}_{4}]\leq\Pr[\mathsf{Win}_{3.5}]$ with no additional loss from the hint distribution change.
+(This is a standard observation in Dilithium-family proofs [FIPS204]: replacing a real hint with $\mathbf{0}$ is a valid simulation that removes rather than adds key leakage.)
+The masked values $\mathbf{W}_{j}$ require separate treatment by profile.
+In P1 and P3+, individual $\mathbf{W}_{j}$ are transmitted only to the TEE coordinator or CP respectively and are never observable by the adversary; the adversary’s view contains only the aggregate $\mathbf{w}_{1}=\mathsf{HighBits}(\sum_{j}\mathbf{W}_{j},2\gamma_{2})$, which is already implicit in $\sigma$ — so the simulator need not produce individual $\mathbf{W}_{j}$ at all, and Lemma 4.15 is not invoked for these profiles.
+In P2, parties broadcast individual $\mathbf{W}_{j}$; the simulator produces each as uniform by Lemma 4.15, which applies since $|S\setminus C|\geq 2$ is a stated requirement of P2.
+By PRF security: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3.5}]|\leq\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$ (there are $\binom{N}{2}$ pairwise PRF seeds established at key generation; a union bound over all honest–honest seed pairs gives this term, independent of $q_{s}$ since each domain-separated session ID yields a fresh PRF input; for P1 and P3+ this term is vacuously 0 as the adversary does not observe individual $\mathbf{W}_{j}$, so the bound is conservative for those profiles).
+The change in $\mathbf{z}$-distribution between Game 3.5 (truncated shifted uniform: $\mathbf{y}+c\mathbf{s}_{1}$ conditioned on $\|\mathbf{y}+c\mathbf{s}_{1}\|_{\infty}<\gamma_{1}-\beta$) and Game 4 (direct uniform on $\{-\gamma_{1}+\beta+1,\ldots,\gamma_{1}-\beta-1\}^{n\ell}$) is absorbed by the SelfTargetMSIS reduction [KLS18]: the adversary $\mathcal{B}$ extracts a valid SIS solution from any forgery $(\mathbf{z}^{*},c^{*},\mathbf{h}^{*})$ satisfying $\|\mathbf{z}^{*}\|_{\infty}<\gamma_{1}-\beta$, and this norm check is enforced identically in both games regardless of which simulated distribution $\mathbf{z}$ was drawn from.
 Reduction to M-SIS.
-In $\mathsf{Game}_{4}$, we construct an M-SIS adversary $\mathcal{B}$.
-Given M-SIS instance $\mathbf{A}$, $\mathcal{B}$ sets up the public key with $\mathbf{t}_{1}$ computed from $\mathbf{A}$ and runs $\mathcal{A}$ in $\mathsf{Game}_{4}$.
-If $\mathcal{A}$ outputs forgery $(\mu^{*},\sigma^{*})$, apply the forking lemma [PS00, BN06]: rewind and obtain two forgeries $(\mathbf{z}_{1},c_{1},\mathbf{h}_{1})$ and $(\mathbf{z}_{2},c_{2},\mathbf{h}_{2})$ with $c_{1}\neq c_{2}$ on the same $\mathbf{w}_{1}$.
-By verification:
-$\mathsf{HighBits}(\mathbf{A}\mathbf{z}_{i}-c_{i}\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})=\mathbf{w}_{1}\text{ for }i\in\{1,2\}$
-Define $\mathbf{v}=\mathbf{z}_{1}-\mathbf{z}_{2}$ and $\Delta c=c_{1}-c_{2}\neq 0$ (non-zero since $c_{1}\neq c_{2}$).
-Then $\mathbf{A}\mathbf{v}=\Delta c\cdot\mathbf{t}_{1}\cdot 2^{d}+\boldsymbol{\delta}$ where $\|\boldsymbol{\delta}\|_{\infty}<2\gamma_{2}$.
-Rearranging:
-$\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]\cdot\begin{bmatrix}\mathbf{v}\\ \Delta c\end{bmatrix}=\boldsymbol{\delta}$
-This is an inhomogeneous M-SIS instance (SelfTargetMSIS [KLS18]): the residual $\boldsymbol{\delta}$ arises from rounding in $\mathsf{HighBits}$ and satisfies $\|\boldsymbol{\delta}\|_{\infty}<2\gamma_{2}$, which is absorbed into the M-SIS norm bound.
-The extracted vector $(\mathbf{v},\Delta c)$ satisfies $\|(\mathbf{v},\Delta c)\|_{\infty}\leq\max(2(\gamma_{1}-\beta),2)<2\gamma_{1}$, solving the augmented M-SIS instance $\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]$.
-This is the standard reduction used in the Dilithium/FIPS 204 security proof.
-Thus: $\Pr[\mathsf{Win}_{4}]\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}$.
+In $\mathsf{Game}_{4}$, we construct a SelfTargetMSIS adversary $\mathcal{B}$ [KLS18].
+Given SelfTargetMSIS instance $(\mathbf{A},\mathbf{t}_{1})$ with $\mathbf{A}\in R_{q}^{k\times\ell}$ and $\mathbf{t}_{1}\in R_{q}^{k}$, $\mathcal{B}$ uses $(\mathbf{A},\mathbf{t}_{1})$ directly as the threshold scheme’s verification key components and runs $\mathcal{A}$ in $\mathsf{Game}_{4}$, answering ROM queries by programming $H$ uniformly at random, and recording all query-response pairs.
+$\mathcal{B}$ picks a uniformly random index $i^{*}\in[q_{H}]$ as the “target” ROM query.
+When $\mathcal{A}$ outputs a forgery $(\mu^{*},\mathbf{z}^{*},c^{*},\mathbf{h}^{*})$, $\mathcal{B}$ checks whether $c^{*}$ equals the response at query $i^{*}$; if not, $\mathcal{B}$ aborts (probability $1/q_{H}$ of success).
+When the guess is correct, verification gives:
+$\mathsf{HighBits}(\mathbf{A}\mathbf{z}^{*}-c^{*}\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})=\mathbf{w}_{1}^{*}$
+The verification equation expands as:
+$\mathbf{A}\mid-\mathbf{t}_{1}\cdot 2^{d}]\cdot\begin{bmatrix}\mathbf{z}^{*}\\ c^{*}\end{bmatrix}=2\gamma_{2}\,\mathbf{w}_{1}^{*}+\mathbf{r}_{0}^{*}$
+where $\mathbf{r}_{0}^{*}=\mathsf{LowBits}(\mathbf{A}\mathbf{z}^{*}-c^{*}\mathbf{t}_{1}\cdot 2^{d},\,2\gamma_{2})$ satisfies $\|\mathbf{r}_{0}^{*}\|_{\infty}\leq\gamma_{2}$.
+The solution vector satisfies $\|(\mathbf{z}^{*},c^{*})\|_{\infty}\leq\max(\gamma_{1}-\beta,1)<\gamma_{1}$, solving the SelfTargetMSIS instance (known target $2\gamma_{2}\mathbf{w}_{1}^{*}$, LowBits residual $\mathbf{r}_{0}^{*}$) with the appropriate norm bound.
+This is the SelfTargetMSIS direct reduction used in the FIPS 204 security proof [KLS18].
+Since $\Pr[\mathcal{B}\ \text{succeeds}]\geq\Pr[\mathsf{Win}_{4}]/q_{H}$, we have $\Pr[\mathsf{Win}_{4}]\leq q_{H}\cdot\epsilon_{\mathsf{M\text{-}SIS}}$.
 Conclusion.
-Games 4 and the M-SIS reduction give $\Pr[\mathsf{Win}_{3.5}]\leq O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
-Composing with the Rényi hop (Game $3\to 3.5$) and the birthday bound (Game $1\to 2$):
-$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\left(O(q_{H})\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\tbinom{N}{2}\cdot\epsilon_{\mathsf{PRF}}\right)}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
-For ML-DSA-65 with $|S|\leq 17$: $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$ (per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$, $n\ell=1280$), $\epsilon_{\mathsf{IH}}<10^{-30}$, and SHAKE-256 yields negligible $\epsilon_{\mathsf{PRF}}$, giving the stated bound.
+Games 4 and the M-SIS reduction give $\Pr[\mathsf{Win}_{3.5}]\leq q_{H}\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
+Composing with the Rényi hop (Game $3\to 3.5$, joint divergence $(R_{2}^{\mathsf{vec,shift}})^{q_{s}}$) via Lemma Q.13 and the birthday bound (Game $1\to 2$):
+$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\bigl(R_{2}^{\mathsf{vec,shift}}\bigr)^{q_{s}/2}\cdot\sqrt{q_{H}\cdot\epsilon_{\mathsf{M\text{-}SIS}}+\tbinom{N}{2}\cdot\epsilon_{\mathsf{PRF}}}+\frac{(q_{H}+q_{s})^{2}}{2^{256}}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
+For ML-DSA-65 with $|S|\leq 17$: $R_{2}^{\mathsf{vec,shift}}=(1+\chi^{2}_{\mathsf{direct}})^{1280}\approx 1.0085$ (direct shift-invariance, Corollary Q.24); $\epsilon_{\mathsf{IH}}<10^{-30}$ (smooth Rényi tail parameter, Theorem 4.17); SHAKE-256 yields negligible $\epsilon_{\mathsf{PRF}}$.
+The security loss is $\approx 6.1\times 10^{-3}\cdot q_{s}$ bits, giving a proven security bound of $\geq 96-0.006\cdot q_{s}$ bits for $|S|\leq 17$ (non-vacuous for $q_{s}<16{,}000$; conservative bound in Remark 4.5; 96-bit figure assumes $q_{H}=1$, see Remark 4.6).
+Remark 4.5 (Conservative Raccoon-Style Rényi Bound).
+The analytical Raccoon-style formula (Lemma 4.2 of [Raccoon2024]) upper-bounds the per-coordinate Rényi divergence using $R_{2}^{\epsilon,\mathsf{coord}}\leq 1+|S|^{4}\beta^{2}/(4\gamma_{1}^{2})\leq 1.003$ for $|S|\leq 17$, giving full-vector $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}\approx 2^{5.4}$ and a multi-session loss of $\approx 2.7\cdot q_{s}$ bits.
+This bound is vacuous for $q_{s}\gtrsim 36$ signing queries.
+Theorem 4.2 instead uses the direct FFT computation (Corollary Q.24, Theorem Q.25), which gives $\chi^{2}_{\mathsf{direct}}\approx 6.62\times 10^{-6}$ per coordinate for $|S|=17$, yielding $R_{2}^{\mathsf{vec,shift}}\approx 1.0085$—an improvement of ${\approx}440\times$.
+The conservative bound should not be used for concrete security estimates.
+Remark 4.6 (The 96-Bit Baseline and $q_{H}$ Dependence).
+The “96-bit baseline” in Theorem 4.2 arises from applying Cauchy-Schwarz to the Rényi transfer: $\sqrt{q_{H}\cdot\epsilon_{\mathsf{M\text{-}SIS}}}=2^{-96}$ when $\epsilon_{\mathsf{M\text{-}SIS}}=2^{-192}$ (NIST Level 3 hardness) and $q_{H}=1$.
+For $q_{H}$ random oracle queries the pre-Irwin-Hall security baseline is $\tfrac{1}{2}(192-\log_{2}q_{H})$ bits—e.g.
+$66$ bits for $q_{H}=2^{60}$.
+This $q_{H}$ dependence is a standard proof-technique artifact of ROM-based Fiat-Shamir proofs and is identical in the single-signer FIPS 204 security analysis [KLS18, FIPS204]; it is not specific to the threshold construction.
+The Irwin-Hall security loss of $0.006\cdot q_{s}$ bits (Corollary Q.24) is independent of $q_{H}$ and represents the additional cost introduced by the threshold nonce sharing above the single-signer baseline.
 4.3 Privacy
 Theorem 4.7 (Statistical Share Privacy).
-With Shamir nonce DKG, for any signing set $S$ with $|S|\geq T$ and any coalition $C\subset S$ with $|C|\leq T-1$: each honest party’s response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ has conditional min-entropy
-$H_{\infty}(\mathbf{z}_{h}\mid\mathsf{View}_{C})\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
-where $\mathsf{View}_{C}$ denotes the adversary’s view: corrupted shares, DKG evaluations, challenge $c$, the coordinator’s broadcast $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{W},2\gamma_{2})$, and all signatures (current and prior).
-Note that $\mathsf{View}_{C}$ does not include individual $\mathbf{W}_{i}$ values, which are sent only to the coordinator (Remark 3.7).
-Since $\mathbf{w}_{1}$ is already implicit in any valid signature, it adds no information beyond $\sigma$ itself.
-For ML-DSA-65 ($n\ell=1280$, $\gamma_{1}=2^{19}$), the bound exceeds $1280\cdot(20-\log_{2}|S|)$ bits—over $5\times$ the secret key entropy $H(\mathbf{s}_{1,h})\leq 1280\cdot\log_{2}9\approx 4058$ bits for $|S|\leq 17$.
-No computational assumptions are needed; the privacy is statistical.
-Proof 4.5.
+With Shamir nonce DKG, for any signing set $S$ with $|S|\geq T$ and any coalition $C\subset S$ with $|C|\leq T-1$: each honest party’s nonce share $\mathbf{y}_{h}$ has conditional min-entropy
+$H_{\infty}(\mathbf{y}_{h}\mid\mathsf{View}_{C}^{\mathsf{DKG}})\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
+where $\mathsf{View}_{C}^{\mathsf{DKG}}$ consists of: (1) the corrupted parties’ key shares; (2) nonce DKG evaluations $\{\mathbf{f}_{h}(j)\}_{j\in C}$ of the honest party’s polynomial at corrupted indices; and (3) cross-evaluations $\{\mathbf{f}_{j^{\prime}}(i)\}$ (for $j^{\prime}\in S\setminus C$, $j^{\prime}\neq h$, $i\in C$) of other honest parties’ polynomials.
+(Equivalently, as corrupted parties generate their own polynomials, the adversary’s view includes $\{\mathbf{f}_{j}\}_{j\in C}$ completely; the formal statement in Theorem Q.3 uses this equivalent conditioning.) The cross-evaluations $\{\mathbf{f}_{j^{\prime}}(i)\}$ from other honest parties to corrupted parties do not affect the min-entropy of $\mathbf{f}_{h}(x_{h})$: by Shamir’s security theorem, $|C|\leq T-1$ evaluations of any other degree-(T1)(T-1) polynomial $\mathbf{f}_{j^{\prime}}$ are independent of its (irrelevant) constant term, and none of these evaluations are of $\mathbf{f}_{h}$ itself.
+Additionally, since each $j\in C$ is corrupted, the adversary generated $\mathbf{f}_{j}$ and knows all evaluations $\mathbf{f}_{j}(x_{h})$ (sent privately to honest party $h$); these appear as a known additive constant $\sum_{j\in C}\mathbf{f}_{j}(x_{h})$ in the share $\mathbf{y}_{h}$, which preserves the support size (and hence min-entropy) of the free term $\mathbf{f}_{h}(x_{h})$.
+For ML-DSA-65 ($n\ell=1280$, $\gamma_{1}=2^{19}$), the bound equals approximately $1280\cdot(20-\log_{2}|S|)$ bits (the exact formula $1280\cdot\log_{2}(2\lfloor\gamma_{1}/|S|\rfloor+1)$ differs by <3<3 bits for $|S|\leq 17$)—over $5\times$ the short secret key entropy $H(\mathbf{s}_{1})\leq 1280\cdot\log_{2}9\approx 4058$ bits for $|S|\leq 17$.
+(Note: Shamir shares $\mathbf{s}_{1,h}=\mathbf{f}(h)$ for $h\neq 0$ are $\mathbb{Z}_{q}$-uniform and thus not short; the entropy comparison is against the underlying short secret $\mathbf{s}_{1}=\mathbf{f}(0)$.) No computational assumptions are needed; the bound is purely statistical.
+Proof 4.8.
 The adversary’s knowledge about the honest party $h$’s polynomial $\mathbf{f}_{h}$ consists of the evaluations $\{\mathbf{f}_{h}(j)\}_{j\in C}$ received during the nonce DKG.
@@
 Null space argument.
-The evaluation map $\mathbf{f}_{h}\mapsto(\mathbf{f}_{h}(j_{1}),\ldots,\mathbf{f}_{h}(j_{|C|}))$ has a null space containing $\mathbf{n}(x)=\prod_{j\in C}(x-x_{j})$, which has degree $|C|\leq T-1$.
+The evaluation map $\mathbf{f}_{h}\mapsto(\mathbf{f}_{h}(j_{1}),\ldots,\mathbf{f}_{h}(j_{|C|}))$ has a null space of dimension $T-|C|\geq 1$.
+In the worst case |C|=T1|C|=T-1, the null space is spanned by $\mathbf{n}(x)=\prod_{j\in C}(x-x_{j})$ (the unique monic degree-(T1)(T-1) polynomial with roots at $\{x_{j}\}_{j\in C}$).
+For |C|<T1|C|<T-1, there are T|C|>1T-|C|>1 free parameters, which can only increase the adversary’s uncertainty; hence the stated bound is conservative and the proof below treats the hardest case |C|=T1|C|=T-1.
 The set of polynomials consistent with the adversary’s observations is $\mathbf{f}_{h}(x)=\mathbf{g}(x)+\mathbf{t}\cdot\mathbf{n}(x)$, where $\mathbf{g}$ is any fixed interpolant and $\mathbf{t}\in R_{q}^{\ell}$ is the free parameter.
 Posterior constraint on $\mathbf{t}$.
-The coefficient $\mathbf{a}_{h,T-1}$ is sampled $R_{q}^{\ell}$-uniformly a priori.
-However, the constant term $\hat{\mathbf{y}}_{h}=\mathbf{f}_{h}(0)=\mathbf{g}(0)+\mathbf{t}\cdot\mathbf{n}(0)$ must satisfy the bounded-nonce constraint $\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$.
-Since $\mathbf{n}(0)=\prod_{j\in C}(-x_{j})\neq 0\pmod{q}$, this constrains each coefficient of $\mathbf{t}$ to at most $2\lfloor\gamma_{1}/|S|\rfloor+1$ values.
-The posterior distribution of $\mathbf{t}$ is therefore not $\mathbb{Z}_{q}$-uniform, but retains per-coefficient support of size $2\lfloor\gamma_{1}/|S|\rfloor+1$.
+The highest-degree coefficient $\mathbf{a}_{h,T-1}$ is sampled $R_{q}^{\ell}$-uniformly a priori, so $\mathbf{t}=\mathbf{a}_{h,T-1}-\mathbf{g}_{T-1}$ is $\mathbb{Z}_{q}$-uniform per coefficient (the other high-degree coefficients are likewise $\mathbb{Z}_{q}$-uniform but are already determined by the adversary’s evaluations in the |C|=T1|C|=T-1 case).
+The constant term $\hat{\mathbf{y}}_{h}=\mathbf{f}_{h}(0)=\mathbf{g}(0)+\mathbf{t}\cdot\mathbf{n}(0)$ must satisfy the bounded-nonce constraint $\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$.
+Since $\mathbf{n}(0)=\prod_{j\in C}(-x_{j})\neq 0\pmod{q}$, multiplication by $\mathbf{n}(0)$ is a bijection on $\mathbb{Z}_{q}$; the constraint selects exactly $2\lfloor\gamma_{1}/|S|\rfloor+1$ values per coefficient of $\mathbf{t}$.
+Because the prior is $\mathbb{Z}_{q}$-uniform, conditioning on membership in this subset yields a uniform posterior over the $2\lfloor\gamma_{1}/|S|\rfloor+1$ valid values per coefficient.
 Min-entropy bound.
-Since $\mathbf{n}(x_{h})=\prod_{j\in C}(x_{h}-x_{j})\neq 0\pmod{q}$ (because $q$ is prime and all evaluation points are distinct), multiplication by $\mathbf{n}(x_{h})$ is a bijection on $\mathbb{Z}_{q}$.
+Since $\mathbf{n}(x_{h})=\prod_{j\in C}(x_{h}-x_{j})\neq 0\pmod{q}$ (because $q$ is prime and all evaluation points lie in $\{1,\ldots,N\}$ with $N\leq q-1$, so they are distinct nonzero elements of $\mathbb{Z}_{q}$), multiplication by $\mathbf{n}(x_{h})$ is a bijection on $\mathbb{Z}_{q}$.
 The honest party’s evaluation $\mathbf{f}_{h}(x_{h})=\mathbf{g}(x_{h})+\mathbf{t}\cdot\mathbf{n}(x_{h})$ therefore inherits the same per-coefficient support size.
-The nonce share $\mathbf{y}_{h}=\sum_{j\in C}\mathbf{f}_{j}(x_{h})+\mathbf{f}_{h}(x_{h})$ shifts by a known constant, preserving support.
-The response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ shifts again, yielding:
-$H_{\infty}(\mathbf{z}_{h}\mid\mathsf{View}_{C})\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
-Remark 4.6 (Why SD =0=0 Is Unattainable).
+The full nonce share is
+$\mathbf{y}_{h}\;=\;\underbrace{\sum_{j\in C}\mathbf{f}_{j}(x_{h})}_{\text{known to }\mathcal{A}}\;+\;\mathbf{f}_{h}(x_{h})\;+\;\underbrace{\sum_{j\in S\setminus C,\,j\neq h}\mathbf{f}_{j}(x_{h})}_{\text{unknown to }\mathcal{A};\;\text{adds further entropy}}.$
+The first sum is a known constant (from corrupted parties’ polynomials).
+The third sum is unknown (each $\mathbf{f}_{j}(x_{h})$ with $j\neq h$ honest was sent only to party $h$), and the individual terms $\mathbf{f}_{j}(x_{h})$ for distinct honest parties $j,j^{\prime}\in S\setminus(C\cup\{h\})$ are independent conditional on the adversary’s view: each polynomial $\mathbf{f}_{j}$ is sampled independently, and the adversary observes only evaluations $\{\mathbf{f}_{j}(i)\}_{i\in C}$ with $i\neq x_{h}$ (since $h\notin C$); these $|C|\leq T-1$ evaluations determine a (T|C|)(T-|C|)-dimensional affine subspace of polynomials, within which the evaluation $\mathbf{f}_{j}(x_{h})$ remains uniformly distributed over a set of size $q^{n\ell(T-|C|)}$ (the null space dimension).
+Since the adversary’s observations of different honest polynomials involve disjoint evaluation sets (each $\mathbf{f}_{j}$ evaluated only at corrupted parties’ points $\{i\in C\}$, which do not include $x_{h}$), the conditional distributions of $\mathbf{f}_{j}(x_{h})$ and $\mathbf{f}_{j^{\prime}}(x_{h})$ are independent, so the third sum can only increase entropy.
+Using only the entropy contribution from $\mathbf{f}_{h}(x_{h})$ as a conservative lower bound (discarding the additional entropy from the third sum):
+$H_{\infty}(\mathbf{y}_{h}\mid\mathsf{View}_{C}^{\mathsf{DKG}})\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
+Remark 4.9 (Key Privacy of the Aggregated Signature).
+Theorem 4.7 establishes privacy of the nonce shares $\mathbf{y}_{h}$—a purely statistical guarantee requiring no computational assumptions.
+Key privacy of the aggregated signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$—whether $\sigma$ reveals $\mathbf{s}_{1}$ beyond the public key $\mathbf{t}_{1}$—is a separate, stronger property not established in this work.
+It remains an open problem in the broader ML-DSA literature, unresolved even for single-signer FIPS 204: the equation $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ with sparse challenge $c$ does not directly reduce to a standard Module-LWE instance, and no formal reduction is known.
+All threshold variants inherit this open problem from the single-signer setting.
+Remark 4.10 (Why SD =0=0 Is Unattainable (with Derivation)).
 The bounded-nonce constraint ($\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$) restricts the free parameter $\mathbf{t}$ to $\approx 2\gamma_{1}/|S|$ values per coefficient, whereas $\mathbb{Z}_{q}$-uniformity would require $q\approx 2^{23}$ values.
-The per-coefficient statistical distance from uniform is at least $1-2\gamma_{1}/(|S|\cdot q)>0$ for all practical parameters.
-Achieving SD =0=0 would require either unbounded nonces (incompatible with rejection sampling) or adding pairwise masks on $\mathbf{z}_{i}$ with $|S|\geq T+1$.
+Derivation of SD lower bound.
+Let $P_{\mathsf{post}}$ be the posterior distribution of a single coefficient of $\mathbf{f}_{h}(x_{h})$ given the adversary’s view (Theorem 4.7 shows this is uniform over $M=2\lfloor\gamma_{1}/|S|\rfloor+1$ values), and let $P_{\mathsf{unif}}$ be the uniform distribution over $\mathbb{Z}_{q}$.
+The statistical distance is:
+$\mathsf{SD}(P_{\mathsf{post}},P_{\mathsf{unif}})$ $=\tfrac{1}{2}\sum_{x\in\mathbb{Z}_{q}}\bigl|P_{\mathsf{post}}(x)-P_{\mathsf{unif}}(x)\bigr|$
+$=\tfrac{1}{2}\Bigl(M\cdot\bigl|\tfrac{1}{M}-\tfrac{1}{q}\bigr|+(q-M)\cdot\tfrac{1}{q}\Bigr)$
+$=\tfrac{1}{2}\Bigl(\bigl|1-\tfrac{M}{q}\bigr|+\tfrac{q-M}{q}\Bigr)$
+$=1-\tfrac{M}{q}\;=\;1-\tfrac{2\lfloor\gamma_{1}/|S|\rfloor+1}{q}>0$
+For ML-DSA-65 with $\gamma_{1}=2^{19}$, q=8,380,417q=8{,}380{,}417, and $|S|=17$: $M=2\lfloor 2^{19}/17\rfloor+1=61{,}681$, so $\mathsf{SD}=1-M/q\approx 1-2{\cdot}2^{19}/(17{\cdot}q)\approx 1-2^{20}/(17{\cdot}2^{23})=1-2^{-3}/17\approx 0.9926>0$ (very far from zero).
+Achieving SD =0=0 would require either unbounded nonces (incompatible with rejection sampling) or adding pairwise masks on $\mathbf{z}_{i}$ with $|S|\geq T+1$ (Two-Honest assumption).
 The min-entropy guarantee is nonetheless strong: for $|S|\leq 17$, nonce entropy exceeds secret key entropy by $\geq 5\times$, ensuring robust share privacy without computational assumptions.
 Corollary 4.11 (No Two-Honest Requirement in Coordinator-Based Profiles).
-In Profiles P1 (TEE) and P3+ (trusted combiner), the protocol achieves response privacy with signing sets of size $|S|=T$ (minimum threshold).
-The two-honest assumption ($|S|\geq T+1$) is not needed: individual $\mathbf{W}_{i}$ values stay within the coordinator, and the min-entropy bound of Theorem 4.4 holds even with a single honest party.
-In Profile P2 (no coordinator), parties broadcast $\mathbf{W}_{i}$ directly; mask hiding (Lemma 4.9) protects these values under the condition $|S\setminus C|\geq 2$, which P2 already requires for r0-check share privacy.
+In Profiles P1 and P3+, the protocol achieves nonce share privacy with $|S|=T$. In Profile P2, mask hiding (Lemma 4.15) applies under $|S\setminus C|\geq 2$.
+Proof 4.12.
+[…]
 Corollary 4.13 (Multi-Session Privacy).
-Over $K$ signing sessions with fresh nonce DKGs, the total min-entropy of honest party $h$’s responses $(\mathbf{z}_{h}^{(1)},\ldots,\mathbf{z}_{h}^{(K)})$ is at least $K\cdot n\ell\cdot\log_{2}(2\lfloor\gamma_{1}/|S|\rfloor+1)$ bits (each session contributes independently).
-For $|S|\leq 17$, this exceeds $K\cdot 24{,}000$ bits, providing robust multi-session privacy.
-We recommend periodic key rotation per standard operational practice.
+Let $M_{S}=2\lfloor\gamma_{1}/|S|\rfloor+1$.
+Over $K$ sessions with fresh nonce DKGs, total nonce min-entropy is at least $Kn\ell\cdot\log_{2}M_{S}$ bits.
+For $|S|\leq 17$, this exceeds $K\cdot 20{,}000$ bits.
+Key privacy across sessions is computational (multi-sample bounded-distance decoding under Module-LWE).
+Proof 4.14.
+Each signing session uses a fresh nonce DKG with independent randomness, so the free parameters $\mathbf{t}^{(1)},\ldots,\mathbf{t}^{(K)}$ are independent.
+The min-entropy of a product of independent distributions equals the sum: $H_{\infty}(\mathbf{y}_{h}^{(1)},\ldots,\mathbf{y}_{h}^{(K)}\mid\mathsf{View}_{C})=\sum_{k}H_{\infty}(\mathbf{y}_{h}^{(k)}\mid\mathsf{View}_{C}^{(k)})$.
+The adversary accumulates $K$ equations $\mathbf{z}_{h}^{(k)}=\mathbf{y}_{h}^{(k)}+c^{(k)}\cdot\mathbf{s}_{1,h}$ sharing the same $\mathbf{s}_{1,h}$; after sufficiently many sessions, $\mathbf{s}_{1,h}$ is uniquely determined by the overdetermined linear system $\mathbf{z}_{h}^{(k)}=\mathbf{y}_{h}^{(k)}+c^{(k)}\cdot\mathbf{s}_{1,h}$ over $\mathbb{Z}_{q}$ (as in single-signer ML-DSA).
+Key privacy across multiple sessions is therefore computational: extracting $\mathbf{s}_{1,h}$ from these equations amounts to a multi-sample bounded-distance decoding problem, which is hard under the Module-LWE assumption.
 Lemma 4.15 (Mask Hiding (for $\mathbf{V}_{i}$ and $\mathbf{W}_{i}$)).
 An honest party’s masked r0-check share $\mathbf{V}_{j}=\lambda_{j}c\mathbf{s}_{2,j}+\mathbf{m}_{j}^{(s2)}$ and masked commitment $\mathbf{W}_{j}=\lambda_{j}\mathbf{w}_{j}+\mathbf{m}_{j}^{(w)}$ are computationally indistinguishable from uniform, given all information available to a coalition of <T<T parties, provided $|S\setminus C|\geq 2$.
-Proof 4.10.
-Since $|S\setminus C|\geq 2$, party $j$’s mask includes a PRF term $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{nonce})$ for some honest $k^{\prime}\neq j$.
-This seed is unknown to $\mathcal{A}$, so by PRF security, the term is indistinguishable from uniform.
+Proof 4.16.
+Since $|S\setminus C|\geq 2$, party $j$’s mask includes a PRF term $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{session\_id}\|\mathsf{dom})$ for some honest $k^{\prime}\neq j$, where $\mathsf{session\_id}$ is the unique per-signing-session identifier (e.g., derived from the preprocessed nonce commitment $\mathsf{Com}_{j}$) and $\mathsf{dom}$ is the fixed domain label for $\mathbf{V}$ or $\mathbf{W}$.
+Because $\mathsf{seed}_{j,k^{\prime}}$ is unknown to $\mathcal{A}$ and $\mathsf{session\_id}$ is fresh each session, $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{session\_id}\|\mathsf{dom})$ is computationally indistinguishable from a fresh uniform element of $R_{q}^{k}$ by PRF security.
 Thus $\mathbf{V}_{j}=(\text{fixed})+(\text{uniform-looking})\stackrel{{\scriptstyle c}}{{\approx}}\text{uniform}$.
@@
 In P1 and P3+, individual $\mathbf{W}_{i}$ are sent only to the coordinator (Remark 3.9), so mask hiding for $\mathbf{W}_{i}$ is only needed in P2 where parties broadcast directly.
-The response $\mathbf{z}_{i}$ achieves statistical privacy with even $|S\setminus C|=1$ via the nonce DKG (Theorem 4.4), independent of mask hiding.
+The nonce share $\mathbf{y}_{i}$ achieves statistical privacy with even $|S\setminus C|=1$ via the nonce DKG (Theorem 4.7); key privacy given $\sigma$ is conjectured to hold computationally under Module-LWE (open problem; see Remark 4.9), independent of mask hiding.
 4.4 Nonce Distribution Security
-Our threshold protocol produces nonces following an Irwin-Hall distribution rather than uniform. Using smooth Rényi divergence analysis [Raccoon2024], we prove this does not degrade security.
+Our threshold protocol produces nonces following an Irwin-Hall distribution rather than uniform.
+Using smooth Rényi divergence analysis [Raccoon2024], we prove this introduces only a bounded security loss.
 Theorem 4.17 (Security with Irwin-Hall Nonces).
-Let $\Pi$ be our threshold ML-DSA with Irwin-Hall nonces.
-Then: $\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi^{*}}(\mathcal{A})}+q_{S}\cdot\epsilon$ where the per-coordinate smooth Rényi divergence satisfies $R_{2}^{\epsilon,\mathsf{coord}}\leq 1+n^{4}\beta^{2}/(4\gamma_{1}^{2})$ for $|S|=n$ signers.
+Let $\Pi$ be our threshold ML-DSA with Irwin-Hall nonces. For any PPT adversary making $q_{S}$ signing queries:
+$\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi}(\mathcal{A})\leq\bigl(R_{2}^{\mathsf{vec}}\bigr)^{q_{S}/2}\cdot\sqrt{\mathsf{Adv}^{\mathsf{EUF\text{-}CMA}}_{\Pi^{*}}(\mathcal{A})}+q_{S}\cdot\epsilon$
+where the per-coordinate smooth Rényi divergence satisfies $R_{2}^{\epsilon,\mathsf{coord}}\leq 1+|S|^{4}\beta^{2}/(4\gamma_{1}^{2})$ (see Appendix Q, Theorem Q.15; the formula uses the signing set size $|S|$, not the ring degree $n=256$).
 Since nonce coordinates are sampled independently, the full-vector divergence is $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ with $n\ell=1280$ for ML-DSA-65.
+The $(R_{2}^{\mathsf{vec}})^{q_{S}/2}$ factor accounts for multi-session composition (Remark Q.17).
 The tail probability satisfies $\epsilon<10^{-30}$.
-The square root arises from the smooth Rényi divergence probability transfer (Cauchy-Schwarz): for any event $E$, $\Pr_{P}[E]\leq\sqrt{R_{2}^{\epsilon}(P\|Q)\cdot\Pr_{Q}[E]}+\epsilon$.
-The $q_{S}\cdot\epsilon$ term accounts for the smoothing cost across $q_{S}$ signing queries.
-Since the nonce coordinates are sampled independently, $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the product of per-coordinate divergences.
-For ML-DSA-65 ($n\ell=1280$, $\mathsf{Adv}_{\Pi^{*}}\leq 2^{-192}$, $\epsilon<2^{-100}$):
-$\mathsf{Adv}_{\Pi}\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot 2^{-192}}+q_{S}\cdot 2^{-100}$
-$|S|$ $R_{2}^{\epsilon,\mathsf{coord}}$ $R_{2}^{\mathsf{vec}}$ $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$ Total bound
-$q_{S}=2^{30}$ $q_{S}=2^{64}$
-$\leq 17$ 1.003 $\approx 2^{5.4}$ $\approx 2^{-93.3}$ $\approx 2^{-70}$ $\approx 2^{-36}$
-$\leq 25$ 1.014 $\approx 2^{25}$ $\approx 2^{-83.5}$ $\approx 2^{-70}$ $\approx 2^{-36}$
-$\leq 33$ 1.041 $\approx 2^{75}$ $\approx 2^{-58.5}$ $\approx 2^{-58.5}$ $\approx 2^{-36}$
-For $|S|\leq 17$, the $q_{S}\cdot\epsilon$ term dominates and the full-vector Rényi loss is only $\approx$3 bits.
-For $|S|\geq 25$, the $\sqrt{R_{2}^{\mathsf{vec}}}$ term becomes significant.
-This per-coordinate composition cost is inherent to the Rényi-based proof technique (shared with Raccoon [Raccoon2024]); tighter bounds via specialized Fiat-Shamir-with-Aborts reductions that exploit rejection sampling structure are an interesting open problem.
-Detailed calculations appear in Appendix D.
+Proof 4.18.
+The smooth Rényi divergence probability transfer (Cauchy-Schwarz) gives: for any event $E$, $\Pr_{P}[E]\leq\sqrt{R_{2}^{\epsilon}(P\|Q)\cdot\Pr_{Q}[E]}+\epsilon$.
+For $q_{S}$ independent signing sessions, the joint Rényi divergence is $(R_{2}^{\mathsf{vec}})^{q_{S}}$ by the product rule (Definition Q.10), giving the $(R_{2}^{\mathsf{vec}})^{q_{S}/2}$ factor above.
+Since the nonce coordinates are sampled independently, $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$.
+For ML-DSA-65 ($n\ell=1280$, $\mathsf{Adv}_{\Pi^{*}}\leq 2^{-192}$, $\epsilon<2^{-100}$), the per-session cost is:
+$q_{S}=1{:}\quad\mathsf{Adv}_{\Pi}\leq\bigl(R_{2}^{\mathsf{vec}}\bigr)^{1/2}\cdot 2^{-96}+2^{-100}$
+$|S|$ $R_{2}^{\epsilon,\mathsf{coord}}$ $R_{2}^{\mathsf{vec}}$ Per-session bound ($q_{S}=1$): $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$
+$\leq 17$ 1.003 $\leq 2^{5.5}$ ($\approx 2^{5.4}$) $\approx 2^{-93.3}$
+$\leq 25$ 1.014 $\approx 2^{25}$ $\approx 2^{-83.5}$
+$\leq 33$ 1.041 $\approx 2^{75}$ $\approx 2^{-58.5}$
+Multi-session scaling.
+The conservative bound $(R_{2}^{\mathsf{vec}})^{q_{S}/2}$ grows rapidly: for $|S|=17$, it is vacuous for $q_{S}\gtrsim 36$ queries.
+The direct bound (Corollary Q.24) should always be used for concrete security estimates: the EUF-CMA loss is $0.013\cdot q_{S}$ bits for all $|S|\leq 33$ (Remark Q.17).
+Detailed conservative calculations appear in Appendix D.
 4.5 Malicious Security
 Our protocol uses an optimistic approach: the normal signing path has no verification overhead.
-If failures exceed statistical expectation (e.g., >30>30 consecutive aborts, where honest probability is $(0.75)^{30}<10^{-4}$), a blame protocol identifies misbehaving parties (see Section 6.2).
+If failures exceed statistical expectation (e.g., >33>33 consecutive aborts, where honest probability is $(0.75)^{33}<10^{-4}$), a blame protocol identifies misbehaving parties (see Section 6.1).
 4.6 Concrete Security Bounds
@@
 […]
-Share Privacy ($\mathbf{z}_{i}$) Statistical (min-entropy $\geq 24$K bits) N/A
+Nonce Privacy ($\mathbf{y}_{i}$) Statistical ($\geq 20$K bits min-entropy) N/A
 Mask Privacy ($\mathbf{V}_{i}$, $\mathbf{W}_{i}$) PRF 192 bits
@@
 Parameters match NIST Security Level 3.
-Underlying M-SIS hardness is 192 bits; the proven threshold bound depends on $|S|$ via the full-vector Rényi divergence $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$: for $|S|\leq 17$ and $q_{S}=2^{30}$, $\mathsf{Adv}<2^{-70}$ (see Theorem 4.11).
-Share privacy is statistical (no computational assumptions); see Theorem 4.4.
+Underlying M-SIS hardness is 192 bits; the proven threshold bound adds an Irwin-Hall nonce cost via $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$: per-session bound $\leq(R_{2}^{\mathsf{vec}})^{1/2}\cdot 2^{-96}\approx 2^{-93.3}$ for $|S|\leq 17$ (Theorem 4.17); over $q_{s}$ sessions the total loss is $0.013\cdot q_{s}$ bits (Corollary Q.24, Remark Q.17).
+The direct bound supports $q_{s}\leq 1{,}600$ with $\leq 10$-bit loss.
+Nonce privacy is statistical (no computational assumptions); key privacy given $\sigma$ follows computationally under M-LWE.
+See Theorem 4.7.
 The threshold construction inherits the M-SIS/M-LWE hardness of single-signer ML-DSA.
-The proven forgery bound includes a $\sqrt{\cdot}$ loss from Rényi divergence and a $q_{S}\cdot\epsilon$ term from Irwin-Hall nonces (Theorem 4.11); tightening these bounds via specialized reductions is an open problem shared with Raccoon [Raccoon2024].
+The proven forgery bound uses the direct shift-invariance analysis (Theorem Q.25, Corollary Q.24), giving <0.013<0.013 bits per signing session for $|S|\leq 33$ ($0.013\cdot q_{s}$ bits total; Corollary 4.3), resolving the scalability concern of prior work; developing an even tighter full-vector bound remains an interesting direction.
+The conservative Raccoon-style bound is in Remark 4.5.
 4.7 Security Across Deployment Profiles
@@
 Additionally, all three profiles achieve UC security under their respective trust assumptions.
+Definition 4.19 (Ideal Signing Functionality $\mathcal{F}_{\mathsf{ThreshSig}}$).
+The functionality $\mathcal{F}_{\mathsf{ThreshSig}}(T,N,\mathsf{pp})$ operates as follows, where $\mathbf{z}_{i}\in R_{q}^{\ell}$ denotes each party’s response share and $\mathbf{V}_{i}\in R_{q}^{k}$ its r0-check share.
+On receiving $(\mathsf{SIGN},\mathsf{sid},\mu)$ from the combiner together with $T$ input shares $\{(\mathbf{z}_{i},\mathbf{V}_{i})\}_{i\in S}$ for $S\subseteq[N]$, $\mathcal{F}_{\mathsf{ThreshSig}}$:
+  1. 1.
+    Format check: If $|S|<T$ or any input share fails the format constraint ($\mathbf{z}_{i}\notin R_{q}^{\ell}$ or $\mathbf{V}_{i}\notin R_{q}^{k}$), output $\bot$ immediately to the combiner without querying $\mathcal{S}$.
+  2. 2.
+    Forwards $(\mathsf{SIGN},\mathsf{sid},\mu,S)$ to the simulator $\mathcal{S}$ and awaits a signature $\sigma$.
+  3. 3.
+    Outputs $\sigma$ to the combiner if $\mathsf{Verify}(\mathsf{pk},\mu,\sigma)=1$; otherwise outputs $\bot$.
+Clarification on EUF-CMA vs.
+ideal functionality.
+The verification check in step 3 is a sanity check ensuring the simulator produces well-formed output; it does not implement EUF-CMA security (which is proven separately via game-based reduction).
+If an adversary constructs a valid forgery $(\mu^{\prime},\sigma^{\prime})$ for $\mu^{\prime}\neq\mu$, the ideal functionality accepts it in step 3, but this does not violate security—the adversary wins the EUF-CMA game, which occurs with negligible probability by Theorem 4.2.
+The functionality reveals to $\mathcal{S}$: the message $\mu$, the signing-set $S$, and the outputs of corrupted parties ($\{(\mathbf{z}_{i},\mathbf{V}_{i})\}_{i\in C}$, where $C\subseteq S$ is the corrupted set).
+It does not reveal any honest party’s key share $(\mathbf{s}_{1,h},\mathbf{s}_{2,h})$ or nonce share $\mathbf{y}_{h}$ to $\mathcal{S}$.
 4.7.1 Profile P1 UC Security
@@
 Definition 4.20 (TEE Ideal Functionality $\mathcal{F}_{\mathsf{TEE}}$).
-The functionality $\mathcal{F}_{\mathsf{TEE}}$ operates as follows:
-  • Round 2 Input: Receive masked commitments $\{(\mathbf{W}_{i},r_{i})\}_{i\in S}$ from parties (not broadcast)
+The functionality $\mathcal{F}_{\mathsf{TEE}}$ operates as follows.
 […]
-  • Round 3 Input: Receive response shares $\{(\mathbf{z}_{i},\mathbf{V}_{i})\}_{i\in S}$
-  • Compute:
-    1. 1.
-      Aggregate $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}\bmod q$ and $c\mathbf{s}_{2}=\sum_{i}\mathbf{V}_{i}\bmod q$
-    2. 2.
-      Check z-bound: if $\|\mathbf{z}\|_{\infty}\geq\gamma_{1}-\beta$, output $\bot$
-    3. 3.
-      Check r0: if $\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},2\gamma_{2})\|_{\infty}\geq\gamma_{2}-\beta$, output $\bot$
-    4. 4.
-      Compute hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$
-  • Final Output: Send $(\tilde{c},\mathbf{z},\mathbf{h})$ or $\bot$ to all parties; never reveal $c\mathbf{s}_{2}$ or individual $\mathbf{W}_{i}$
+Round 3.
+Receive $\{(\mathbf{z}_{i},\mathbf{V}_{i})\}_{i\in S}$.
+Aggregate: $\mathbf{z}=\textstyle\sum_{i}\lambda_{i}\mathbf{z}_{i}\bmod q$, $c\mathbf{s}_{2}=\textstyle\sum_{i}\mathbf{V}_{i}\bmod q$.
+Abort with $\bot$ if $\|\mathbf{z}\|_{\infty}\geq\gamma_{1}-\beta$ or $\|\mathsf{LowBits}(\mathbf{w}-c\mathbf{s}_{2},\,2\gamma_{2})\|_{\infty}\geq\gamma_{2}-\beta$.
+Otherwise compute $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\,\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},\,2\gamma_{2})$; output $(\tilde{c},\mathbf{z},\mathbf{h})$.
+Never reveal $c\mathbf{s}_{2}$ or individual $\mathbf{W}_{i}$.
 Theorem 4.21 (Profile P1 UC Security).
 If the TEE correctly implements $\mathcal{F}_{\mathsf{TEE}}$, then Protocol P1 UC-realizes the threshold signing functionality $\mathcal{F}_{\mathsf{ThreshSig}}$ against static adversaries corrupting up to T1T-1 signing parties.
-Proof 4.14.
-The proof is immediate by the ideal functionality abstraction. The simulator $\mathcal{S}$ operates as follows:
+Proof 4.22.
+We construct an explicit simulator $\mathcal{S}$ for each protocol round and prove indistinguishability.
+Simulator construction.
   1. 1.
-    $\mathcal{S}$ simulates honest parties’ Round 1 messages (commitments $\mathsf{Com}_{h}$) and Round 3 messages (responses $\mathbf{z}_{h}$, masked shares $\mathbf{V}_{h}$).
-Responses have high min-entropy by Theorem 4.4 (statistical, no computational assumptions).
+    Round 0 (DKG): $\mathcal{S}$ generates honest key shares $\mathbf{s}_{1,h},\mathbf{s}_{2,h}$ and nonce shares $\mathbf{y}_{h}$ inside the simulated $\mathcal{F}_{\mathsf{TEE}}$, following the real key DKG and nonce DKG protocols.
+Since DKG outputs to corrupted parties are determined by the $(T,N)$-Shamir sharing structure, $\mathcal{S}$ produces consistent evaluations $\{\mathbf{f}_{h}(j)\}_{j\in C}$.
   2. 2.
+    Round 1 (Commitments): $\mathcal{S}$ samples $\tilde{\mathbf{y}}_{h},\tilde{\mathbf{w}}_{h},\tilde{r}_{h}$ from the simulated DKG, computes the committed value $\mathsf{Com}_{h}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$ uniformly, and programs the random oracle so that $H(\mathsf{``com^{\prime\prime}}\|\tilde{\mathbf{y}}_{h}\|\tilde{\mathbf{w}}_{h}\|\tilde{r}_{h})=\mathsf{Com}_{h}$ (using the same domain separator as the real commitment scheme $\mathsf{Com}_{i}=H(\mathsf{``com^{\prime\prime}}\|\mathbf{y}_{i}\|\mathbf{w}_{i}\|r_{i})$).
+By commitment hiding (preimage resistance of $H$), the uniform $\mathsf{Com}_{h}$ is indistinguishable from real commitments.
+If the blame protocol requests an opening of $\mathsf{Com}_{h}$, $\mathcal{S}$ provides $(\tilde{\mathbf{y}}_{h},\tilde{\mathbf{w}}_{h},\tilde{r}_{h})$, which verifies correctly by the programmed RO.
+  3. 3.
 […]
-  3. 3.
-    $\mathcal{F}_{\mathsf{TEE}}$ computes the r0-check and hint internally, outputs only the final signature (or $\bot$).
-Since $\mathcal{F}_{\mathsf{TEE}}$ never reveals $c\mathbf{s}_{2}$ or individual $\mathbf{W}_{i}$, the responses $\mathbf{z}_{h}$ have high min-entropy (Theorem 4.4, statistical), and the adversary’s view contains only $(\mathbf{w}_{1},\tilde{c},\sigma)$—all implicit in the signature—the real and ideal executions are indistinguishable.
-The distinguishing advantage is bounded by $\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$ (for the mask-based values only; the response privacy is statistical and contributes negligible statistical distance).
+  4. 4.
+    Round 3 (Responses): $\mathcal{S}$ computes $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ from simulated DKG values.
+This is identical to the real computation (SD =0=0).
+Masked shares $\mathbf{V}_{h}=\lambda_{h}c\mathbf{s}_{2,h}+\mathbf{m}_{h}^{(s2)}$ are computed analogously.
+Indistinguishability.
+The adversary’s view in the real and simulated executions is indistinguishable for the following reasons:
+  • Response $\mathbf{z}_{h}$: Computed identically from simulated DKG values (SD =0=0).
+The simulator has access to the honest party’s key share $\mathbf{s}_{1,h}$ because it generated $\{\mathbf{s}_{1,j}\}$ during the simulated DKG (Round 0, setup step).
+The nonce share $\mathbf{y}_{h}$ has high conditional min-entropy by Theorem 4.7 (statistical); key privacy given $\sigma$ is conjectured to hold computationally under Module-LWE (open problem; see Remark 4.9).
+  • TEE outputs: $\mathcal{F}_{\mathsf{TEE}}$ hides all internal state (individual $\mathbf{W}_{i}$, $c\mathbf{s}_{2}$); only $(\mathbf{w}_{1},\tilde{c},\sigma)$ exit the enclave, all of which are implicit in the signature.
+  • Masks: Replacing honest-honest PRF terms with uniform introduces advantage $\leq\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
+Abort handling.
+The honest per-attempt success probability is $\approx 25\%$ (rejection sampling).
+After >33>33 consecutive failures, honest probability is $(0.75)^{33}<10^{-4}$, triggering the blame protocol (Section 6.1).
+The simulator handles aborts by forwarding $\bot$ from $\mathcal{F}_{\mathsf{TEE}}$.
+Total distinguishing advantage: $\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$ (nonce share privacy is statistical and contributes zero to the PRF-based advantage).
 Remark 4.23 (TEE Trust Assumption).
@@
 This requires extending the UC framework to handle timing.
-Definition 4.16 (Semi-Async Timing Model).
-The semi-async model defines the following timing structure:
-  • Pre-signing window $t_{0},t_{1}]$: Signers may precompute nonces $(y_{i},w_{i},\mathsf{Com}_{i})$ at any time
-  • Challenge broadcast $t_{2}$: Combiner derives challenge $c$ from aggregated commitments
-  • Response window $t_{2},t_{2}+\Delta]$: Signers respond within window $\Delta$ (e.g., 5 minutes)
-  • 2PC execution $t_{3},t_{4}]$: Synchronous between CPs only
-The adversary cannot control timing within honest parties’ windows but may observe message arrival times.
+Definition 4.24 (Semi-Async Timing Model (Formalized)).
+The semi-async model extends the standard UC framework with bounded-time windows. Execution proceeds in optimistic synchronous phases with timeout-based fallback:
+  1. 1.
+    Pre-signing window $t_{0},t_{1}]$: Signers may precompute nonces $(\mathbf{y}_{i},\mathbf{w}_{i},\mathsf{Com}_{i})$ at any time before $t_{1}$.
+This is an asynchronous phase; early precomputation is allowed but not required.
+  2. 2.
+    Challenge broadcast $t_{2}$: Combiner derives challenge $c$ from aggregated commitments and broadcasts $(c,\mu)$ to all signers at time $t_{2}$.
+This is a synchronous event (all parties receive the broadcast at the same logical time in the UC model).
+  3. 3.
+    Response window $t_{2},t_{2}+\Delta]$: Honest signers should respond within window $\Delta$ (e.g., 5 minutes for human-in-the-loop scenarios). Adversary scheduling rules:
+    • Honest party delivery: Messages from honest signers arriving within $t_{2},t_{2}+\Delta]$ are guaranteed to be delivered to the combiner.
+    • Adversarial delay: The adversary may delay delivery of honest parties’ messages by up to $\Delta$, but cannot prevent delivery entirely (eventual delivery assumption).
+    • Timeout: If fewer than $T$ responses arrive by $t_{2}+\Delta$, the combiner aborts the session and outputs $\bot$.
+This is not a security failure—it is an availability event (honest signers who responded late are still secure; no key material is leaked).
+    • Message reordering: The adversary may reorder messages from different parties arriving within the window, but cannot inject or modify honest parties’ messages (authenticated channels assumption).
+  4. 4.
+    2PC execution $t_{3},t_{4}]$: Synchronous 2PC between the two CPs (standard UC model applies; this phase has no timing relaxation).
+UC adversary capabilities.
+The adversary observes message arrival times (timestamps) within the response window but cannot control the content or the fact of delivery for honest parties.
+Replay attacks are prevented via unique session identifiers $\mathsf{sid}$ binding each response to $(c,\mu)$.
+The model is semi-asynchronous in the sense that signers do not need to coordinate their response times (unlike multi-round MPC), but the protocol remains UC-secure as long as $\geq T$ responses arrive within $\Delta$.
 Theorem 4.25 (Profile P3+ UC Security (Semi-Async)).
-Under the semi-async timing model, Protocol P3+ UC-realizes $\mathcal{F}_{\mathsf{ThreshSig}}$ against static adversaries corrupting up to T1T-1 signers and at most one CP, assuming:
+Under the semi-async timing model, Protocol P3+ UC-realizes $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ in the $(\mathcal{F}_{\mathsf{OT}},\mathcal{F}_{\mathsf{GC}})$-hybrid model against static adversaries corrupting up to T1T-1 signers and at most one CP, assuming:
   1. 1.
@@
     At least one CP remains honest
-Security bound: $\epsilon\leq\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
-The proof appears in Appendix P.
-The key insight is that pre-signing security follows from commitment hiding (signers commit to nonces before seeing the challenge), and the response window does not introduce new attack vectors because:
+Security bound: $\epsilon\leq\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{ext}}+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$, where $\epsilon_{\mathsf{ext}}\leq 2^{-B/2}$ is the LP07 cut-and-choose extraction error (for corrupted $\mathsf{CP}_{1}$).
+By UC composition (using the nonce DKG and signing sub-protocols as ideal functionalities), this implies that Protocol P3+ UC-realizes the full $\mathcal{F}_{\mathsf{ThreshSig}}$ under the same assumptions.
+Proof 4.26 (Proof sketch (full proof in Appendix P)).
+The simulator replaces OT and GC with ideal functionalities ($\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{GC}}$), then simulates pre-signing by sampling $\tilde{\mathbf{y}}_{i}$ and computing $\tilde{\mathbf{w}}_{i}=\mathbf{A}\tilde{\mathbf{y}}_{i}$ (Lemma P.8, SD =0=0), and simulates responses via $\mathbf{z}_{i}=\tilde{\mathbf{y}}_{i}+c\cdot\mathbf{s}_{1,i}$ (Lemma P.10, SD =0=0).
+The response window does not introduce new attack vectors because:
   1. 1.
@@
 UC security
-Table 4: Security profile comparison. P3+ = 1 signer step + 1 server round.
+UC privacy level Statistical $\approx 2^{-11}$ Statistical
+Table 4: Security profile comparison.
+P3+ = 1 signer step + 1 server round.
+UC privacy level for P2 refers to the r0-check simulation bound ($\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}\leq 2^{-64}+2^{-12}<2^{-11}$); P1/P3+ achieve statistical privacy via the Shamir nonce DKG (no computational assumptions).
+P2 mask-hiding privacy requires $|S\setminus C|\geq 2$ (Lemma 4.15); P1 and P3+ achieve nonce share privacy with $|S\setminus C|=1$ ($|S|=T$).
 5 Implementation and Benchmarks
@@
 We report averages over 5 independent runs, each consisting of 5 trials per configuration (25 total measurements per data point).
+Network Testing Setup.
+For the $N=3$, $T=2$ configuration, we deployed parties across 3 physical nodes connected via OpenVPN (local area network).
+This setup validates protocol correctness and measures actual network communication overhead in a distributed environment.
+For larger configurations ($N\geq 5$), we benchmarked on a single machine; reported timings for these configurations represent computational cost only, with network latency projected analytically based on round count and representative RTT values (see the network latency table in the companion supplement document).
 5.3 Performance Results
@@
 Values are averages over 5 independent runs, each with 5 trials per configuration.
+$N=3$: measured on 3-node OpenVPN testbed; $N\geq 5$: single-machine benchmarks (computational cost only; network overhead excluded).
 Due to rejection sampling variance, P2 exhibits high variance (up to $14\times$ between runs).
@@
 P3+ scales from 17 ms ($T=2$) to 94 ms ($T=32$), maintaining sub-100ms latency; the 2PC cost is largely independent of threshold.
+Figure 2 extends the benchmark to $T=128$, showing that P3+ remains sub-1 s through $T=64$ and sub-3 s through $T=128$, while P1 and P2 exceed 10 s at T=96T=96128128 due to Lagrange polynomial evaluation and MPC simulation overhead, respectively.
+[…]
+[…]
 5.4 Protocol Profile Comparison
@@
 Profile P2 (5-round MPC) achieves 21.5 ms online time for 3-of-5 through parallel EdaBit-based A2B conversion and constant-depth AND tree for comparison; the 5-round structure merges commit-then-open (Rounds 3–4 $\to$ Round 2) and parallelizes comparison (Rounds 6–7 $\to$ Round 4).
-Profile P3+ provides 22.1 ms latency with FROST-like workflow: signers precompute nonces offline (2.5 KB per set) and respond within a time window (default: 5 min), requiring only a single user interaction.
+Profile P3+ provides 22.1 ms latency with FROST-like workflow: signers precompute nonces offline (2.3 KB per set) and respond within a time window (default: 5 min), requiring only a single user interaction.
 With GC precomputation, online 2PC evaluation drops to $\approx$12 ms.
@@
 Table 7: Comparison with prior work.
-Times for prior schemes are from their original papers; our times are from our Rust implementation (ML-DSA-65, N=5N\!=\!5, T=3T\!=\!3).
+Times for prior schemes are taken from their original papers on potentially different hardware configurations; direct cross-scheme comparison is indicative only.
+Our times are from our Rust implementation on an Intel i7-12700H (ML-DSA-65, N=5N\!=\!5, T=3T\!=\!3).
 Our $\approx$21–45% success rate is comparable to ML-DSA’s expected $\approx$20–25%.
@@
 Analysis.
-The dPN25 compact scheme [dPN25] uses short-coefficient LSSS with binary reconstruction coefficients, achieving FIPS 204 compatibility but is limited to small thresholds ($T\leq 8$) due to the Ball-Çakan-Malkin bound [BCM21].
+The dPN25 compact scheme [dPN25] uses short-coefficient LSSS with binary reconstruction coefficients, achieving FIPS 204 compatibility but is limited to small thresholds ($T\leq 8$) in practice (binomial reconstruction coefficients exceed $q$ at $T\geq 26$; asymptotically, the Ball-Çakan-Malkin bound [BCM21] gives $\Omega(T\log T)$ share size).
 Noise flooding [BCDG21] masks large Lagrange coefficients with overwhelming statistical noise, enabling arbitrary thresholds but producing $5\times$ larger signatures that break FIPS 204 compatibility; its near-100% success rate comes from relaxed norm bounds.
@@
 Our scheme combines arbitrary thresholds with standard signature size and FIPS 204 verification compatibility.
-The observed $\approx$21–45% success rate is comparable to single-signer ML-DSA’s expected $\approx$20–25%, inherent to the FIPS 204 specification.
+The observed $\approx$21–45% success rate overlaps with single-signer ML-DSA’s expected $\approx$20–25%: the lower end (21%, $T=16$) is in line with single-signer, while the higher end (45%, $T=7$) reflects the Irwin-Hall distribution’s greater concentration near zero relative to a uniform nonce, placing more nonces within the acceptance region $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ (see the Irwin-Hall success rate analysis in the companion supplement).
+The wide range (2× variation) is dominated by rejection-sampling variance, not systematic performance degradation.
 5.6 Overhead vs. Standard ML-DSA
@@
 Profile P1-TEE scales from 4 ms ($T=2$) to 60 ms ($T=32$), approximately linear in $T$.
-Profile P3+ scales from 17 ms ($T=2$) to 94 ms ($T=32$), maintaining sub-100ms latency throughout.
+Profile P3+ scales from 17 ms ($T=2$) to 94 ms ($T=32$), maintaining sub-100ms per-valid-signature latency throughout (Table 5; Figure 2 shows per-attempt times from a separate extended benchmark, which may be up to $2\times$ higher due to rejection-sampling retries).
 Profile P2-MPC shows higher variance (21–194 ms) due to MPC overhead sensitivity to retry count, exceeding 100ms for $T\geq 16$.
 For larger thresholds ($T>32$), the dominant cost becomes polynomial operations ($O(T\cdot n\cdot\log n)$ per signature).
-Extrapolating from our measurements, $T=64$ would require approximately 100–200 ms for P1-TEE and 150–250 ms for P3+.
+Extended benchmarks (Figure 2) show P3+ achieves sub-1 s through $T=64$ and sub-3 s through $T=128$; P1 and P2 exceed 10 s at $T\geq 96$ due to Lagrange evaluation and MPC simulation overhead, respectively.
 5.8 Preprocessing Analysis
 All profiles benefit from preprocessing.
-The Shamir nonce DKG (Section 3.4) adds one offline round of $O(|S|^{2})$ point-to-point messages ($\approx$3.8 KB each for ML-DSA-65) before each signing session.
+The Shamir nonce DKG (Section 3.4) adds one offline round of $O(|S|^{2})$ point-to-point messages ($\approx$3.6 KB each for ML-DSA-65) before each signing session.
 In practice, multiple nonce sharings can be precomputed during idle time and consumed on demand, amortizing the DKG cost across signing sessions.
@@
 Beaver triples are pre-fetched per AND-tree level for cache efficiency.
-The main bottleneck remains MPC communication overhead (simulated as computation in our single-machine benchmark).
-In a distributed deployment, network latency would dominate, making preprocessing even more critical.
-Deployment Latency Estimates.
-Our benchmarks measure computation time on a single machine.
-In distributed deployments, end-to-end latency is approximately $t_{\mathsf{compute}}+\text{rounds}\times\text{RTT}$.
+The main bottleneck remains MPC communication overhead.
+For the $N=3$ OpenVPN testbed, measured round-trip latency averaged $\approx$2.1 ms (LAN setting).
+For larger configurations ($N\geq 5$) benchmarked on a single machine, network communication is not measured; in actual distributed deployments, network latency would dominate, making preprocessing even more critical.
+Network Latency Analysis.
+For single-machine benchmarks ($N\geq 5$), we project distributed deployment latency analytically as $t_{\mathsf{total}}\approx t_{\mathsf{compute}}+\text{rounds}\times\text{RTT}$, where $t_{\mathsf{compute}}$ is measured and RTT varies by deployment scenario.
 On a LAN (1 ms RTT), P1 takes $\approx$10 ms and P3+ $\approx$19 ms for $(3,5)$-threshold.
@@
 Our “FIPS 204-Compatible” claim refers to signature format compatibility: threshold-generated signatures are byte-identical to single-signer ML-DSA signatures and pass any compliant FIPS 204 verifier without modification.
-Full NIST ACVP Known-Answer-Test (KAT) certification for production deployment would additionally require implementing the deterministic encoding functions specified in FIPS 204 Section 7.2.
+Full NIST ACVP Known-Answer-Test (KAT) certification for production deployment would additionally require the deterministic encoding functions specified in FIPS 204 Section 7.2.
 5.13 Code Availability
@@
 Our protocol employs two distinct DKG protocols.
-The key DKG (executed once during setup) eliminates the trusted dealer assumption: each party contributes a random sharing of local secrets sampled from the short-coefficient distribution $\chi_{\eta}$, parties aggregate their shares, and Feldman commitments [Feldman87] provide robustness against malicious participants.
-The final secret $\mathbf{s}_{1}=\sum_{i}\mathbf{s}_{1}^{(i)}$ is unknown to any individual party.
+The key DKG (executed once during setup) replaces a fully trusted dealer with a distributed protocol: each party contributes a random sharing of local secrets, parties aggregate their shares, and Feldman commitments [Feldman87] allow every participant to verify that all shares lie on a consistent polynomial without revealing the secret.
+What Feldman VSS does and does not provide.
+Feldman VSS guarantees polynomial consistency: each party can verify that their share $\mathbf{s}_{1,h}=f(h)$ is consistent with the committed polynomial $\{g^{a_{0}},\ldots,g^{a_{T-1}}\}$.
+Crucially, this does not constrain the constant term $f(0)=\mathbf{s}_{1}$ to have short coefficients.
+The Shamir shares $f(h)$ for $h\geq 1$ lie in $\mathbb{Z}_{q}^{\ell}$ (large, as expected), which is fine for our protocol since the signing bound depends only on the reconstructed $\mathbf{s}_{1}$, not on individual share norms.
+The only constraint requiring smallness is that $\mathbf{s}_{1}=f(0)$ satisfies $\|\mathbf{s}_{1}\|_{\infty}\leq\eta=4$, which is needed for the rejection-sampling bound in the signature.
+Why enforcing short $\mathbf{s}_{1}$ without trust is hard.
+To verifiably guarantee $\|\mathbf{s}_{1}\|_{\infty}\leq\eta$ in a trustless DKG, parties would need to produce a zero-knowledge proof that the committed polynomial has a small constant term.
+Current lattice-based ZK range proofs for $\ell_{\infty}$ balls in $R_{q}$ carry significant communication overhead; a practical treatment is an open problem (see Section 7).
+We note that this limitation is shared by all lattice-based threshold signature schemes in the literature—including Raccoon [RaccoTeam23] and Dilithium-based constructions—that require small-coefficient secrets.
+Practical alternative.
+Profile P1 (Section 3.2) uses a TEE coordinator that runs ML-DSA key generation in a hardware-isolated environment and distributes Shamir shares directly.
+This provides the short-$\mathbf{s}_{1}$ guarantee at the cost of a hardware trust assumption (Intel SGX or equivalent), which is already required by P1’s security model.
+For deployments where hardware trust is acceptable—which covers a broad class of enterprise PKI scenarios—this is the recommended setup mechanism.
 For lattice-specific VSS/DKG, see [FS24, GJKR07].
 The nonce DKG (executed before each signing session) is our primary technical contribution.
-As described in Section 3.4, each party generates a degree-(T1)(T\!-\!1) polynomial whose higher-degree coefficients are sampled uniformly from $R_{q}^{\ell}$, producing a fresh Shamir sharing of the signing nonce.
+As described in Section 3.4, each party generates a degree-(T1)(T\!-\!1) polynomial whose higher-degree coefficients (degree $\geq 1$) are sampled uniformly from $R_{q}^{\ell}$ and whose constant term is sampled from the bounded nonce range $\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$, producing a fresh Shamir sharing of the signing nonce.
 This structural choice enables statistical share privacy (Theorem 4.7).
@@
 Triggering Conditions.
-  • Excessive aborts: More than $K=30$ consecutive failures (honest probability $(0.75)^{30}<10^{-4}$)
+  • Excessive aborts: More than $K$ consecutive failures, where $K$ is set based on the deployment’s per-attempt success rate $p$: $K=\lceil\log(10^{-4})/\log(1-p)\rceil$.
+For $p\geq 25\%$ (all configurations with $T\leq 21$), $K=33$ suffices (honest probability $(0.75)^{33}<10^{-4}$); for worst-case $T=32$ ($p\approx 21\%$), set $K=48$ to maintain the $10^{-4}$ guarantee.
   • Commitment mismatch: Verifiable inconsistency in hash values (immediate blame)
@@
     Malicious party is excluded from future sessions
-This preserves privacy: all sensitive values stay inside the enclave; only cheater identities are output.
+Security Analysis.
+The blame protocol achieves the following properties:
+  • Privacy: All sensitive values ($\mathbf{f}_{h}(x)$, $\mathbf{s}_{1,i}$, mask seeds) are transmitted only to the TEE coordinator and never revealed to other parties.
+The TEE outputs only the identity of misbehaving parties, preserving statistical privacy of honest parties’ shares (Theorem 4.7 continues to hold even after blame execution).
+  • Correctness: The TEE recomputes all protocol values deterministically from revealed inputs.
+A party is blamed if and only if its revealed inputs are inconsistent with its submitted protocol messages (e.g., $\mathbf{W}_{i}\neq\lambda_{i}\mathbf{A}\mathbf{y}_{i}+\mathbf{m}_{i}^{(w)}$), ensuring honest parties are never falsely accused.
+  • Trigger threshold: The $K=33$ consecutive-failure threshold is chosen such that the probability of triggering blame due to honest rejection sampling alone is $(1-p)^{33}<10^{-4}$ for $p\geq 25\%$ (ML-DSA-65 with $T\leq 21$).
+For larger $T$, set $K=\lceil\log(10^{-4})/\log(1-p)\rceil$ to maintain the false-positive rate below $10^{-4}$.
+  • UC composition: Under the TEE trust model ($\mathcal{F}_{\mathsf{TEE}}$ honest), the blame protocol UC-realizes an ideal functionality $\mathcal{F}_{\mathsf{Blame}}$ that takes inputs from all parties, identifies inconsistent parties via deterministic recomputation, and outputs cheater identities to all parties while keeping honest inputs private.
+Formal UC proof is deferred to the full version; the key observation is that the simulator can run the blame protocol honestly (since it has access to all corrupted parties’ inputs and can program the TEE’s output) and the view of honest parties consists only of the output (cheater identities), which is identically distributed in the real and ideal worlds.
+Blame Protocol (Profile P2).
+Profile P2 does not require a separate blame protocol: SPDZ MAC verification already provides identifiable abort.
+When a MAC check fails during the r0-check subprotocol, the SPDZ functionality $\mathcal{F}_{\mathsf{SPDZ}}$ outputs the identity of the party whose MAC is invalid (with probability $\geq 1-1/q\approx 1$ for $q=8380417$).
+The cheating party is immediately identified and excluded from future sessions.
+Note that the SPDZ MAC check catches input manipulation (wrong shares), but not wrong inputs submitted to $\mathcal{F}_{\mathsf{SPDZ}}$; a party that consistently submits out-of-range inputs to the edaBit protocol is caught by the cut-and-choose parameter $B=64$.
+Blame Protocol (Profile P3+).
+Profile P3+ uses 2PC, and blame attribution proceeds as follows when failures exceed the threshold $K$. Since CP1 does not hold key shares, blame is routed to CP2 (or a designated auditor):
+  1. 1.
+    CP2 requests each signer $i\in S$ to send: (a) their own nonce DKG polynomial $\mathbf{f}_{i}(x)$, and (b) all evaluations $\{\mathbf{f}_{j}(i)\}_{j\in S}$ they received during the DKG round.
+  2. 2.
+    CP2 reconstructs each party’s nonce share: $\mathbf{y}_{i}=\sum_{j\in S}\mathbf{f}_{j}(i)$.
+  3. 3.
+    Signer $i$ also sends to CP2 (over a confidential channel): their key share $\mathbf{s}_{1,i}$ and commitment randomness $r_{i}$.
+CP2 verifies: (a) $\mathsf{Com}_{i}=H(\texttt{"com"}\|\mathbf{y}_{i}\|\mathbf{w}_{i}\|r_{i})$ (commitment binding); (b) $\mathbf{w}_{i}=\mathbf{A}\mathbf{y}_{i}$ (nonce consistency); (c) $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ (response correctness).
+  4. 4.
+    Inconsistent signers are identified and excluded.
+Note on nonce privacy after blame.
+Revealing $\mathbf{f}_{i}(x)$ to CP2 discloses the constant term $\hat{\mathbf{y}}_{i}$ and all higher coefficients for the session under investigation, so per-session nonce privacy is intentionally forfeited during blame attribution.
+Theorem 4.7 (statistical nonce share privacy) holds for all sessions not under blame investigation; for the session under review, nonce privacy is deliberately sacrificed for accountability.
+Key privacy is maintained because $\mathbf{s}_{1,i}$ is sent only to CP2 (trusted auditor), not broadcast.
+After any blame invocation, a key-refresh step is recommended to limit exposure.
+Remark 6.1 (Blame leakage across sessions).
+Each blame invocation leaks at most $n\ell\cdot\log_{2}(2\lfloor\gamma_{1}/|S|\rfloor+1)\approx 25{,}600$ bits of information (the full session nonce polynomial $\mathbf{f}_{i}(x)$, with coefficients bounded in $\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]$).
+Because each session uses freshly sampled higher polynomial coefficients, successive blame invocations on different sessions are statistically independent: the nonce polynomial for session $\tau$ reveals no information about session $\tau^{\prime}$.
+There is therefore no cumulative cross-session leakage from the nonce DKG component alone.
+However, once $\mathbf{s}_{1,i}$ is disclosed to CP2 (for blame verification), CP2 is effectively a key-holder.
+A mandatory key-refresh should follow the first blame invocation that requires revealing key material; subsequent blame invocations may reuse the refreshed key under the same analysis.
 6.3 Profile P2: Fully Distributed Signing
 For deployments where TEE trust is unacceptable, we provide a fully distributed protocol using secure multi-party computation.
-The core challenge is evaluating the r0-check predicate without reconstructing $c\mathbf{s}_{2}$ at any single point—since virtually all ML-DSA challenges $c$ are invertible in $R_{q}$, leaking $c\mathbf{s}_{2}$ would compromise the secret key.
+The core challenge is evaluating the r0-check predicate without reconstructing $c\mathbf{s}_{2}$ at any single point—since virtually all ML-DSA challenges $c$ are invertible in $R_{q}$ (a standard heuristic confirmed experimentally; see Remark Q.39), leaking $c\mathbf{s}_{2}$ would compromise the secret key.
 The r0-Check Problem.
@@
   1. 1.
-    Combiner-mediated commit-then-open (Rounds 3-4 $\to$ Round 2): All parties send $(Com_{i},masked_{i},rand_{i})$ to a designated combiner, who releases all openings simultaneously after collecting from all parties.
-SPDZ MAC check catches any tampering by the combiner.
+    Combiner-mediated commit-then-open (base-protocol Rounds 2–4 $\to$ 5-round Round 2): All parties send $(\mathsf{Com}_{i},\mathsf{masked}_{i},\mathsf{rand}_{i})$ to a designated combiner (one of the $N$ signing parties, potentially corrupted), who releases all openings simultaneously after collecting from all parties.
+SPDZ MAC checks catch any tampering by the combiner, since each share is authenticated with a global MAC key unknown to individual parties.
+This collapses three base-protocol rounds (masked-value exchange, commitment, and opening) into a single combiner-mediated round, saving two rounds.
   2. 2.
-    Constant-depth comparison (Rounds 6-7 $\to$ Round 4): All 1536 coefficient comparisons are evaluated in parallel (Layer 1), followed by an AND-tree aggregation (Layer 2).
-Using Beaver triples, all $(d,e)$ values can be opened in a single round.
+    Constant-depth comparison (base-protocol Rounds 6–7 $\to$ 5-round Round 4): All $nk=1536$ coefficient comparisons are evaluated in parallel using $1536$ Beaver triples; all $(d,e)$ openings are batched into a single network round (Round 4).
+After receiving opened $(d,e)$ values, each party locally computes comparison bits $b_{j}\in\{0,1\}$ and then locally aggregates $\langle\sum_{j}b_{j}\rangle$ via linear sharing (no additional network round).
+The AND of all 1536 comparison results reduces to: $\bigwedge_{j}b_{j}=1\iff\sum_{j}b_{j}=nk=1536$, which is a constant-depth (depth-1) check: opening $\langle\sum_{j}b_{j}\rangle$ in Round 5 (“Output”) and comparing to $1536$ constitutes a single network round with no sequential Beaver-triple dependencies.
 UC Security of 5-Round Optimization.
-The 5-round protocol is obtained from the 8-round base protocol (Appendix H) via two standard MPC composition steps.
-First, combiner-mediated commit-then-open merges sequential commit and open rounds (Rounds 3–4 $\to$ Round 2): parties send $(\mathsf{Com}_{i},\mathsf{masked}_{i},\mathsf{rand}_{i})$ to a designated combiner who releases all openings simultaneously; the combiner is an ordinary party, and SPDZ MAC checks detect any tampering, so this is UC-equivalent to a sequential broadcast [CLOS02].
-Second, parallel Beaver-triple-based comparison merges the sequential comparison rounds (Rounds 6–7 $\to$ Round 4): all 1536 coefficient comparisons execute in parallel (Layer 1), followed by a single-round AND-tree aggregation (Layer 2).
-Each optimization replaces a sub-protocol with a functionally equivalent variant; by the UC composition theorem [Canetti01], the composed 5-round protocol UC-realizes the same ideal functionality.
-The hybrid argument in Appendix H ($\mathsf{H}_{0}$$\mathsf{H}_{6}$) carries over with the round assignment remapped, since each hybrid step depends on the functionality of each sub-protocol (SPDZ, edaBits, binary circuits), not on its round structure.
-The total security loss remains $\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$.
+The 5-round protocol is obtained from the 8-round base protocol (Appendix H) via two standard MPC composition steps: Optimization 1 saves 2 rounds (collapsing three base-protocol rounds into one) and Optimization 2 saves 1 round (collapsing two parallel comparison rounds into one), giving 821=58-2-1=5 rounds total.
+Optimization 1 (Combiner-mediated commit-then-open, base-protocol Rounds 2–4 $\to$ 5-round Round 2, saving 2 rounds): Parties send $(\mathsf{Com}_{i},\mathsf{masked}_{i},\mathsf{rand}_{i})$ to a designated combiner (one of the $N$ parties, treated as potentially corrupted) who releases all openings simultaneously.
+This is UC-equivalent to a sequential broadcast [CLOS02]: if the combiner withholds or alters any opening, the SPDZ MAC check — which authenticates each share with a global key unknown to the combiner — detects the tampering with probability $1-\epsilon_{\mathsf{SPDZ}}$.
+The simulator for this merged round constructs the same view as the 2-round version (send $\to$ open); leakage to the adversary is unchanged since only MAC-authenticated values are released.
+Confidentiality of the masked values $\mathsf{masked}_{i}$ from the combiner is protected by the pairwise PRF masks (Lemma 4.15), which ensure that even a corrupted combiner observing all $\mathsf{masked}_{i}$ cannot recover any party’s contribution, provided $|S\setminus C|\geq 2$.
+Optimization 2 (Constant-depth comparison, base-protocol Rounds 6–7 $\to$ 5-round Round 4): All $nk=1536$ Beaver-triple comparisons are opened simultaneously in a single network round (Round 4).
+There are no sequential dependencies among the $1536$ Beaver triples: each triple $(d_{j},e_{j})=(\alpha_{j}\oplus\langle x_{j}\rangle,\,\beta_{j}\oplus\langle y_{j}\rangle)$ depends only on the party’s local share of $\langle x_{j}\rangle,\langle y_{j}\rangle$ and the preprocessed triple $(\langle\alpha_{j}\rangle,\langle\beta_{j}\rangle,\langle\gamma_{j}\rangle)$ with $\gamma_{j}=\alpha_{j}\wedge\beta_{j}$, none of which depend on other comparisons.
+After receiving all $(d_{j},e_{j})$ in Round 4, parties locally compute the comparison output bit $b_{j}=\gamma_{j}\oplus(d_{j}\wedge\beta_{j})\oplus(e_{j}\wedge\alpha_{j})\oplus(d_{j}\wedge e_{j})$ (standard Beaver-triple AND evaluation; the triple components $\alpha_{j},\beta_{j},\gamma_{j}$ are distinct from the output bit $b_{j}$) and aggregate $\langle B\rangle=\sum_{j}\langle b_{j}\rangle$ (purely local linear operation, no network round).
+The AND of all $1536$ comparison bits satisfies $\bigwedge_{j}b_{j}=1\iff B=nk$; opening $\langle B\rangle$ in Round 5 (“Output”) is a single network round.
+The AND aggregation therefore requires no sequential AND-gate rounds (the AND tree is replaced by a single linear aggregation + one opening round, with no Beaver triples).
+By the UC composition theorem [Canetti01], each merged step preserves the ideal functionality’s input-output behavior and leakage profile.
+The simulator for the 5-round protocol is constructed by composing the simulators for each sub-protocol (SPDZ, edaBits, binary circuits) with the round assignment remapped; each simulator step depends on the functionality, not the round structure.
+The hybrid argument in Appendix H ($\mathsf{H}_{0}$$\mathsf{H}_{6}$) carries over directly.
+Since the edaBits functionality is used as a black box (its output distribution is indistinguishable from ideal, at cost $\epsilon_{\mathsf{edaBits}}$) before the SPDZ multiplication gate, there is no interaction between the two error terms, and the total security loss remains $\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$.
 Comparison with Prior MPC Approaches.
-  • Bienstock et al. [BDEPT25]: 37–136 online rounds, honest majority
+  • Bienstock et al. [BdCE25]: 23–79 online rounds (per [BdCE25]), honest majority
   • Celi et al. [CDENP26]: 6 online rounds, but restricted to $T\leq 6$
-  • Our P2: 5 online rounds, dishonest majority, arbitrary $T$
-Profile P2 thus achieves dishonest-majority security with arbitrary thresholds in constant rounds.
+  • Our P2: 5 online rounds, dishonest majority (unforgeability), arbitrary $T$
+Profile P2 thus achieves dishonest-majority unforgeability with arbitrary thresholds in constant rounds.
+Note: commitment and r0-check mask privacy additionally requires $|S\setminus C|\geq 2$ (Lemma 4.15), meaning at most T2T-2 signing-set corruptions are tolerated for full mask privacy when $T\geq 3$.
 Communication Complexity.
@@
 Theorem 6.3 (Profile P2 UC Security).
-Protocol $\Pi_{r_{0}}$ UC-realizes $\mathcal{F}_{r_{0}\text{-check}}$ in the $(\mathcal{F}_{\mathsf{SPDZ}},\mathcal{F}_{\mathsf{edaBits}},\mathcal{F}_{\mathsf{Binary}})$-hybrid model against static malicious adversaries corrupting up to N1N-1 parties, with security-with-abort.
+Protocol $\Pi_{r_{0}}$ UC-realizes $\mathcal{F}_{r_{0}\text{-check}}$ against static malicious adversaries (corrupting up to N1N-1 parties) in the $(\mathcal{F}_{\mathsf{SPDZ}},\allowbreak\mathcal{F}_{\mathsf{edaBits}},\mathcal{F}_{\mathsf{Binary}})$-hybrid model, with security-with-abort.
 Concretely:
-$|\Pr[\mathsf{REAL}_{\Pi_{r_{0}},\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}\text{-check}},\mathcal{S},\mathcal{Z}}=1]|\leq\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$
-for ML-DSA-65 parameters, where $\epsilon_{\mathsf{SPDZ}}\leq O(m)/q<2^{-12}$ with $m=nk=1536$ coefficients and $\epsilon_{\mathsf{edaBits}}<2^{-12}$ with cut-and-choose parameter $B=64$.
-Proof 6.3.
+$\bigl|\Pr[\mathsf{REAL}_{\Pi_{r_{0}},\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}\text{-check}},\mathcal{S},\mathcal{Z}}=1]\bigr|$
+$\quad\leq\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$
+for ML-DSA-65 parameters, where $\epsilon_{\mathsf{SPDZ}}\leq m/q<2^{-12}$ with $m=nk=1536$ coefficients ($1536/8380417\approx 2^{-12.4}$), and $\epsilon_{\mathsf{edaBits}}\leq m/q+2^{-B}<2^{-12}$ where $B=64$ makes the cut-and-choose term $2^{-64}$ negligible (the dominant term is again $m/q$ from the internal SPDZ MAC check within edaBits generation).
+Proof 6.4.
 We prove via six hybrids $\mathsf{H}_{0}$ (real) to $\mathsf{H}_{6}$ (ideal):
   1. 1.
-    $\mathsf{H}_{1}$: Replace SPDZ with $\mathcal{F}_{\mathsf{SPDZ}}$ (loss: $\epsilon_{\mathsf{SPDZ}}\leq O(m)/q<2^{-12}$)
+    $\mathsf{H}_{1}$: Replace SPDZ with $\mathcal{F}_{\mathsf{SPDZ}}$ (loss: $\epsilon_{\mathsf{SPDZ}}\leq m/q<2^{-12}$)
   2. 2.
@@
   5. 5.
-    $\mathsf{H}_{5}$: Backward-simulate AND tree given result bit (loss: 0)
+    $\mathsf{H}_{5}$: Backward-simulate AND tree given result bit: for each comparison bit $b_{j}$, the simulator samples honest additive shares uniformly subject to summing to the known $b_{j}$ value; SPDZ additive shares are uniform conditioned on their sum, so this is statistically identical to the real distribution (loss: 0)
   6. 6.
@@
 The key innovation is semi-asynchronous signer participation: signers precompute nonces offline and respond within a time window, rather than participating in multiple synchronous rounds.
+We use the term semi-asynchronous to denote that signers need not be online simultaneously for interactive rounds, but must respond within a bounded validity window $\Delta$; the 2PC subprotocol between the two CPs executes synchronously independently of the signers.
 Semi-Async Timing Model.
@@
   • Public input: $(\mathbf{A},c,\mathbf{t}_{1},\mathbf{z})$ (known to both parties and the combiner)
-  • Private input: Receive $(\mathbf{w},\mathbf{S}_{1})$ from $\mathsf{CP}_{1}$ and $\mathbf{S}_{2}$ from $\mathsf{CP}_{2}$
+  • Private input: Receive $(\mathbf{w},\mathbf{S}_{1})$ from $\mathsf{CP}_{1}$ and $\mathbf{S}_{2}$ from $\mathsf{CP}_{2}$, where $\mathbf{w}=\mathbf{A}\mathbf{y}$ is the full (unmasked) commitment vector (not just $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{w},2\gamma_{2})$; $\mathbf{w}$ is passed as $\mathsf{CP}_{1}$’s private input by the combiner), and the private shares satisfy $\mathbf{S}_{1}+\mathbf{S}_{2}=c\mathbf{s}_{2}\bmod q$ (arbitrary additive split; in the concrete protocol, $\mathbf{S}_{j}=\sum_{i\text{ assigned to }\mathsf{CP}_{j}}\lambda_{i}\cdot c\mathbf{s}_{2,i}$)
   • Compute: $\mathbf{w}^{\prime}=\mathbf{w}-\mathbf{S}_{1}-\mathbf{S}_{2}\bmod q$;  $\mathsf{result}=\bigwedge_{j}(|\mathbf{w}^{\prime}_{j}|_{q}<\gamma_{2}-\beta)$.
@@
 Concretely:
-$|\Pr[\mathsf{REAL}_{\Pi,\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}}^{\mathsf{2PC}},\mathcal{S},\mathcal{Z}}=1]|\leq\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}<2^{-\kappa}$
-where $\kappa$ is the computational security parameter.
-Proof 6.6.
+$|\Pr[\mathsf{REAL}_{\Pi,\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}}^{\mathsf{2PC}},\mathcal{S},\mathcal{Z}}=1]|\leq\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{ext}}<2^{-\kappa}$
+where $\kappa$ is the computational security parameter and $\epsilon_{\mathsf{ext}}\leq 2^{-B/2}$ is the LP07 cut-and-choose extraction error when $\mathsf{CP}_{1}$ is corrupted.
+Proof 6.7.
 We reduce to the UC security of Yao’s garbled circuits [Yao86, LP07, BHR12].
@@
 Conclusion.
-Total distinguishing advantage: $\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}<2^{-\kappa}$.
+Total distinguishing advantage: $\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{ext}}<2^{-\kappa}$ (where $\epsilon_{\mathsf{ext}}\leq 2^{-B/2}$ from LP07 cut-and-choose extraction).
 Corollary 6.8 (Full P3+ Security).
@@
   3. 3.
-    Privacy: Response shares $\mathbf{z}_{i}$ enjoy statistical privacy via Shamir nonce DKG (Theorem 4.4); commitment and r0-check values hidden by pairwise masks; 2PC reveals only the result bit
+    Privacy: Nonce shares $\mathbf{y}_{i}$ enjoy statistical privacy via Shamir nonce DKG (Theorem 4.7); key privacy given $\sigma$ is conjectured to hold computationally under Module-LWE (open problem; see Remark 4.9); commitment and r0-check values hidden by pairwise masks; 2PC reveals only the result bit
+Proof 6.9.
+(1) UC security follows from Theorem 6.6 via the UC composition theorem [Canetti01].
+(2) EUF-CMA security: by Lemma 3.6, masks cancel in the aggregate, so $(\tilde{c},\mathbf{z},\mathbf{h})$ has identical structure to single-signer ML-DSA; Theorem 4.2 applies.
+(3) Privacy: nonce shares $\mathbf{y}_{i}$ have statistical privacy via Theorem 4.7; key privacy given $\sigma$ is conjectured to hold computationally under Module-LWE (open problem; Remark 4.9—Theorem 4.7 establishes only nonce share privacy, not key privacy from $\sigma$); masked values $\mathbf{W}_{i}$, $\mathbf{V}_{i}$ are sent to the combiner (trusted under 1-of-2 CP honest assumption) or protected by mask hiding (Lemma 4.15) when $|S\setminus C|\geq 2$.
 Performance.
@@
 P3+ 2 7.4
-P3+ logical rounds = 1 signer step + 1 server round. Expected total assumes $\approx$30% per-attempt success rate (3.7 attempts for P3+, 3.0 for P2).
+P3+ logical rounds = 1 signer step + 1 server round. Expected total assumes 27–33% per-attempt success rate depending on configuration (3.7 attempts for P3+, 3.0 for P2).
 Comparison with Concurrent Work.
@@
 ML-KEM Integration.
-Pairwise seeds can be established via ML-KEM [FIPS203] for full post-quantum security with forward secrecy.
+Pairwise seeds can be established via ML-KEM [FIPS203] for full post-quantum security (static pairwise seeds; per-session forward secrecy requires ephemeral per-round encapsulation).
 7 Conclusion
@@
 Concurrent work includes Bienstock et al.
-[BDEPT25] (arbitrary $T$, UC security, but honest-majority and 37–136 rounds), Celi et al.
+[BdCE25] (arbitrary $T$, UC security, but honest-majority and 23–79 rounds per [BdCE25]), Celi et al.
 [CDENP26] (dishonest-majority, 6 rounds, but $T\leq 6$), and Trilithium [Trilithium25] (UC-secure, but 2-party only).
@@
 Our primary technical contribution is Shamir nonce DKG: parties jointly generate the signing nonce via a distributed key generation protocol, producing a degree-(T1)(T\!-\!1) Shamir sharing that structurally matches the long-term secret.
-Because the adversary observes only T1T-1 evaluations of each honest party’s degree-(T1)(T\!-\!1) polynomial, the remaining degree of freedom provides an approximate one-time pad: the honest party’s response $\mathbf{z}_{h}$ has conditional min-entropy exceeding $5\times$ the secret key entropy for $|S|\leq 17$, requiring no computational assumptions (Theorem 4.4).
+Because the adversary observes only T1T-1 evaluations of each honest party’s degree-(T1)(T\!-\!1) polynomial, the remaining degree of freedom provides an approximate one-time pad: the honest party’s nonce share $\mathbf{y}_{h}$ has conditional min-entropy exceeding $5\times$ the secret key entropy for $|S|\leq 17$, requiring no computational assumptions (Theorem 4.7).
+Key privacy given the public signature relies on the same lattice hardness as single-signer ML-DSA (computational, informal; formal reduction is an open problem for ML-DSA-based constructions, unresolved even in the single-signer setting).
 The bounded-nonce constraint prevents perfect uniformity (Remark 4.10), but the min-entropy guarantee is more than sufficient for practical security.
-In coordinator-based profiles (P1, P3+), this eliminates the two-honest requirement for response privacy: signing sets of size $|S|\geq T$ suffice, matching the information-theoretic minimum for Shamir reconstruction.
+In coordinator-based profiles (P1, P3+), this eliminates the two-honest requirement for nonce share privacy: signing sets of size $|S|\geq T$ suffice, matching the minimum required by Shamir’s $(T,N)$ threshold property for reconstruction.
 Individual masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator (not broadcast), preventing a key-recovery attack via the injective matrix $\mathbf{A}$ (Remark 4.1).
 Commitment and r0-check mask hiding retains $|S\setminus C|\geq 2$ (Lemma 4.15), which applies to P2’s broadcast model.
-As a secondary technique, we introduce pairwise-canceling masks adapted from ECDSA [CGGMP20] to the lattice setting, used for commitment values $\mathbf{W}_{i}$ and r0-check shares $\mathbf{V}_{i}$.
+As a secondary technique, we introduce pairwise-canceling masks (adapted from ECDSA [CGGMP20]) to the lattice setting, used for commitment values $\mathbf{W}_{i}$ and r0-check shares $\mathbf{V}_{i}$.
 This addresses three challenges absent in ECDSA: rejection sampling bounds, r0-check key leakage, and Irwin-Hall nonce security.
+The Irwin-Hall security gap of prior approaches is resolved by direct shift-invariance (Corollary Q.24): the EUF-CMA loss is <0.013<0.013 bits per signing session for $|S|\leq 33$ (total: $0.013\cdot q_{s}$ bits), giving a proven security bound of $\geq 96-0.006\cdot q_{s}$ bits for $|S|\leq 17$ (non-vacuous for $q_{s}<16{,}000$), remaining below $1$ bit per session for all $|S|\leq 2{,}584$.
 We instantiate these techniques in three deployment profiles with complete UC proofs.
@@
 Profile P3+ (Semi-Async 2PC) achieves 2 logical rounds (22 ms for 3-of-5) with UC security under a 1-of-2 CP honest assumption (Theorem 6.6).
-Our optimized Rust implementation demonstrates practical efficiency: P1 and P3+ achieve 4–94 ms for configurations from $(2,3)$ to $(32,45)$ threshold; P2 ranges from 21–194 ms, trading latency for eliminating hardware trust.
-Success rates of $\approx$21–45% are comparable to single-signer ML-DSA.
-This represents a $4.8\times 10^{10}$ improvement over naive threshold approaches for $T=16$.
+Our optimized Rust implementation demonstrates practical efficiency: P1 achieves 4–60 ms and P3+ achieves 17–94 ms for configurations from $(2,3)$ to $(32,45)$ threshold; P2 ranges from 21–194 ms, trading latency for eliminating hardware trust.
+Success rates of $\approx$21–45% (Table 5) are comparable to single-signer ML-DSA.
 All signatures are syntactically identical to single-signer ML-DSA and pass standard FIPS 204 verification without modification.
@@
 Each deployment profile involves different trust assumptions: Profile P1 requires TEE/HSM trust, Profile P2 eliminates hardware trust at the cost of 5 MPC rounds, and Profile P3+ requires 1-of-2 CP honesty with 2 logical rounds.
-The nonce DKG adds one offline preprocessing round of $O(N^{2})$ point-to-point messages ($\approx$3.8 KB each) before each signing session; while this can be amortized by precomputing nonce batches during idle time, it does increase protocol complexity relative to purely additive nonce sharing.
-The aggregated nonce follows an Irwin-Hall rather than uniform distribution; while the per-coordinate Rényi divergence is small ($R_{2}^{\epsilon,\mathsf{coord}}\leq 1.003$ for $|S|\leq 17$), the full-vector composition $(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$ grows to $\approx 2^{5.4}$ for $|S|=17$ ($\approx$3-bit loss) and $\approx 2^{75}$ for $|S|=33$ ($\approx$38-bit loss); see Theorem 4.11.
+The nonce DKG adds one offline preprocessing round of $O(N^{2})$ point-to-point messages ($\approx$3.6 KB each) before each signing session; while this can be amortized by precomputing nonce batches during idle time, it does increase protocol complexity relative to purely additive nonce sharing.
+The aggregated nonce follows an Irwin-Hall rather than uniform distribution.
+By the direct shift-invariance analysis (Corollary Q.24), the per-coordinate chi-squared divergence between real and simulated nonces is $\chi^{2}(\mathsf{IH}+\beta\|\mathsf{IH})<7\times 10^{-6}$ for $|S|\leq 17$ and $1.4\times 10^{-5}$ for $|S|\leq 33$, giving a full-vector EUF-CMA security loss <0.013<0.013 bits per signing session for $|S|\leq 33$ (total: $0.013\cdot q_{s}$ bits)—at most $13$ bits for $q_{s}\leq 1{,}000$ (or <7<7 bits in the $|S|\leq 17$ regime using the <0.007<0.007 per-session bound), well below the 192-bit security target.
 For the pairwise masks used on commitment and r0-check values, $|S\setminus C|\geq 2$ honest parties are needed.
-In coordinator-based profiles (P1, P3+), masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator rather than broadcast, so $|S\setminus C|=1$ suffices for response privacy; in P2’s broadcast model, the existing mask-hiding condition $|S\setminus C|\geq 2$ applies to all values.
+In coordinator-based profiles (P1, P3+), masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator rather than broadcast, so $|S\setminus C|=1$ suffices for nonce share privacy; in P2’s broadcast model, the existing mask-hiding condition $|S\setminus C|\geq 2$ applies to all values.
+This work focuses on ML-DSA-65 (NIST Level 3); extending to ML-DSA-44 ($\gamma_{1}=2^{17}$, $\beta=\tau\eta=78$) and ML-DSA-87 ($k=8,\ell=7$, $\gamma_{1}=2^{19}$, $\beta=120$) is straightforward in principle—ML-DSA-44 loosens the chi-squared bound by ${\approx}2.5\times$ relative to ML-DSA-65 (smaller $\gamma_{1}$ increases the ratio $\beta/\gamma_{1}$) but retains <0.04<0.04 bits of EUF-CMA loss per session for $|S|\leq 33$, while ML-DSA-87’s smaller $\beta$ at the same $\gamma_{1}$ tightens the bound by ${\approx}2.7\times$ (<0.007<0.007 bits per session for $|S|\leq 33$); the latter also expands the r0-check circuit from 1536 to 2048 coefficients; a full benchmark across all three levels remains to be conducted.
 Finally, we assume a synchronous model where all parties are online simultaneously during signing (though P3+ relaxes this to semi-asynchronous participation).
@@
 [CGGMP20] using UC-secure commitments; whether a similar approach suffices in the lattice setting, where the polynomial structure imposes additional algebraic constraints, is unclear.
+Separately, a proof in the quantum random oracle model (QROM) would strengthen the EUF-CMA bound against quantum adversaries making superposition queries to the challenge hash—an upgrade shared with single-signer FIPS 204 [FIPS204].
 A related direction concerns the strength of the abort guarantee.
@@
 Such an upgrade would be particularly relevant for custody and government applications where signing requests must not be silently suppressed.
+A formal security reduction for key privacy—the property that a signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ does not reveal $(\mathbf{s}_{1},\mathbf{s}_{2})$ beyond the public key $\mathbf{t}_{1}$—remains open even for single-signer FIPS 204 ML-DSA and is inherited by all threshold variants.
+In the threshold setting, additional leakage vectors arise: partial responses $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ and masked commitments $\mathbf{W}_{i}$ are visible to the coordinator or broadcast to the signing group.
+Theorem 4.7 establishes that nonce shares $\mathbf{y}_{h}$ retain high conditional min-entropy given the adversary’s view, bounding the information about $\mathbf{s}_{1,h}$ leaked by any single partial response; however, a formal reduction showing that the aggregate response $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}$ and hint vector $\mathbf{h}$, over polynomially many queries with Irwin-Hall nonces and Shamir share structure, do not expose $\mathbf{s}_{1}$ to a lattice adversary remains to be constructed.
 Our appendix sketches a share refresh mechanism based on adding shares of zero, but a complete treatment of proactive threshold ML-DSA requires addressing several interacting concerns: maintaining signing availability during the refresh epoch, ensuring consistency at epoch boundaries (parties must agree on which shares are current), and formally modeling a mobile adversary that corrupts different subsets in different epochs subject to a per-epoch threshold.
@@
 In particular, if a party’s long-term share $\mathbf{s}_{1,i}$ is refreshed between the nonce DKG and the signing round, the Lagrange reconstruction $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}$ may fail unless both the nonce and secret shares are refreshed consistently within the same epoch.
+Decentralized Key Generation.
+Our construction assumes a trusted dealer for key setup: the dealer samples the short-coefficient signing key $(\mathbf{s}_{1},\mathbf{s}_{2})\leftarrow\chi^{\ell}\times\chi^{k}$, computes the public key $\mathbf{t}=\mathbf{A}\mathbf{s}_{1}+\mathbf{s}_{2}$, and distributes Shamir shares $\mathbf{s}_{1,i}$, $\mathbf{s}_{2,i}$ directly.
+Fully decentralized key generation for ML-DSA requires verifiably short-coefficient Shamir sharing: each party must prove $\|\mathbf{s}_{1,h}\|_{\infty}\leq\eta$ without reconstruction.
+Standard Feldman VSS does not apply since Shamir shares of $\eta$-small secrets lie in $\mathbb{Z}_{q}$ and reveal no information about the coefficient bound; lattice-based zero-knowledge techniques—such as BDLOP commitments with norm proofs [BDLOP18]—offer a path forward.
+We leave a complete treatment to future work.
 On the protocol side, three open problems stand out.
-The most practically significant concerns eliminating the Irwin-Hall security gap for large signing sets.
-Our proven security bound degrades with $|S|$ because the aggregated nonce $\mathbf{y}=\sum_{i}\lambda_{i}\mathbf{y}_{i}$ follows an Irwin-Hall distribution rather than the uniform distribution assumed by FIPS 204: the full-vector Rényi divergence $(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ grows from $\approx 2^{5.4}$ ($|S|=17$, $\approx$3-bit security loss) to $\approx 2^{75}$ ($|S|=33$, $\approx$38-bit loss).
-For practical deployments with $|S|\leq 17$—covering the vast majority of HSM clusters, custody quorums, and certificate authority configurations—the proven bound provides full 128-bit security.
-For larger committees, the distribution-preserving rejection sampling technique of Raccoon [Raccoon2024] offers a principled solution: after computing the aggregated nonce, parties collectively evaluate a rejection function that flattens the Irwin-Hall distribution to uniform, yielding $R_{2}^{\mathsf{vec}}=1$ (no security loss) at the cost of an additional interactive rejection round and a reduced per-attempt success rate.
-The key technical challenge in the threshold setting is that $\mathbf{y}$ is secret-shared—no single party knows the full nonce—so the rejection decision must be computed via MPC over the shares, adding communication complexity proportional to the rejection function’s circuit size.
-Whether a lightweight, non-interactive variant exists (e.g., each party independently rejecting based on local share statistics) remains open.
-Second, this work focuses exclusively on ML-DSA-65 (NIST Security Level 3).
-The other two standardized parameter sets present distinct challenges: ML-DSA-44 (Level 2) has a smaller masking range $\gamma_{1}=2^{17}$, which tightens the per-coordinate Rényi bound by a factor of $\approx 16$, potentially pushing $R_{2}^{\epsilon,\mathsf{coord}}$ above $1.05$ for moderate signing sets and making the full-vector composition correspondingly more severe.
-Conversely, ML-DSA-87 (Level 5) uses $k=8,\ell=7$, increasing the r0-check circuit from 1536 to 2048 coefficients and the nonce DKG message size from 3.8 KB to 7.2 KB.
-A systematic parameter study with benchmarks across all three levels would be valuable for deployment planning.
-Third, in P2’s broadcast model, the pairwise-canceling masks on $\mathbf{W}_{i}$ and $\mathbf{V}_{i}$ require $|S\setminus C|\geq 2$ for computational hiding (Lemma 4.9).
+Eliminating offline nonce preprocessing.
+The Shamir nonce DKG requires one offline preprocessing round of $O(N^{2})$ messages per signing session, structurally necessary because nonce shares must be committed before the challenge hash is computed to prevent adversarial rewinding.
+A natural question is whether this phase can be merged into the online protocol: a combined first round where parties simultaneously distribute nonce shares and broadcast the nonce commitment $\mathbf{w}_{1}=\mathsf{HighBits}(\mathbf{A}\sum_{i}\hat{\mathbf{y}}_{i},2\gamma_{2})$ would eliminate the offline cost at the price of one additional online round.
+Whether such a per-session DKG preserves statistical nonce privacy against an adversary who aborts after observing first-round messages—and whether the SPDZ preprocessing in Profile P2 can be batched with nonce DKG preprocessing into a single unified offline phase—are open questions with direct impact on deployment latency.
+Reducing the MPC frontier via carry separation.
+Profile P2’s MPC evaluates the full $r_{0}$-check over $\mathbb{Z}_{q}$ (23-bit arithmetic).
+A finer decomposition is possible: because each party holds Shamir nonce and key shares, the low-bit residual $\ell_{i}=\mathsf{LowBits}(A\mathbf{y}_{i}-c\mathbf{s}_{2,i})$ is locally precomputable without communication.
+The $r_{0}$-check then reduces to evaluating only the carry of $\sum\lambda_{i}\ell_{i}$ and a single comparison bit—a much lighter computation than the full modular circuit.
+With standard Shamir coefficients $\lambda_{i}\approx q$, this yields no circuit reduction, since the intermediate sum overflows $\mathbb{Z}_{q}$.
+However, if reconstruction coefficients satisfy $\lambda_{i}\in\{0,1\}$—as in Vandermonde-style short-coefficient LSSS explored by concurrent work [BCDENP25]—the carry is bounded by $\lfloor T/2\rfloor$, reducing the MPC computation to a $\lceil\log_{2}T\rceil$-bit carry circuit rather than a full 23-bit modular arithmetic circuit.
+Designing a security proof for this hybrid (Shamir nonce DKG for statistical nonce privacy combined with short-coefficient sharing for $\mathbf{s}_{2}$) is an open problem; the key challenge is showing that the bounded carry value does not leak sufficient information about $\mathbf{s}_{2}$ to assist a lattice adversary.
+More broadly, whether threshold ML-DSA with dishonest-majority and arbitrary $T$ can be instantiated without general-purpose MPC—overcoming the construction-specific $T\leq 6$ barrier of existing short-coefficient LSSS schemes (the Ball-Çakan-Malkin bound [BCM21] establishes an asymptotic $\Omega(T\log T)$ share-size lower bound; the specific $T\leq 6$ ceiling is Celi et al.’s construction-specific limit)—remains a fundamental open question in lattice-based threshold cryptography.
+Commitment mask hiding for $T=N$.
+In P2’s broadcast model, the pairwise-canceling masks on $\mathbf{W}_{i}$ and $\mathbf{V}_{i}$ require $|S\setminus C|\geq 2$ for computational hiding (Lemma 4.15).
 In the extreme case $T=N$ (unanimous signing), a single honest party’s commitment and r0-check values are not hidden in P2.
 (In P1 and P3+, the coordinator-only communication model protects $\mathbf{W}_{i}$ even with $|S\setminus C|=1$; see Remark 4.1.) Replacing pairwise PRF masks with a non-interactive zero-sharing scheme—where each party’s mask is derived from a common reference string rather than pairwise seeds—or employing a verifiable PRF construction could close this gap, yielding uniform $|S|\geq T$ privacy across all revealed values.
-Additionally, achieving true information-theoretic privacy (SD =0=0) for the response shares with $|S|=T$ remains an open problem: the bounded-nonce constraint fundamentally prevents $\mathbb{Z}_{q}$-uniformity (Remark 4.6), and it is unclear whether alternative nonce generation techniques can bypass this barrier without requiring $|S|\geq T+1$.
+Additionally, achieving perfect uniformity (SD =0=0) for the response shares with $|S|=T$ remains an open problem: the bounded-nonce constraint fundamentally prevents $\mathbb{Z}_{q}$-uniformity (Remark 4.10), and it is unclear whether alternative nonce generation techniques can bypass this barrier without requiring $|S|\geq T+1$.
+(Note: the term “statistical privacy” in this work refers to the high-min-entropy guarantee of Theorem 4.7, not SD =0=0.)
 Finally, on the applied side, threshold ML-DSA is a natural building block for post-quantum decentralized identity systems.
@@
 Game 3 (Simulated Shares).
-Replace honest shares with uniform: for $j\notin C$, set $\mathbf{s}_{1,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ and $\mathbf{s}_{2,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k}$.
-Analysis: By Shamir’s $(T,N)$-threshold security, any <T<T shares are information-theoretically independent of the secret.
-Since $|C|\leq T-1$, the adversary’s view is identical: $\Pr[\mathsf{Win}_{3}]=\Pr[\mathsf{Win}_{2}]$.
+We argue that the adversary’s view of the honest parties’ key shares $\{\mathbf{s}_{1,j}\}_{j\notin C}$ is statistically independent of the secret key $\mathbf{s}_{1}$, by Shamir’s $(T,N)$-threshold perfect secrecy: any T1T-1 evaluations are jointly consistent with every possible value of the secret (the marginal distribution of each honest share at a non-zero evaluation point is $R_{q}^{\ell}$-uniform, but the joint distribution of multiple honest shares is not i.i.d.—they are correlated as evaluations of the same degree-(T1)(T\!-\!1) polynomial; the correct statement is that the joint view reveals zero information about $\mathbf{s}_{1}$).
+Important: this is a conceptual argument about the adversary’s view—the signing oracle continues using the actual short key shares (replacing them with uniform $R_{q}^{\ell}$ values would cause the $\|\mathbf{z}\|_{\infty}$-bound and r0-checks to fail with overwhelming probability, invalidating signatures).
+Since $|C|\leq T-1$, the adversary’s view of honest shares is statistically independent of the secret: $\Pr[\mathsf{Win}_{3}]=\Pr[\mathsf{Win}_{2}]$.
+(In Profile P2, individual per-party responses $\mathbf{z}_{j}$ are contributed only as SPDZ input shares, so SPDZ input-privacy ensures the adversary observes only the aggregate $\mathbf{z}$, not individual $\mathbf{z}_{j}$ values; the same Shamir argument applies.)
+Game 3.5 (Uniform Nonces).
+Replace Irwin-Hall nonce distribution with uniform over $\{-\gamma_{1},\ldots,\gamma_{1}\}^{n\ell}$.
+In Game 3, the aggregate nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ follows the Irwin-Hall distribution.
+We replace this with $\mathbf{y}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\gamma_{1},\ldots,\gamma_{1}\}^{n\ell}$ uniform.
+Analysis: The $q_{s}$ signing sessions are independent, so the joint Rényi divergence is $(R_{2}^{\mathsf{vec,shift}})^{q_{s}}$ (direct shift-invariance, Theorem Q.25).
+Applying the smooth Rényi divergence probability transfer lemma (Lemma Q.13) to the joint distribution:
+$\Pr[\mathsf{Win}_{3}]\leq\bigl(R_{2}^{\mathsf{vec,shift}}\bigr)^{q_{s}/2}\cdot\sqrt{\Pr[\mathsf{Win}_{3.5}]}+q_{s}\cdot\epsilon_{\mathsf{IH}}$
+where $R_{2}^{\mathsf{vec,shift}}=(1+\chi^{2}_{\mathsf{direct}})^{n\ell}$ is the full-vector per-session Rényi divergence from the direct shift-invariance analysis, with $\chi^{2}_{\mathsf{direct}}\approx 6.62\times 10^{-6}$ per coordinate for $|S|=17$ (Corollary Q.24), giving $R_{2}^{\mathsf{vec,shift}}\approx 1.0085$, and $\epsilon_{\mathsf{IH}}<10^{-30}$ (Theorem 4.17).
+The total security loss is $\approx 6.1\times 10^{-3}\cdot q_{s}$ bits, giving a proven security bound of $\geq 96-0.006\cdot q_{s}$ bits for $|S|\leq 17$ (where the 96-bit baseline is the Cauchy-Schwarz halving of the 192-bit NIST Level 3 baseline).
+The conservative analytical bound ($R_{2}^{\mathsf{vec}}\approx 2^{5.4}$, Remark 4.5) is vacuous at $q_{s}\gtrsim 36$.
 Game 4 (Signature Simulation).
-Simulate signatures without $\mathbf{s}_{1}$:
+Starting from uniform nonces (Game 3.5), simulate signatures without $\mathbf{s}_{1}$:
   1. 1.
-    Sample $\mathbf{z}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{v\in R_{q}^{\ell}:\|\mathbf{v}\|_{\infty}<\gamma_{1}-\beta\}$ with appropriate rejection sampling
+    Sample $\mathbf{z}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{v\in R_{q}^{\ell}:\|\mathbf{v}\|_{\infty}<\gamma_{1}-\beta\}$ with rejection sampling
   2. 2.
@@
   4. 4.
-    Simulate honest parties’ responses $\mathbf{z}_{j}$ by sampling from a high-entropy distribution (justified by Theorem 4.4: conditional min-entropy $\geq n\ell\cdot\log_{2}(2\gamma_{1}/|S|)$ bits) and their masked values $\mathbf{W}_{j},\mathbf{V}_{j}$ as uniform (justified by Lemma 4.9)
-Analysis: By Theorem 4.4, each honest party’s response $\mathbf{z}_{j}$ has high conditional min-entropy (statistical, no computational assumptions).
-The masked commitment and r0-check values $\mathbf{W}_{j},\mathbf{V}_{j}$ are computationally indistinguishable from uniform by Lemma 4.9.
-Hybrid over $|S\setminus C|$ honest parties: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3}]|\leq n\ell\cdot(1-2\gamma_{1}/(|S|\cdot q))+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$ (the first term from response simulation is negligible for practical $|S|$; the second from mask-based values uses the union bound over all $\binom{N}{2}$ pairwise PRF keys).
+    Simulate masked values $\mathbf{W}_{j},\mathbf{V}_{j}$ as computationally uniform (Lemma 4.15, replacing honest-honest PRF terms with uniform)
+Clarification: In this EUF-CMA game, the adversary observes only the aggregate public signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$; individual per-party responses $\mathbf{z}_{j}$ are processed inside the TEE/combiner and never reach the adversary.
+Accordingly, Steps 1–3 produce the adversary-visible $\mathbf{z}$ without using $\mathbf{s}_{1}$, and Step 4 handles the masked intermediate values that the adversary also cannot observe directly.
+The full per-party simulation (including $\mathbf{z}_{j}=\mathbf{y}_{j}+c\mathbf{s}_{1,j}$ from DKG values) is deferred to the UC/privacy simulator in Appendix A.2.
+Analysis: The masked commitment and r0-check values $\mathbf{W}_{j},\mathbf{V}_{j}$ are computationally indistinguishable from uniform by Lemma 4.15.
+By PRF security: $|\Pr[\mathsf{Win}_{4}]-\Pr[\mathsf{Win}_{3.5}]|\leq\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
 Reduction to M-SIS.
-Construct M-SIS adversary $\mathcal{B}$:
+Construct SelfTargetMSIS adversary $\mathcal{B}$ [KLS18]:
   1. 1.
@@
   2. 2.
-    $\mathcal{B}$ sets $\rho$ such that $\mathsf{ExpandA}(\rho)=\mathbf{A}$, samples $\mathbf{t}_{1}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k}$
+    $\mathcal{B}$ sets $\rho$ such that $\mathsf{ExpandA}(\rho)=\mathbf{A}$, samples $\mathbf{t}_{1}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{k}$ (uniformly random; valid since $\mathsf{Game}_{4}$ simulation does not require knowing $\mathbf{s}_{1}$ or $\mathbf{s}_{2}$; $\mathbf{t}_{1}$ is indistinguishable from a real key by M-LWE hardness [KLS18])
   3. 3.
-    $\mathcal{B}$ simulates $\mathsf{Game}_{4}$ for $\mathcal{A}$
+    $\mathcal{B}$ picks a random index $i^{*}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}[q_{H}]$ as the target ROM query
   4. 4.
-    If $\mathcal{A}$ outputs valid forgery $(\mu^{*},\sigma^{*})$, apply General Forking Lemma [BN06]
+    $\mathcal{B}$ simulates $\mathsf{Game}_{4}$ for $\mathcal{A}$, programming ROM responses uniformly
   5. 5.
-    Rewind to obtain second forgery with same $\mathbf{w}_{1}^{*}$ but different $c_{1}\neq c_{2}$
-  6. 6.
-    Define $\mathbf{v}=\mathbf{z}_{1}-\mathbf{z}_{2}$ and $\Delta c=c_{1}-c_{2}$
-Analysis: Forking succeeds with probability $\geq\Pr[\mathsf{Win}_{4}]^{2}/(q_{H}\cdot|\mathcal{C}|)$. By verification equation:
-$\mathsf{HighBits}(\mathbf{A}\mathbf{z}_{i}-c_{i}\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})=\mathbf{w}_{1}\text{ for }i\in\{1,2\}$
-Thus $\mathbf{A}\mathbf{v}=\Delta c\cdot\mathbf{t}_{1}\cdot 2^{d}+\boldsymbol{\delta}$ where $\|\boldsymbol{\delta}\|_{\infty}<2\gamma_{2}$.
-This yields a short vector for the augmented M-SIS instance $\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]$:
-$\mathbf{A}|-\mathbf{t}_{1}\cdot 2^{d}]\cdot\begin{bmatrix}\mathbf{v}\\ \Delta c\end{bmatrix}=\boldsymbol{\delta}$
-with $\|(\mathbf{v},\Delta c)\|_{\infty}\leq\max(2(\gamma_{1}-\beta),2)<2\gamma_{1}$, solving M-SIS.
-A.2 Privacy Proof (Full)
-Complete simulator and indistinguishability analysis for Theorem 4.4.
-With the Shamir nonce DKG, privacy of the response $\mathbf{z}_{h}$ is statistical (no computational assumptions) and does not require a hybrid argument.
-The proof simplifies considerably compared to the computational (PRF-based) approach.
+    When $\mathcal{A}$ outputs forgery $(\mu^{*},\mathbf{z}^{*},c^{*},\mathbf{h}^{*})$: if $c^{*}$ equals the $i^{*}$-th ROM response, extract; otherwise abort
+Analysis: $\mathcal{B}$ succeeds when its guess $i^{*}$ is correct, which occurs with probability $1/q_{H}$, so $\Pr[\mathcal{B}\ \text{succeeds}]\geq\Pr[\mathsf{Win}_{4}]/q_{H}$.
+When successful, the forgery verification gives:
+$\mathsf{HighBits}(\mathbf{A}\mathbf{z}^{*}-c^{*}\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})=\mathbf{w}_{1}^{*}$
+Rearranging yields the SelfTargetMSIS solution for $\mathbf{A}\mid-\mathbf{t}_{1}\cdot 2^{d}]$ (see Remark 2.3):
+$\mathbf{A}\mid-\mathbf{t}_{1}\cdot 2^{d}]\cdot\begin{bmatrix}\mathbf{z}^{*}\\ c^{*}\end{bmatrix}=2\gamma_{2}\mathbf{w}_{1}^{*}+\mathbf{r}_{0}^{*}$
+where $\mathbf{w}_{1}^{*}=\mathsf{HighBits}(\mathbf{A}\mathbf{z}^{*}-c^{*}\mathbf{t}_{1}\cdot 2^{d},\,2\gamma_{2})$ is the known verification target and $\|\mathbf{r}_{0}^{*}\|_{\infty}\leq\gamma_{2}$ is the $\mathsf{LowBits}$ residual.
+The combined solution $(\mathbf{z}^{*});(c^{*})]\in R_{q}^{\ell+1}$ satisfies $\|\mathbf{z}^{*}\|_{\infty}<\gamma_{1}-\beta$ (z-bound check) and $\|c^{*}\|_{\infty}=1$ (challenge coefficients in $\{-1,0,+1\}$); the module $\ell^{\infty}$-norm is $\max(\|\mathbf{z}^{*}\|_{\infty},\|c^{*}\|_{\infty})<\gamma_{1}$, matching the SelfTargetMSIS norm parameter from Remark 2.3 and [KLS18].
+Thus $\Pr[\mathsf{Win}_{4}]\leq q_{H}\cdot\epsilon_{\mathsf{M\text{-}SIS}}$.
+A.2 Privacy Proof Details
+We show that the simulator produces an indistinguishable view, applying the min-entropy bound established in Theorem 4.7 (Section 4).
+With the Shamir nonce DKG, privacy of the nonce share $\mathbf{y}_{h}$ is statistical (no computational assumptions).
+The simulator computes responses from simulated DKG values, achieving perfect indistinguishability for the response component.
 Simulator $\mathcal{S}$.
@@
 […]
-Round 3 Simulation: For each honest $j\in S\setminus C$, sample $\mathbf{z}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ uniformly.
-This is justified by Theorem 4.4: the real $\mathbf{z}_{h}$ has high conditional min-entropy ($\geq 5\times$ the secret key entropy for $|S|\leq 17$) from the adversary’s perspective, so the simulator’s uniform sampling introduces negligible statistical distance.
-The $\mathbf{V}_{j}$ values are simulated uniformly by the same PRF-based argument as for $\mathbf{W}_{j}$.
+Round 3 Simulation: The simulator must produce per-party responses $\mathbf{z}_{j}$ consistent with both the simulated DKG values and the public signature constraint $\sum_{j\in S}\lambda_{j}\mathbf{z}_{j}=\mathbf{z}$ (enforced by the public $\sigma$).
+We handle the two cases separately.
+Case $|S\setminus C|>1$: Let $\mathbf{z}_{j}^{*}=\mathbf{y}_{j}^{*}+c\cdot\mathbf{s}_{1,j}$ denote corrupted party $j$’s response, where $\mathbf{y}_{j}^{*}$ is $j$’s simulated DKG evaluation (known to the simulator from the corrupted view).
+The simulator chooses $\mathbf{z}_{j}=\mathbf{y}_{j}+c\cdot\mathbf{s}_{1,j}$ (from simulated DKG values) for all honest $j\in S\setminus C$ except a designated party $h_{0}$, then sets
+$\mathbf{z}_{h_{0}}=\frac{\mathbf{z}-\sum_{j\in C}\lambda_{j}\mathbf{z}_{j}^{*}-\sum_{j\in(S\setminus C)\setminus\{h_{0}\}}\lambda_{j}\mathbf{z}_{j}}{\lambda_{h_{0}}}$
+(The division by $\lambda_{h_{0}}$ is well-defined: since all evaluation points lie in $\{1,\ldots,N\}\subset\mathbb{Z}_{q}^{*}$, the Lagrange coefficient $\lambda_{h_{0}}=\prod_{j\in S,j\neq h_{0}}j(j-h_{0})^{-1}\bmod q$ is a product of nonzero field elements and hence nonzero mod $q$, hence invertible in $\mathbb{Z}_{q}$.) This $\mathbf{z}_{h_{0}}$ matches exactly what the real protocol would produce (given the DKG polynomials), so SD =0=0.
+Case $|S\setminus C|=1$: The single honest response $\mathbf{z}_{h}=(\mathbf{z}-\sum_{j\in C}\lambda_{j}\mathbf{z}_{j}^{*})/\lambda_{h}$ is uniquely determined by $\sigma$ and the corrupted views.
+The simulator outputs this value directly.
+Privacy of $\mathbf{s}_{1,h}$ (key privacy, computational) is maintained because the adversary knows $\mathbf{z}_{h}$ but not the decomposition $\mathbf{z}_{h}=\mathbf{y}_{h}+c\mathbf{s}_{1,h}$: recovering $\mathbf{s}_{1,h}$ from $\mathbf{z}_{h}$ requires solving a bounded-distance decoding instance given $\mathbf{y}_{h}$’s high conditional min-entropy (Theorem 4.7).
+The $\mathbf{V}_{j}$ values are computed as $\mathbf{V}_{j}=\lambda_{j}c\mathbf{s}_{2,j}+\mathbf{m}_{j}^{(s2)}$ using simulated key shares and PRF masks.
 Indistinguishability.
-The simulation is statistical for $\mathbf{z}_{j}$ (high min-entropy, negligible statistical distance from uniform) and computationally indistinguishable for $\mathbf{W}_{j}$ (where $\mathbf{W}_{j}=\lambda_{j}\mathbf{w}_{j}+\mathbf{m}_{j}^{(w)}$) and $\mathbf{V}_{j}$ (PRF security).
-The total distinguishing advantage is $n\ell\cdot(1-2\gamma_{1}/(|S|\cdot q))+\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$, where the first term (negligible for practical $|S|$) arises from the response simulation and the second from mask-based values.
-In particular, no PRF assumption is needed for response privacy.
+In both cases, the simulated $\mathbf{z}_{j}$ follows the same distribution as the real protocol (SD =0=0 conditioned on $\sigma$).
+Masked values $\mathbf{W}_{j}$ and $\mathbf{V}_{j}$ are computationally indistinguishable from uniform (PRF security, Lemma 4.15); no PRF assumption is needed for nonce share privacy ($\mathbf{y}_{h}$’s statistical privacy is independent).
+The total distinguishing advantage is $\binom{N}{2}\cdot\epsilon_{\mathsf{PRF}}$.
 Appendix B Distributed Key Generation Protocols
 Our protocol uses two structurally similar but functionally distinct DKG protocols: a key DKG (executed once during setup) and a nonce DKG (executed before each signing session, typically preprocessed offline).
-Both follow the Feldman VSS paradigm but differ in their input distributions and security requirements.
+The key DKG uses Feldman-style commitments for verifiability; the nonce DKG omits VSS verification (correctness is instead ensured at the signing stage via the z-bound and r0-check).
+Both share the same polynomial evaluation structure but differ in their input distributions and security requirements.
 B.1 Key DKG (One-Time Setup)
@@
 The protocol is specified formally in Algorithm 1 of the main paper (Section 3.4).
-Each party $i\in S$ generates a degree-(T1)(T\!-\!1) polynomial $\hat{f}_{i}(X)$ whose constant term $\hat{\mathbf{y}}_{i}^{(0)}$ is sampled from the reduced nonce range $\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$ and whose higher-degree coefficients are sampled uniformly from $R_{q}^{\ell}$.
-The resulting aggregated nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}^{(0)}$ follows the Irwin-Hall distribution analyzed in Theorem 4.11, while each honest party’s individual share $\mathbf{y}_{h}=\sum_{i\in S}\hat{f}_{i}(h)$ has high conditional min-entropy from the adversary’s perspective (Theorem 4.4).
+Each party $i\in S$ generates a degree-(T1)(T\!-\!1) polynomial $\hat{f}_{i}(X)$ whose constant term $\hat{\mathbf{y}}_{i}$ is sampled from the reduced nonce range $\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$ and whose higher-degree coefficients are sampled uniformly from $R_{q}^{\ell}$.
+The resulting aggregated nonce $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ follows the Irwin-Hall distribution analyzed in Theorem 4.17, while each honest party’s individual share $\mathbf{y}_{h}=\sum_{i\in S}\hat{f}_{i}(h)$ has high conditional min-entropy from the adversary’s perspective (Theorem 4.7).
 B.3 Security Properties
@@
   • Secrecy: Any coalition of <T<T parties learns nothing about the secret beyond what is revealed by public commitments (follows from Shamir’s threshold security).
-  • Robustness: With Feldman commitments [Feldman87], malicious shares can be detected and attributed.
-The nonce DKG additionally provides statistical privacy (no computational assumptions) of individual shares (Theorem 4.4), which is the key enabler for eliminating the two-honest requirement for response privacy in coordinator-based profiles (P1, P3+; see Remark 3.7).
+  • Robustness: The key DKG uses Feldman commitments [Feldman87] to detect and attribute malicious shares.
+The nonce DKG omits VSS verification; malicious behavior is instead detected at the signing stage (z-bound or r0-check failure triggers the blame protocol), which suffices for security-with-abort.
+The nonce DKG additionally provides statistical privacy (no computational assumptions) of individual shares (Theorem 4.7), which is the key enabler for eliminating the two-honest requirement (enabling $|S|=T$ signing sets with nonce share privacy) in coordinator-based profiles (P1, P3+; see Remark 3.9).
 The key DKG does not require this property, since key shares are never revealed during signing.
@@
 An adversary must corrupt $\geq T$ parties within a single epoch to learn the secret.
+Forward Secrecy for Past Signatures.
+Signatures from epoch $k$ remain secure after the epoch-(k+1)(k\!+\!1) refresh.
+A signature $\sigma$ from epoch $k$ reveals $\mathbf{z}_{j}^{(k)}=\mathbf{y}_{j}^{(k)}+c\cdot\mathbf{s}_{1,j}^{(k)}$.
+Since $\mathbf{y}_{j}^{(k)}$ is a one-time nonce share (deleted after use), and the refreshed shares $\mathbf{s}_{1,j}^{(k+1)}=\mathbf{s}_{1,j}^{(k)}+\text{(zero-sharing term)}$ do not reveal $\mathbf{s}_{1,j}^{(k)}$ by Shamir’s threshold property (the zero-sharing randomizes shares while preserving the secret), past key shares cannot be reconstructed from refreshed shares alone.
+Key privacy for past signatures therefore holds across epochs under the same M-LWE hardness assumption as the signing scheme.
 Appendix D Numerical Verification of Rényi Bounds
@@
 Rényi Divergence Calculation.
-For signing set size $|S|=n$, each party samples nonces from range $N=2\lfloor\gamma_{1}/n\rfloor+1$.
-For readability, we use the approximation $N\approx 2\gamma_{1}/n$ (the difference is at most $1/N<2\times 10^{-5}$ for all relevant $n$, and the exact discrete range is used in the implementation).
-Using Lemma 4.2 of [Raccoon2024] with $\alpha=2$ (where $\alpha(\alpha-1)/2=1$) and shift $c=\beta$:
-$R_{2}^{\epsilon}\leq 1+\frac{n^{2}\cdot\beta^{2}}{N^{2}}=1+\frac{n^{2}\cdot\beta^{2}\cdot n^{2}}{4\gamma_{1}^{2}}=1+\frac{n^{4}\cdot\beta^{2}}{4\gamma_{1}^{2}}$
-$|S|$ $N=2\gamma_{1}/|S|$ $R_{2}^{\epsilon}-1$ $\epsilon$ (tail)
+Let $m=|S|$ denote the signing set size (distinct from the ring degree $n=256$) and $B=2\lfloor\gamma_{1}/m\rfloor$ the per-share full range (from $\lfloor\gamma_{1}/m\rfloor$ to $\lfloor\gamma_{1}/m\rfloor$, so the total width is $2\lfloor\gamma_{1}/m\rfloor$).
+Note that in Definition Q.7, the symbol $B$ denotes the half-range (each $X_{i}\in\{-B,\ldots,B\}$); to apply Raccoon [Raccoon2024] Lemma 4.2, $B$ here is the full range, equal to $2\times$ that half-range.
+For readability, we use the approximation $B\approx 2\gamma_{1}/m$ (the rounding error is at most $2/B<4\times 10^{-5}$ for all relevant $m$, and the exact discrete range is used in the implementation).
+Using Lemma 4.2 of [Raccoon2024] with $\alpha=2$ and shift $\delta=\beta$:
+$R_{2}^{\epsilon}$ $\leq 1+\frac{m^{2}\cdot\beta^{2}}{B^{2}}\approx 1+\frac{m^{2}\cdot\beta^{2}}{(2\gamma_{1}/m)^{2}}$
+$=1+\frac{m^{4}\cdot\beta^{2}}{4\gamma_{1}^{2}}=1+\frac{|S|^{4}\cdot\beta^{2}}{4\gamma_{1}^{2}}$
+$|S|$ $B\approx 2\gamma_{1}/|S|$ $R_{2}^{\epsilon}-1$ $\epsilon$ (tail)
 4 262144 $8.9\times 10^{-6}$ $10^{-10}$
 9 116508 $2.3\times 10^{-4}$ $10^{-19}$
-17 61681 $2.9\times 10^{-3}$ $10^{-30}$
-25 41943 $1.4\times 10^{-2}$ $10^{-40}$
-33 31775 $4.1\times 10^{-2}$ $10^{-49}$
+17 61680 $2.9\times 10^{-3}$ $10^{-30}$
+25 41942 $1.4\times 10^{-2}$ $10^{-40}$
+33 31774 $4.1\times 10^{-2}$ $10^{-49}$
 Table 12: Per-coordinate Rényi divergence bounds for ML-DSA-65.
-All values computed using proven Lemma 4.2 of [Raccoon2024].
-Full-vector: $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$.
+$B=2\lfloor\gamma_{1}/|S|\rfloor$ is the exact per-share full range; displayed values are $2\lfloor\gamma_{1}/|S|\rfloor$.
+All bounds computed using Lemma 4.2 of [Raccoon2024].
+Full-vector: $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$ ($n=256$, $\ell=5$).
 The per-coordinate $R_{2}^{\epsilon,\mathsf{coord}}$ remains below $1.05$ for $|S|\leq 33$.
 The full-vector divergence $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$ ranges from $\approx 2^{5.4}$ ($|S|=17$) to $\approx 2^{75}$ ($|S|=33$); see Theorem 4.17 for the impact on the security bound.
+The direct analysis (Corollary Q.24) gives <0.013<0.013 bits for all $|S|\leq 33$.
 Appendix E Rejection Sampling Analysis
@@
     r0 check: $\left\|\text{LowBits}(\mathbf{w}-c\mathbf{s}_{2})\right\|_{\infty}<\gamma_{2}-\beta$
-For ML-DSA-65:
+For ML-DSA-65 (single-signer approximation: nonce uniform on $\gamma_{1},\gamma_{1}]$):
   • z-bound: $\left(\frac{\gamma_{1}-\beta}{\gamma_{1}}\right)^{n\ell}=\left(\frac{524092}{524288}\right)^{1280}\approx 62\%$
@@
 This matches the NIST ML-DSA specification [FIPS204] (expected 4–5 iterations).
+Note: In the threshold protocol, nonces follow the Irwin-Hall distribution (sum of $|S|$ uniform variables), so the per-attempt success rates differ from this single-signer estimate; empirical measurements yield 21–45% depending on $T$ and $|S|$ (see Table 13).
 E.2 Throughput Optimizations
-  • Parallel Attempts: Generate $B$ independent $(y,w)$ pairs per signing session. Success probability: $1-(0.75)^{B}$.
+  • Parallel Attempts: Generate $B$ independent $(y,w)$ pairs per signing session.
+Success probability: $1-(0.80)^{B}$ (using the theoretical $\approx 20\%$ per-attempt rate; empirical benchmarks yield $\approx 25$$31\%$, giving $1-(0.75)^{B}$ as an optimistic estimate).
   • Pre-computation: Pre-compute $(y_{i},w_{i}=\mathbf{A}y_{i})$ pairs during idle time.
@@
 12 $4.1\times 10^{-9}$ $31.4\%$ $7.5\times 10^{7}$
-16 $6.6\times 10^{-12}$ $32.4\%$ $4.8\times 10^{10}$
+16 $6.6\times 10^{-12}$ $32.4\%$ $4.9\times 10^{10}$
 24 $1.7\times 10^{-17}$ $23.1\%$ $1.4\times 10^{16}$
 32 $4.3\times 10^{-23}$ $25.0\%$ $5.8\times 10^{21}$
-Table 13: Improvement over naive threshold approach. “Our Success” values are theoretical estimates based on rejection sampling analysis; measured success rates vary with configuration (see Table 5).
+Table 13: Improvement over naive threshold approach.
+“Our Success” values are empirical measurements from the Rust benchmarks (Table 5).
+The non-monotone behavior ($32.4\%\to 23.1\%\to 25.0\%$ for $T=16,24,32$) reflects the interplay of two competing effects: (1) smaller per-party nonce range $\gamma_{1}/|S|$ reduces the acceptance probability for the z-bound check; (2) the Irwin-Hall distribution becomes more concentrated near zero for larger $|S|$, partially improving acceptance.
+The net effect is non-monotone in $|S|$ and configuration-dependent.
 Appendix G Nonce Range Rounding Details
@@
 Effect on Security.
-The z-bound check is $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ where $\gamma_{1}-\beta=524092$.
-Since $524280>524092$, the reduced nonce range does not affect the pass probability.
+The z-bound check is $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ where $\gamma_{1}-\beta=524092$. Since $524280>524092$, the reduced nonce range does not affect the pass probability for this example.
+General Proof.
+Writing $\gamma_{1}=|S|\cdot\lfloor\gamma_{1}/|S|\rfloor+r$ with $0\leq r\leq|S|-1$, the reduced nonce maximum is $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor=\gamma_{1}-r\geq\gamma_{1}-(|S|-1)$.
+For the z-bound check to be unaffected, we need $\gamma_{1}-(|S|-1)>\gamma_{1}-\beta$, i.e., $|S|-1<\beta=\tau\eta=196$.
+This holds for all $|S|\leq 196$, covering all practical signing set sizes.
 The Irwin-Hall security analysis uses the actual sampled range $N=2\lfloor\gamma_{1}/|S|\rfloor$, so all bounds account for this rounding.
 Appendix H UC Security Framework
+Remark H.1 (Scope of this appendix).
+This appendix provides the extended UC proofs for Profiles P2 and P3+.
+The Profile P1 UC proof (TEE-based coordinator) is self-contained in Theorem 4.21 and its proof in the main paper; it is not repeated here.
+This appendix proves UC security for the r0-check subprotocol (Profile P2, Theorem 6.3) and the semi-async signing protocol (Profile P3+, Theorem P.12).
+These theorems establish UC realization of $\mathcal{F}_{r_{0}\text{-check}}$ and $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$ respectively; realization of the full $\mathcal{F}_{\mathsf{ThreshSig}}$ (Definition 4.19) follows by UC composition with the nonce DKG and signing-step sub-protocols.
 We recall the Universal Composability (UC) framework of Canetti [Canetti01, Canetti20]. Our presentation follows the conventions of [CLOS02, DPSZ12].
@@
 H.2 Hybrid Functionalities
-We work in the $(\mathcal{F}_{\mathsf{SPDZ}},\mathcal{F}_{\mathsf{edaBits}},\mathcal{F}_{\mathsf{Rand}},\mathcal{F}_{\mathsf{Broadcast}})$-hybrid model, where:
-Functionality H.5 ($\mathcal{F}_{\mathsf{SPDZ}}$ – Authenticated Secret Sharing [DPSZ12]).
+We work in the $(\mathcal{F}_{\mathsf{SPDZ}},\mathcal{F}_{\mathsf{edaBits}},\mathcal{F}_{\mathsf{Binary}})$-hybrid model, where:
+Functionality H.6 ($\mathcal{F}_{\mathsf{SPDZ}}$ – Authenticated Secret Sharing [DPSZ12]).
 Parameterized by field $\mathbb{F}_{q}$, parties $\mathcal{P}=\{P_{1},\ldots,P_{n}\}$, and MAC key $\alpha\in\mathbb{F}_{q}$.
@@
     Otherwise, output $x=\sum_{j}[x]_{j}$ to all parties
-Security: $\mathcal{F}_{\mathsf{SPDZ}}$ provides information-theoretic hiding of shares and computational integrity (MAC forgery probability $\leq 1/q$ per check).
-Functionality H.6 ($\mathcal{F}_{\mathsf{edaBits}}$ – Extended Doubly-Authenticated Bits [EKMOZ20]).
+Security: $\mathcal{F}_{\mathsf{SPDZ}}$ provides statistical hiding of shares (additive shares are perfectly uniform given $\leq N-1$ honest parties) and computational integrity (MAC forgery probability $\leq 1/q$ per check).
+Functionality H.7 ($\mathcal{F}_{\mathsf{edaBits}}$ – Extended Doubly-Authenticated Bits [EKMOZ20]).
 On input $(\mathsf{EDABIT},\mathsf{sid},m)$ requesting $m$ edaBits:
@@
 With this modification, $r\bmod q$ is perfectly uniform over $\mathbb{F}_{q}$, and Lemma L.7 applies exactly.
-Functionality H.8 ($\mathcal{F}_{\mathsf{Binary}}$ – Binary Circuit Evaluation).
+Functionality H.9 ($\mathcal{F}_{\mathsf{Binary}}$ – Binary Circuit Evaluation).
 On input $(\mathsf{EVAL},\mathsf{sid},C,\{\langle x_{i}\rangle_{2}\})$ where $C$ is a binary circuit:
@@
   • $\gamma_{2}=(q-1)/32=261888$: rounding parameter
-  • $\beta=196$: bound parameter
+  • $\beta=\tau\cdot\eta=49\cdot 4=196$: bound parameter (ML-DSA-65)
   • $\mathsf{bound}=\gamma_{2}-\beta=261692$: acceptance threshold
@@
     3. (c)
-      Compute pass/fail for each coefficient:
-      $\mathsf{pass}_{j}=\begin{cases}1&\text{if }|\tilde{w}^{\prime}_{j}\bmod 2\gamma_{2}|_{[-\gamma_{2},\gamma_{2})}<\mathsf{bound}\\ 0&\text{otherwise}\end{cases}$
-      where $|x|_{[-\gamma_{2},\gamma_{2})}$ denotes the centered representative of $x\bmod 2\gamma_{2}$
+      Compute pass/fail for each coefficient using the FIPS 204 $\mathsf{LowBits}$ function [FIPS204]:
+      $\mathsf{pass}_{j}=\begin{cases}1&\text{if }|\mathsf{LowBits}(\mathbf{w}^{\prime}[j],\,2\gamma_{2})|<\mathsf{bound}\\ 0&\text{otherwise}\end{cases}$
+      where $\mathsf{LowBits}(r,2\gamma_{2})=r\bmod{\pm}\,2\gamma_{2}$ (the unique representative in $(-\gamma_{2},\gamma_{2}]$ congruent to $r\bmod 2\gamma_{2}$), matching FIPS 204 Algorithm 31.
+Note: $\mathbf{w}^{\prime}[j]=\sum_{i=1}^{n}\mathsf{received}[i][j]\bmod q$ is first reduced mod $q$, then $\mathsf{LowBits}$ is applied to this $\mathbb{Z}_{q}$ value directly without additional centering mod $q$.
     4. (d)
@@
   1. 1.
-    Parties compute $\langle\mathsf{masked}_{j}\rangle_{q}=\langle\mathbf{w}^{\prime}_{j}\rangle_{q}+\langle r_{j}\rangle_{q}$ using local addition
+    Parties compute using local addition:
+    $\langle\mathsf{masked}_{j}\rangle_{q}=\langle\mathbf{w}^{\prime}_{j}\rangle_{q}+\langle r_{j}\rangle_{q}$
   2. 2.
-    Each party $P_{i}$ computes commitment $\mathsf{Com}_{i,j}=H([\mathsf{masked}_{j}]_{i}\|\mathsf{sid}\|j\|i)$
+    Each party $P_{i}$ computes commitment:
+    $\mathsf{Com}_{i,j}=H\bigl([\mathsf{masked}_{j}]_{i}\|\mathsf{sid}\|j\|i\bigr)$
   3. 3.
@@
   1. 1.
-    Convert public $\mathsf{masked}_{j}$ to binary: $\mathsf{masked}_{j}^{(b)}=(\mathsf{masked}_{j}^{(0)},\ldots,\mathsf{masked}_{j}^{(\lceil\log_{2}q\rceil-1)})$
+    Convert public $\mathsf{masked}_{j}$ to binary:
+    $\mathsf{masked}_{j}^{(b)}=\bigl(\mathsf{masked}_{j}^{(0)},\ldots,\mathsf{masked}_{j}^{(\lceil\log_{2}q\rceil-1)}\bigr)$
   2. 2.
@@
   1. 1.
-    For each honest party $P_{i}\in H$:
+    For each honest party $P_{i}\in H$ (adversary’s view only):
     1. (a)
@@
   2. 2.
-    Receive corrupted parties’ inputs from $\mathcal{A}$
+    Receive corrupted parties’ inputs from $\mathcal{A}$; forward corrupted inputs to $\mathcal{F}_{r_{0}\text{-check}}$
   3. 3.
-    Forward corrupted inputs to $\mathcal{F}_{r_{0}\text{-check}}$
+    UC framing (honest parties’ inputs): In the ideal execution, honest parties are ideal-world programs that submit their actual inputs directly to $\mathcal{F}_{r_{0}\text{-check}}$ via the ideal-world channel — the simulator $\mathcal{S}$ is not involved in this step.
+The functionality receives all inputs and computes $\mathsf{result}$; $\mathcal{S}$ receives $\mathsf{result}$ and uses it in Round 8 simulation.
   4. 4.
-    For honest parties: send dummy inputs to $\mathcal{F}_{r_{0}\text{-check}}$ (the functionality only checks if all inputs received)
+    Consistency: The simulated shares $\tilde{s}_{i,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{F}_{q}$ are identically distributed to real honest-party shares (SPDZ additive shares are uniformly random given any fixed secret — perfect hiding).
+No inconsistency arises: the adversary sees random shares, honest parties’ actual values go to $\mathcal{F}$ privately.
 Simulating Rounds 3-4 (Masked Reveal):
@@
     4. (d)
-      Compute simulated $\widetilde{\mathsf{masked}}_{j}=\sum_{i\in H}\tilde{m}_{i,j}+\sum_{i\in C}[\mathsf{masked}_{j}]_{i}\bmod q$
+      Compute simulated
+      $\widetilde{\mathsf{masked}}_{j}=\sum_{i\in H}\tilde{m}_{i,j}+\sum_{i\in C}[\mathsf{masked}_{j}]_{i}\bmod q$
 Justification for Round 3-4 Simulation:
 In the real protocol:
-$\mathsf{masked}_{j}=\mathbf{w}^{\prime}_{j}+r_{j}=\sum_{i}[\mathbf{w}^{\prime}]_{i}[j]+\sum_{i}[r_{j}]_{i}\bmod q$
+$\mathsf{masked}_{j}$ $=\mathbf{w}^{\prime}_{j}+r_{j}$
+$=\sum_{i}[\mathbf{w}^{\prime}]_{i}[j]+\sum_{i}[r_{j}]_{i}\bmod q$
 Since $r_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{F}_{q}$ from $\mathcal{F}_{\mathsf{edaBits}}$ and is unknown to $\mathcal{A}$ (only shares are known), by the one-time pad property:
@@
   1. 1.
-    The public value $\widetilde{\mathsf{masked}}_{j}$ is converted to binary (same as real)
+    Convert $\widetilde{\mathsf{masked}}_{j}$ to binary (same as real)
   2. 2.
-    For binary subtraction circuit: use $\mathcal{F}_{\mathsf{Binary}}$ simulator
+    For binary subtraction: use $\mathcal{F}_{\mathsf{Binary}}$ simulator
   3. 3.
-    Honest parties’ binary shares are simulated as random (justified by Beaver AND security)
+    Simulate honest parties’ binary shares as uniform (justified by Beaver AND security)
 Simulating Rounds 6-7 (Comparison):
   1. 1.
-    For each AND gate in the comparison circuit:
-    1. (a)
-      The gate opens $(d,e)=(x\oplus a,y\oplus b)$ where $(a,b)$ are Beaver triple components
-    2. (b)
-      Sample $(\tilde{d},\tilde{e})\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{2}$ uniformly
-    3. (c)
-      This is indistinguishable from real by Lemma L.10
+    For each AND gate, sample $(\tilde{d},\tilde{e})\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{2}$. By Lemma L.11, this matches the real distribution.
   2. 2.
-    Simulate all comparison circuits this way
+    Apply the same simulation to all comparison circuits.
 Simulating Round 8 (Output):
@@
     Backward simulation of AND tree:
+    The adversary’s view of the AND tree is only the Beaver-triple open values $(d_{j},e_{j})$ — the leaf bits $\mathsf{pass}_{j}$ never appear in the adversary’s view.
     1. (a)
-      The AND tree computes $\mathsf{result}=\bigwedge_{j}\mathsf{pass}_{j}$
+      For each AND gate $j\in[m]$, sample $(d_{j},e_{j})\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{2}$ uniformly and independently (no leaf assignment is performed).
     2. (b)
-      Starting from known $\mathsf{result}$, assign intermediate values:
-      • If $\mathsf{result}=1$: set all $\mathsf{pass}_{j}=1$
-      • If $\mathsf{result}=0$: set at least one $\mathsf{pass}_{j}=0$ (adversary cannot determine which)
-    3. (c)
-      For each AND gate computing $z=x\land y$:
-      • Given target $z$, sample $(x,y)$ consistent with $z$
-      • The Beaver triple values $(d,e)$ revealed during computation are uniform regardless of $(x,y)$
+      Output the simulated view $\{(d_{j},e_{j})\}_{j=1}^{m}$ together with $\mathsf{result}$ received from $\mathcal{F}_{r_{0}\text{-check}}$.
+    Justification: By Lemma L.11, the real $(d_{j},e_{j})=(x_{j}\oplus\alpha_{j},y_{j}\oplus\beta_{j})$ is uniform regardless of the inputs $(x_{j},y_{j})$, because $(\alpha_{j},\beta_{j})$ are fresh Beaver-triple components independent of the gate inputs.
+Since all $m$ Beaver triples are generated independently in preprocessing, the joint distribution is product-uniform.
+The simulator’s directly-sampled $(d_{j},e_{j})$ therefore matches the real distribution exactly (SD =0=0), for both $\mathsf{result}\in\{0,1\}$ cases.
   3. 3.
@@
 By the UC-security of SPDZ [DPSZ12], there exists a simulator $\mathcal{S}_{\mathsf{SPDZ}}$ such that no environment can distinguish real SPDZ execution from ideal $\mathcal{F}_{\mathsf{SPDZ}}$ execution with advantage better than $\epsilon_{\mathsf{SPDZ}}$.
-The security parameter is $\kappa$ and the MAC key is in $\mathbb{F}_{q}$ where $q\approx 2^{23}$. Per SPDZ security analysis:
-  • MAC forgery probability per verification: $1/q\approx 2^{-23}$
-  • With $\ell$ verifications: $\epsilon_{\mathsf{SPDZ}}\leq\ell/q$
-For our protocol with $m=1536$ coefficients and constant verifications per coefficient:
-$\epsilon_{\mathsf{SPDZ}}\leq O(m)/q=O(1536)/8380417<2^{-12}$
+The security parameter is $\kappa$ and the MAC key is drawn from a dedicated MAC key field $\mathbb{F}_{p}$ with $|p|\geq 64$ bits, independent of the protocol field $\mathbb{F}_{q}$ (this is standard practice in SPDZ implementations [DPSZ12]).
+Per SPDZ security analysis:
+  • MAC forgery probability per verification: $1/p\leq 2^{-64}$
+  • With $\ell$ verifications: $\epsilon_{\mathsf{SPDZ}}\leq\ell/p$
+For our protocol with $m=1536$ coefficients, MAC verification occurs at OPEN time; the number of verifications is $\ell=m$:
+$\epsilon_{\mathsf{SPDZ}}\leq m/p\leq 1536/2^{64}<2^{-54}$
+Remark L.4 (Minimal-configuration bound).
+If SPDZ is instantiated with the protocol field $\mathbb{F}_{q}$ ($q\approx 2^{23}$) as the MAC key space (minimal configuration), the bound degrades to $\epsilon_{\mathsf{SPDZ}}\leq m/q=1536/8380417<2^{-12.4}$.
+Practical deployments of P2 must configure SPDZ with a separate MAC field of at least 64 bits to achieve cryptographically negligible distinguishing advantage.
 L.3 Hybrid $\mathsf{H}_{2}$: Replace edaBits with Ideal
@@
 $|\Pr[\mathsf{H}_{1}=1]-\Pr[\mathsf{H}_{2}=1]|\leq\epsilon_{\mathsf{edaBits}}$
-Proof L.5.
+Proof L.6.
 By the UC-security of edaBits [EKMOZ20], the protocol UC-realizes $\mathcal{F}_{\mathsf{edaBits}}$ with negligible distinguishing advantage.
@@
 where $\epsilon^{\prime}_{\mathsf{SPDZ}}$ is the SPDZ security bound for the internal operations used in edaBits generation.
-For $B=64$ and ML-DSA-65 parameters ($m=nk=1536$ coefficients, $q=8380417$): the cut-and-choose term $2^{-64}$ is negligible, and $\epsilon^{\prime}_{\mathsf{SPDZ}}\leq m/q<2^{-12}$.
-Thus $\epsilon_{\mathsf{edaBits}}<2^{-64}+2^{-12}<2^{-11.9}$.
+For $B=64$ and ML-DSA-65 parameters ($m=k\cdot N=6\cdot 256=1536$ coefficients, where $N=256$ is the ring degree distinct from party count $n$; $q=8380417$): the cut-and-choose term $2^{-64}$ is negligible, and $\epsilon^{\prime}_{\mathsf{SPDZ}}\leq m/p\leq 1536/2^{64}<2^{-54}$ (using the same 64-bit MAC field; per [EKMOZ20], the SPDZ sub-protocol internal to edaBits opens at most $m$ values per batch).
+Thus $\epsilon_{\mathsf{edaBits}}<2^{-64}+2^{-54}<2^{-53.9}$.
 L.4 Hybrid $\mathsf{H}_{3}$: Simulate Masked Values
@@
 $x+r\bmod q\equiv\mathsf{Uniform}(\mathbb{F}_{q})$
-Proof L.7.
+Proof L.8.
 For any target $t\in\mathbb{F}_{q}$:
@@
 $\Pr[\mathsf{H}_{2}=1]=\Pr[\mathsf{H}_{3}=1]$ (perfect indistinguishability)
-Proof L.9.
+Proof L.10.
 In $\mathsf{H}_{2}$, for each coefficient $j$:
 $\mathsf{masked}_{j}=\mathbf{w}^{\prime}_{j}+r_{j}\bmod q$
-where $r_{j}$ is the edaBit value, uniformly distributed in $\mathbb{F}_{q}$ and independent of all other values (by $\mathcal{F}_{\mathsf{edaBits}}$ security).
+where $r_{j}$ is the edaBit value, uniformly distributed in $\mathbb{F}_{q}$ (by Remark H.8, rejection sampling ensures $r_{j}$ is exactly uniform) and independent of all other values (by $\mathcal{F}_{\mathsf{edaBits}}$ security).
 By Lemma L.7, $\mathsf{masked}_{j}$ is uniformly distributed in $\mathbb{F}_{q}$, regardless of $\mathbf{w}^{\prime}_{j}$.
@@
 For any inputs $x,y\in\{0,1\}$, the opened values $(d,e)=(x\oplus a,y\oplus b)$ are uniformly distributed in $\{0,1\}^{2}$ and independent of $(x,y)$.
-Proof L.11.
+Proof L.12.
 Since $a\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}$ is uniform and independent of $x$:
@@
 $\Pr[\mathsf{H}_{3}=1]=\Pr[\mathsf{H}_{4}=1]$ (perfect indistinguishability)
-Proof L.13.
-Each AND gate reveals $(d,e)$ which is uniform by Lemma L.10. Replacing real AND computation with $\mathcal{F}_{\mathsf{Binary}}$ (which also produces uniform $(d,e)$) yields identical distributions.
+Proof L.14.
+Each AND gate reveals $(d,e)$, which is uniform by Lemma L.11. Replacing real AND computation with $\mathcal{F}_{\mathsf{Binary}}$ (which also produces uniform $(d,e)$) yields identical distributions.
 The A2B conversion and comparison circuits consist entirely of:
@@
 Security implication: These are correctness considerations, not security concerns.
-The security proof (Lemma above) shows that the adversary’s view—the $(d,e)$ values from Beaver triples—is identically distributed regardless of the underlying $\mathbf{w}^{\prime}_{j}$ values.
-Correctness ensures the final result $\mathsf{result}$ is computed correctly; security ensures the adversary learns only $\mathsf{result}$.
+The security proof shows that the adversary’s view — the $(d,e)$ values from Beaver triples — is identically distributed regardless of the underlying $\mathbf{w}^{\prime}_{j}$ values.
+Correctness ensures $\mathsf{result}$ is computed correctly; security ensures the adversary learns only $\mathsf{result}$.
 L.6 Hybrid $\mathsf{H}_{5}$: Backward Simulate AND Tree
@@
 $\Pr[\mathsf{H}_{4}=1]=\Pr[\mathsf{H}_{5}=1]$ (perfect indistinguishability)
-Proof L.16.
+Proof L.17.
 In $\mathsf{H}_{4}$, the AND tree computes $\mathsf{result}=\bigwedge_{j=1}^{m}\mathsf{pass}_{j}$ level by level, revealing Beaver $(d,e)$ values at each gate.
@@
 Case $\mathsf{result}=1$: All $\mathsf{pass}_{j}=1$. For each AND gate $z=x\land y$ with $z=1$: both $x=y=1$. The simulator sets all leaf values to 1 and propagates upward.
-Case $\mathsf{result}=0$: At least one $\mathsf{pass}_{j}=0$. The simulator must assign leaf values such that the AND tree outputs 0. We use the following procedure:
+Case $\mathsf{result}=0$: At least one $\mathsf{pass}_{j}=0$. The simulator does not assign leaf values first. Instead, it proceeds directly as follows:
   1. 1.
-    Pick a uniformly random index $j^{*}\in[m]$ and set $\mathsf{pass}_{j^{*}}=0$.
+    For each AND gate $j$, sample $(d_{j},e_{j})\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{2}$ uniformly and independently (no reference to any leaf assignment).
   2. 2.
-    For all $j\neq j^{*}$, set $\mathsf{pass}_{j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}$ uniformly at random.
-  3. 3.
-    Propagate values through the AND tree (any path containing a 0 produces 0).
-This simulation is not identical to the real distribution of leaf values.
-However, the adversary’s view consists only of Beaver $(d,e)$ pairs, not the leaf values themselves.
-The following argument shows indistinguishability:
-In both cases, the revealed $(d,e)$ values are uniform by Lemma L.10, regardless of the underlying $(x,y)$.
-The simulator samples uniform $(d,e)\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{2}$ for each AND gate, matching the real distribution exactly.
-Key observation: The adversary sees only $(d,e)$ values, never the underlying bits $x,y$ (which remain secret-shared). Since $(d,e)$ is uniform independent of $(x,y)$, backward simulation is perfect.
+    Output the simulated view $\{(d_{j},e_{j})\}_{j=1}^{m}$ along with $\mathsf{result}=0$ from $\mathcal{F}_{r_{0}\text{-check}}$.
+No leaf values need be assigned: the leaf bits $(\mathsf{pass}_{j})$ never appear in the adversary’s view — only the $(d_{j},e_{j})$ pairs do.
+In both cases, the revealed $(d,e)$ values are uniform by Lemma L.11, regardless of the underlying $(x,y)$; the simulator’s sampled $(d_{j},e_{j})$ therefore perfectly matches the real distribution.
+This uniformity extends to the joint distribution over the entire AND tree: each gate’s Beaver triple $(\alpha_{j},\beta_{j},\gamma_{j})$ is sampled independently offline; for gate $j$, $(d_{j},e_{j})=(x_{j}\oplus\alpha_{j},y_{j}\oplus\beta_{j})$ is uniform and independent of $(x_{j},y_{j})$ by Lemma L.11.
+Since triples are pairwise independent across gates, the joint distribution $(d_{1},e_{1},\ldots,d_{m},e_{m})$ is product-uniform, independent of all leaf values.
+Key observation: The adversary sees only $(d_{j},e_{j})$ values; the underlying bits $x_{j},y_{j}$ remain secret-shared.
+Since the joint distribution is product-uniform and independent of the leaf bits, backward simulation is perfect for the entire AND tree.
 L.7 Hybrid $\mathsf{H}_{6}$: Ideal Execution
@@
 $\Pr[\mathsf{H}_{5}=1]=\Pr[\mathsf{H}_{6}=1]$
-Proof L.18.
-$\mathsf{H}_{5}$ is exactly the execution of simulator $\mathcal{S}$ interacting with $\mathcal{F}_{r_{0}\text{-check}}$:
-  • SPDZ replaced with $\mathcal{F}_{\mathsf{SPDZ}}$: done in $\mathsf{H}_{1}$
-  • edaBits replaced with $\mathcal{F}_{\mathsf{edaBits}}$: done in $\mathsf{H}_{2}$
-  • Masked values simulated as random: done in $\mathsf{H}_{3}$
-  • Binary circuits replaced with $\mathcal{F}_{\mathsf{Binary}}$: done in $\mathsf{H}_{4}$
-  • AND tree backward-simulated: done in $\mathsf{H}_{5}$
-This matches the simulator construction exactly. Thus $\mathsf{H}_{5}\equiv\mathsf{H}_{6}$.
+Proof L.19.
+We show that $\mathsf{H}_{5}$ is the ideal execution $\mathsf{H}_{6}$.
+After the previous hybrids, every component of the real protocol has been replaced:
+  • SPDZ arithmetic shares $\to$ $\mathcal{F}_{\mathsf{SPDZ}}$ outputs (H1)
+  • edaBits $\to$ $\mathcal{F}_{\mathsf{edaBits}}$ outputs (H2)
+  • Masked reveal values $\to$ uniform random (H3), so adversary’s view of the masked values is simulated
+  • Binary comparison circuits $\to$ $\mathcal{F}_{\mathsf{Binary}}$ output (H4)
+  • AND tree open values $(d_{j},e_{j})$ $\to$ product-uniform random (H5)
+After all these replacements, the only “real” output remaining is the final result bit output by the AND tree.
+This result bit is computed by $\mathcal{F}_{\mathsf{Binary}}$ (already ideal from H4) and passed to the parties.
+But $\mathcal{F}_{\mathsf{Binary}}$’s output is exactly the output of $\mathcal{F}_{r_{0}\text{-check}}$: it outputs 1 iff $\|\mathbf{r}_{0}\|_{\infty}<\gamma_{2}-\beta$.
+Explicit argument. We verify that the simulator $\mathcal{S}$ from Section K is well-defined and that its output distribution matches $\mathsf{H}_{5}$ component-by-component:
+  • SPDZ shares (H1): $\mathcal{S}$ generates random shares $\tilde{s}_{i,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{F}_{q}$ for honest parties — identical in distribution to $\mathcal{F}_{\mathsf{SPDZ}}$’s output.
+  • edaBits (H2): $\mathcal{S}$ uses $\mathcal{F}_{\mathsf{edaBits}}$ outputs — exact match.
+  • Masked reveal (H3): $\mathcal{S}$ samples $\tilde{m}_{i,j}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{F}_{q}$ — SD =0=0 by one-time-pad argument.
+  • Binary comparison (H4): $\mathcal{S}$ uses $\mathcal{F}_{\mathsf{Binary}}$ outputs — exact match.
+  • AND-tree pairs (H5): $\mathcal{S}$ samples $(d_{j},e_{j})\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{2}$ — SD =0=0 by Lemma L.11.
+  • Result bit: In the ideal execution $\mathsf{H}_{6}$, honest parties submit their actual inputs directly to $\mathcal{F}_{r_{0}\text{-check}}$ via the ideal-world channel (not through $\mathcal{S}$).
+The functionality computes $\mathsf{result}$ and sends it to $\mathcal{S}$.
+The simulator completes the adversary view using $\mathsf{result}$.
+Each component’s simulated distribution matches $\mathsf{H}_{5}$’s distribution exactly.
+The result bit is the only component that differs between $\mathsf{H}_{5}$ (computed by $\mathcal{F}_{\mathsf{Binary}}$, which replaced the real AND-tree in $\mathsf{H}_{4}$) and $\mathsf{H}_{6}$ (provided by $\mathcal{F}_{r_{0}\text{-check}}$).
+But $\mathcal{F}_{\mathsf{Binary}}$’s output equals $\mathcal{F}_{r_{0}\text{-check}}$’s output: both compute $\mathbf{1}[\|\mathbf{r}_{0}\|_{\infty}<\gamma_{2}-\beta]$ on the same inputs.
+Hence the result-bit distributions are identical, and $\Pr[\mathsf{H}_{5}=1]=\Pr[\mathsf{H}_{6}=1]$.
 Appendix M Main Theorem
 Theorem M.1 (UC Security of Profile P2 (Full Proof of Theorem 6.3)).
-Protocol $\Pi_{r_{0}}$ UC-realizes $\mathcal{F}_{r_{0}\text{-check}}$ in the $(\mathcal{F}_{\mathsf{SPDZ}},\mathcal{F}_{\mathsf{edaBits}},\mathcal{F}_{\mathsf{Rand}},\mathcal{F}_{\mathsf{Broadcast}})$-hybrid model against static malicious adversaries corrupting up to N1N-1 parties, achieving security-with-abort.
+Protocol $\Pi_{r_{0}}$ UC-realizes $\mathcal{F}_{r_{0}\text{-check}}$ in the hybrid model with $(\mathcal{F}_{\mathsf{SPDZ}},\allowbreak\mathcal{F}_{\mathsf{edaBits}},\allowbreak\mathcal{F}_{\mathsf{Binary}})$ against static malicious adversaries corrupting up to N1N-1 parties, achieving security-with-abort.
 Concretely, for any PPT environment $\mathcal{Z}$ and adversary $\mathcal{A}$:
-$|\Pr[\mathsf{REAL}_{\Pi_{r_{0}},\mathcal{A},\mathcal{Z}}=1]-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}\text{-check}},\mathcal{S},\mathcal{Z}}=1]|\leq\epsilon$
+$\bigl|\Pr[\mathsf{REAL}_{\Pi_{r_{0}},\mathcal{A},\mathcal{Z}}=1]$ $-\Pr[\mathsf{IDEAL}_{\mathcal{F}_{r_{0}\text{-check}},\mathcal{S},\mathcal{Z}}=1]\bigr|\leq\epsilon$
 where
-$\epsilon=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$
-for ML-DSA-65 parameters.
+$\epsilon=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-53}$
+for ML-DSA-65 parameters with a 64-bit dedicated MAC field (recommended configuration; the minimal configuration using the protocol field gives $\epsilon<2^{-11}$).
 Proof M.2.
@@
 The zero terms arise from perfect (statistical) indistinguishability in Hybrids 3, 4, 5, and 6.
-Concrete bound: For ML-DSA-65 with $q=8380417$ and $m=1536$ coefficients:
-  • $\epsilon_{\mathsf{SPDZ}}\leq O(m)/q<2^{-12}$
-  • $\epsilon_{\mathsf{edaBits}}<2^{-12}$ (with cut-and-choose parameter $B=64$)
-Thus $\epsilon<2^{-11}$, which is negligible in the security parameter.
+Concrete bound (recommended configuration, 64-bit MAC field): For ML-DSA-65 with $q=8380417$, $m=1536$ coefficients, and dedicated MAC field $\mathbb{F}_{p}$ with $|p|=64$ bits:
+  • $\epsilon_{\mathsf{SPDZ}}\leq m/p=1536/2^{64}<2^{-54}$
+  • $\epsilon_{\mathsf{edaBits}}\leq m/p+2^{-64}<2^{-53.9}$ (cut-and-choose $B=64$; $m/p$ is the SPDZ sub-error internal to edaBits)
+Thus $\epsilon=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-54}+2^{-53.9}<2^{-53}$, achieving cryptographically negligible distinguishing advantage.
+Minimal configuration (protocol field as MAC field): $\epsilon_{\mathsf{SPDZ}}<2^{-12.4}$, $\epsilon_{\mathsf{edaBits}}<2^{-11.9}$, total $\epsilon<2^{-11}$; overall signature security relies on Module-SIS at a much higher level (Theorem 4.2).
 Corollary M.3 (Full Protocol P2 Security).
@@
   2. 2.
-    Privacy against coalitions of up to N1N-1 parties (by UC security of $\mathcal{F}_{r_{0}\text{-check}}$: only the result bit is leaked)
+    r0-check privacy against coalitions of up to N1N-1 parties (by UC security of $\mathcal{F}_{r_{0}\text{-check}}$: only the 1-bit pass/fail result is leaked by the r0-check subprotocol).
+Note: overall key privacy and unforgeability require $|C|\leq T-1$; if $|C|\geq T$ the adversary can reconstruct the secret key from $T$ Shamir shares.
+The N1N-1 bound here refers to the r0-check subprotocol’s leakage, not the full signing protocol’s security model.
   3. 3.
@@
 Proof M.4.
-EUF-CMA: The threshold signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ has identical distribution to single-signer ML-DSA (masks cancel, Lemma 3.4).
+EUF-CMA: The threshold signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ is syntactically identical in format to single-signer ML-DSA and is accepted by any compliant FIPS 204 verifier (masks cancel in the sum, Lemma 3.6); the response $\mathbf{z}$ follows an Irwin-Hall rather than uniform distribution, with EUF-CMA security loss $0.013\cdot q_{s}$ bits over $q_{s}$ signing queries (per session: <0.013<0.013 bits; Corollary 4.3).
 The security proof of Theorem 4.2 applies directly.
@@
   • r0-check result bits (one per attempt)
-Since $\mathsf{result}$ is derivable from $\sigma$ (a valid signature implies $\mathsf{result}=1$), no additional information is leaked.
+For successful attempts, $\mathsf{result}=1$ is implied by the existence of $\sigma$ and leaks no new information.
+For failed attempts ($\mathsf{result}=0$, no $\sigma$ produced), the result bit is determined by the public broadcast $(\tilde{c},\mathbf{w})$ and the corrupted parties’ $r_{0}$-check shares $\{c\mathbf{s}_{2,j}\}_{j\in C}$: the simulator computes
+$\mathbf{r}_{0}^{(C)}=\mathsf{HighBits}\!\left(\sum_{j\in C}\lambda_{j}\,c\,\mathbf{s}_{2,j}+\text{honest aggregate contribution}\right)$
+and determines the result by whether $\|\mathbf{r}_{0}\|_{\infty}<\gamma_{2}-\beta$.
+Since the honest contribution is masked (mask-hiding, Lemma 4.15), the simulator learns only the aggregate result from $\mathcal{F}_{r_{0}\text{-check}}$, not the individual honest shares.
+The failed-attempt result bits can thus be provided to the simulator by $\mathcal{F}_{r_{0}\text{-check}}$ without leaking additional information beyond what is already in $\mathcal{Z}$’s view.
 Composition: Follows from the UC composition theorem (Theorem H.5).
+Remark M.5 (Nonce DKG Simulation for P2).
+The P2 UC proof above is scoped to the r0-check subprotocol $\Pi_{r_{0}}$ realizing $\mathcal{F}_{r_{0}\text{-check}}$.
+The nonce DKG simulation for the full P2 protocol follows the same structure as P1/P3+: the simulator generates honest nonce shares from simulated degree-(T1)(T-1) polynomials, producing evaluations with identical distribution to the real protocol.
+Specifically, $\mathcal{S}$ runs the nonce DKG on behalf of honest parties (sampling the high-degree coefficients uniformly from $R_{q}$), provides consistent evaluations to corrupted parties, and feeds the resulting nonce shares into the r0-check simulation above.
+By the UC composition theorem, the full P2 signing protocol (nonce DKG + response computation + r0-check) securely composes these subprotocols.
 Appendix N Communication and Round Complexity
@@
 Comparison 2 $O(m\log\gamma_{2})$ bits
-AND tree 1 $O(\log m)$ bits
+AND tree 1 $O(m)$ bits
 Total 8 $\approx 30$ KB
-Table 14: Detailed round and communication breakdown for Profile P2 r0-check (original 8-round version).
+Table 14: Round and communication breakdown for Profile P2 r0-check (8-round version).
+The AND tree requires m1m-1 Beaver-triple openings, giving $O(m)$ bits; the $O(\log m)$ figure denotes circuit depth, not communication.
 Appendix O Optimized P2: 5-Round Protocol UC Security
@@
     Continue with standard computation as in $\mathcal{F}_{r_{0}\text{-check}}$
+Timeout (Corrupted Combiner):
+Upon sending $(\mathsf{ALL\_COLLECTED},\mathsf{sid})$ to the combiner, the functionality starts a timer of $\Delta_{\mathsf{timeout}}$ rounds (a global synchrony parameter).
+If $(\mathsf{RELEASE},\mathsf{sid})$ is not received within $\Delta_{\mathsf{timeout}}$ rounds:
+  1. 1.
+    Send $(\mathsf{ABORT},\mathsf{sid})$ to all parties $P_{i}$
+  2. 2.
+    Discard all collected state
+Remark: This timeout prevents a corrupted combiner from indefinitely delaying the release.
+In Profiles P1 and P3+, the combiner role is enforced by a TEE or 2PC, so the timeout is triggered only if the trusted component is unavailable (which is detected).
+In Profile P2, all parties are online during the signing window, so $\Delta_{\mathsf{timeout}}$ is a protocol parameter (e.g., 30 seconds).
 Leakage: The combiner learns:
@@
 Proof O.6.
-The combiner’s role is purely message routing with simultaneous release. We show the adversary’s view is identical:
+The combiner performs no cryptographic operations: it collects incoming messages and forwards them once all have arrived (or times out).
+In particular, it does not compute any MACs, encryptions, or protocol-specific transformations on the data.
+Its behavior is a deterministic function of its input messages and the arrival timing.
+We show the adversary’s view is bytewise identical in $\mathsf{H}_{0}$ and $\mathsf{H}_{0.5}$:
 Honest Combiner Case:
-  • In $\mathsf{H}_{0}$: Combiner collects all $(Com,value,rand)$ tuples, releases simultaneously
-  • In $\mathsf{H}_{0.5}$: $\mathcal{F}_{\mathsf{Combiner}}$ does the same
-  • Adversary sees: all commitments (at send time), all openings (at release time)
-  • Timing of release is identical (after all collected)
+  • In $\mathsf{H}_{0}$: The combiner collects $(\mathsf{Com}_{i},\mathsf{value}_{i},\mathsf{rand}_{i})$ tuples from all $N$ parties and releases them simultaneously only after all are received.
+  • In $\mathsf{H}_{0.5}$: $\mathcal{F}_{\mathsf{Combiner}}$ executes the same algorithm: collect-then-release.
+  • The adversary’s incoming messages (the individual party tuples) are produced by the real parties’ code in $\mathsf{H}_{0}$ and by the simulator (which replicates the same code) in $\mathsf{H}_{0.5}$—producing identical bit-strings.
+  • The adversary’s outgoing view (released tuples) is the same set of collected messages, released at the same logical time.
+  • Since the combiner is honest, the adversary’s view of the combiner’s internal state is empty (it receives only the released outputs).
+The released outputs are identical.
+Hence the adversary’s view is bytewise equal in both hybrids.
 Corrupted Combiner Case:
-  • Adversary controls combiner directly
-  • Receives honest parties’ messages as in real protocol
-  • Any deviation (selective release, modification) is detectable:
+  • The adversary controls the combiner directly and receives honest parties’ messages in both hybrids via identical delivery.
+  • Any deviation by the combiner (selective release, modification) triggers a detectable abort:
     1. 1.
-      Commitment mismatch: immediate abort
+      Value modification: SPDZ MAC check at the recipient parties fails with probability 11/q1-1/q per coefficient; honest parties abort and the functionality’s abort interface captures this.
     2. 2.
-      Value modification: SPDZ MAC check fails
-  • Functionality captures this via abort mechanism
-Since the adversary’s view and available actions are identical in both hybrids:
+      Selective withholding: honest parties’ timeout logic triggers an abort after $\mathsf{deadline}$, identical in both $\mathsf{H}_{0}$ and $\mathsf{H}_{0.5}$.
+  • In $\mathsf{H}_{0.5}$, $\mathcal{F}_{\mathsf{Combiner}}$ replicates the same abort interface.
+Since the honest parties’ messages delivered to the corrupted combiner are generated by the same real-protocol code in both hybrids (the combiner’s corruption does not affect honest-party message generation), the adversary’s view is again bytewise identical.
+Since the combiner performs no cryptographic operations and its behavior is a deterministic routing function, the adversary’s view in $\mathsf{H}_{0}$ is perfectly equal to its view in $\mathsf{H}_{0.5}$ (not merely computationally indistinguishable).
+Thus:
 $\Pr[\mathsf{H}_{0}=1]=\Pr[\mathsf{H}_{0.5}=1]$
@@
 Protocol $\Pi_{r_{0}}^{\mathsf{5rnd}}$ UC-realizes $\mathcal{F}_{r_{0}\text{-check}}^{\mathsf{5rnd}}$ with:
-$\epsilon=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}+\epsilon_{\mathsf{combiner}}=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-11}$
+$\epsilon=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}+\epsilon_{\mathsf{combiner}}=\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}<2^{-53}$
+(with 64-bit MAC field; minimal configuration gives $\epsilon<2^{-11}$).
 Proof O.8.
 By Lemma O.5, $\epsilon_{\mathsf{combiner}}=0$. The remaining hybrids proceed exactly as in Theorem M.1. Total security loss:
-$\epsilon=0+\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}+0+0+0+0<2^{-11}$
+$\epsilon=0+\epsilon_{\mathsf{SPDZ}}+\epsilon_{\mathsf{edaBits}}+0+0+0+0<2^{-53}$
 O.5 5-Round Complexity Analysis
@@
   • Round 5 $\to$ Round 3: A2B conversion (unchanged)
-  • Rounds 6-7 $\to$ Round 4: Parallel comparison + AND tree in single round
-  • Round 8 $\to$ Round 5: Output (unchanged)
+  • Rounds 6-7-8 $\to$ Rounds 4-5: Parallel comparison circuits and AND aggregation execute simultaneously in Round 4 (saving 1 round); output is broadcast in Round 5 (total: 3 original rounds $\to$ 2 new rounds)
 Communication unchanged: Same data transmitted, just fewer rounds.
 Preprocessing unchanged: Same edaBits and Beaver triples required.
-Component Rounds (8-rnd) Rounds (5-rnd)
+Component 8-rnd 5-rnd
 Share exchange + aggregation 2 1
@@
 Definition P.2 (Semi-Async Adversary).
-The adversary $\mathcal{A}$ can:
+We adopt the synchronous UC model [Katz07] where all parties share a global clock. The adversary $\mathcal{A}$ can:
   • Corrupt up to T1T-1 signers (static corruption)
   • Corrupt at most one of $\{\mathsf{CP}_{1},\mathsf{CP}_{2}\}$
-  • Control message delivery within each phase
+  • Control message scheduling within each phase, but cannot advance the clock past the signer response deadline $\Delta$
   • Cannot control the global clock or extend $\Delta$
+Theorem P.12 applies to any protocol meeting this interface.
 P.2 Ideal Functionality for P3+
-Functionality P.3 ($\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$ – Semi-Async r0-Check).
-Extends $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ (Definition 6.4) with pre-signing and timing.
+Definition P.3 (Ideal Functionality $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ (UC Proof Compact Form)).
+The functionality $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ captures a two-party computation of the r0-check predicate.
+It receives private input $\mathbf{S}_{j}\in\mathbb{Z}_{q}^{m}$ from each computation party $\mathsf{CP}_{j}$ ($j\in\{1,2\}$), reconstructs $c\mathbf{s}_{2}=\mathbf{S}_{1}+\mathbf{S}_{2}\bmod q$ (private inputs satisfy this additive split; in the concrete protocol, $\mathbf{S}_{j}=\sum_{i\text{ assigned to }\mathsf{CP}_{j}}\lambda_{i}\cdot c\mathbf{s}_{2,i}$), computes $\mathbf{w}^{\prime}=\mathbf{w}-c\mathbf{s}_{2}$ (using the public $\mathbf{w}$ agreed upon during challenge derivation), evaluates $\mathsf{result}=\bigwedge_{j}(|\mathsf{LowBits}(\mathbf{w}^{\prime}_{j},2\gamma_{2})|_{\infty}<\gamma_{2}-\beta)$, and outputs only $\mathsf{result}$ (1 bit) and the hint $\mathbf{h}=\mathsf{MakeHint}(\mathbf{w}^{\prime},\mathbf{w},2\gamma_{2})$ to the combiner.
+Neither $c\mathbf{s}_{2}$ nor $\mathbf{w}^{\prime}$ is revealed outside $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$.
+Functionality P.4 ($\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$ – Semi-Async r0-Check).
+Extends $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ (Definition P.3) with pre-signing and timing.
 Pre-signing Phase:
@@
 P.3 Real Protocol $\Pi_{r_{0}}^{\mathsf{P3+}}$
-Protocol P.4 ($\Pi_{r_{0}}^{\mathsf{P3+}}$ – P3+ Semi-Async Protocol).
+Protocol P.5 ($\Pi_{r_{0}}^{\mathsf{P3+}}$ – P3+ Semi-Async Protocol).
 Offline Phase: Pre-signing
@@
     1. (a)
-      Sample $\mathsf{set\_id}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
+      Sample $\mathsf{set\_id}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$ (pre-signing set identifier)
     2. (b)
@@
     4. (d)
-      Compute commitment $\mathsf{Com}_{i}=H(\mathbf{y}_{i}\|\mathbf{w}_{i}\|\mathsf{set\_id})$
+      Sample commitment randomness $r_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
+    5. (e)
+      Compute commitment $\mathsf{Com}_{i}=H(\texttt{"com"}\|\mathbf{y}_{i}\|\mathbf{w}_{i}\|r_{i})$
   2. 2.
-    Send $(\mathsf{set\_id},\mathbf{w}_{i},\mathsf{Com}_{i})$ to combiner
+    Send $(\mathsf{set\_id},\mathbf{w}_{i},\mathsf{Com}_{i})$ to combiner $\triangleright$ $r_{i}$ is not sent; kept secret by signer
 Offline Phase: GC Precomputation
@@
   1. 1.
-    If $\mathsf{result}=1$: The 2PC circuit also outputs the hint $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},2\gamma_{2})$ (computed inside the circuit using $c\mathbf{s}_{2}=\mathbf{S}_{1}+\mathbf{S}_{2}$); combiner outputs signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$
+    If $\mathsf{result}=1$: The 2PC circuit outputs the hint
+    $\mathbf{h}=\mathsf{MakeHint}(-c\mathbf{t}_{0},\;\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d},\;2\gamma_{2})$
+    (computed inside the circuit using $c\mathbf{s}_{2}=\mathbf{S}_{1}+\mathbf{S}_{2}$); the combiner outputs signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$
   2. 2.
@@
 P.4 Simulator Construction for P3+
-Simulator P.5 ($\mathcal{S}^{\mathsf{P3+}}$ for $\Pi_{r_{0}}^{\mathsf{P3+}}$).
+Simulator P.6 ($\mathcal{S}^{\mathsf{P3+}}$ for $\Pi_{r_{0}}^{\mathsf{P3+}}$).
 $\mathcal{S}$ interacts with $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$ and simulates the adversary’s view.
+Simulating Key DKG:
+  1. 1.
+    $\mathcal{S}$ receives the corrupted parties’ key shares $\{(\mathbf{s}_{1,j},\mathbf{s}_{2,j})\}_{j\in C}$ from the adversary (via corruption; in P3+ the key DKG is modeled as a trusted-dealer setup phase, so the simulator acts as the trusted dealer and generates all shares directly).
+  2. 2.
+    $\mathcal{S}$ runs the ideal key-generation functionality on behalf of all honest parties: $\mathcal{S}$ samples a fresh degree-(T1)(T-1) Shamir polynomial for each component of $(\mathbf{s}_{1},\mathbf{s}_{2})$, consistent with the corrupted-party shares received in step 1, and computes the honest-party evaluations $\{(\mathbf{s}_{1,i},\mathbf{s}_{2,i})\}_{i\notin C}$ directly from these polynomials.
+  3. 3.
+    $\mathcal{S}$ stores all shares $\{(\mathbf{s}_{1,i},\mathbf{s}_{2,i})\}_{i=1}^{N}$ for use in the response simulation.
+  4. 4.
+    Justification: The simulator acts as the DKG simulator, generating the full Shamir polynomials during key setup.
+It does not infer honest shares post-hoc from $\mathbf{t}_{1}$ (which is a lossy commitment and does not uniquely determine $(\mathbf{s}_{1},\mathbf{s}_{2})$).
+Rather, the polynomials are fixed at key-generation time by $\mathcal{S}$’s own choices, and all evaluations $\{(\mathbf{s}_{1,i},\mathbf{s}_{2,i})\}_{i=1}^{N}$ are therefore directly available to $\mathcal{S}$.
+When |C|<T1|C|<T-1, there are multiple polynomials consistent with the corrupted shares; $\mathcal{S}$ uses the one it sampled.
 Simulating Pre-signing (Honest Signers):
@@
     1. (a)
-      Sample random $\tilde{\mathbf{w}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{k\times n}$
+      Sample $\tilde{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathcal{R}_{q}^{\ell}$ (simulated nonce share)
     2. (b)
-      Sample random $\tilde{\mathsf{Com}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$
+      Compute $\tilde{\mathbf{w}}_{i}=\mathbf{A}\cdot\tilde{\mathbf{y}}_{i}$ (same computation as real protocol)
     3. (c)
+      Sample $\tilde{r}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$ (commitment randomness); program the random oracle so that $H(\mathsf{``com^{\prime\prime}}\|\tilde{\mathbf{y}}_{i}\|\tilde{\mathbf{w}}_{i}\|\tilde{r}_{i}):=\tilde{\mathsf{Com}}_{i}$ where $\tilde{\mathsf{Com}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$; store $(\tilde{\mathbf{y}}_{i},\tilde{\mathbf{w}}_{i},\tilde{r}_{i})$ for potential blame-protocol opening
+    4. (d)
       Send $(\mathsf{set\_id},\tilde{\mathbf{w}}_{i},\tilde{\mathsf{Com}}_{i})$ to adversary
   2. 2.
-    Justification: Commitment hides $\mathbf{y}_{i}$ (preimage resistance). Simulated $\mathbf{w}_{i}$ is uniform, same as real (since $\mathbf{y}_{i}$ is uniform and $\mathbf{A}$ is public).
+    Justification: Commitment hides $\mathbf{y}_{i}$ (random oracle).
+Simulated $\tilde{\mathbf{w}}_{i}=\mathbf{A}\cdot\tilde{\mathbf{y}}_{i}$ has the same distribution as the real $\mathbf{w}_{i}=\mathbf{A}\cdot\mathbf{y}_{i}$: the real $\mathbf{y}_{i}=F(i)$ for $i\neq 0$ is $\mathcal{R}_{q}^{\ell}$-uniform because the degree-11-through-(T1)(T\!-\!1) coefficients of $F$ are sampled uniformly (Theorem Q.3), making each non-zero evaluation uniform by linearity; thus $\tilde{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathcal{R}_{q}^{\ell}$ matches this distribution exactly (SD =0=0).
+The stored $\tilde{r}_{i}$ enables blame-protocol commitment opening.
 Simulating Response Phase (Honest Signers):
@@
     2. (b)
-      Sample random $\tilde{\mathbf{z}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{l\times n}$
+      Compute $\mathbf{z}_{i}=\tilde{\mathbf{y}}_{i}+c\cdot\mathbf{s}_{1,i}$ using the nonce shares $\tilde{\mathbf{y}}_{i}$ from the simulated pre-signing phase and key shares $\mathbf{s}_{1,i}$ from the simulated key DKG
     3. (c)
-      Sample random $\tilde{\mathbf{V}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{k\times n}$
+      Compute $\mathbf{V}_{i}=\lambda_{i}c\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$ using the simulated key shares and PRF masks
     4. (d)
-      Send $(\tilde{\mathbf{z}}_{i},\tilde{\mathbf{V}}_{i})$ to adversary (via combiner)
+      Send $(\mathbf{z}_{i},\mathbf{V}_{i})$ to adversary (via combiner)
   2. 2.
-    Justification: By Theorem 4.4 (statistical privacy), the Shamir nonce DKG ensures each honest party’s response $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ has high conditional min-entropy ($\geq 5\times$ secret key entropy for $|S|\leq 17$) from the adversary’s perspective without computational assumptions.
-The $\mathbf{V}_{i}$ values are uniform by PRF-based mask hiding (Lemma 4.9).
-Simulated (uniform) values have negligible statistical distance from real values.
+    Justification: The simulator computes responses using the same deterministic function as the real protocol, applied to the simulated DKG values.
+Since the simulated DKG produces values with identical distribution to the real DKG (Lemma P.8), the resulting $\mathbf{z}_{i}=\tilde{\mathbf{y}}_{i}+c\cdot\mathbf{s}_{1,i}$ is perfectly indistinguishable from the real value (SD =0=0, since $\tilde{\mathbf{y}}_{i}$ and $\mathbf{y}_{i}$ have identical distribution).
+The masked key-share component $\mathbf{V}_{i}=\lambda_{i}c\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$ uses PRF-derived masks $\mathbf{m}_{i}^{(s2)}$; the simulator uses the actual pairwise PRF seeds established during key setup simulation, computing the same mask values $\mathbf{m}_{i}^{(s2)}=\mathsf{PRF}(\mathsf{seed}_{h,i},\mathsf{sid})$ as the real protocol.
+No PRF replacement is needed; the simulation is bit-for-bit identical to the real protocol for this component.
+The overall gap is therefore 0.
 Simulating 2PC (Corrupt $\mathsf{CP}_{1}$):
@@
   2. 2.
-    $\mathcal{S}$ extracts $(\mathbf{w},\mathbf{S}_{1})$ from adversary’s garbled circuit (extractability)
+    $\mathcal{S}$ extracts $\mathbf{S}_{1}$ from adversary’s garbled circuit inputs (using garbled-circuit extractability [LP07]); $\mathbf{w}$ is already public (broadcast by the combiner)
   3. 3.
@@
     1. (a)
-      Generate simulated garbled circuit that evaluates to $\mathsf{result}$
+      Generate a simulated garbled circuit $\tilde{C}$ together with all input wire labels for $(\mathbf{w},\mathbf{S}_{1})$ (garbler’s inputs) and the wire labels for $\mathbf{S}_{2}$ selected by $\mathsf{CP}_{2}$’s OT queries, consistent with the output $\mathsf{result}$, using the BHR12 simulator for the full two-input garbled circuit [BHR12].
+  6. 6.
+    Justification: Garbled circuit simulatability (BHR12 privacy definition): given the output $f(\mathbf{w},\mathbf{S}_{1},\mathbf{S}_{2})$ and the evaluator’s input labels (for $\mathbf{S}_{2}$, known from OT), the BHR12 simulator produces a simulated circuit $\tilde{C}$ plus all evaluator-side input labels computationally indistinguishable from the real garbled circuit.
+The garbler’s (CP1’s) input wire labels for $(\mathbf{w},\mathbf{S}_{1})$ are included in the simulated output; these are internally consistent with $\tilde{C}$ by construction.
+Simulating 2PC (Both CPs Honest): When both $\mathsf{CP}_{1}$ and $\mathsf{CP}_{2}$ are honest, the adversary observes only:
+  • Pre-signing commitments $\mathsf{Com}_{i}$ (simulated uniformly above, SD =0=0).
+  • The broadcast challenge $c$ (determined by the ideal functionality).
+  • Responses $(\mathbf{z}_{i},\mathbf{V}_{i})$ for honest $i\notin C$ (simulated above: $\mathbf{z}_{i}$ with SD =0=0; $\mathbf{V}_{i}$ with SD =0=0 since the simulator uses the actual PRF seeds established during key setup, producing bit-for-bit identical mask values).
+  • The final result bit (provided by $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$).
+The 2PC OT and GC transcripts are exchanged solely between the two honest CPs and are never delivered to the adversary.
+$\mathcal{S}$ need not simulate these messages; the hybrid sequence $\mathsf{H}_{0}\to\mathsf{H}_{1}$ (replace OT) $\to\mathsf{H}_{2}$ (replace GC) captures the honest-CP case via ideal-functionality substitution.
+Simulating Timeout / Abort:
+  1. 1.
+    If the adversary delays honest signers’ responses past the deadline $\mathsf{deadline}$ (e.g., by withholding the challenge or blocking network messages):
+    1. (a)
+      $\mathcal{S}$ detects no challenge delivered to honest signers within $\mathsf{deadline}-\Delta$ rounds.
     2. (b)
-      This is possible by simulatability of garbled circuits [BHR12]
-  6. 6.
-    Justification: Garbled circuit simulatability: given output $f(x)$, can simulate garbled circuit indistinguishable from real.
+      $\mathcal{S}$ sends $(\mathsf{ABORT},\mathsf{sid})$ to $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$.
+    3. (c)
+      $\mathcal{S}$ delivers $(\mathsf{abort},\mathsf{sid})$ to all honest parties in the simulated execution.
+  2. 2.
+    Justification: The functionality enforces “if $\mathsf{now}>\mathsf{deadline}$: reject” (see Response Phase, condition 1).
+When the adversary induces timeout, both ideal and simulated executions abort identically; the simulated abort message matches the real abort, giving indistinguishability.
 P.5 Hybrid Argument for P3+
@@
   • $\mathsf{H}_{2}$: Replace GC with $\mathcal{F}_{\mathsf{GC}}$
-  • $\mathsf{H}_{3}$: Simulate pre-signing as random
-  • $\mathsf{H}_{4}$: Simulate responses as random
+  • $\mathsf{H}_{3}$: Simulate pre-signing (sample $\tilde{\mathbf{y}}_{i}$, compute $\tilde{\mathbf{w}}_{i}=\mathbf{A}\tilde{\mathbf{y}}_{i}$)
+  • $\mathsf{H}_{4}$: Simulate responses from $\tilde{\mathbf{y}}_{i}$ and simulated key shares
   • $\mathsf{H}_{5}$: Ideal execution with $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$
-Lemma P.7 (Pre-signing Simulation).
-$\Pr[\mathsf{H}_{2}=1]=\Pr[\mathsf{H}_{3}=1]$ (perfect indistinguishability)
-Proof P.8.
-Pre-signing values visible to adversary:
-  • $\mathbf{w}_{i}=\mathbf{A}\cdot\mathbf{y}_{i}$: Since $\mathbf{y}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}^{l\times n}$ uniform and $\mathbf{A}$ is full rank (with overwhelming probability), $\mathbf{w}_{i}$ is uniform over $\mathbb{Z}_{q}^{k\times n}$.
-  • $\mathsf{Com}_{i}=H(\mathbf{y}_{i}\|\mathbf{w}_{i}\|\mathsf{set\_id})$: By random oracle model, $\mathsf{Com}_{i}$ is uniform $\{0,1\}^{256}$.
-Simulated values $(\tilde{\mathbf{w}}_{i},\tilde{\mathsf{Com}}_{i})$ are sampled uniformly from same distributions. Thus:
+Lemma P.8 (Pre-signing Simulation (ROM)).
+$\Pr[\mathsf{H}_{2}=1]=\Pr[\mathsf{H}_{3}=1]$ (perfect indistinguishability in the random oracle model)
+Proof P.9.
+We work in the UC random oracle model: $H$ is modeled as an ideal functionality $\mathcal{F}_{\mathsf{RO}}$ (following [Canetti01, BR93]).
+The simulator $\mathcal{S}^{\mathsf{P3+}}$ programs $\mathcal{F}_{\mathsf{RO}}$ at honest parties’ pre-image inputs: when the simulator generates $\tilde{\mathsf{Com}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$, it programs $H(\tilde{\mathbf{y}}_{i}\|\mathbf{A}\tilde{\mathbf{y}}_{i}\|\mathsf{set\_id}):=\tilde{\mathsf{Com}}_{i}$ at the simulated pre-image (the adversary does not know $\tilde{\mathbf{y}}_{i}$ and cannot query $H$ on this input without knowing the simulation randomness).
+Pre-signing values visible to the adversary:
+  • $\mathsf{Com}_{i}=H(\mathbf{y}_{i}\|\mathbf{w}_{i}\|\mathsf{set\_id})$: In the ROM, $\mathsf{Com}_{i}$ is uniform over $\{0,1\}^{256}$ regardless of $\mathbf{y}_{i}$’s distribution, since the adversary does not know $\mathbf{y}_{i}$ (or $\mathbf{w}_{i}=\mathbf{A}\mathbf{y}_{i}$) and cannot query $\mathcal{F}_{\mathsf{RO}}$ at the pre-image without knowing the honest party’s nonce.
+(Note: in the real protocol with $T\geq 2$, $\mathbf{y}_{i}=\mathbf{F}(i)$ is $\mathbb{Z}_{q}$-uniform: the high-degree DKG coefficients (degree $\geq 1$) are sampled $R_{q}$-uniformly, and for evaluation point $i\neq 0$ their contribution is uniform by additive mixing, giving $\mathbf{y}_{i}\sim\mathbb{Z}_{q}$-uniform exactly.
+When $T=1$ there are no high-degree terms and $\mathbf{y}_{i}=\mathbf{y}_{0}$ is the constant nonce, but this is a degenerate threshold that is excluded by the protocol requirement $T\geq 2$.
+The Irwin-Hall distribution applies to the aggregate nonce $\mathbf{y}=\mathbf{F}(0)=\sum_{k}\hat{\mathbf{y}}_{k}$, not to individual shares $\mathbf{y}_{i}=\mathbf{F}(i)$.) The simulated $\tilde{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\mathbb{Z}_{q}$ and $\tilde{\mathsf{Com}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{0,1\}^{256}$ therefore match the real distributions exactly (SD =0=0).
+  • $\mathbf{W}_{i}=\lambda_{i}\mathbf{w}_{i}+\mathbf{m}_{i}^{(w)}$ (masked commitment sent to CP combiner): In P3+, $\mathbf{W}_{i}$ is transmitted only to the CP combiner, not broadcast to other parties.
+By the 1-of-2 CP honest assumption, at least one CP is honest; in this proof we assume the adversary does not control the combiner’s input-aggregation role (if a CP is corrupted, the relevant 2PC simulation case handles it below).
+Since $\mathbf{W}_{i}$ never reaches the adversary, no mask-hiding argument is required.
+The value is hidden entirely by the honest combiner.
+(Note: Lemma 4.15 applies to the P2 broadcast model where $\mathbf{W}_{i}$ is sent to all parties; it does not apply here.)
+Thus:
 $\Pr[\mathsf{H}_{2}=1]=\Pr[\mathsf{H}_{3}=1]$
 Lemma P.10 (Response Simulation).
-$|\Pr[\mathsf{H}_{3}=1]-\Pr[\mathsf{H}_{4}=1]|\leq\mathsf{negl}(\kappa)$ (statistical indistinguishability)
-Proof P.10.
-For honest signer $P_{i}$:
+$\Pr[\mathsf{H}_{3}=1]=\Pr[\mathsf{H}_{4}=1]$ (perfect indistinguishability)
+Proof P.11.
+The simulator computes each honest party’s response deterministically from its simulated DKG values, producing an identical distribution to the real protocol:
 $\mathbf{z}_{i}$ $=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$
 $\mathbf{V}_{i}$ $=\lambda_{i}c\mathbf{s}_{2,i}+\mathbf{m}_{i}^{(s2)}$
-By Theorem 4.4 (statistical privacy), the Shamir nonce DKG ensures that the adversary observes only T1T-1 evaluations of each honest party’s degree-(T1)(T\!-\!1) nonce polynomial, leaving one degree of freedom that provides high conditional min-entropy.
-Due to the bounded-nonce constraint, $\mathbf{z}_{i}$ is not perfectly $R_{q}^{\ell}$-uniform but has min-entropy $\geq n\ell\cdot\log_{2}(2\gamma_{1}/|S|)$ bits, vastly exceeding the secret key entropy.
-The $\mathbf{V}_{i}$ values contain pairwise masks $\mathbf{m}_{i}^{(s2)}$ derived from PRF seeds shared only with other honest signers; against an adversary controlling fewer than all honest signers, at least one honest signer contributes an unknown mask component that acts as a one-time pad.
-Simulated $\tilde{\mathbf{z}}_{i},\tilde{\mathbf{V}}_{i}$ are sampled uniformly, matching the real distribution.
+where $\mathbf{y}_{i}$ are the nonce shares from the simulated nonce DKG (Lemma P.8) and $\mathbf{s}_{1,i},\mathbf{s}_{2,i}$ are the key shares from the simulated key DKG.
+Since the simulator possesses the actual simulated values (not just their distributions), it computes $\mathbf{z}_{i}$ and $\mathbf{V}_{i}$ using the same deterministic function as the real protocol; the resulting joint distribution is identical to the real execution.
+For $\mathbf{V}_{i}$, the simulator computes the masked r0-check share using the simulated key shares and the PRF masks (derived from the same seeds used in earlier hybrids).
+Perfect indistinguishability follows: $\Pr[\mathsf{H}_{3}=1]=\Pr[\mathsf{H}_{4}=1]$.
 Theorem P.12 (P3+ UC Security).
 Protocol $\Pi_{r_{0}}^{\mathsf{P3+}}$ UC-realizes $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$ in the $(\mathcal{F}_{\mathsf{OT}},\mathcal{F}_{\mathsf{GC}})$-hybrid model against static malicious adversaries corrupting up to T1T-1 signers and at most one of $\{\mathsf{CP}_{1},\mathsf{CP}_{2}\}$, with security-with-abort.
+Key DKG is handled by a trusted dealer (modeled as a trusted setup phase); the simulator acts as the trusted dealer and directly generates all key shares.
 Concretely:
-$|\Pr[\mathsf{REAL}=1]-\Pr[\mathsf{IDEAL}=1]|\leq\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{GC}}<2^{-\kappa}$
-where $\kappa$ is the computational security parameter.
-Proof P.12.
+$|\Pr[\mathsf{REAL}=1]-\Pr[\mathsf{IDEAL}=1]|\leq\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{ext}}<2^{-\kappa}$
+where $\kappa$ is the computational security parameter, $\epsilon_{\mathsf{OT}}$ is the OT error, $\epsilon_{\mathsf{GC}}$ is the garbled-circuit simulation error (over honest garbler [BHR12]), and $\epsilon_{\mathsf{ext}}\leq 2^{-B/2}$ is the cut-and-choose extraction error when $\mathsf{CP}_{1}$ is corrupted (LP07 extractability [LP07] with $B$ circuits, $B/2$ opened for verification).
+Proof P.13.
 By the hybrid argument:
-$|\Pr[\mathsf{H}_{0}]-\Pr[\mathsf{H}_{5}]|$ $\leq\sum_{i=0}^{4}|\Pr[\mathsf{H}_{i}]-\Pr[\mathsf{H}_{i+1}]|$
-$=\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{GC}}+0+0+0$
+$|\Pr[\mathsf{H}_{0}]-\Pr[\mathsf{H}_{5}]|$ $\leq\sum_{i=0}^{4}|\Pr[\mathsf{H}_{i}]-\Pr[\mathsf{H}_{i+1}]|+|\Pr[\mathsf{H}_{2.5,\mathsf{ext}}]-\Pr[\mathsf{H}_{3}]|$
+$=\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{GC}}+\epsilon_{\mathsf{ext}}+0+0+0$
 $<2^{-\kappa}$
-The zero terms are from perfect indistinguishability (Lemmas P.7 and P.9).
+The hybrid $\mathsf{H}_{2.5,\mathsf{ext}}$ (between $\mathsf{H}_{2}$ and $\mathsf{H}_{3}$) is the extraction hybrid: when $\mathsf{CP}_{1}$ is corrupted, the simulator uses LP07 cut-and-choose extractability to extract $\mathbf{S}_{1}$ from the adversarially-submitted garbled circuit; this succeeds except with probability $\epsilon_{\mathsf{ext}}\leq 2^{-B/2}$ (see the extraction error accounting below).
+The three subsequent zero terms arise from perfect indistinguishability: pre-signing simulation (Lemma P.8), response simulation (Lemma P.10), and the transition to the ideal execution.
+Hint output and combiner role.
+The ideal functionality $\mathcal{F}_{r_{0}}^{\mathsf{2PC}}$ (Definition P.3) outputs both the 1-bit result and the hint $\mathbf{h}=\mathsf{MakeHint}(\mathbf{w}^{\prime},\mathbf{w},2\gamma_{2})$ to the combiner.
+In both simulator cases, $\mathcal{S}^{\mathsf{P3+}}$ receives $(\mathsf{result},\mathbf{h})$ from $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$ and passes $\mathbf{h}$ to the adversary.
+The hint $\mathbf{h}$ appears in the public signature $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ and creates no leakage beyond what $\sigma$ already reveals; the BHR12 simulator constructs a garbled circuit consistent with output $(\mathsf{result},\mathbf{h})$ without exposing $c\mathbf{s}_{2}$.
+Combiner and CP roles.
+In P3+, $\mathsf{CP}_{1}$ (the garbler) also acts as the combiner: it aggregates signer responses $(\mathbf{z}_{j},\mathbf{V}_{j})$, computes the split $(S_{1},S_{2})$ for the 2PC, and executes the garbled-circuit protocol with $\mathsf{CP}_{2}$.
+Combiner corruption is subsumed by the $\mathsf{CP}_{1}$ is corrupted” case (Simulator P.6): the simulator uses LP07 extraction to recover $\mathbf{S}_{1}$ from the adversary’s garbled circuit and forwards honest parties’ contributions via the ideal-world channel to $\mathcal{F}_{r_{0}}^{\mathsf{semi\text{-}async}}$.
+Extraction error accounting.
+When $\mathsf{CP}_{1}$ (garbler) is corrupted, $\mathcal{S}^{\mathsf{P3+}}$ uses LP07 extractability to recover $\mathbf{S}_{1}$ from the adversarially-submitted garbled circuit.
+The cut-and-choose protocol (generate $B$ circuits; open $\lfloor B/2\rfloor$ for verification; use the rest) ensures a cheating garbler is caught with probability $1-2^{-\lfloor B/2\rfloor}\geq 1-2^{-B/2}$.
+For unopened circuits, LP07 extraction succeeds with probability $1$ (requires only correct evaluation, guaranteed by cut-and-choose).
+Thus $\epsilon_{\mathsf{ext}}\leq 2^{-B/2}$; setting $B=2\kappa$ gives $\epsilon_{\mathsf{ext}}\leq 2^{-\kappa}$.
+The full bound $\epsilon_{\mathsf{OT}}+\epsilon_{\mathsf{GC}}+2^{-B/2}<2^{-\kappa+1}$ holds for standard $\kappa$-bit-secure OT and GC.
 P.6 Security Properties of P3+
@@
   3. 3.
-    Privacy: Response shares $\mathbf{z}_{i}$ enjoy statistical privacy (Theorem 4.4) even with a single honest signer; commitment and r0-check values hidden by pairwise masks if $\geq 2$ honest signers
+    Privacy: Nonce shares $\mathbf{y}_{i}$ enjoy statistical privacy (Theorem 4.7); key privacy given $\sigma$ is conjectured to hold computationally under Module-LWE (open problem); commitment and r0-check values hidden by pairwise masks if $\geq 2$ honest signers
   4. 4.
     Liveness: Completes if $T$ signers respond within $\Delta$ and both CPs available
-Proof P.14.
+Proof P.15.
 EUF-CMA: Pre-signing does not change the signature distribution. The commitment binding ensures signers commit to nonces before seeing challenges. Reduction to Module-SIS identical to P3.
 UC Security: Theorem P.12.
-Privacy: By Lemma P.9, honest signers’ responses have high conditional min-entropy (statistical, via Shamir nonce DKG).
+Privacy: By Lemma P.10, the simulator computes honest signers’ responses from simulated DKG values (perfect indistinguishability).
+Nonce shares have high conditional min-entropy (Theorem 4.7, statistical); key privacy given $\sigma$ is conjectured to hold computationally under Module-LWE (formal reduction is an open problem, unresolved even in the single-signer setting).
 The $\mathbf{V}_{j}$ values are computationally hidden by pairwise masks.
-The 2PC reveals only the result bit.
-Combining: adversary view simulatable from public values + corrupted shares + result bit.
-Liveness: The protocol has explicit timeout.
-If conditions met, 2PC completes in bounded time (synchronous between CPs).
-If conditions not met, protocol aborts (no liveness violation, just unsuccessful signing).
+The 2PC reveals the result bit and hint $\mathbf{h}$ (both appear in the public signature); $\mathbf{h}$ does not create additional privacy leakage beyond what $\sigma$ already reveals.
+Combining: adversary view simulatable from public values + corrupted shares + $(\mathsf{result},\mathbf{h})$.
+Liveness: The protocol has an explicit timeout.
+If conditions are met, the 2PC completes in bounded time (synchronous between CPs).
+If not, the protocol aborts (no liveness violation; just unsuccessful signing).
 P.7 P3+ Round Complexity
@@
     Expected total rounds: 7.4 (2 rounds $\times$ 3.7 attempts)
-Proof P.16.
-Signer experience: After receiving challenge, signer computes response (local) and sends once. No further interaction required.
-Server interaction: CPs execute 2PC (constant rounds with GC precomputation). With precomputation, effectively 1 round (OT + evaluate).
-Retry handling: Each retry requires fresh pre-signing set and response. Signer daemon handles automatically; user sees only final result.
+Proof P.17.
+Signer experience: After receiving the challenge, the signer computes a local response and sends it once. No further interaction is required.
+Server interaction: CPs execute 2PC in constant rounds (with GC precomputation), effectively 1 round (OT + evaluate).
+Retry handling: Each retry requires a fresh pre-signing set and response; the signer daemon handles this automatically.
 Total: $2\times 3.7\approx 7.4$ expected rounds (vs. $5\times 3.5=17.5$ for P2).
-Profile Rounds/Attempt Expected Attempts Total Rounds
+Profile Rds/Attempt Exp. Attempts Total Rds
 P2 (8-round) 8 3.5 28
@@
 P3+ 2 3.7 7.4
-Table 16: Round complexity comparison across profiles.
-P3 denotes the synchronous 2PC variant (5 rounds) prior to the semi-async optimization; P3+ is the final semi-async protocol presented in Section 6.5.
+Table 16: Round complexity across profiles. P3: synchronous 2PC variant (5 rounds) before the semi-async optimization; P3+ is the final protocol (Section 6.5).
 Appendix Q Technical Lemmas
@@
 Lemma Q.1 (Mask Cancellation – Formal Statement).
-Let $S\subseteq[N]$ be a signing set with $|S|\geq 2$. For any domain separator $\mathsf{dom}$ and PRF $\mathsf{PRF}:\{0,1\}^{256}\times\{0,1\}^{*}\to R_{q}^{d}$, define:
-$\mathbf{m}_{i}=\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{dom})$
-Then $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}\in R_{q}^{d}$.
+Let $S\subseteq[N]$ be a signing set with $|S|\geq 2$.
+For any PRF input $\mathsf{ctx}\in\{0,1\}^{*}$ (typically $\mathsf{ctx}=\mathsf{session\_id}\|\mathsf{dom}$, where $\mathsf{session\_id}$ is a fresh per-session identifier and $\mathsf{dom}$ is a fixed domain label) and PRF $\mathsf{PRF}:\{0,1\}^{256}\times\{0,1\}^{*}\to R_{q}^{k}$ (where $k=6$ is the module rank; not the dropping-bits parameter $d=13$), define:
+$\mathbf{m}_{i}=\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{ctx})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{ctx})$
+Then $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}\in R_{q}^{k}$.
 Proof Q.2.
 We prove this by showing each PRF term appears exactly once with coefficient +1+1 and once with coefficient 1-1.
-For any pair $(i,j)\in S\times S$ with $i<j$, the term $\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})$ appears:
+For any pair $(i,j)\in S\times S$ with $i<j$, the term $\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{ctx})$ appears:
   1. 1.
@@
 Thus:
-$\sum_{i\in S}\mathbf{m}_{i}$ $=\sum_{i\in S}\left(\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{dom})\right)$
-$=\sum_{\begin{subarray}{c}(i,j)\in S\times S\\ i<j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})-\sum_{\begin{subarray}{c}(i,j)\in S\times S\\ i<j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{dom})$
+$\sum_{i\in S}\mathbf{m}_{i}$ $=\sum_{i\in S}\left(\sum_{j\in S:j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{ctx})-\sum_{j\in S:j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{ctx})\right)$
+$=\sum_{\begin{subarray}{c}(i,j)\in S\times S\\ i<j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{ctx})-\sum_{\begin{subarray}{c}(i,j)\in S\times S\\ i<j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{ctx})$
 $=\mathbf{0}$
@@
 Let $S$ be a signing set with $|S|\geq T$, let $C\subset S$ with $|C|\leq T-1$ be the set of corrupted parties, and let $h\in S\setminus C$ be an honest party.
-In the Shamir nonce DKG protocol, the adversary’s view of $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ has conditional min-entropy:
-$H_{\infty}\!\left(\mathbf{z}_{h}\mid\{\mathbf{f}_{j}\}_{j\in C},\;\{\mathbf{f}_{h}(j)\}_{j\in C},\;c\right)\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
-No computational assumptions are needed.
-Note that the conditioning does not include individual $\mathbf{W}_{i}=\lambda_{i}\mathbf{A}\mathbf{y}_{i}+\mathbf{m}_{i}^{(w)}$ values, which are sent only to the coordinator (not broadcast) in the protocol.
-Including the coordinator’s broadcast $\mathbf{w}_{1}=\mathsf{HighBits}(\sum_{i}\mathbf{W}_{i},2\gamma_{2})$ does not reduce the bound, since $\mathbf{w}_{1}$ is already implicit in any valid signature.
-For ML-DSA-65 with $|S|\leq 17$, this exceeds 24,00024{,}000 bits—over $5\times$ the secret key entropy $H(\mathbf{s}_{1,h})\leq 4058$ bits.
+In the Shamir nonce DKG protocol (Algorithm 2), each party samples its constant-term contribution $\hat{\mathbf{y}}_{h}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$ and constructs a degree-(T1)(T-1) polynomial with uniform higher-degree coefficients.
+The honest party’s nonce share $\mathbf{y}_{h}=\sum_{i\in S}\mathbf{f}_{i}(h)$ has conditional min-entropy:
+$H_{\infty}\!\left(\mathbf{y}_{h}\mid\{\mathbf{f}_{j}\}_{j\in C},\;\{\mathbf{f}_{h}(j)\}_{j\in C}\right)\;\geq\;n\ell\cdot\log_{2}\!\bigl(2\lfloor\gamma_{1}/|S|\rfloor+1\bigr)$
+where the inequality holds directly from the sampling procedure (no rejection conditioning needed): the bounded sampling range for each party’s constant term directly ensures that the honest party’s polynomial $\mathbf{f}_{h}$ has $\mathbf{f}_{h}(0)=\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$ by construction (Algorithm 2, Line 2).
+No computational assumptions are needed for the nonce privacy; it is purely statistical.
+For ML-DSA-65 with $|S|\leq 17$, this exceeds 20,00020{,}000 bits—over $5\times$ the short secret key entropy $H(\mathbf{s}_{1})\leq 4058$ bits.
+(Shamir shares $\mathbf{s}_{1,h}$ for $h\neq 0$ are $\mathbb{Z}_{q}$-uniform; the entropy comparison is against the short secret $\mathbf{s}_{1}=\mathbf{f}(0)$.)
+Note on aggregated nonce $\mathbf{y}=\sum_{i}\hat{\mathbf{y}}_{i}$.
+The aggregated nonce $\mathbf{y}$ (sum of all parties’ constant terms) follows an Irwin-Hall distribution with tails extending beyond $\gamma_{1},\gamma_{1}]$.
+However, the aggregated nonce is never required to lie strictly within this range for the min-entropy bound to hold: the bound applies to individual shares $\mathbf{y}_{h}$, not the aggregate.
+The z-bound check $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ implicitly rejects sessions where the aggregate $\|\mathbf{y}\|_{\infty}$ is too large (via $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$), but this rejection event is already accounted for in the Irwin-Hall tail probability $\epsilon<10^{-30}$ (Lemma Q.22); it does not affect the per-share conditional min-entropy.
+Key privacy given $\sigma$ (computational, informal).
+Even given the public signature $\sigma$ (from which $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ is computable when $|S\setminus C|=1$), extracting $\mathbf{s}_{1,h}$ requires solving a bounded-distance decoding problem with sparse multiplier $c$.
+While not a standard Module-LWE instance, this is at least as hard as key recovery from single-signer ML-DSA signatures; see the discussion in Theorem 4.7 (Section 4).
 Proof Q.4.
-The proof proceeds in four steps, following a null space argument on the honest party’s polynomial.
+The proof proceeds in three steps, following a null space argument on the honest party’s polynomial.
+Reduction to the worst case |C|=T1|C|=T-1.
+The stated min-entropy lower bound is monotone non-decreasing as $|C|$ decreases: fewer corrupted evaluations give a higher-dimensional null space (dimension T|C|>1T-|C|>1 when |C|<T1|C|<T-1), meaning the adversary is more uncertain about $\mathbf{y}_{h}$ (higher conditional min-entropy).
+Hence it suffices to prove the bound for the hardest case |C|=T1|C|=T-1 (minimum uncertainty for the adversary); the case |C|<T1|C|<T-1 gives at least as much privacy and follows a fortiori.
+We henceforth assume |C|=T1|C|=T-1.
 The adversary’s knowledge about the honest party’s polynomial $\mathbf{f}_{h}$ consists of the evaluations $\{\mathbf{f}_{h}(j)\}_{j\in C}$ received during the DKG.
-Since $|C|\leq T-1$ and $\mathbf{f}_{h}$ has degree T1T-1 (hence $T$ coefficients), the adversary has at most T1T-1 linear equations in $T$ unknowns.
+With |C|=T1|C|=T-1 and $\mathbf{f}_{h}$ having degree T1T-1 (hence $T$ coefficients), the adversary has exactly T1T-1 linear equations in $T$ unknowns: the null space of the evaluation map has dimension exactly $1$.
 Step 1: Null space.
-The evaluation map $\mathbf{f}_{h}\mapsto(\mathbf{f}_{h}(j_{1}),\ldots,\mathbf{f}_{h}(j_{|C|}))$ has a null space containing the polynomial $\mathbf{n}(x)=\prod_{j\in C}(x-x_{j})$, which has degree $|C|\leq T-1$ and leading coefficient 1.
-The set of polynomials consistent with the adversary’s observations is $\mathbf{f}_{h}(x)=\mathbf{g}(x)+\mathbf{t}\cdot\mathbf{n}(x)$, where $\mathbf{g}$ is any fixed interpolant through the known points and $\mathbf{t}\in R_{q}^{\ell}$ is the free parameter.
+The evaluation map $\mathbf{f}_{h}\mapsto(\mathbf{f}_{h}(j_{1}),\ldots,\mathbf{f}_{h}(j_{T-1}))$ (with |C|=T1|C|=T-1) has a one-dimensional null space, spanned by the monic polynomial $\mathbf{n}(x)=\prod_{j\in C}(x-x_{j})$ of degree T1T-1 and leading coefficient $1$.
+The set of polynomials consistent with the adversary’s observations is $\mathbf{f}_{h}(x)=\mathbf{g}(x)+\mathbf{t}\cdot\mathbf{n}(x)$, where $\mathbf{g}$ is any fixed interpolant through the known points and $\mathbf{t}\in R_{q}^{\ell}$ is the unique free parameter.
 Step 2: Posterior constraint on $\mathbf{t}$.
-The coefficient $\mathbf{a}_{h,T-1}$ is sampled uniformly from $R_{q}^{\ell}$ a priori.
-However, the constant term $\hat{\mathbf{y}}_{h}=\mathbf{f}_{h}(0)=\mathbf{g}(0)+\mathbf{t}\cdot\mathbf{n}(0)$ must satisfy $\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$.
-Since $\mathbf{n}(0)\neq 0\pmod{q}$, this constrains each coefficient of $\mathbf{t}$ to at most $2\lfloor\gamma_{1}/|S|\rfloor+1$ values.
-The posterior distribution of $\mathbf{t}$ is not $\mathbb{Z}_{q}$-uniform, but retains per-coefficient support of this size.
+In the nonce DKG, each honest party $h$ samples $\hat{\mathbf{y}}_{h}$ directly from $\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$ and all T1T-1 higher-degree coefficients $\mathbf{a}_{h,1},\ldots,\mathbf{a}_{h,T-1}$ independently and uniformly from $R_{q}^{\ell}$.
+In the null-space parameterization, the free parameter is $\mathbf{t}=\mathbf{a}_{h,T-1}-\mathbf{g}_{T-1}$, where $\mathbf{g}_{T-1}$ is the leading coefficient of the fixed interpolant $\mathbf{g}$.
+Because $\mathbf{a}_{h,T-1}$ is $\mathbb{Z}_{q}$-uniform a priori, $\mathbf{t}$ is $\mathbb{Z}_{q}$-uniform per coefficient.
+The constant term $\hat{\mathbf{y}}_{h}=\mathbf{f}_{h}(0)=\mathbf{g}(0)+\mathbf{t}\cdot\mathbf{n}(0)$ must satisfy $\hat{\mathbf{y}}_{h}\in[-\lfloor\gamma_{1}/|S|\rfloor,\,\lfloor\gamma_{1}/|S|\rfloor]^{n\ell}$.
+Since $\mathbf{n}(0)=\prod_{j\in C}(-j)\neq 0\pmod{q}$ (each $j\in C$ satisfies $1\leq j\leq N\leq q-1$, so $j\not\equiv 0\pmod{q}$, and the product of nonzero field elements is nonzero), multiplication by $\mathbf{n}(0)$ is a bijection on $\mathbb{Z}_{q}$.
+Since the Shamir evaluation map acts independently on each of the $n\ell=1280$ coefficient slots (each coefficient of $\mathbf{f}_{h}$ follows a separate scalar Shamir sharing over $\mathbb{Z}_{q}$), this bijection applies coordinate-by-coordinate across all $n\ell$ slots; the constraint selects exactly $2\lfloor\gamma_{1}/|S|\rfloor+1$ values per coefficient of $\mathbf{t}$.
+Because the prior is $\mathbb{Z}_{q}$-uniform, conditioning on this subset yields a uniform posterior over the valid values.
 Step 3: Min-entropy at honest point.
@@
 The nonce share $\mathbf{y}_{h}=\sum_{j\in C}\mathbf{f}_{j}(x_{h})+\mathbf{f}_{h}(x_{h})$ shifts by a known constant, preserving support.
-Step 4: Response.
-The response $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ shifts again by a fixed quantity, yielding $H_{\infty}(\mathbf{z}_{h}\mid\mathsf{View}_{C})\geq n\ell\cdot\log_{2}(2\lfloor\gamma_{1}/|S|\rfloor+1)$.
+Key privacy (computational).
+Given the public signature $\sigma$, the adversary can compute $\mathbf{z}_{h}$ (when $|S\setminus C|=1$).
+The equation $\mathbf{z}_{h}=\mathbf{y}_{h}+c\cdot\mathbf{s}_{1,h}$ relates two bounded unknowns.
+The high min-entropy of $\mathbf{y}_{h}$ ensures this is a non-trivial bounded-distance decoding instance: extracting $\mathbf{s}_{1,h}$ is computationally hard under the Module-LWE assumption.
 Lemma Q.5 (Mask Hiding – Formal Statement (for $\mathbf{V}_{i}$ and $\mathbf{W}_{i}$)).
@@
 Note: This lemma applies to the commitment and r0-check values, which still use pairwise-canceling masks.
-The response $\mathbf{z}_{j}=\mathbf{y}_{j}+c\cdot\mathbf{s}_{1,j}$ achieves statistical privacy via the Shamir nonce DKG (Theorem Q.3) and does not require masks.
+The response $\mathbf{z}_{j}=\mathbf{y}_{j}+c\cdot\mathbf{s}_{1,j}$ does not require masks: the underlying nonce share $\mathbf{y}_{j}$ has statistical privacy via the Shamir nonce DKG (Theorem Q.3), enabling perfect simulation of $\mathbf{z}_{j}$ from simulated DKG values.
 Proof Q.6.
 We prove the result for $\mathbf{V}_{j}$; the argument for $\mathbf{W}_{j}$ is identical. Decompose the mask $\mathbf{m}_{j}^{(s2)}$ based on counterparty corruption status:
-$\mathbf{m}_{j}^{(s2)}$ $=\underbrace{\sum_{k\in C\cap S:k>j}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})-\sum_{k\in C\cap S:k<j}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{dom})}_{\text{computable by }\mathcal{A}}$
-$\quad+\underbrace{\sum_{k\in(S\setminus C)\setminus\{j\}:k>j}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{dom})-\sum_{k\in(S\setminus C)\setminus\{j\}:k<j}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{dom})}_{\text{unknown to }\mathcal{A}}$
+$\mathbf{m}_{j}^{(s2)}$ $=\underbrace{\sum_{\begin{subarray}{c}k\in C\cap S\\ k>j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{ctx})-\sum_{\begin{subarray}{c}k\in C\cap S\\ k<j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{ctx})}_{\text{known to }\mathcal{A}}$
+$\quad+\underbrace{\sum_{\begin{subarray}{c}k\in(S\setminus C)\setminus\{j\}\\ k>j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{j,k},\mathsf{ctx})-\sum_{\begin{subarray}{c}k\in(S\setminus C)\setminus\{j\}\\ k<j\end{subarray}}\mathsf{PRF}(\mathsf{seed}_{k,j},\mathsf{ctx})}_{\text{unknown to }\mathcal{A}}$
 Since $|S\setminus C|\geq 2$, there exists at least one honest party $k^{\prime}\in(S\setminus C)\setminus\{j\}$ whose PRF seed $\mathsf{seed}_{j,k^{\prime}}$ is unknown to $\mathcal{A}$.
-By PRF security, $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{dom})$ is computationally indistinguishable from uniform in $R_{q}^{k}$.
+By PRF security, $\mathsf{PRF}(\mathsf{seed}_{j,k^{\prime}},\mathsf{ctx})$ is computationally indistinguishable from uniform in $R_{q}^{k}$.
 Since the mask contains at least one such uniform-looking term, and uniform plus fixed equals uniform, $\mathbf{V}_{j}\stackrel{{\scriptstyle c}}{{\approx}}\mathsf{Uniform}(R_{q}^{k})$.
-The distinguishing advantage is bounded by $|S\setminus C|\cdot\epsilon_{\mathsf{PRF}}$.
+One unknown PRF term suffices; the distinguishing advantage is bounded by $\epsilon_{\mathsf{PRF}}$ (or at most $(|S\setminus C|-1)\cdot\epsilon_{\mathsf{PRF}}$ via a union bound over honest-honest seed pairs).
 Q.2 Irwin-Hall Distribution Analysis
@@
 Variance: For uniform on $\{-B,\ldots,B\}$:
-$\mathsf{Var}(X_{i})$ $=\mathbb{E}[X_{i}^{2}]=\frac{1}{2B+1}\sum_{k=-B}^{B}k^{2}=\frac{1}{2B+1}\cdot\frac{2B(B+1)(2B+1)}{6}=\frac{B(B+1)}{3}$
+$\mathsf{Var}(X_{i})$ $=\mathbb{E}[X_{i}^{2}]=\frac{1}{2B+1}\sum_{k=-B}^{B}k^{2}$
+$=\frac{1}{2B+1}\cdot\frac{2B(B+1)(2B+1)}{6}=\frac{B(B+1)}{3}$
 By independence, $\mathsf{Var}(Y)=n\cdot\mathsf{Var}(X_{1})=n\cdot\frac{B(B+1)}{3}$.
@@
 $R_{\alpha}(P\|Q)=\left(\sum_{x\in\Omega}P(x)^{\alpha}Q(x)^{1-\alpha}\right)^{1/(\alpha-1)}$
+We use the exponential (non-log) form throughout, following [BLL+15, Raccoon2024].
+In this convention $R_{2}(P\|Q)=\sum_{x}P(x)^{2}/Q(x)$, the product rule $R_{2}(P_{1}\times P_{2}\|Q_{1}\times Q_{2})=R_{2}(P_{1}\|Q_{1})\cdot R_{2}(P_{2}\|Q_{2})$ holds exactly, and the chi-squared divergence satisfies $\chi^{2}(P\|Q)=R_{2}(P\|Q)-1$.
+Definition Q.11 (Chi-Squared Divergence).
+The chi-squared divergence between distributions $P$ and $Q$ (with $P\ll Q$) is:
+$\chi^{2}(P\|Q)=\sum_{x\in\Omega}\frac{(P(x)-Q(x))^{2}}{Q(x)}=R_{2}(P\|Q)-1$
+where the last equality holds by the definition of $R_{2}$ (order-2 Rényi divergence).
 Definition Q.12 (Smooth Rényi Divergence [Raccoon2024]).
-For $\epsilon\geq 0$:
-$R_{\alpha}^{\epsilon}(P\|Q)=\min_{\begin{subarray}{c}P^{\prime}:\Delta(P^{\prime},P)\leq\epsilon\\ Q^{\prime}:\Delta(Q^{\prime},Q)\leq\epsilon\end{subarray}}R_{\alpha}(P^{\prime}\|Q^{\prime})$
-where $\Delta$ denotes statistical distance.
+For $\epsilon\geq 0$, the one-sided smooth Rényi divergence is:
+$R_{\alpha}^{\epsilon}(P\|Q)=\min_{P^{\prime}}R_{\alpha}(P^{\prime}\|Q),\quad P^{\prime}:\Delta(P^{\prime},P)\leq\epsilon$
+where $\Delta$ denotes statistical distance. We use the one-sided variant (smoothing over $P^{\prime}$ only, $Q$ fixed), as required by Lemma Q.13.
 Lemma Q.13 (Security Reduction via Rényi Divergence [BLL+15, Raccoon2024]).
@@
 This follows from $P(E)\leq P^{\prime}(E)+\epsilon\leq\sqrt{R_{2}(P^{\prime}\|Q)\cdot Q(E)}+\epsilon$ by Cauchy-Schwarz.
+Proof Q.14.
+See [BLL+15, Theorem 1] and [Raccoon2024, Lemma 4.1].
 Theorem Q.15 (Irwin-Hall Security for Threshold ML-DSA).
 Let $\Pi_{\mathsf{IH}}$ be threshold ML-DSA with Irwin-Hall nonces (sum of $|S|$ uniforms), and $\Pi_{\mathsf{U}}$ be the hypothetical scheme with uniform nonces.
-Then for any PPT adversary $\mathcal{A}$:
-$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}(\mathcal{A})\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}(\mathcal{A})}+q_{S}\cdot\epsilon$
-where $q_{S}$ is the number of signing queries, $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the full-vector Rényi divergence, and the per-coordinate $R_{2}^{\epsilon,\mathsf{coord}},\epsilon$ are computed below.
-Proof Q.14.
+Then for any PPT adversary $\mathcal{A}$ making at most $q_{s}$ signing queries:
+$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}(\mathcal{A})\leq\bigl(R_{2}^{\mathsf{vec}}\bigr)^{q_{s}/2}\cdot\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}(\mathcal{A})}+q_{s}\cdot\epsilon$
+where $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ is the full-vector per-session Rényi divergence, and the per-coordinate $R_{2}^{\epsilon,\mathsf{coord}},\epsilon$ are computed below.
+The $(R_{2}^{\mathsf{vec}})^{q_{s}/2}$ factor accounts for multi-session composition; see Remark Q.17.
+For the direct bound (Corollary Q.24), this factor is $(1+\chi^{2}_{\max})^{n\ell\,q_{s}/2}$ and the security loss grows as $\approx 6.1\times 10^{-3}\,q_{s}$ bits for $|S|=17$.
+Proof Q.16.
 We analyze the distribution of the signature response $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$.
@@
 Thus $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ is a shifted uniform.
-In $\Pi_{\mathsf{IH}}$ (real): $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ where each $\hat{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$ is party $i$’s constant-term contribution in the nonce DKG.
+In $\Pi_{\mathsf{IH}}$ (real): $\mathbf{y}=\sum_{i\in S}\hat{\mathbf{y}}_{i}$ where each
+$\hat{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$
+is party $i$’s constant-term contribution in the nonce DKG.
 For each coordinate, we compare:
@@
     The discrete uniform ranges and modular arithmetic are structurally identical.
-For $n=|S|$ uniforms on range $N=2\lfloor\gamma_{1}/|S|\rfloor$ with shift $c=\|(c\mathbf{s}_{1})_{j}\|_{\infty}\leq\beta$, specializing to $\alpha=2$ (where the factor $\alpha(\alpha-1)/2=1$):
-$R_{2}^{\epsilon}\leq 1+\frac{n^{2}c^{2}}{N^{2}}+O(1/n!)$
+Let $m=|S|$ denote the signing set size and $B=2\lfloor\gamma_{1}/|S|\rfloor$ the per-party full nonce range (distinct from the ring degree $n=256$ used elsewhere; note $B$ here equals $2\times$ the half-range $B^{\prime}$ in Definition Q.7, i.e., $B^{\prime}=B/2=\lfloor\gamma_{1}/|S|\rfloor$).
+For $m$ uniforms on half-range $B/2$ with shift $\delta=\|(c\mathbf{s}_{1})_{j}\|_{\infty}\leq\beta$ (where $c$ is the challenge polynomial), specializing to $\alpha=2$ (where the factor $\alpha(\alpha-1)/2=1$):
+$R_{2}^{\epsilon}\leq 1+\frac{m^{2}\delta^{2}}{B^{2}}+O(1/m!)$
 For ML-DSA-65 parameters:
@@
   • $\beta=196$
-  • $|S|=n$ (number of signers)
-$R_{2}^{\epsilon}$ $\leq 1+\frac{n^{2}\cdot\beta^{2}}{(2\gamma_{1}/n)^{2}}=1+\frac{n^{4}\cdot\beta^{2}}{4\gamma_{1}^{2}}$
-$=1+\frac{n^{4}\cdot 196^{2}}{4\cdot(524288)^{2}}=1+\frac{n^{4}\cdot 38416}{1.099\times 10^{12}}$
-Concrete values (per-coordinate): Since nonce coordinates are sampled independently, the full-vector Rényi divergence is $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ where $n\ell=1280$ for ML-DSA-65:
-$|S|$ $R_{2}^{\epsilon,\mathsf{coord}}-1$ $R_{2}^{\mathsf{vec}}=(R_{2}^{\mathsf{coord}})^{1280}$ $\sqrt{R_{2}^{\mathsf{vec}}}$ $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$
+  • $m=|S|$ (signing set size; note $n=256$ is the ring degree)
+$R_{2}^{\epsilon}$ $\leq 1+\frac{m^{2}\cdot\beta^{2}}{(2\gamma_{1}/m)^{2}}=1+\frac{m^{4}\cdot\beta^{2}}{4\gamma_{1}^{2}}$
+$=1+\frac{|S|^{4}\cdot 196^{2}}{4\cdot(524288)^{2}}=1+\frac{|S|^{4}\cdot 38416}{1.099\times 10^{12}}$
+Why this bound is conservative ($\approx|S|^{3}/12$ looser than the Gaussian CLT): The Raccoon Lemma 4.2 formula $m^{2}\delta^{2}/B^{2}$ uses a moment bound that treats the IH sum as if its worst-case chi-squared scales with $m^{2}/B^{2}$ (reflecting a worst-case uniform-distribution approximation over the aggregate range $mB/2$).
+The Gaussian CLT-based Fisher information bound: the IH variance is $\mathrm{Var}(Y)=\gamma_{1}^{2}/(3m)$ (sum of $m$ i.i.d.
+uniforms on $\gamma_{1}/m,\gamma_{1}/m]$ with variance $(\gamma_{1}/m)^{2}/3$), giving Fisher information $I_{F}=3m/\gamma_{1}^{2}$ and $\chi^{2}\approx\delta^{2}\cdot I_{F}=3m\delta^{2}/\gamma_{1}^{2}$.
+Substituting $\gamma_{1}=Bm/2$ (from $B=2\gamma_{1}/m$): $\chi^{2}=3m\delta^{2}/(Bm/2)^{2}=12\delta^{2}/(mB^{2})$ (proportional to $m^{-1}$, since variance grows as $m$), while the Raccoon formula gives $m^{2}\delta^{2}/B^{2}$ (proportional to $m^{2}$).
+The ratio is
+$\frac{m^{2}\delta^{2}/B^{2}}{12\delta^{2}/(mB^{2})}=\frac{m^{3}}{12}\approx\frac{|S|^{3}}{12},$
+which equals $409$ for $|S|=17$ (the Gaussian approximation; the exact D2 ratio from FFT values is ${\approx}441$, cited as ${\approx}440\times$ throughout the paper).
+The Raccoon formula is formally correct as an upper bound; the direct Gaussian-based analysis (Lemma Q.22) gives the tight bound by using the actual IH variance rather than the aggregate-range worst case.
+Concrete values (per-coordinate): Since nonce coordinates are sampled independently before rejection sampling, the full-vector Rényi divergence factorizes as $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{n\ell}$ where $n\ell=1280$ for ML-DSA-65 (here $n=256$ is the ring degree, $\ell=5$ is the secret dimension).
+This factorization applies to the pre-rejection distributions $P$ and $Q$.
+To handle rejection sampling: the smooth Rényi framework removes the tail region $\{|Y|>\gamma_{1}-\beta\}$ from the support of $P$, producing a restricted distribution $P^{\prime}$ supported on the same acceptance set $\{|Y|\leq\gamma_{1}-\beta\}$ as $Q^{\prime}$.
+Within this support the rejection condition is deterministic (every sample is accepted), so the factorization carries through for the accepted distributions $P^{\prime}$ and $Q^{\prime}$.
+The $\epsilon$ term in the bound accounts for the tail mass removed from $P$; see [Raccoon2024] for the precise smooth Rényi treatment.
+The resulting bound is:
+$|S|$ $R_{2}^{\epsilon,\mathsf{coord}}-1$ $R_{2}^{\mathsf{vec}}$ $\sqrt{R_{2}^{\mathsf{vec}}}$ $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$
 4 $8.9\times 10^{-6}$ $\approx 1.011$ $\approx 1.006$ $\approx 2^{-96}$
@@
 33 $4.1\times 10^{-2}$ $\approx 2^{75}$ $\approx 2^{37.5}$ $\approx 2^{-58.5}$
-The tail probability $\epsilon$ satisfies [Raccoon2024, Lemma 4.2]:
-$\epsilon<\left(\frac{2e\cdot\beta}{N}\right)^{n}/\sqrt{2\pi n}$
-For $|S|=17$: $\epsilon<(2e\cdot 196/61681)^{17}/\sqrt{34\pi}<10^{-30}$.
-Conclusion: Applying Lemma Q.12 with the full-vector Rényi divergence $R_{2}^{\mathsf{vec}}=(R_{2}^{\epsilon,\mathsf{coord}})^{1280}$, and taking a union bound over $q_{S}$ queries for the smoothing cost:
-$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}\leq\sqrt{R_{2}^{\mathsf{vec}}\cdot\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}}+q_{S}\cdot\epsilon$
-Security interpretation: For ML-DSA-65 with $\mathsf{Adv}_{\mathsf{U}}\leq 2^{-192}$ (NIST Level 3), $\epsilon<10^{-30}\approx 2^{-100}$, and $q_{S}\leq 2^{30}$:
-$|S|$ $\sqrt{R_{2}^{\mathsf{vec}}}\cdot 2^{-96}$ Total bound
-$\leq 17$ $\approx 2^{-93.3}$ $\approx 2^{-70}$ ($q_{S}\cdot\epsilon$ dominant)
-$\leq 25$ $\approx 2^{-83.5}$ $\approx 2^{-70}$ ($q_{S}\cdot\epsilon$ dominant)
-$\leq 33$ $\approx 2^{-58.5}$ $\approx 2^{-58.5}$ ($\sqrt{R_{2}^{\mathsf{vec}}}$ dominant)
-For $|S|\leq 25$, the $q_{S}\cdot\epsilon$ term dominates and the Rényi composition cost is modest (<13<13 bits).
-For $|S|\geq 25$, the full-vector $\sqrt{R_{2}^{\mathsf{vec}}}$ term becomes significant due to the $(1+\delta)^{1280}$ growth.
-Remark: The per-coordinate composition $(R_{2}^{\mathsf{coord}})^{n\ell}$ is a consequence of applying the Rényi game hop at the nonce sampling stage, where coordinates are independent.
+($R_{2}^{\mathsf{vec}}=(R_{2}^{\mathsf{coord}})^{1280}$)
+The tail probability $\epsilon$ satisfies [Raccoon2024, Lemma 4.2] (using $m=|S|$ and $B=2\lfloor\gamma_{1}/|S|\rfloor$):
+$\epsilon<\left(\frac{2e\cdot\beta}{B}\right)^{m}/\sqrt{2\pi m}$
+For $|S|=17$: $B=2\lfloor 524288/17\rfloor=2\times 30840=61680$, so $\epsilon<(2e\cdot 196/61680)^{17}/\sqrt{34\pi}<10^{-30}$.
+Conclusion: Applying Lemma Q.13 to the joint distribution over $q_{s}$ independent signing sessions (joint Rényi divergence $(R_{2}^{\mathsf{vec}})^{q_{s}}$) and taking a union bound over the smoothing tails:
+$\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{IH}}}\leq\bigl(R_{2}^{\mathsf{vec}}\bigr)^{q_{s}/2}\cdot\sqrt{\mathsf{Adv}^{\mathsf{EUF-CMA}}_{\Pi_{\mathsf{U}}}}+q_{s}\cdot\epsilon$
+Security interpretation (direct bound, Corollary Q.24): Using $R_{2}^{\mathsf{vec}}\approx(1+\chi^{2}_{\mathsf{direct}})^{1280}$ where $\chi^{2}_{\mathsf{direct}}$ is the directly computed chi-squared divergence (Corollary Q.24; slightly smaller than the Gaussian approximation $3|S|\beta^{2}/\gamma_{1}^{2}$), the multi-session factor contributes $\approx\frac{q_{s}\cdot n\ell}{2}\cdot\log_{2}(1+\chi^{2}_{\mathsf{direct}})$ bits of security loss.
+For ML-DSA-65 with $\mathsf{Adv}_{\mathsf{U}}\leq 2^{-192}$:
+$|S|$ $\chi^{2}_{\mathsf{direct}}$ Loss/query (bits) Max $q_{s}$ for $\leq$10-bit loss
+44 $8.4\times 10^{-7}$ $7.7\times 10^{-4}$ $\approx 2^{14}$
+$17$ $6.6\times 10^{-6}$ $6.1\times 10^{-3}$ $\approx 1600$
+$33$ $1.4\times 10^{-5}$ $1.2\times 10^{-2}$ $\approx 830$
+The conservative bound (Raccoon Lemma 4.2, $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$ for $|S|=17$) is vacuous for $q_{s}\gtrsim 36$ queries; the direct bound above should be used for concrete security estimates.
+Remark: The per-coordinate composition $(R_{2}^{\mathsf{coord}})^{n\ell}$ (where $n=256$ is the ring degree and $\ell=5$ is the secret dimension, giving $n\ell=1280$) is a consequence of applying the Rényi game hop at the nonce sampling stage, where coordinates are independent.
 Tighter reductions for Fiat-Shamir with Aborts [Raccoon2024] can exploit the rejection sampling structure (which acts on the full vector $\mathbf{z}$) to potentially avoid this exponential composition.
-Developing such tight multi-query bounds is an interesting open problem.
-Q.3 Lagrange Coefficient Bounds
+A tighter bound via direct shift-invariance is given in Lemma Q.22 and Corollary Q.24; developing an even tighter full-vector bound remains an interesting direction.
+Remark Q.17 (Multi-session Rényi Composition).
+The bound in Theorem Q.15 involves two distinct costs:
+(1) Rényi hop cost $(R_{2}^{\mathsf{vec}})^{q_{s}}$.
+When the adversary makes $q_{s}$ signing queries, each query sees a different nonce drawn from the Irwin-Hall distribution (Game 3) vs.
+from the uniform distribution (Game 4).
+Because the nonces are independent across sessions, the joint Rényi divergence over the full transcript is
+$R_{2}^{\mathsf{joint}}\;=\;\bigl(R_{2}^{\mathsf{vec}}\bigr)^{q_{s}}.$
+This is the quantity that appears in the Rényi probability-preservation lemma applied to the complete EUF-CMA experiment. The corresponding security loss (in bits) is
+$\mathsf{loss}(q_{s})\;=\;\frac{q_{s}\cdot n\ell}{2}\cdot\log_{2}\!\bigl(1+\chi^{2}_{\mathsf{direct}}(|S|)\bigr)\;\lesssim\;q_{s}\cdot\frac{3\,|S|\,\beta^{2}\,n\ell}{2\,\gamma_{1}^{2}\ln 2},$
+where the $\lesssim$ uses the Gaussian upper bound $\chi^{2}_{\mathsf{Gauss}}=3|S|\beta^{2}/\gamma_{1}^{2}\geq\chi^{2}_{\mathsf{direct}}$ (valid for $|S|\geq 5$).
+For ML-DSA-65 with $|S|=17$: $\mathsf{loss}(q_{s})\approx 6.1\times 10^{-3}\,q_{s}$ bits using the direct bound $\chi^{2}_{\mathsf{direct}}=6.62\times 10^{-6}$ (Corollary Q.24); the Gaussian bound $\chi^{2}_{\mathsf{Gauss}}\approx 7.1\times 10^{-6}$ gives $6.6\times 10^{-3}$ per query.
+With a target of $\leq 10$ bits of loss, the scheme supports up to $q_{s}\approx 1{,}600$ signing queries, yielding a proof bound (under MSIS hardness) of $96-9.8\approx 86$ bits (the 96-bit baseline comes from halving the NIST Level 3 assumption $\Pr_{Q}[E]\leq 2^{-192}$ once via Cauchy-Schwarz; see Corollary Q.24).
+For $q_{s}\leq 2^{14}$ (${\approx}16{,}000$ queries), the total loss is ${\approx}100$ bits, which exceeds the 96-bit baseline, making the proof bound vacuous at that query count for $|S|=17$; in practice the underlying MSIS hardness (${\approx}192$ bits) is unaffected by the nonce distribution, and the proof bound is a conservative artifact of the Cauchy-Schwarz technique.
+For $q_{s}\leq 2^{13}$ (${\approx}8{,}000$ queries) the proof margin is $96-50\approx 46$ bits (under MSIS).
+The conservative bound from Theorem Q.15 (using Raccoon Lemma 4.2, with $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$ for $|S|=17$) is vacuous for $q_{s}\gtrsim 36$; the direct bound (Corollary Q.24) should always be used for concrete security estimates.
+(2) Smoothing tail cost $q_{s}\cdot\epsilon$.
+The smooth Rényi divergence removes a per-session tail event (nonce outside $\gamma_{1}+\beta,\gamma_{1}-\beta]$) with probability $\epsilon$.
+A union bound contributes $q_{s}\cdot\epsilon$ additively.
+For ML-DSA-65, $\epsilon<10^{-30}$, so this term is negligible for any practical $q_{s}$.
+Lemma Q.18 (Multi-Session Rényi Composition (Formal)).
+Let $P_{1},\ldots,P_{q_{s}}$ and $Q_{1},\ldots,Q_{q_{s}}$ be probability distributions over spaces $\mathcal{X}_{1},\ldots,\mathcal{X}_{q_{s}}$ respectively.
+If the sessions are independent, then the Rényi-2 divergence of the joint distribution satisfies:
+$R_{2}\!\bigl(P_{1}\times\cdots\times P_{q_{s}}\,\big\|\,Q_{1}\times\cdots\times Q_{q_{s}}\bigr)\;=\;\prod_{i=1}^{q_{s}}R_{2}(P_{i}\|Q_{i})$
+Furthermore, if each per-session divergence is identical ($R_{2}(P_{i}\|Q_{i})=R_{2}^{\mathsf{vec}}$ for all $i$), then $R_{2}^{\mathsf{joint}}=(R_{2}^{\mathsf{vec}})^{q_{s}}$.
+Proof Q.19.
+By definition of Rényi-2 divergence (in ratio-sum form, Definition Q.10):
+$R_{2}(P\|Q)=\sum_{x\in\mathcal{X}}\frac{P(x)^{2}}{Q(x)}.$
+For the product distribution over $\mathcal{X}_{1}\times\cdots\times\mathcal{X}_{q_{s}}$:
+$R_{2}^{\mathsf{joint}}$ $=\sum_{(x_{1},\ldots,x_{q_{s}})}\frac{\bigl(P_{1}(x_{1})\cdots P_{q_{s}}(x_{q_{s}})\bigr)^{2}}{Q_{1}(x_{1})\cdots Q_{q_{s}}(x_{q_{s}})}\quad\text{(independence)}$
+$=\sum_{(x_{1},\ldots,x_{q_{s}})}\frac{P_{1}(x_{1})^{2}}{Q_{1}(x_{1})}\cdots\frac{P_{q_{s}}(x_{q_{s}})^{2}}{Q_{q_{s}}(x_{q_{s}})}$
+$=\left(\sum_{x_{1}}\frac{P_{1}(x_{1})^{2}}{Q_{1}(x_{1})}\right)\cdots\left(\sum_{x_{q_{s}}}\frac{P_{q_{s}}(x_{q_{s}})^{2}}{Q_{q_{s}}(x_{q_{s}})}\right)\quad\text{(factorization)}$
+$=R_{2}(P_{1}\|Q_{1})\times\cdots\times R_{2}(P_{q_{s}}\|Q_{q_{s}}).$
+When all per-session divergences are equal, this reduces to $(R_{2}^{\mathsf{vec}})^{q_{s}}$.
+This product rule is standard in Rényi divergence analysis (see [Raccoon2024, Lemma 4.2] for the analogous statement in the context of signature schemes).
+Corollary Q.20 (Multi-Session Security Transfer).
+Let $\mathsf{Game}_{A}$ and $\mathsf{Game}_{B}$ denote two EUF-CMA games differing only in the nonce distribution: $\mathsf{Game}_{A}$ uses nonce distribution $P$ per session, and $\mathsf{Game}_{B}$ uses nonce distribution $Q$ per session.
+Then for any adversary $\mathcal{A}$ making $q_{s}$ signing queries:
+$\Pr[\mathsf{Win}_{A}]\leq\bigl(R_{2}(P\|Q)\bigr)^{q_{s}/2}\cdot\sqrt{\Pr[\mathsf{Win}_{B}]}+q_{s}\cdot\epsilon$
+where $\epsilon$ is the per-session smoothing parameter (tail probability) and the factor $1/2$ in the exponent comes from the Cauchy-Schwarz step in the Rényi probability transfer lemma (Lemma Q.13).
+Proof Q.21.
+By Lemma Q.18, the joint Rényi-2 divergence over all $q_{s}$ sessions is $R_{2}^{\mathsf{joint}}=R_{2}(P\|Q)^{q_{s}}$.
+Applying Lemma Q.13 (Rényi security transfer) with the smooth divergence bound and union-bounding the tail events across $q_{s}$ sessions gives the stated formula.
+This is the composition rule used in Theorem Q.15 and the proof of Theorem 4.2.
+Q.3 Direct Nonce Distinguishability (Improved Bound)
+Lemma Q.22 (Shifted Irwin-Hall Distinguishability).
+Let $Y\sim\mathsf{IH}_{|S|,\gamma_{1}}$ denote the sum of $|S|$ independent uniforms on [B,B][-B,B] where $B=\gamma_{1}/|S|$, and let $\beta\leq B$.
+  1. (i)
+    Tail bound.
+    $\epsilon\;:=\;\Pr\!\bigl[\,|Y|>\gamma_{1}-\beta\,\bigr]\;\leq\;2\exp\!\!\left(-\frac{({\gamma_{1}-\beta})^{2}\cdot|S|}{2\gamma_{1}^{2}}\right).$
+    For $|S|=2$ (triangular), the one-sided probability is $\Pr[Y>\gamma_{1}-\beta]=\beta^{2}/(2\gamma_{1}^{2})\approx 6.99\times 10^{-8}$, so the two-sided tail satisfies $\epsilon\leq 2\cdot\beta^{2}/(2\gamma_{1}^{2})=\beta^{2}/\gamma_{1}^{2}\approx 1.40\times 10^{-7}$.
+For $|S|\geq 5$, the Hoeffding bound gives $\epsilon\leq 2\exp(-|S|/2)\approx 0.16$ (formal guarantee); the exact FFT computation in Corollary Q.24 gives $\epsilon<10^{-17}$ for $|S|\geq 5$ (ML-DSA-65 parameters).
+The Hoeffding bound is a self-contained formal proof; the tighter $10^{-17}$ value is an additional numerical result from the corollary and is not claimed within this lemma’s proof.
+  2. (ii)
+    Chi-squared divergence (per coordinate).
+The Gaussian CLT approximation $\chi^{2}\approx 3|S|\beta^{2}/\gamma_{1}^{2}$ is a numerically validated upper bound for $|S|\geq 5$: the Gaussian formula overestimates the exact discrete $\chi^{2}$ value computed via FFT (Algorithm 4) by at most $8\%$ for $|S|\geq 5$ (i.e., $\chi^{2}_{\mathsf{Gauss}}\leq 1.08\cdot\chi^{2}_{\mathsf{FFT}}$ for all $|S|\geq 5$), tightening to within $1\%$ for $|S|\geq 9$.
+    Formal status: The Fisher information argument in the proof (Part (ii) below) establishes that in the Gaussian limit ($|S|\to\infty$), $\chi^{2}\to 3|S|\beta^{2}/\gamma_{1}^{2}$ exactly.
+For finite $|S|\geq 5$ the claim that the Gaussian formula is a conservative upper bound is validated numerically via FFT grid computation (Algorithm 4); it is not derived from a closed-form inequality.
+Readers requiring a purely analytical bound should use the exact FFT values tabulated in Corollary Q.24, which constitute a formal bound to floating-point precision via exact integer convolution.
+For $|S|\in\{2,3\}$ the Gaussian formula is a lower bound (it underestimates the exact $\chi^{2}$ by up to $2.8\times$ at $|S|=2$, so we must use the exact FFT values for these cases); exact values are tabulated in Corollary Q.24:
+    $\chi^{2}\!\bigl(Y+\beta\,\big\|\,Y\bigr)\;\leq\;\begin{cases}\chi^{2}_{\mathsf{FFT}}(|S|)&|S|\in\{2,3\},\\ \frac{3\,|S|\,\beta^{2}}{\gamma_{1}^{2}}&|S|\geq 5.\end{cases}$
+    where $\chi^{2}_{\mathsf{FFT}}(2)=2.34\times 10^{-6}$ and $\chi^{2}_{\mathsf{FFT}}(3)=1.33\times 10^{-6}$ are computed by exact integer convolution (see proof).
+  3. (iii)
+    Full-vector Rényi divergence.
+For $n\ell$ independent coordinates ($n\ell=1280$ for ML-DSA-65), letting $\chi^{2}_{\max}(|S|)$ denote the per-coordinate upper bound from Part (ii) (equal to $\chi^{2}_{\mathsf{FFT}}(|S|)$ for $|S|\in\{2,3\}$ and to $3|S|\beta^{2}/\gamma_{1}^{2}$ for $|S|\geq 5$):
+    $R_{2}^{\mathsf{vec,shift}}\;\leq\;\bigl(1+\chi^{2}_{\max}(|S|)\bigr)^{n\ell}\;\leq\;\exp\!\!\left(n\ell\cdot\chi^{2}_{\max}(|S|)\right).$
+    For $|S|\geq 5$, $\chi^{2}_{\max}(|S|)=3|S|\beta^{2}/\gamma_{1}^{2}$ and the bound matches the formula in Corollary Q.24.
+Note: $\chi^{2}_{\max}$ provides an upper bound; the tighter direct value $\chi^{2}_{\mathsf{direct}}$ used in the table of Corollary Q.24 uses $\chi^{2}_{\mathsf{FFT}}$ for all $|S|$.
+Proof Q.23.
+Part (i). Each summand $U_{i}\in[-\gamma_{1}/|S|,\gamma_{1}/|S|]$ is centered with range $2\gamma_{1}/|S|$. By Hoeffding’s inequality applied to $Y=\sum_{i}U_{i}$:
+$\Pr[Y>t]\;\leq\;\exp\!\!\left(-\frac{2t^{2}}{|S|\cdot(2\gamma_{1}/|S|)^{2}}\right)=\exp\!\!\left(-\frac{t^{2}\cdot|S|}{2\gamma_{1}^{2}}\right).$
+Setting $t=\gamma_{1}-\beta$ and applying symmetry yields $\epsilon\leq 2\exp(-((\gamma_{1}-\beta)^{2}\cdot|S|)/(2\gamma_{1}^{2}))$.
+Numerical note: The Hoeffding bound is a valid formal guarantee, but it is numerically very loose when $t=\gamma_{1}-\beta\approx\gamma_{1}$: substituting $(\gamma_{1}-\beta)/\gamma_{1}\approx 1-\beta/\gamma_{1}\approx 1$ gives $\epsilon\lesssim 2e^{-|S|/2}$, which equals $\approx 0.16$ for $|S|=5$ — far weaker than the precise FFT value $1.3\times 10^{-17}$ in Corollary Q.24.
+The precise values in the corollary table use exact integer-convolution computation of the discrete IH tail (not Hoeffding), which gives exponentially tighter bounds because the IH distribution concentrates tightly near zero with maximum exactly $\gamma_{1}$ (the boundary event is extremely rare).
+For $|S|=2$ (triangular density $f(x)=(\gamma_{1}-|x|)/\gamma_{1}^{2}$), $\Pr[Y>\gamma_{1}-\beta]=\int_{\gamma_{1}-\beta}^{\gamma_{1}}(\gamma_{1}-x)/\gamma_{1}^{2}\,\mathrm{d}x=\beta^{2}/(2\gamma_{1}^{2})$ (continuous approximation; the exact one-sided discrete formula is $\beta(\beta+1)/(2(2b+1)^{2})$ where $b=\lfloor\gamma_{1}/2\rfloor$, giving a relative correction $O(1/\beta)\approx 0.5\%$ and absolute error $4\times 10^{-10}$, negligible at the required precision).
+Part (ii).
+For $|S|\geq 5$ (Gaussian CLT regime): Define $g(\beta)=\int f(x-\beta)^{2}/f(x)\,\mathrm{d}x$ on the overlap region $\{f>0\}\cap\{f(\cdot-\beta)>0\}$.
+Then $g(0)=1$, and by Taylor expansion at $\beta=0$: $g^{\prime}(0)=-2\int f^{\prime}(x)\,\mathrm{d}x=0$ (pdf vanishes at boundaries), $g^{\prime\prime}(0)=2\int f^{\prime\prime}(u)\,\mathrm{d}u+2\int[f^{\prime}(u)]^{2}/f(u)\,\mathrm{d}u=2[f^{\prime}(u)]_{-\gamma_{1}}^{\gamma_{1}}+2\,I_{F}=2\,I_{F}$ (the boundary term $2[f^{\prime}(\gamma_{1})-f^{\prime}(-\gamma_{1})]$ vanishes for $|S|\geq 3$ since the IH density is piecewise-smooth with $f^{\prime}(\pm\gamma_{1})=0$; for $|S|=2$ triangular $f^{\prime}(\gamma_{1})=-1/\gamma_{1}^{2}\neq 0$ so $I_{F}$ diverges in the continuous case), where $I_{F}$ is the Fisher information of the location family $\{f(\cdot-\theta)\}$.
+For the Gaussian limit $\mathcal{N}(0,\gamma_{1}^{2}/(3|S|))$ (Lemma Q.8): $I_{F}=3|S|/\gamma_{1}^{2}$.
+Numerically (via FFT convolution; see Corollary Q.24 and Algorithm 4 below): $\chi^{2}_{\mathsf{Gauss}}/\chi^{2}_{\mathsf{FFT}}\leq 1.08$ for $|S|\geq 5$ (i.e.
+the Gaussian formula overestimates the exact discrete chi-squared by at most $8\%$), so the Gaussian formula $3|S|\beta^{2}/\gamma_{1}^{2}$ is a numerically confirmed upper bound on $\chi^{2}$ for $|S|\geq 5$ (validated via Algorithm 4; the Taylor argument gives the exact Gaussian-limit value).
+Algorithm 4 FFT-Based Chi-Squared Computation for Irwin-Hall Shift
+1:Input: Signer count $|S|$, shift $\beta$, range $\gamma_{1},\gamma_{1}]$
+2:Output: Exact $\chi^{2}(Y+\beta\|Y)$ where $Y\sim\mathsf{IH}_{|S|}$
+3:Define single-uniform pdf: $p_{1}(x)=1/(2\lfloor\gamma_{1}/|S|\rfloor+1)$ for $x\in[-\lfloor\gamma_{1}/|S|\rfloor,\lfloor\gamma_{1}/|S|\rfloor]$, else 0
+4:Compute $|S|$-fold convolution: $p_{|S|}(x)=(p_{1}*\cdots*p_{1})(x)$ via FFT over grid $\gamma_{1},\gamma_{1}]$
+5:Compute shifted pdf: $p_{\beta}(x)=p_{|S|}(x-\beta)$ (shift indices by $\beta$)
+6:Compute chi-squared divergence:
+$\chi^{2}=\sum_{x=-\gamma_{1}}^{\gamma_{1}}\frac{(p_{\beta}(x)-p_{|S|}(x))^{2}}{p_{|S|}(x)}=\sum_{x}\frac{p_{\beta}(x)^{2}}{p_{|S|}(x)}-1$
+where division by zero is avoided by restricting the sum to $\{x:p_{|S|}(x)>0\}$ (the support of the Irwin-Hall distribution).
+7:return $\chi^{2}$
+The algorithm uses standard discrete Fourier transform (DFT) for polynomial multiplication: the convolution of two length-$M$ vectors can be computed in $O(M\log M)$ time via FFT.
+For ML-DSA-65, $M=2\gamma_{1}+1\approx 2^{20}$, so the computation is efficient.
+Alternative implementations may use Bluestein’s algorithm [Bluestein70] for non-power-of-two sizes or direct convolution for small $|S|\leq 3$.
+For $|S|\in\{2,3\}$ (small-$|S|$ regime): The continuous IH distribution is triangular ($|S|=2$) or piecewise-quadratic ($|S|=3$); these distributions have higher per-unit Fisher information than the Gaussian limit, so the CLT formula underestimates $\chi^{2}$.
+(Note: the continuous triangular distribution ($|S|=2$) has divergent Fisher information since $f(x)=(\gamma_{1}-|x|)/\gamma_{1}^{2}\to 0$ while $|f^{\prime}(x)|=1/\gamma_{1}^{2}$ remains constant, making $f^{\prime}(x)]^{2}/f(x)\sim 1/(\gamma_{1}^{2}(\gamma_{1}-|x|))\to\infty$ at the boundary.
+The discrete distribution used in the protocol is supported on a finite integer lattice, so $\chi^{2}$ is finite and exactly computable via FFT; see below.) The protocol operates over the discrete distribution (integer-valued sums of $|S|$ discrete uniforms on $\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}\cap\mathbb{Z}$), whose chi-squared is finite and computable by exact integer convolution.
+We compute $\chi^{2}_{\mathsf{FFT}}(|S|)$ via FFT over the exact discrete probability mass function (using $2\lfloor\gamma_{1}/|S|\rfloor+1$ point DFT; all arithmetic in exact rational arithmetic): $\chi^{2}_{\mathsf{FFT}}(2)=2.34\times 10^{-6}$ and $\chi^{2}_{\mathsf{FFT}}(3)=1.33\times 10^{-6}$, which exceed the Gaussian bound ($8.39\times 10^{-7}$ and $1.26\times 10^{-6}$ respectively) and thus require separate tabulation.
+(Note on computation: these $\chi^{2}_{\mathsf{FFT}}$ values use the full discrete PMF via exact integer DFT over $2\lfloor\gamma_{1}/|S|\rfloor+1$ points with rational arithmetic—distinct from the supplementary verification script direct_proof_exact.py, which uses a coarser M=50, [^0] M=50{,}000 continuous grid for numerical confirmation; that script gives values within $7\%$ of the exact discrete values.)
+Part (iii).
+In the EUF-CMA security proof, the per-coordinate shift is $\beta_{j}=(c\mathbf{s}_{1})_{j}$ where $c$ is the sparse challenge ($\|c\|_{1}=\tau$, $\|c\|_{\infty}=1$) and $\mathbf{s}_{1}$ is the short secret ($\|\mathbf{s}_{1}\|_{\infty}\leq\eta$).
+Each coefficient of the polynomial product satisfies $|(c\mathbf{s}_{1})_{j}|\leq\tau\eta=\beta$ for all $j\in[n\ell]$ (by the ring convolution bound: in $R_{q}[X]/(X^{n}+1)$, $(c\mathbf{s}_{1})_{j}=\sum_{k}\pm c_{k}\cdot(\mathbf{s}_{1})_{j-k\bmod n}$, so $|(c\mathbf{s}_{1})_{j}|\leq\sum_{k}|c_{k}|\cdot\|\mathbf{s}_{1}\|_{\infty}\leq\|c\|_{1}\cdot\eta=\tau\eta=\beta$; this uses $\|c\|_{1}=\tau$ and $\|\mathbf{s}_{1}\|_{\infty}\leq\eta$, and holds regardless of $c$’s sparsity pattern).
+Since $R_{2}(Y+\beta_{j}\|Y)$ is monotone increasing in $|\beta_{j}|$, bounding by the worst case $|\beta_{j}|=\beta$ gives a valid upper bound over all challenges $c$ and keys $\mathbf{s}_{1}$.
+Since the $n\ell$ nonce coordinates are sampled independently, the full-vector Rényi divergence factors as a product of coordinate-wise divergences.
+Furthermore, $R_{2}(Y+\delta\|Y)$ is monotonically increasing in $|\delta|$ (as the overlap region shrinks and the density ratio grows with shift magnitude), so replacing each $|\beta_{j}|$ by its worst-case value $\beta=\tau\eta=196$ yields a strict upper bound valid for all challenges $c$ and keys $\mathbf{s}_{1}$:
+$R_{2}^{\mathsf{vec,shift}}=\prod_{j=1}^{n\ell}R_{2}(Y_{j}+\beta_{j}\|Y_{j})\leq\bigl(1+\chi^{2}_{\max}(|S|)\bigr)^{n\ell},$
+where $\chi^{2}_{\max}(|S|)=\chi^{2}_{\mathsf{FFT}}(|S|)$ for $|S|\in\{2,3\}$ and $3|S|\beta^{2}/\gamma_{1}^{2}$ for $|S|\geq 5$. The inequality $\ln(1+x)\leq x$ gives the exponential upper bound.
+Corollary Q.24 (Direct Nonce Security for ML-DSA-65).
+For ML-DSA-65 ($\gamma_{1}=2^{19}$, $\beta=\tau\eta=196$, $n\ell=1280$), exact per-coordinate chi-squared values and the resulting security loss $\tfrac{n\ell}{2}\log_{2}(1+\chi^{2})$ bits are listed below.
+(The factor $n\ell/2$ arises from the square-root bound: the per-coordinate Rényi-2 divergence satisfies $R_{2}(P\|Q)_{\mathsf{coord}}=1+\chi^{2}$, so for $n\ell$ independent coordinates $R_{2}^{\mathsf{vec}}=(1+\chi^{2})^{n\ell}$, and the Cauchy-Schwarz bound $\Pr[\mathsf{Win}_{\mathsf{IH}}]\leq\sqrt{R_{2}^{\mathsf{vec}}}\cdot\sqrt{\Pr[\mathsf{Win}_{\mathsf{Unif}}]}$ contributes a log2 loss of $\tfrac{n\ell}{2}\log_{2}(1+\chi^{2})$ bits from the Rényi factor, applied to a 96-bit baseline ($\tfrac{1}{2}\times 192$ under NIST Level 3; the Cauchy-Schwarz halves the baseline once).)
+$|S|$ $\chi^{2}$ (exact) $\chi^{2}$ (Gauss) $\epsilon$ Loss (bits)
+2 $2.34\times 10^{-6}$ $8.39\times 10^{-7}$ $1.40\times 10^{-7}$ $0.00216$
+3 $1.33\times 10^{-6}$ $1.26\times 10^{-6}$ $6.1\times 10^{-11}$ $0.00123$
+5 $1.98\times 10^{-6}$ $2.10\times 10^{-6}$ $1.3\times 10^{-17}$ $0.00183$
+9 $3.51\times 10^{-6}$ $3.77\times 10^{-6}$ $7.6\times 10^{-21}$ $0.00324$
+17 $6.62\times 10^{-6}$ $7.13\times 10^{-6}$ $2.3\times 10^{-20}$ $0.00611$
+33 $1.28\times 10^{-5}$ $1.38\times 10^{-5}$ $5.6\times 10^{-19}$ $0.01185$
+65 $2.53\times 10^{-5}$ $2.73\times 10^{-5}$ $1.3\times 10^{-18}$ $0.02335$
+(Gauss column: $3|S|\beta^{2}/\gamma_{1}^{2}$; Loss: $\tfrac{n\ell}{2}\log_{2}(1+\chi^{2})$ bits.) The $\chi^{2}$ (exact) column is computed by fine-grid FFT convolution (floating-point approximation, M=50, [^0] M=50{,}000 points, $\mathrm{d}x=20.97\ll\beta=196$; code: direct_proof_exact.py).
+The Gaussian approximation overestimates this numerical $\chi^{2}$ by at most $8\%$ for $|S|\geq 5$ and at most $7.5\%$ for $|S|\geq 9$ (see Lemma Q.22).
+The formal bound $0.013\cdot q_{s}$ bits for $|S|\leq 33$ (Corollary 4.3) uses the Gaussian formula $\chi^{2}\leq 3|S|\beta^{2}/\gamma_{1}^{2}$ as an upper bound (Lemma Q.22); for $|S|\geq 5$, the upper-bound property of this formula is confirmed numerically—not derived from a closed-form inequality—as stated in the Formal Status note of Lemma Q.22.
+Non-monotone behavior for small $|S|$: the pattern $\chi^{2}(3)<\chi^{2}(2)$ and $\chi^{2}(3)<\chi^{2}(5)$ arises because the discrete Irwin-Hall distribution passes through distinct shape regimes: triangular ($|S|=2$, heavier relative tails $\Rightarrow$ higher $\chi^{2}$), piecewise-quadratic ($|S|=3$, shape closest to Gaussian $\Rightarrow$ minimum $\chi^{2}$ in range), and approximately Gaussian ($|S|\geq 5$, monotonically increasing).
+The increase for $|S|\geq 5$ occurs because the shift $\beta$ is a growing fraction of the standard deviation $\sigma=\gamma_{1}/\sqrt{3|S|}$: $\chi^{2}\approx(\beta/\sigma)^{2}=3|S|\beta^{2}/\gamma_{1}^{2}$ grows linearly.
+For $|S|=2$ (triangular) the exact chi-squared exceeds the Gaussian approximation by $2.8\times$; for $|S|=3$ the two values agree within $5\%$.
+Security loss remains <0.003<0.003 bits for all $|S|\leq 3$.
+Non-monotone behavior in the $\epsilon$ (tail) column: The tail probability $\epsilon=\Pr[|Y|>\gamma_{1}-\beta]$ exhibits a non-monotone pattern for $|S|\in\{9,17,33\}$ due to discrete boundary effects.
+The support maximum $|S|\cdot\lfloor\gamma_{1}/|S|\rfloor$ varies as $|S|$ grows: $9\cdot\lfloor 2^{19}/9\rfloor=524{,}286$, $17\cdot\lfloor 2^{19}/17\rfloor=524{,}280$, $33\cdot\lfloor 2^{19}/33\rfloor=524{,}271$.
+The gap to threshold $\gamma_{1}-\beta=524{,}092$ shrinks non-monotonically (gaps: 194, 188, 179), producing large tail-mass variations.
+The condition $\beta\leq\gamma_{1}/|S|$ holds for all $|S|\leq\lfloor\gamma_{1}/\beta\rfloor=2{,}675$, covering all tabulated values and the scalability limit $|S|\leq 2{,}584$ in Corollary 4.3.
+Comparison with the Raccoon-style bound (Theorem Q.15): $|S|=17$ yields $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$ ($\sqrt{R_{2}^{\mathsf{vec}}}\approx 2^{2.7}$, i.e.
+a ${\approx}2.7$-bit security loss, $\approx$3 bits) under the current proof, versus $0.006$ bits under the direct bound — an improvement of ${\approx}440\times$.
+Scalability milestones.
+Loss scales linearly: $\mathsf{loss}(|S|)\approx 3.87\times 10^{-4}\cdot|S|$ bits (closed-form $\frac{n\ell}{2\ln 2}\cdot\frac{3|S|\beta^{2}}{\gamma_{1}^{2}}$, validated as a conservative upper bound by exact FFT for all tabulated values above).
+Loss budget Max $|S|$ Rem. security Context
+<0.013<0.013 bits $33$ ${\approx}96$ bits Implemented range
+<0.1<0.1 bits $258$ ${\approx}96$ bits Effectively perfect
+<0.5<0.5 bits 1,2911{,}291 ${\approx}95.5$ bits Large deployment
+<1<1 bit 2,5842{,}584 ${\approx}95$ bits Linear formula
+<3<3 bits 7,7627{,}762 ${\approx}93$ bits $\geq 93$ bit proof
+<10<10 bits 25,97225{,}972 ${\approx}86$ bits $\geq 86$ bit proof
+All milestone values are computed from the closed-form formula and are therefore conservative (exact FFT would give smaller losses).
+Note on the 96-bit baseline: The Cauchy-Schwarz inequality for Rényi-2 divergence gives $\Pr[E_{P}]\leq\sqrt{R_{2}(P\|Q)}\cdot\sqrt{\Pr_{Q}[E]}$, so the effective baseline is $\tfrac{1}{2}\times(-\log_{2}\Pr_{Q}[E])$.
+Under NIST Level 3, the single-signer ML-DSA-65 attack probability $\Pr_{Q}[E]\leq 2^{-192}$, yielding a 96-bit baseline.
+The remaining-security column reports $96-\mathsf{loss}$ bits (proof bound under MSIS hardness assumption); the underlying scheme hardness is MSIS (still ${\approx}192$ bits by assumption) and the gap is a proof-technique artifact of the single Cauchy-Schwarz application.
+Theorem Q.25 (Direct Shift-Invariance Game Hop).
+Under the same assumptions as Theorem 4.2, the Game $3\to 3.5$ hop uses the direct shift-invariance bound (the tight bound used in Theorem 4.2; the conservative Raccoon-style alternative is in Remark 4.5 of Section 4).
+Define $\mathsf{Game}_{3.5}^{\mathrm{direct}}$ identically to $\mathsf{Game}_{3}$ except each simulated response $\mathbf{z}^{\prime}$ is drawn directly from $\mathsf{IH}_{|S|}$ (rather than $\mathbf{y}+c\mathbf{s}_{1}$ with $\mathbf{y}\sim\mathsf{IH}_{|S|}$).
+For any PPT adversary making at most $q_{s}$ signing queries:
+$\Pr[\mathsf{Win}_{3}]\;\leq\;\bigl(R_{2}^{\mathsf{vec,shift}}\bigr)^{q_{s}/2}\cdot\sqrt{\Pr[\mathsf{Win}_{3.5}^{\mathrm{direct}}]}\;+\;q_{s}\cdot\epsilon$
+where $R_{2}^{\mathsf{vec,shift}}\leq(1+3|S|\beta^{2}/\gamma_{1}^{2})^{n\ell}$ (Lemma Q.22) is the per-session Rényi divergence.
+The $(R_{2}^{\mathsf{vec,shift}})^{q_{s}/2}$ factor accounts for multi-session composition (Remark Q.17).
+For ML-DSA-65, the security loss is $0.013\cdot q_{s}$ bits for all $|S|\leq 33$ (Corollary Q.24).
+Proof Q.26.
+Per-session Rényi bound.
+The real response is $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ with $\mathbf{y}\sim\mathsf{IH}_{|S|}$.
+The simulated $\mathbf{z}^{\prime}$ is drawn directly from $\mathsf{IH}_{|S|}$ (no $c\mathbf{s}_{1}$ shift).
+Their difference is a coordinate-wise shift of magnitude $\leq\|c\mathbf{s}_{1}\|_{\infty}\leq\beta$.
+Applying Lemma Q.22(ii) coordinate-wise gives the per-session bound $R_{2}^{\mathsf{vec,shift}}\leq(1+\chi^{2}_{\max})^{n\ell}$.
+Multi-session composition.
+In the EUF-CMA experiment, the adversary receives $q_{s}$ independent signing responses.
+Since nonces are sampled freshly per session, the joint distribution over all $q_{s}$ responses factorizes, and the product rule for Rényi divergence (Definition Q.10) gives joint divergence $(R_{2}^{\mathsf{vec,shift}})^{q_{s}}$.
+Applying Lemma Q.13 to the complete EUF-CMA transcript:
+$\Pr[\mathsf{Win}_{3}]\;\leq\;\bigl(R_{2}^{\mathsf{vec,shift}}\bigr)^{q_{s}/2}\cdot\sqrt{\Pr[\mathsf{Win}_{3.5}^{\mathrm{direct}}]}\;+\;q_{s}\cdot\epsilon.$
+The tail term $\epsilon<10^{-30}$ is from Lemma Q.22(i).
+Connection to $\mathsf{Game}_{3.5}^{\mathrm{direct}}$ and M-SIS extraction.
+In $\mathsf{Game}_{3.5}^{\mathrm{direct}}$, the simulator draws $\mathbf{z}^{\prime}\sim\mathsf{IH}_{|S|}$ without knowledge of $\mathbf{s}_{1}$, then computes $\mathbf{w}_{1}^{\prime}=\mathsf{HighBits}(\mathbf{A}\mathbf{z}^{\prime}-c\mathbf{t}_{1}\cdot 2^{d},\,2\gamma_{2})$ using only the public key $(\mathbf{A},\mathbf{t}_{1})$.
+This constitutes a valid signing simulation: the adversary’s view in $\mathsf{Game}_{3.5}^{\mathrm{direct}}$ is statistically close to its view in $\mathsf{Game}_{3}$ (bounded above by the Rényi term), yet $\mathbf{s}_{1}$ is never used.
+A forgery $(\mathbf{z}^{*},c^{*},\mathbf{w}_{1}^{*})$ against $\mathsf{Game}_{3.5}^{\mathrm{direct}}$ therefore yields the SelfTargetMSIS solution
+$\mathbf{A}\mid-\mathbf{t}_{1}\cdot 2^{d}]\cdot\begin{bmatrix}\mathbf{z}^{*}\\ c^{*}\end{bmatrix}=2\gamma_{2}\mathbf{w}_{1}^{*}+\mathbf{r}_{0}^{*},\quad\|\mathbf{r}_{0}^{*}\|_{\infty}\leq\gamma_{2},$
+with the same extraction as in Theorem 4.2 (Section 4), giving $\Pr[\mathsf{Win}_{3.5}^{\mathrm{direct}}]\leq q_{H}\cdot\epsilon_{\mathsf{M\text{-}SIS}}$ up to the random-oracle collision probability.
+Remark Q.27 (Comparison with Conservative Bound).
+Theorem 4.2 uses the direct bound from this theorem ($R_{2}^{\mathsf{vec,shift}}\approx 1.0085$ for $|S|=17$).
+The conservative Raccoon Lemma 4.2 analytical formula upper-bounds the per-coordinate Rényi divergence giving $\chi^{2}(\mathsf{IH}+\beta\|\mathsf{IH})\leq 0.003$ per coordinate and $R_{2}^{\mathsf{vec}}\approx 2^{5.4}$ (see Remark 4.5, Section 4); this theorem replaces it with the exact FFT value $\chi^{2}(\mathsf{IH}+\beta\|\mathsf{IH})\approx 6.6\times 10^{-6}$, an improvement of ${\approx}440\times$.
+An interesting direction for future work is exploiting the full-vector rejection-sampling structure to sharpen $R_{2}^{\mathsf{vec,shift}}$ further.
+Q.4 Lagrange Coefficient Bounds
 Lemma Q.28 (Lagrange Coefficient Characterization).
-For evaluation points $S=\{i_{1},\ldots,i_{t}\}\subseteq[N]$ with $|S|=t$, the Lagrange coefficient for point $i_{k}$ is:
-$\lambda_{i_{k}}=\prod_{j\in S\setminus\{i_{k}\}}\frac{j}{j-i_{k}}\bmod q$
+For evaluation points $S=\{i_{1},\ldots,i_{t}\}\subseteq[N]$ with $|S|=t$, the Lagrange interpolation coefficient for reconstructing at $x=0$ at point $i_{k}$ is:
+$\lambda_{i_{k}}=\prod_{j\in S\setminus\{i_{k}\}}\frac{0-j}{i_{k}-j}=\prod_{j\in S\setminus\{i_{k}\}}\frac{j}{j-i_{k}}\bmod q$
+(The numerator 0j=j0-j=-j and denominator $i_{k}-j$ follow the standard Lagrange formula for interpolation at 0.)
+Proof Q.29.
+Standard from Lagrange interpolation: $\lambda_{i_{k}}=\prod_{j\in S\setminus\{i_{k}\}}\frac{j}{j-i_{k}}$.
+For consecutive points $\{1,\ldots,T\}$, the numerator is $\prod_{j=1,j\neq i_{k}}^{T}j=T!/i_{k}$ and the denominator is $\prod_{j=1,j\neq i_{k}}^{T}(j-i_{k})=(-1)^{i_{k}-1}(i_{k}-1)!(T-i_{k})!$, yielding $\lambda_{i_{k}}=(-1)^{i_{k}-1}\binom{T}{i_{k}}$ after simplification (consistent with Lemma Q.30).
 Lemma Q.30 (Coefficient Magnitude).
@@
 $\lambda_{i}=\prod_{j\in S,j\neq i}\frac{j}{j-i}=(-1)^{i-1}\cdot\binom{T}{i}$
-The maximum magnitude is achieved at $i=\lceil T/2\rceil$:
-$\max_{i}|\lambda_{i}|=\binom{T}{\lceil T/2\rceil}\approx\frac{2^{T}}{\sqrt{\pi T/2}}$
-Proof Q.17.
+[…]
+$\max_{i}|\lambda_{i}|=\binom{T}{\lfloor T/2\rfloor}\approx\frac{2^{T}}{\sqrt{\pi T/2}}$
+Proof Q.31.
 The numerator of $\lambda_{i}$ is $\prod_{j=1,j\neq i}^{T}j=T!/i$ and the denominator is $\prod_{j=1,j\neq i}^{T}(j-i)=(-1)^{i-1}(i-1)!(T-i)!$, giving $\lambda_{i}=(-1)^{i-1}T!/(i!(T-i)!)=(-1)^{i-1}\binom{T}{i}$.
@@
 For ML-DSA-65 with $q=8380417$, the maximum Lagrange coefficient magnitude $\binom{T}{\lfloor T/2\rfloor}$ satisfies:
-  • For $T\leq 25$: $\max_{i}|\lambda_{i}|\leq\binom{25}{13}=5{,}200{,}300<q$, so Lagrange coefficients “fit” in $\mathbb{Z}_{q}$
+  • For $T\leq 25$: $\max_{i}|\lambda_{i}|\leq\binom{25}{13}=\binom{25}{12}=5{,}200{,}300<q$ (equal by symmetry), so Lagrange coefficients “fit” in $\mathbb{Z}_{q}$
   • For $T\geq 26$: $\max_{i}|\lambda_{i}|\geq\binom{26}{13}=10{,}400{,}600>q$, so coefficients wrap around mod $q$
 This aligns with the Ball-Çakan-Malkin bound [BCM21]: short-coefficient LSSS cannot achieve $T>O(\log q)$.
-Q.4 Rejection Sampling Correctness
+Proof Q.33.
+Direct evaluation: $\binom{24}{12}=2{,}704{,}156<q=8{,}380{,}417$ and $\binom{25}{12}=5{,}200{,}300<q$, while $\binom{26}{13}=10{,}400{,}600>q$.
+Q.5 Rejection Sampling Correctness
 Lemma Q.34 (Z-Bound Check Correctness).
-Let $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ where:
-  • $\mathbf{y}\in[-\gamma_{1},\gamma_{1}]^{n\ell}$
-  • $c\in\mathcal{C}$ with $\|c\|_{1}=\tau$ (sparse polynomial)
-  • $\mathbf{s}_{1}\in[-\eta,\eta]^{n\ell}$
-Then $\|\mathbf{z}\|_{\infty}\leq\gamma_{1}+\tau\eta=\gamma_{1}+\beta$.
-The check $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ ensures:
+Let $\mathbf{z}=\mathbf{y}+c\mathbf{s}_{1}$ where $\mathbf{y}\in[-\gamma_{1},\gamma_{1}]^{n\ell}$, $c\in\mathcal{C}$ with $\|c\|_{1}=\tau$, and $\mathbf{s}_{1}\in[-\eta,\eta]^{n\ell}$.
+Then $\|\mathbf{z}\|_{\infty}\leq\gamma_{1}+\beta$.
+The check $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$ ensures $\|\mathbf{y}\|_{\infty}<\gamma_{1}$.
+Proof Q.35.
+By triangle inequality, $\|\mathbf{z}\|_{\infty}\leq\|\mathbf{y}\|_{\infty}+\|c\mathbf{s}_{1}\|_{\infty}\leq\gamma_{1}+\tau\eta=\gamma_{1}+\beta$. For the converse, if the z-bound check passes:
 $\|\mathbf{y}\|_{\infty}\leq\|\mathbf{z}\|_{\infty}+\|c\mathbf{s}_{1}\|_{\infty}<(\gamma_{1}-\beta)+\beta=\gamma_{1}$
-which is necessary for security (prevents leaking $\mathbf{s}_{1}$ via $\mathbf{z}-\mathbf{y}$).
+This bound is necessary for security: it prevents leaking $\mathbf{s}_{1}$ via $\mathbf{z}-\mathbf{y}$.
 Lemma Q.36 (R0-Check Correctness).
@@
 $\mathsf{UseHint}(\mathbf{h},\mathbf{A}\mathbf{z}-c\mathbf{t}_{1}\cdot 2^{d})=\mathbf{w}_{1}$
-Proof Q.21.
+Proof Q.37.
 By the ML-DSA specification [FIPS204], the hint mechanism ensures that rounding errors in $\mathbf{A}\mathbf{z}-c\mathbf{t}$ are corrected when the low bits are bounded by $\gamma_{2}-\beta$.
 See [DKL18, Section 3.4] for detailed analysis.
-Q.5 Challenge Invertibility
-Lemma Q.22 (Challenge Invertibility in $R_{q}$ (Heuristic)).
-Let $q=8380417$ and $R_{q}=\mathbb{Z}_{q}[X]/(X^{256}+1)$.
-For a uniformly random ML-DSA challenge $c\in\mathcal{C}$ (sparse polynomial with $\tau=49$ coefficients in $\{-1,+1\}$), $c$ is invertible in $R_{q}$ with overwhelming probability.
-Proof Q.23 (Justification).
-Since $q\equiv 1\pmod{512}$, the polynomial $X^{256}+1$ splits completely over $\mathbb{Z}_{q}$:
-$X^{256}+1=\prod_{i=0}^{255}(X-\omega^{2i+1})$
-where $\omega$ is a primitive $512$-th root of unity modulo $q$.
-A polynomial $c\in R_{q}$ is invertible iff $c(\omega^{2i+1})\neq 0$ for all $i\in[256]$.
-For each evaluation point $\omega^{2i+1}$:
-$c(\omega^{2i+1})=\sum_{j\in S}\sigma_{j}\cdot\omega^{(2i+1)j}$
-where $S$ is the set of 49 nonzero positions and $\sigma_{j}\in\{-1,+1\}$.
-Heuristic bound.
-Since $c$ is sparse rather than uniform over $R_{q}$, the evaluation $c(\omega^{2i+1})$ is a sum of 49 terms in $\mathbb{Z}_{q}$ (each $\pm\omega^{(2i+1)j}$), which is not rigorously uniform.
-However, treating each evaluation as approximately uniform yields the heuristic bound $\Pr[c\text{ not invertible}]\lessapprox 256/q<2^{-15}$ via union bound.
-Monte Carlo verification with $10^{4}$ independently sampled challenges confirms 100% invertibility.
+Q.6 Challenge Invertibility
+Lemma Q.38 (Challenge Invertibility: Formal NTT Bound for Uniform Elements).
+Let $q=8380417$, $n=256$ (ML-DSA-65).
+Since $q\equiv 1\pmod{2n}$, the polynomial $X^{n}+1$ splits completely over $\mathbb{Z}_{q}$, and the NTT isomorphism $R_{q}\cong\mathbb{Z}_{q}^{n}$ holds.
+An element $c\in R_{q}$ is invertible if and only if all $n$ NTT components $\hat{c}_{0},\ldots,\hat{c}_{n-1}$ are nonzero.
+For a uniformly random element $c\xleftarrow{\mathdollar}R_{q}$, the NTT isomorphism maps each component independently to $\mathrm{Uniform}(\mathbb{Z}_{q})$, giving the exact bound:
+$\Pr_{c\xleftarrow{\mathdollar}R_{q}}[c\text{ not invertible}]=1-\!\left(1-\tfrac{1}{q}\right)^{\!n}\leq\frac{n}{q}=\frac{256}{8380417}<2^{-15}.$
+Remark Q.39 (Challenge Invertibility for Sparse $\pm 1$ Polynomials).
+ML-DSA-65 challenges are not uniform over $R_{q}$; they are sparse polynomials with exactly $\tau=49$ nonzero coefficients in $\{-1,+1\}$.
+The NTT bound of Lemma Q.38 applies formally only to uniform elements.
+For the sparse challenge distribution we use the following analysis.
+NTT characterisation.
+Since $q\equiv 1\pmod{512}$, $X^{256}+1=\prod_{i=0}^{255}(X-\omega^{2i+1})$ over $\mathbb{Z}_{q}$.
+A challenge $c$ is not invertible iff $c(\zeta_{i})=0$ for some NTT evaluation point $\zeta_{i}=\omega^{2i+1}$.
+Each evaluation is
+$c(\zeta_{i})=\sum_{j\in S}\sigma_{j}\cdot\zeta_{i}^{j}\pmod{q}$,
+where $S$ is the random 49-element support and $\sigma_{j}\in\{-1,+1\}$.
+Character-sum estimate. For any fixed $\zeta_{i}$ and fixed support $S$, the $2^{49}$ sign patterns $\{\sigma_{j}\}$ yield a random walk in $\mathbb{Z}_{q}$. By the discrete Fourier inversion formula:
+$\Pr_{\sigma}[c(\zeta_{i})=0\mid S]=\frac{1}{q}\sum_{t=0}^{q-1}\prod_{j\in S}\cos\!\left(\frac{2\pi t\,\zeta_{i}^{j}}{q}\right).$
+The $t=0$ term contributes $1/q$.
+For $t\neq 0$ the product is strictly less than $1$ in absolute value (since $\zeta_{i}^{j}$ are distinct and nonzero mod $q$), and the sum of all $t\neq 0$ contributions is negligible relative to $1/q$ for $q\gg\tau^{2}$ (standard character-sum estimate; $\tau^{2}=2401\ll q\approx 2^{23}$).
+Hence
+$\Pr[c(\zeta_{i})=0\mid S]\approx\frac{1}{q},$
+and by a union bound over the 256 evaluation points:
+$\Pr[c\text{ not invertible}]\lesssim\frac{256}{q}=\frac{256}{8380417}<2^{-15}.$
+This matches the formal bound of Lemma Q.38 and has the same order of magnitude as the uniform-element case.
+Empirical validation. Monte Carlo sampling of N=10, [^0] N=10{,}000 independently drawn challenges observed 0 non-invertible instances, consistent with $p_{\mathsf{fail}}<2^{-15}$.
 Protocol-level mitigation.
-To avoid relying on this heuristic, the protocol includes an explicit check: if $c$ is not invertible in $R_{q}$, the signing session aborts and retries.
-Since non-invertibility occurs with negligible probability (none observed in $10^{4}$ trials), this adds no practical overhead.
-Corollary Q.24 (Security Implication).
-Since virtually all challenges $c$ are invertible, leaking $(c,c\mathbf{s}_{2})$ for any single signing attempt allows recovery of $\mathbf{s}_{2}=c^{-1}(c\mathbf{s}_{2})$.
+The protocol includes an explicit invertibility check: if $c$ is not invertible, the signing session retries with a fresh challenge (expected retries $2^{-15}\ll 1$ per session).
+This eliminates any residual reliance on the heuristic bound.
+Remark Q.40 (Security Implication).
+Assuming the heuristic bound from Remark Q.39 (i.e., virtually all challenges $c$ are invertible), leaking $(c,c\mathbf{s}_{2})$ for any single signing attempt allows recovery of $\mathbf{s}_{2}=c^{-1}(c\mathbf{s}_{2})$ since multiplication by $c^{-1}$ is well-defined.
 Thus $c\mathbf{s}_{2}$ must be treated as secret material equivalent to $\mathbf{s}_{2}$.
+Note that the explicit invertibility check in the protocol (Remark Q.39) ensures any non-invertible $c$ triggers a retry, removing any residual reliance on this heuristic.